A Complete Bibliography of International Conferences on Very Large Data Bases

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

06 November 2012 Version 2.10

Title word cross-reference

0 [670]. 1 [1248]. 1: N [1290]. 2 [1126]. 20 [1171]. 20/20 [1511]. 80 [1171]. ^{2} [1572]. ^{Π} [1092]. A [1488]. K [1402]. L_{p} [1476]. N [1403, 1125]. ± [567].

-Based [1037]. -Complete [670]. -D [1126]. -Dimensional [1125]. -Safe [1248]. -Tree [637, 1092, 1488]. -Trees [515].

1.0 [1859]. 10-Year [1464]. 11th [1892]. 13th [1894]. 14th [1909, 1895]. 16th [1897]. 18th [1899]. 19th [1900].

2002 [1910]. 2003 [1911]. 2020 [1511]. 20th [1901]. 21st [1902]. 26th [1908]. 29th [1911]. 2nd [1882]. 2PL [502]. 2Q [1012].

3 [691, 1718]. 370 [46]. 3W [1443].

548-559 [1805].

7.0 [1387]. 7.3 [1198]. 7.5 [1386].

’86 [1893, 556].

’92 [1899]. ’93 [1900]. ’95 [1902]. 9th [1890].

= [1773].

A-TOPSS [1750]. Abduction [784].

abridged [1761]. Abstract

[943, 1542, 189, 447, 125, 849, 154, 554, 810, 830, 797, 450, 198, 396, 1222, 418, 432, 191, 119, 576, 173, 118, 821, 421, 1223, 1282, 158, 393, 569, 1600, 1538, 1539]. Abstraction [255, 1044, 9, 1028, 58]. Accelerating [1623]. Acceleration [1614, 1732, 1867]. Accelerator [1032]. Acceptable [743]. Access [23, 1185, 1367, 1653, 1873, 469, 360, 696, 1259, 101, 766, 830, 93, 267, 731, 662, 15, 998, 914, 49, 1720, 368, 92, 686, 51, 926, 1482, 326, 755, 1690, 1030, 1834, 1320, 1495, 615, 98, 1135, 669, 779, 1217, 58, 457, 1582, 1825, 1149, 1682, 1580]. Accesses [541, 1793]. Accessibility [1682]. Accessible [1564]. Accessing [347, 401, 1111, 1101, 995]. Accommodating [491]. Accounts [499]. Accurate [1697, 1593, 1835]. Accurately [1251]. Achievement [110]. Achieving [704]. Acquired [1648]. Across [1493, 1244, 1735, 1211]. ACS [419]. Active [1374, 1436, 1745, 826, 1420, 1026, 810, 825, 882, 946, 838, 948, 329, 390, 1287, 729, 1749, 839, 881, 1103, 1855]. ACTIVIEW [1621]. Activities [802, 640]. Activity [29, 768, 1278, 896, 1062, 1603]. Actual [1648]. Acyclic [303, 302, 451]. Ada [566]. ADABAS [1347]. ADAPT [218, 1199]. ADAPT/X [1199]. Adaptable [1299, 1646, 750, 618, 1267]. Adaptation [1808]. Adaptive [1819, 531, 938, 947, 1794, 795, 772, 808, 1790, 1067, 1621, 891, 960, 1806, 1562, 1469, 1249, 1677, 1375, 1468, 746]. ADD [187]. Adding [923]. Additional [1535]. Addressable [510, 482]. Addressing [268]. Adjustment [1254]. Administration [654, 54, 230]. ADMS [567]. ADSM [1096]. Advanced [1738, 755, 634, 1862, 1417, 714, 1543, 446, 1718]. Advantage [1339]. Aerospace [1730]. Affinity [555]. After [940, 1084]. Against [365]. Agencies [968]. Agent [1739, 1622]. Agent-based [1622]. AgFlow [1622]. Aggregate [1490, 257, 1122, 1612, 543, 1547, 1076, 1809, 711, 860, 1709, 679, 1364, 1813]. Aggregate-Query [1076, 1813].

Aggregates [1188, 605, 506, 1710, 1322, 753, 1496, 712, 1366, 1189, 1456, 609]. Aggregation [1348, 1711, 1824, 1833, 1708, 1424, 1454, 1425, 1172, 843, 1075]. Agora [1502]. Ahead [820, 717, 1326]. Aided [631, 370]. AIM [404]. Alert [839]. Algebra [462, 1443, 713, 1647, 732]. Algebraic [811, 1015, 1815, 572, 588, 384, 184, 1010, 463, 595, 911]. Algebras [422]. Algorithm [1251, 310, 1444, 1667, 947, 1290, 927, 695, 625, 789, 1012, 236, 1665, 1708, 694, 829, 1319, 1485, 1082, 452, 548, 746, 451]. Algorithmic [204]. Algorithms [611, 674, 758, 1016, 275, 442, 639, 498, 836, 742, 1576, 1301, 749, 121, 1178, 1650, 978, 1255, 1248, 671, 645, 1315, 867, 1129, 1424, 511, 747, 549, 1314, 689, 860, 1373, 1328, 995, 1234, 1018, 379, 561, 1230, 302, 660]. ALIAS [1749]. alignment [1835]. Allocation [938, 816, 52, 939, 1700, 1466, 1634, 412]. Almost [1701, 1588, 1823]. Almost-Serial [1701]. alone [1276]. Along [1291]. Alpha [436]. Alternating [959]. Alternative [1614, 740, 165, 890]. Ambient [1538]. Ambiguity [238]. America [1339]. Among [1807, 582, 195]. Amsterdam [1896]. Analogical [570]. Analogue [932]. Analysis [943, 192, 287, 1015, 484, 667, 1500, 1669, 33, 1186, 428, 632, 254, 1125, 1752, 1301, 4, 124, 438, 509, 173, 298, 1362, 17, 329, 744, 190, 102, 829, 925, 1742, 1796, 1841, 1176, 1208, 1298]. Analytic [413, 292, 14]. Analytical [620, 1600, 1425, 312]. Analyzing [1548, 1580, 29]. Anatomy [657]. Angeles [1895]. AniPQO [1823]. Anomalies [603]. Anonymity [1281]. ANSI [279, 117, 150]. ANSI/SPARC [117, 150]. Answer [1214]. Answering [1457, 181, 1130, 1563, 589, 1485, 1172]. Answers [1439, 1001, 1593, 1377, 1179, 706, 792]. Any [1101]. Anymore [1533]. Anyone [581]. Anyway [1430]. Application [1327, 1517, 192, 31, 197, 296, 336, 626, 185, 1116, 260, 768, 326, 295, 1732, 24, 655, 1226, 74, 335, 664, 1754, 546, 242, 1862, 1417, 388, 1183, 1866, 1467, 430, 62, 282, 557, 1821]. Applications [575, 1598, 904, 1276, 857, 1509, 1684, 1420, 1660, 1421, 1464, 1880, 630, 547, 374, 502, 685, 629, 1747, 1795, 83, 1690, 1722, 1847, 1867, 604, 1520, 1641, 516, 738, 204, 1137, 902, 912, 264, 87, 446, 709]. Applied [243, 112]. Applying [1096, 951, 1169]. Approach [486, 751, 1203, 218, 1015, 1559, 449, 405, 387, 475, 492, 279, 493, 345, 572, 1449, 354, 605, 889, 824, 1104, 1758, 113, 383, 1308, 619, 957, 903, 103, 759, 948, 977, 1780, 739, 432, 1706, 1335, 1810, 413, 316, 896, 1067, 1086, 425, 1102, 952, 1625, 1029, 61, 284, 994, 1492, 416, 933, 842, 944, 905, 863, 239, 595, 1764, 1318, 1796, 804, 259, 409, 1257, 1543, 446, 1235, 324, 613, 1048, 70]. Approaches [1693, 897, 1477]. Approximate [1439, 1451, 1493, 1457, 1638, 1263, 1547, 1588, 1130, 1833, 1750, 1671, 1304]. Approximating [1490]. Approximation [1824, 1377, 1488, 1329]. Approximations [1299]. Aqua [1439]. AQuery [1788]. Arbitrary [1476, 1797]. Architectural [85, 77, 476]. Architecture [1346, 1345, 1380, 586, 1367, 494, 1387, 1441, 702, 36, 73, 954, 1587, 173, 1089, 1616, 1874, 1518, 21, 1243, 567, 96, 11, 839, 106, 187, 1375, 1440, 1605]. Architectures [667, 1689, 313, 740, 908, 907, 690, 1160, 1755, 1211, 1608]. Archival [92, 81]. Archives [1241]. Archiving [1498]. Area [969, 1278]. Argument [699]. ARIEL [512]. ARIES [764, 1363, 717]. ARIES/KVL [764]. ARIES/NT [717]. Array [24, 301, 1135]. Arrays [743, 1221, 1053, 744, 1137]. Art [1537, 821, 1041]. ASE [1542]. Asera [1518]. Aspects [128, 1145, 144, 1392, 180]. Assembly [368]. Assertions [257]. Assessment [348, 342]. Assignment [239].

Assignments [750]. Assistant [1498, 1632]. Assisted [368]. Associated [1754]. Association [1016, 1143, 1081, 1140, 1314, 1313, 1699, 1082, 1080, 1376, 1142, 569]. Associative [24, 87]. Assumption [1265]. Astronomical [1126]. Asynchronous [1690, 1319]. Athens [1905]. ATLAS [1868]. Atmospheric [1155]. Atomic [879]. Atomicity [1281]. Attachment [785]. Attribute [1712, 1072, 903, 736, 38, 1265]. Attribute-Oriented [903]. Attributes [541]. Auditing [123]. Augmented [752]. August [1900, 1896, 1895, 1910, 1891, 1906, 1905, 1897, 1892, 1899]. Australia [1897]. Authentic [1604]. Author [1534]. Authority [1791]. Authorization [914, 346, 235]. Authorizations [1185]. Automata [1780]. Automata-based [1780]. Automated [1733, 1486, 981, 524, 457]. Automatic [1386, 1815, 1550, 1789, 323, 16, 459, 248, 13, 483]. Automating [12, 832]. Automation [1521, 1756]. Automotive [1741]. Autonomic [1715]. Autonomous [1089, 879, 841]. Autonomously [579]. Autonomy [654, 993]. Auxiliary [26]. Availability [352, 1198, 1085, 940]. Available [1380, 1584, 1515, 807]. Avoidance [891, 1319]. Avoidance-Based [1319]. Avoiding [1070, 121, 1828]. Award [1464]. Away [1066]. Axiomatic [196]. Axioms [590]. Axis [1803].

B

[886, 885, 735, 1114, 325, 764, 515, 710, 995]. B-Tree [885, 764]. B-Trees [735, 1114, 325, 995]. B2B [1518]. Back [299]. Back-End [299]. Backed [1558]. Backend [551]. Backout [214]. Backup [941]. Bad [1473]. Baes [830]. Bag [811]. Balanced [88]. Balances [1804]. Balancing [1182, 1792, 829, 1183, 925, 1079, 1376]. Bandwidth [1687]. Bank [1339, 1338]. BANKS [1744]. Barcelona [1898]. Base [186, 143, 299, 192, 339, 426, 130, 321, 163, 69, 101, 45, 296, 313, 144, 122, 227, 184, 254, 66, 972, 174, 501, 126, 22, 29, 314, 2, 99, 55, 185, 4, 111, 183, 331, 260, 556, 160, 168, 95, 74, 59, 208, 232, 335, 10, 156, 109, 16, 17, 534, 61, 108, 81, 213, 165, 388, 230, 77, 328, 330, 96, 366, 11, 779, 312, 39, 132, 342, 278, 13, 264, 62, 320, 131, 133, 82, 162, 327, 112, 249, 288, 319, 135, 136, 50, 324, 134, 41, 1886]. Based [1064, 1330, 675, 375, 1875, 1853, 275, 1395, 1311, 493, 1620, 760, 1024, 1444, 498, 1470, 46, 1657, 1308, 1037, 1297, 1072, 258, 1377, 877, 789, 196, 668, 38, 1315, 1383, 1241, 875, 747, 1805, 1506, 1661, 1558, 1450, 1854, 952, 1200, 628, 1029, 1862, 1303, 170, 1319, 1492, 1247, 1829, 466, 717, 899, 1818, 1796, 1841, 282, 1767, 955, 391, 532, 585, 1468, 472, 555, 709, 775, 1606, 1702, 1667, 1758, 1178, 1780, 1625, 1572, 1635, 529, 974, 1098, 1816, 1622]. Bases [1909, 225, 120, 1910, 252, 849, 1901, 159, 31, 1898, 28, 44, 971, 200, 1902, 334, 1908, 35, 311, 1911, 217, 245, 138, 124, 577, 1905, 820, 755, 859, 954, 1011, 668, 1881, 1423, 203, 119, 350, 54, 9, 562, 21, 281, 922, 155, 1890, 123, 25, 47, 1904, 1886, 340, 483, 528, 1883, 1900, 1896, 1895, 1887, 359, 1891, 1885, 1893, 1882, 1897, 1892, 1894, 1889, 1884, 1899, 1888, 1046]. Basic [399]. Basis [623, 530, 213, 732]. Be [1453, 1030]. BEA [1843]. BEA/XQRL [1843]. Became [1204]. Behaved [420]. Behavior [33, 896, 1580]. Behavioral [984]. Behaviour [604]. Behavioural [297]. Belgium [1882]. Bellcore [1199]. Benchmark [1336, 1726, 1821]. Benchmarking [387, 1097]. Berlin [1911, 1884, 1886]. Between [1132, 61, 677, 412, 41]. Beyond [837, 1363]. BHUNT [1815]. BibFinder [1864]. BibFinder/StatMiner [1864]. Bibliographic [335, 67]. BigSur [1115]. Binary [104, 191, 43, 1314]. Binding [1306]. Bio [1841]. Bio-Molecular [1841].

Biodiversity [1523, 1525, 1532, 1524]. Bioinformatics [1526, 1642, 1527]. Biological [1059, 1553, 1150, 1835]. Biotech [1609]. Bisimulation [1773]. Bit [452, 526]. Bit-Sliced [452]. Bitemporal [1311]. Bitmap [1391]. Bitmaps [1471]. Black [969, 1432]. Blackboard [954]. Block [668, 1053, 190, 541]. blocking [1667]. Bo [132]. Bombay [1904]. Bonananzas [1212]. Boolean [238, 859]. Bottleneck [1367]. Bottlenecks [1212]. Bottom [762, 1810, 761, 1799]. Bottom-Up [762, 1810, 761, 1799]. Bound [1074, 504]. Bounding [1452]. Boyce [272, 361]. Boyce-Codd [272, 361]. Brain [1742]. Braking [1295]. Branched [1482]. Brazil [1885, 230]. Breadth [1256]. Breadth-First [1256]. Breaking [1411]. Brick [1612]. Bridges [1132]. Bridging [1354]. Brighton [1894]. Brisbane [1897]. Broadcast [1175, 1596, 1372, 1249]. Broadening [1514, 1513]. Broker [1052]. Brokering [1781]. Browser [1156, 349]. Browsing [1744, 1498, 647]. Brussels [1882]. BT [1482]. BT-tree [1482]. Bubble [163]. Bubbles [1797]. Bucket [708, 748]. Buddy [779]. Buddy-Tree [779]. Buffer [1279, 892, 938, 496, 707, 766, 816, 789, 1012, 634, 1708, 1069, 1013, 372, 615, 1347]. Buffering [1321, 1396, 559, 1011, 1326, 1793]. Build [1456]. Building [709, 1509, 1037, 977, 1138, 1316, 1435, 391, 1398, 1536]. Bulk [1327, 1838, 1257, 1585, 985, 1048]. Bulk-Inserts [1838]. Bulk-Loading [1327]. Bureau [27]. Bushy [951]. Business [1521, 1537, 1519, 1555, 1639, 1640, 1520, 1716, 1869, 1613]. Bypassing [1065].

C [775, 1037, 1572]. C-based [775]. Cache [1556, 1819, 1557, 1397, 1794, 1197, 1618, 1623, 1368, 1319, 815, 1369, 1127, 1713, 1018]. Cache-Conscious [1557, 1794]. Cached [741]. CachePortal [1867]. Caches [1618].

Caching [1689, 1173, 1614, 962, 1690, 1732, 1558, 1641, 1319, 1375, 1458, 1765]. Cactis [742]. CAD [467, 486, 487, 494, 565, 630, 364, 466]. CAD/CAM [466]. CAD/VLSI [467]. Cairo [1908]. Calculus [244, 250, 382]. Calibrating [1177]. Call [1417]. Call-Level [1417]. Calls [932]. CAM [466]. Can [1352, 1030, 1413, 1382]. Canada [1887, 1899]. Cannes [1888]. Capabilities [986, 1475, 1242]. Capability [104, 65]. Capacity [474, 920]. Caprera [1278]. Capture [713]. Capturing [1392, 896, 1842, 862]. Card [1605]. Cardinality [1814]. Cares [1040]. Cascade [479]. Case [818, 1339, 716, 1492, 230, 1678, 1223]. Case-Based [1492]. CASSM [25, 87]. Catalog [233, 926]. Catalonia [1898]. Categorical [1308, 1314]. Category [160]. Causal [1332]. Cause [569]. Cause-Effect [569]. CB [453]. CB-trees [453]. Cellular [301, 25]. Censoring [1120]. Census [50, 27]. Center [1867]. Center-Hosted [1867]. Centered [906, 1155]. Centralization [1104]. Centralized [400]. Centric [1722, 1597, 1760, 1880]. Certification [468]. Chain [1241, 752]. Challenge [1519, 1484, 1532, 876, 1031]. Challenges [1207, 1610, 1431, 972, 1043, 1275, 964, 1730]. Chameleon [1861]. Champagne [1748]. Change [1685, 631, 1597, 1510, 1748, 873]. Change-Centric [1597]. Changes [62]. Changing [1118, 1790, 193]. Channel [1616]. Characteristics [1818, 367, 445]. Characterization [1162, 420]. Charge [1352]. Checking [530, 244, 591, 377, 1187, 1226, 844, 437, 256]. Checkpointing [1344]. Checks [1804]. CheeTah [1460]. Chile [1901]. China [1910]. Chip [1653, 1873]. Chip-Secured [1653, 1873]. Choice [443, 18]. Choices [363]. Choosing [580, 57]. CIM [974]. City [1889]. Class [310, 1660, 1398]. Classical [704]. Classification [192, 1697, 1475, 1068, 483]. Classifier [904, 1316, 1191]. Classifiers [1398]. Classroom [1501]. Clause [659]. Cleaning [1576, 1577]. Client [1049, 1022, 908, 907, 962, 1319, 1494, 955, 1375, 741]. Client-Based [955]. Client-Cached [741]. Client-Server [1049, 1022, 908, 907, 962]. Clinical [1147]. Closed [642]. Closest [1572]. Closure [611, 758, 760, 759, 927, 573, 671, 612]. Closures [698]. Clues [80]. Cluster [4, 1705, 411, 1376]. Clustered [1840, 1073, 736, 1703]. Clusteredness [1814]. Clustering [1766, 989, 551, 760, 1174, 1309, 547, 1657, 1308, 1411, 894, 1053, 1572, 987, 1318, 1797]. Clusters [1618]. ClustRa [1085]. CLV [833]. CMD [852]. Coalescing [1151, 1152]. Coarse [1800]. Coarse-Grained [1800]. Cockpit [1716]. CODASYL [148, 211, 359, 149, 150, 164, 220, 151]. CODASYL-DBTG-Type [359]. CODASYL-Like [220]. Codd [272, 361]. Coding [668]. Coexistence [100, 41]. cognizant [1056]. Coherence [632]. Coherency [1083, 1197, 809, 1686]. Coherency-Control [809]. Collaborative [1498, 1600]. Collection [1049, 1253, 1648, 1252, 1176, 1624]. Collections [930]. Coloring [1066]. Colours [1432]. Column [658]. COMA [1693]. COMBI [1795]. COMBI-Operator [1795]. Combination [1693]. Combinatorial [89]. Combining [957, 1404]. Command [1273]. Comments [81, 132, 133]. Commerce [1374, 1554, 1737, 1876, 1789, 1282, 1601, 1281, 1623]. Commercial [299, 1116, 1194, 1848]. Commit [961, 446]. Commit LSN [765]. Commitment [879]. Committees [1535]. Common [1109]. Commons [1525]. Communication [1542, 1352, 978, 1066, 582, 896, 694, 61]. Communications [489, 138]. Community [1525]. Commutativity [697]. Compact [1462, 670]. Compacted [373]. Compacting [995]. Company [1353]. Comparative [1614, 945]. Comparing [1370, 1670, 1595]. Comparison [606, 662, 400, 398]. Compensating [739]. Compensation [1297]. Compensation-Based [1297]. Compilers [1195]. Complementary [318, 333]. Complete [692, 1179, 670, 1868]. Completeness [546, 732]. Complex [649, 1560, 981, 45, 470, 687, 97, 733, 1532, 43, 713, 1762, 770, 248, 539, 1272, 527]. Complexity [757, 1247, 524]. Compliant [1262]. Component [66, 286, 566]. Components [332, 1705]. Composing [1807]. Composite [1026, 1070, 1288, 882, 915]. Composition [1736, 115, 1860]. Comprehensive [449, 206]. Compressed [1471, 1856, 1771, 267, 1391, 1424, 215, 561, 1682, 934]. Compression [334, 1034, 1378, 1871, 616, 335, 1570, 1837]. Compromise [365]. Computable [579]. Computation [1188, 573, 1228, 1569, 873]. Computational [1164]. Computations [759, 911]. Computed [1496]. Computer [141, 772, 1215, 631, 748, 582, 10, 370, 108, 21, 1538, 929]. Computer-Aided [631]. Computers [845, 432, 74, 747]. Computing [611, 375, 1637, 1491, 1307, 1714, 969, 508, 661, 612, 1759, 1362, 390, 1799]. Concept [116, 488, 790, 308, 396, 429, 1044, 389, 955]. Concepts [175, 301]. Conceptual [410, 255, 212, 692, 1148, 793, 797, 175, 149, 91, 53, 246, 894, 54, 213, 520, 315, 40, 256, 682]. Concerning [971]. Concurrency [675, 287, 274, 275, 355, 1056, 442, 639, 1557, 497, 397, 632, 718, 1092, 1058, 794, 1057, 234, 963, 764, 515, 621, 398, 395, 306, 995, 1531, 356, 676]. Concurrency-Coherence [632]. Concurrent [553, 191, 226, 30, 1252, 342, 1054]. Conference [1883, 1900, 1535, 1896, 1909, 1907, 1895, 1910, 1901, 1887, 1898, 1891, 1902, 1908, 1911, 1885, 1906, 1905, 1881, 1882, 1897, 1892, 1890, 1894, 1903, 1889, 1904, 1886, 1884, 1899, 1888, 730, 1893]. Confidential [1653]. Configuration [1229]. Configurations [1248, 26]. Configuring [720]. Conflicts [804]. Conjunctive [251, 215]. Connected [119]. Connectionist [1086]. Connectivity [1741]. Conscious [1557, 1794, 1674, 1369, 1018]. Consensus [674]. Consequences [429]. conservation [1603]. Consideration [114]. Considerations [44, 648]. Consistency [984, 530, 767, 228, 1747, 737, 364, 1319, 1373, 741]. Consistent [1205, 1453, 635, 519]. Consolidation [1092]. Constrained [1817, 1571]. Constraint [956, 777, 948, 1187, 459, 922, 529, 627]. Constraints [602, 1815, 530, 1464, 1193, 1481, 1407, 460, 1384, 825, 304, 1824, 737, 599, 220, 459, 706, 844, 1700, 813, 437, 1181, 1445]. Construct [1625]. Constructing [1880, 596, 1143]. Construction [1317, 1233]. Constructors [505]. Constructs [838, 1195]. Consultable [200]. Contain [605]. Contained [606]. Containing [715]. Containment [1473]. Content [1620, 1689, 1024, 1614, 1269, 510, 482, 1468]. Content-Based [1620, 1024, 1468]. Content-Specific [1269]. Contention [1070, 766, 1007, 891]. Context [1395, 1489, 73, 763]. Context-Based [1395]. Contexts [1026]. Continuous [1268, 1801, 1802, 418, 1296, 1069, 1872, 1666]. Continuously [1802]. Contrast [1480]. Control [675, 287, 274, 275, 355, 1185, 1056, 442, 639, 1557, 565, 19, 497, 397, 718, 816, 1058, 957, 795, 794, 1273, 976, 323, 234, 952, 891, 764, 809, 515, 621, 874, 395, 306, 974, 995, 42, 1682, 676]. Controlling [1834, 1528]. Controls [632]. Conventional [602]. Conversational [180]. Conversion [218, 428, 103, 248, 151, 62, 320, 229]. Converter [39]. Cooperating [877, 1686, 709]. Cooperative [830, 691, 634, 738, 1062, 823]. Coordination [1537, 1705]. Coping [524]. CORAL [874, 923]. Correct [305]. Correctly [720]. Correctness [718]. Correlated [541]. Correlation [1507, 1071, 506]. Correspondences [1829]. Cost [1231, 1073, 1178, 1177, 743, 1297, 1656, 1823, 1658, 940, 1303, 1025, 1419]. Cost-Based [1303, 1178]. Cost-Driven [1231]. Cost-Effective [940]. Costly [437]. Costs [34, 1768, 677]. Counting [1253]. Countries [374, 217]. Counts [1671]. Coupled [613]. Coupling [483]. Cover [326]. Cover-Trees [326]. Coverage [1864]. Covering [1773]. CPU [1472]. Crash [475]. Crashes [626, 1084]. Crawler [1459]. Crawling [1489, 1552, 1866]. Creation [1740, 769]. Criteria [53, 75]. Criterion [718]. Critical [1277]. Cross [1622]. Cross-Enterprise [1622]. Crossroads [966]. Cryptography [1834]. CS4 [125]. Cube [1591, 1190, 1707, 1871, 1462]. Cubes [1386, 1426, 1571, 1799]. Cubetree [1355]. Cubing [1799]. Cumulative [1020]. Cupid [1544]. Curio [1433]. Currency [1339]. Current [348, 293, 167, 821]. Curse [1411]. Curve [1404]. Customer [970]. Customizable [1090, 1861]. Customization [1863]. Customizing [1746, 1536]. Cyber [1382]. Cycles [1409]. Cyclic [1253, 495].

D [352, 1207, 866, 1126]. Dalí [979]. Dark [1852]. Data [722, 1285, 1745, 186, 143, 322, 467, 1766, 751, 865, 1554, 1596, 913, 811, 1721, 1251, 943, 1327, 1691, 1132, 1517, 1909, 1856, 299, 1346, 192, 218, 116, 225, 120, 1077, 447, 1271, 417, 339, 1123, 1853, 257, 1910, 426, 1122, 252, 489, 1311, 1901, 1610, 318, 194, 1438, 1653, 1873, 159, 1688, 154, 255, 493, 408, 1115, 1815, 31, 130, 494, 909, 1059, 1612, 321, 1898, 1508, 28, 1660, 1830, 1500, 790, 1421, 827, 1751, 866, 360, 424, 1711, 935, 1659, 163, 1636, 1370, 8, 69, 705, 1669, 1637, 1444, 1240, 44, 856, 1876, 1036]. Data [971, 1573, 1670, 1798, 1550, 1586, 348, 1209, 200, 359, 101, 766, 1173, 45, 1433, 1902, 1431, 296, 1148, 313, 1877, 144, 830, 1519, 1407, 1337, 1571, 122, 227, 1190, 1392, 267, 334, 184, 1250, 1908, 1758, 731, 254, 66, 1309, 100, 1220, 1752, 35, 1564, 15, 80, 311, 972, 1238, 1422, 174, 261, 1911, 501, 1832, 1576, 126, 383, 1324, 1753, 1357, 217, 1308, 1144, 1801, 245, 22, 1455, 636, 1076, 1214, 1244, 910, 29, 1845, 314, 1381, 2, 1794, 99, 55, 32, 185, 97, 198, 1740, 1795, 1033, 988, 4, 111, 103, 138, 772, 845, 1273]. Data [124, 1219, 331, 1643, 1474, 1034, 1218, 1413, 1879, 505, 1905, 1484, 260, 1385, 1443, 438, 808, 556, 733, 196, 104, 820, 179, 1704, 65, 533, 1150, 295, 363, 668, 1881, 38, 56, 385, 1160, 1274, 1729, 748, 160, 86, 436, 1780, 323, 1331, 1710, 1804, 1587, 1878, 1690, 1241, 418, 1849, 1129, 1578, 1792, 1707, 1871, 1739, 1532, 168, 316, 1788, 852, 1424, 1575, 1867, 203, 1674, 95, 663, 119, 1004, 350, 74, 59, 1086, 1805, 208, 232, 335, 431, 1807, 1046, 114, 10, 1671, 1117, 156, 564, 1563, 1570, 1627, 109, 1002, 54, 1834, 1292, 16, 17]. Data [118, 891, 1322, 940, 1202, 329, 1625, 1069, 1029, 1320, 290, 61, 190, 9, 284, 1529, 108, 987, 81, 1864, 1696, 1454, 1312, 213, 518, 1196, 1727, 165, 1460, 973, 540, 21, 281, 416, 1681, 1692, 1837, 453, 787, 1811, 230, 155, 1211, 77, 1429, 1577, 1748, 864, 5, 863, 328, 1287, 330, 283, 1699, 1031, 1243, 1419, 593, 1611, 98, 1469, 1756, 1592, 1528, 96, 366, 1127, 421, 1890, 11, 1726, 123, 779, 819, 1713, 1101, 48, 1223, 312, 1121, 1191, 1686, 1764, 1447, 1565, 161, 523, 1760, 1796, 918, 39, 537, 132, 342, 210, 1841]. Data [1249, 278, 13, 1466, 1634, 78, 912, 264, 25, 62, 320, 932, 1161, 807, 131, 1785, 1531, 1767, 1229, 133, 82, 162, 47, 110, 1861, 1239, 1006, 327, 42, 532, 1527, 393, 112, 1341, 1904, 1169, 270, 249, 288, 846, 1235, 1868, 1870, 319, 135, 136, 340, 50, 324, 741, 18, 483, 1536, 1458, 1230, 569, 134, 555, 1816, 1797, 1672, 41, 528, 63, 1883, 1613, 1619, 1900, 1580, 1896, 1895, 1887, 1880, 1891, 1885, 1784, 1893, 1882, 1621, 1897, 1603, 1892, 1435, 1894, 1889, 1886, 1884, 1899, 1888, 660, 337]. Data-Bases [334]. Data-Centric [1760, 1880]. Data-Contention [891]. Data-Driven [1421, 329]. Data-Intensive [1421, 1536, 1458, 1625]. Data-Parallel [988]. data-sources [1435]. Data-Structure [453]. Data-World [932]. Data/Metadata [1144]. Database [602, 253, 611, 904, 1737, 850, 1819, 943, 1049, 1032, 775, 176, 601, 287, 1380, 594, 967, 172, 1526, 1015, 1232, 1014, 1212, 1077, 1348, 274, 1045, 417, 145, 449, 586, 125, 275, 1042, 1185, 704, 387, 727, 381, 623, 831, 1522, 1701, 1442, 468, 1367, 490, 475, 531, 1182, 492, 818, 1584, 989, 1514, 530, 394, 237, 345, 1096, 442, 538, 1420, 1500, 376, 1110, 495, 883, 84, 424, 935, 696, 303, 1441, 141, 473, 554, 938, 1427, 474, 496, 606, 734, 1193, 986, 1648, 596, 707, 404, 1186, 448, 481, 1340, 197, 247, 810, 552, 740, 1145]. Database [908, 428, 1104, 460, 3, 113, 1108, 374, 997, 1060, 353, 1351, 907, 962, 626, 993, 550, 305, 228, 266, 1177, 825, 996, 1321, 771, 946, 422, 1720, 901, 406, 1588, 1008, 685, 559, 91, 90, 289, 368, 36, 1501, 1093, 1795, 926, 129, 219, 73, 1738, 772, 845, 1553, 1085, 976, 1675, 221, 948, 1215, 1371, 968, 768, 645, 977, 438, 624, 784, 837, 83, 20, 631, 1453, 805, 385, 1850, 1668, 965, 693, 748, 1630, 1768, 191, 440, 75, 1792, 1163, 1408, 85, 724, 180, 870, 1133, 1258, 560, 887, 1147, 1270, 1623, 1867, 338, 24]. Database [604, 663, 1842, 1275, 471, 941, 694, 576, 612, 1558, 226, 1277, 1658, 150, 164, 461, 60, 1078, 1102, 286, 205, 549, 68, 1089, 346, 1651, 1520, 1200, 897, 583, 242, 414, 292, 102, 362, 1529, 364, 635, 1417, 67, 1368, 684, 1111, 1353, 170, 518, 821, 157, 562, 280, 398, 933, 262, 944, 1010, 776, 925, 1079, 520, 1465, 216, 728, 700, 72, 142, 7, 567, 832, 411, 372, 1210, 1159, 536, 1513, 1358, 1818, 188, 306, 770, 1848, 974, 6, 964, 1136, 243, 880, 1757, 524, 841, 248, 79, 590, 1103, 872, 823, 106, 259]. Database [923, 1466, 430, 756, 807, 1208, 1105, 1903, 285, 115, 307, 229, 999, 42, 409, 525, 557, 858, 1543, 1349, 1181, 420, 1717, 446, 585, 774, 1644, 929, 617, 351, 235, 1054, 1039, 302, 570, 240, 1718, 1633, 451, 800, 1354, 1294, 676, 1213]. Database-Backed [1558]. Database-driven [1623, 1867]. Database/Data [385]. Databases [435, 485, 535, 1744, 486, 673, 1016, 1087, 1486, 1654, 1655, 726, 1327, 1548, 806, 407, 1253, 1203, 182, 1907, 1857, 332, 172, 826, 1386, 1151, 1152, 491, 791, 1291, 255, 279, 297, 956, 469, 494, 1119, 1156, 592, 767, 1261, 571, 1026, 347, 1170, 651, 1636, 705, 1055, 276, 1245, 1851, 397, 1201, 1148, 644, 1269, 824, 384, 947, 1342, 788, 834, 888, 914, 317, 1083, 121, 1073, 1178, 1268, 882, 1260, 1284, 304, 642, 434, 1052, 1306, 1497, 1479, 1906, 1000, 1227, 1130, 903, 1081, 1116, 1698, 1831, 838, 1822, 1034, 895, 1583, 1405]. Databases [1574, 236, 1850, 400, 465, 650, 725, 343, 563, 1194, 429, 1158, 1804, 737, 591, 1478, 867, 1849, 1335, 1532, 1187, 413, 365, 1179, 616, 894, 975, 1561, 896, 664, 658, 711, 1117, 597, 803, 603, 369, 389, 701, 1283, 980, 515, 390, 1314, 370, 455, 1224, 844, 545, 1312, 1025, 441, 1647, 1373, 1826, 842, 519, 905, 822, 1225, 1100, 1495, 729, 1137, 712, 868, 871, 1095, 1146, 522, 1082, 703, 1267, 584, 595, 786, 1393, 1318, 841, 367, 445, 1162, 1538, 992, 1549, 1694, 169, 427, 843, 457, 307, 443, 548, 391, 1142, 187, 231, 873, 1106]. Databases [358, 681, 911, 561, 1019, 412, 1041, 884, 672, 709, 680, 1300, 652]. DataBlitz [1356]. Datacomputer [93, 51]. Datacubes [1228]. Datadict [438]. Dataflow [552, 689]. DataGuides [1260].

DataJoiner [1349]. Datalanguage [51]. Datalog [588, 587, 1222]. Datasets [1317, 1315, 1827, 1323]. DataSpot [1340]. Datawarehousing [1432]. Datbase [425]. Date [1882, 388]. DB [1342, 1177, 927, 1844, 900]. DB1 [282]. DB2 [1619, 1840, 1414, 1427, 1108, 1109, 1715, 1196, 1358, 1719, 1541, 1349, 1107]. DBGraph [787]. DBMS [654, 147, 723, 293, 878, 336, 687, 1037, 1058, 1803, 910, 444, 629, 1515, 221, 647, 564, 173, 220, 940, 1069, 466, 839, 881, 27, 566, 152, 966, 282, 162, 231, 900, 70]. DBMS-Based [282]. DBMS-Processable [221]. DBMSs [575, 1390, 512, 1319]. DBTG [359, 110]. DDB [470]. De-Centralization [1104]. De-Clustering [547]. Deadlock [310, 236, 357, 509]. Dealing [1046, 603, 676]. Decade [1119]. Decentralized [235]. Decide [1409]. Decision [1439, 1143, 126, 1317, 1475, 1369, 1316, 207]. Decision-Support [1369]. Declarative [1193, 1576, 957, 769, 896, 459]. Declustering [917, 1055, 771, 852]. Decoding [95]. Decomposing [470]. Decomposition [113, 406, 205, 1161, 115]. Decompositions [1064]. Decompression [368]. Dedicated [32]. Deductive [1014, 571, 200, 683, 65, 591, 1187, 661, 980, 844, 922, 842, 843, 548, 873, 527, 672, 682, 652]. Deduplication [1749]. Deep [853]. Deferred [497, 959]. Deferring [321]. Define [1205]. Defined [1350, 1415, 504, 1410, 663, 664, 454, 1134]. Defining [315]. Definition [381, 15, 580, 382, 210, 1]. Delay [1849]. Delayed [377, 536]. Deletions [1723]. Deliver [1519]. Delivery [1688, 1713]. DELTA [353]. Demand [1596, 1759, 1013, 1025, 1239]. Demonstration [1436, 1499, 1504]. Demonstrations [1535, 1358]. Denodo [1727]. Denotational [382]. Dependability [1821]. Dependencies [578, 921, 601, 641, 715, 1124, 377, 203, 205, 514, 177, 204, 1813, 273]. Dependency [380, 434, 219, 599, 600, 437]. Deploying [1722]. Deployment [1608]. Derivation [653, 333]. Derived [322, 579, 705, 910, 323, 508]. Deriving [777, 851]. Describing [1242]. Description [790, 1333, 160, 209]. Descriptions [1165]. Descriptive [111]. Descriptor [668]. Design [225, 449, 125, 831, 492, 408, 130, 212, 69, 474, 1444, 44, 200, 359, 45, 296, 632, 122, 1157, 793, 113, 499, 174, 175, 126, 1301, 1144, 314, 91, 1381, 4, 1418, 219, 12, 260, 438, 631, 363, 1728, 465, 693, 75, 1258, 316, 24, 663, 232, 286, 298, 16, 980, 1320, 292, 370, 362, 738, 67, 213, 1494, 416, 223, 933, 520, 1192, 567, 832, 315, 1159, 188, 669, 166, 1136, 590, 132, 823, 259, 614, 665, 131, 999, 133, 409, 525, 112, 324, 358, 88, 1230, 70, 134, 682]. Designer [221]. Designing [558]. Designs [37, 34, 662]. Desired [48]. Desk [1090]. Desk-Top [1090]. Desktop [1168]. Destaging [678]. Detecting [579]. Detection [310, 1026, 1685, 236, 201]. Determining [500, 1283]. Deterministic [672]. Developing [1337, 374, 217, 97, 1722, 1624]. Development [456, 577, 260, 768, 582, 1532, 1805, 1494, 230, 72, 243, 340]. Deviants [1371]. Device [49, 1818, 58]. Device-specific [1818]. Devices [1817, 1286, 1538, 1816]. Diabetic [1738]. Diag [1290]. Diag-Join [1290]. Diagram [520]. Diagrama [557]. Dialog [580]. Dialogue [264]. DIAM [48, 58]. Dictionaries [348]. Dictionary [385, 283, 161]. Difference [195]. Differences [1366]. Different [238, 61, 1106, 41]. Differential [1129]. Digital [1606, 1030, 325, 1626]. Dimension [1071, 1125]. Dimensional [1123, 1449, 1669, 1571, 1125, 1293, 1411, 1487, 1400, 1114, 1312, 1488, 637, 1796, 1466, 1298, 1227, 1234, 1841]. Dimensionality [1449, 1411]. Dimensionally [1840]. Dimensions [1412]. Direct [611, 92, 691]. Directed [1712, 1135]. Directions [1514, 84, 1431, 293, 1109, 970, 1513, 1524, 1041, 730]. Directory [347, 44, 1747, 480]. Disaster [749]. DISCO [233]. Disconnection [759]. DISCOVER [1698]. Discovering [1607]. Discovery [1815, 1399, 903, 1081, 1448, 1313, 1549]. Discrete [1389, 1811]. Discrete-Event [1389]. Discriminant [1697]. Discriminants [1261]. Disjoint [396]. Disjoint-Interval [396]. Disjunctive [1306, 1065, 613]. Disk [722, 1251, 666, 1055, 962, 743, 1618, 85, 1053, 744, 1013, 26, 959, 958, 1054]. Disk-Resident [1251]. Disks [1175, 662, 833, 809, 1372]. DISNIC [1200]. DISNIC-PLAN [1200]. Dissecting [1472]. Disseminating [1175, 1764]. Dissemination [1446, 1711, 1051]. Distance [1295, 1315, 1383, 780, 1581]. Distance-Based [1315, 1383]. Distinct [1593, 1072]. Distributed [654, 1855, 186, 143, 322, 1276, 352, 287, 274, 586, 275, 355, 426, 252, 623, 831, 468, 309, 475, 233, 639, 1615, 495, 883, 1399, 424, 76, 354, 473, 554, 276, 44, 19, 630, 254, 353, 1832, 266, 185, 1501, 926, 759, 690, 855, 1675, 236, 1150, 1594, 1160, 1630, 75, 477, 1004, 471, 74, 511, 425, 544, 52, 1038, 286, 980, 1200, 414, 108, 1862, 821, 21, 77, 330, 72, 142, 306, 584, 786, 841, 169, 307, 1787, 162, 187, 1281, 288, 412, 1769, 676]. Distributing [1735]. Distribution [639, 866, 1171, 658, 241, 1183]. Distributions [885]. Distributive [1709]. Diverse [1244, 1419]. DMS [1357]. Do [272, 1430, 1413, 146, 1419]. Document [488, 1045, 1241, 1624, 800, 1154]. Documents [1446, 1836, 1746, 1508, 1568, 1664, 1782, 1777, 1210, 1393, 1447, 1068, 483, 1603]. Does [1390, 393, 272, 1147]. Domain [1648, 1609, 805, 107, 382, 944, 1527]. Domain-Oriented [107]. Domain-specific [1648]. Domain-theoretic [944]. Domains [1467]. Domino [1520]. Domino/Notes [1520]. Done [1414]. Don’t [1453, 1243]. Double [760, 1044]. Down [1442, 1336, 994, 1799]. downfall [1539]. DPLS [42]. DRC [382]. Dream [1600]. Drill [1336]. Driven [1421, 347, 1231, 90, 1404, 329, 1303, 999, 1863, 1623, 1867, 1406, 1872]. DSIS [338]. DTD [1712]. DTD-Directed [1712]. DTL [1340]. Dual [636, 1011]. Dual-Buffering [1011]. Dublin [1900]. due [62]. Duplicate [714, 1051]. Duplicates [1691, 753]. During [489, 1472, 940, 288, 1325]. DVSS [630]. DWMS [1202]. Dynamic [649, 750, 1182, 1689, 866, 1591, 1614, 460, 1680, 976, 1594, 1676, 708, 894, 1450, 939, 1053, 478, 678, 653, 635, 401, 518, 1811, 925, 1079, 1429, 637, 1686, 813, 1376, 42, 1589, 454, 936, 1254]. Dynamical [1308]. Dynamically [554, 1790, 193, 1112]. Dynamics [224].

E-Business [1521]. e-Commerce

[1876, 1601, 1623, 1554, 1789]. E-R [797]. E-Service [1734]. e-Services [1611, 1735]. E.piphany [1530]. Eager [1075]. Earth [1115, 1875]. Ease [984]. Easier [1455]. Easily [1205]. Easy [1497, 1120]. Easy-to-Use [1497]. eBusiness [1755]. ECRINS [556]. ECRINS/86 [556]. Eddies [1787]. Edinburgh [1907]. Edition [1108, 1196]. Education [1501]. Educational [1510]. EEV [280]. Effect [766, 708, 569]. Effective [1014, 1266, 1685, 940, 688, 987, 1154]. Effectively [1864]. Effectiveness [950]. Effects [287, 818, 1472, 846]. Efficacious [1871]. Efficiency [510, 584, 800]. Efficient [673, 751, 865, 1446, 1049, 1840, 1663, 1231, 171, 1779, 1568, 1664, 1259, 986, 1433, 267, 1143, 501, 1650, 1824, 1831, 671, 1574, 357, 506, 1780, 762, 1129, 1114, 1723, 1233, 987, 787, 627, 1082, 669, 779, 1267, 1764, 1166, 992, 152, 1694, 1569, 47, 57, 1543, 319, 1280, 1019, 1765, 1452, 1581, 1682, 884, 1154]. Efficiently [1307, 1478, 1408, 1447, 1325]. Egypt [1908]. Eighth [1910, 1903, 1889]. Electronic [1374, 1737, 1604, 1605, 647, 1281, 1603]. Elevation [1239]. Eliminating [1691]. Elimination [714, 756]. Ellipsoid [1562]. Embedded [606, 1538]. Embedding [78]. empire [1539]. Empirical [1032, 29, 398]. Enabled [1510]. Enabling [1504, 1260, 1587, 1625]. Encipherment [23, 735]. Encoding [95]. ENCOMPASS [309]. Encryption [1873]. End [299, 1209, 146, 179, 512, 1625, 106, 457]. End-User [146, 106]. End-users [1625]. Ended [640, 42]. Ending [1353]. Endowment [1535]. Enduser [66]. Energy [1721, 1277, 1580]. Enforcement [460, 459]. Enforcing [921, 641]. Engine [1522, 1295, 171, 1763, 1362]. Engineering [629, 1215, 465, 389, 1029, 562, 135, 136, 1644, 70, 134]. Engines [1269, 1564, 1673, 1674, 1770, 1651]. England [1894]. English [1386, 827]. English-like [827]. Enhanced [1876, 1268, 390, 1223]. Enhancements [1110, 625]. Enhancing [1410]. ENIAM [692]. Enity [8]. Enity-Relationship [8]. Ensembles [1742]. Enterprise [1516, 1622]. Entire [519]. Entity [339, 683, 556, 222, 599, 869]. Entity-Relationship [339, 683, 556, 599]. enTrans [1747]. Enumeration [757, 1247]. Environment [1461, 704, 565, 456, 1648, 596, 707, 766, 1309, 906, 1116, 1168, 808, 68, 809, 516, 1494, 879, 1736, 1579, 457, 1230]. Environmental [1271]. Environments [375, 849, 883, 1741, 1076, 444, 1219, 855, 390, 1713, 1821, 613, 746]. EOS [1734]. Epicenter [1530]. Epoch [749]. Equations [588, 459]. Equijoin [836]. Equivalence [194, 448, 195]. EROC [1138]. Error [201, 1452]. Escrow [867]. Estimating [1598, 1251, 1071, 1814, 620, 698, 26, 541]. Estimation [250, 1072, 1400, 1833, 1404, 1335, 1679, 658, 1183, 1265, 1189]. EUFID [179]. Evaluability [711]. Evaluate [1389]. Evaluating [1291, 588, 1401, 1402, 444, 574]. Evaluation [638, 127, 1560, 1779, 1683, 496, 313, 836, 1047, 947, 15, 1812, 796, 1255, 945, 927, 794, 89, 762, 633, 544, 344, 891, 280, 761, 320, 1585, 585]. Evaluations [1732]. Event [1507, 1348, 1389, 882, 1593, 703, 999]. Event-Driven [999]. Event-Join [703]. Events [1026, 296]. Everything [1548]. Everywhere [1643, 1217]. Evidence [730]. Evolution [984, 806, 1459, 783, 1060, 73, 1520]. Evolutionary [856]. Evolvable [1273]. Evolving [1766, 1044, 1808]. Exact [1676]. Exactly [1734]. Exactly-Once [1734]. Example [418, 1625, 466, 1]. Example-based [1625]. Examples [1399, 1300]. Exception [1555]. Exceptions [984, 491]. Exclusive [277]. Execution [1817, 1701, 1350, 853, 951, 35, 946, 950, 1770, 847, 688, 1146, 677]. Executive [209]. Exegesis [1420]. Exhibit [1360, 1361]. EXODUS [538]. Expansion [1731]. Expansions [269, 378]. Expensive [1350]. Experience [1628, 1859, 20, 1718]. Experiences [1431, 1342, 906, 647, 1728, 1294]. Experiment [32, 1168]. Experimental [46, 1501, 794, 511]. Experiments [1133, 1788, 1030, 1731, 1757]. Expert [1015, 586, 492, 455, 635, 1866, 393, 483]. Expiring [1324]. Explaining [1366]. Exploiting [978, 1704, 1690, 1139, 941, 1372, 1718]. Exploration [1340, 1871, 1270, 1469, 523, 1866, 1694, 351]. Exploring [1475]. Explosion [1099]. Expressed [244, 250, 319]. Expression [1178, 1384, 436, 459]. Expressions [1598, 1663, 998, 859, 1575, 568, 195, 248]. Expressive [1780, 691]. Extended [578, 1063, 554, 71, 801, 797, 587, 450, 396, 556, 418, 432, 870, 664, 756, 557, 393, 613, 454]. Extender [1619]. Extenders [1456]. Extendible [553, 1561, 480, 1137]. Extending [624, 1618, 828, 1743, 713, 1425, 463]. Extensibility [1410, 1624]. Extensible [486, 538, 892, 935, 1702, 687, 685, 1428, 1335, 663, 952, 1861]. Extension [271, 196, 436, 85, 1868, 817]. Extensions [1617, 64]. External [1370, 148, 211, 101, 1174, 350, 960, 1254, 1326]. Externally [1131]. Extinction [1099]. Extracting [1155, 1423, 1196]. Extraction [1551, 1386, 1550, 1378, 1756]. Extranet [1518]. Extreme [1503].

FaceKit [693]. Facilitating [1270]. Facilities [146, 157, 161, 427]. Facility [211, 90, 106, 13]. Facing [972]. Fact [1703]. Factor [190]. Factoring [699, 752]. Factors [281]. Facts [661]. FAD [594]. Failure [940, 744]. Failures [489]. FALCON [1468]. Family [1740, 1004, 684]. FAS [1705]. Fascicles [1378]. Fast [1439, 1016, 1087, 257, 1560, 991, 1697, 1573, 662, 982, 1317, 1263, 1158, 1331, 616, 1570, 809, 1312, 1373, 1204, 1228, 1476]. Fault [807]. Faulting [890]. Feasibility [502, 10]. Feasible [1512]. Feature [1479]. Features [1170, 290, 40, 1718]. Federated [1342, 1177, 976, 1194, 1419, 841]. Federations [1497, 977]. Feedback [938, 1404, 1364, 1468]. Feedback-Driven [1404]. FeedbackBypass [1559]. Fertile [969]. Fetch [815]. Fetches [341]. Fido [815]. Field [721]. Fifteenth [1896]. Fifth [1885, 1907]. File [916, 233, 37, 538, 44, 34, 499, 996, 12, 89, 86, 378, 1258, 510, 268, 414, 517, 401, 745, 88, 14, 412, 936]. Files [1132, 189, 23, 171, 668, 1842, 215, 526, 412, 934].

Filestore [482]. Filing [386]. Fill [521]. Fill-in-the-Form [521]. Filter [1121, 936]. Filtering [1446, 1590, 1764, 1304]. Filters [951]. Finance [1282]. Financial [935, 1036, 582]. Find [80]. Finding [1798, 1250, 695, 1383, 701]. Fine [1751, 1396, 809]. Fine-Grained [1751]. Fine-Granularity [809]. Finite [615]. Firm [1321, 1058]. First [407, 1256]. Fittest [1067]. Fitting [1404]. Five [818]. Fixed [660]. Fixed-point [660]. Fixpoint [761]. Flat [598]. Flexible [857, 1626, 1702, 1693, 816, 1747, 1091, 283]. Florence [1890]. Flows [1069]. Fly [519]. Focused [1489, 1866]. Footprint [1717]. FORAL [48, 166]. Foreign [953]. Foreword [1535]. Forget [754]. Form [407, 406, 782, 618, 1558, 114, 521, 158, 263, 613, 272, 361]. Form-Based [1558]. Formal [1064, 225, 381, 408, 1545, 739, 933, 677, 356]. Formalism [801]. Formalisms [206]. Formalization [381]. format [833]. Forms [649, 1]. Formulation [1260]. Foundation [1521, 1227]. Foundations [1668]. Four [238]. Fourteenth [1895]. Fourth [406, 1886, 1884, 1906]. FQI [589]. Fractal [1071, 1125]. Fractals [1017]. Fractured [1678]. Fragmentation [1077]. Framework [1766, 447, 147, 376, 565, 1645, 788, 1317, 117, 323, 1278, 440, 150, 463, 832, 1700, 1866, 151, 566, 229, 652, 1624]. Framingham [1881]. France [1888]. Free [33, 1588, 509]. Free-Space [33]. Freeform [618]. French [252, 1603]. Frequencies [499]. Frequency [1671]. Frequent [1810, 1445]. Frequently [1549]. Freshness [1792, 1705]. Freshness-Sensitive [1705]. Friendly [179]. FROG [1132]. Front [512]. Front-End [512]. Full [991, 691, 884]. Full-Text [991]. Fully [215]. Function [1302, 480, 1649, 454]. Functional [601, 1291, 3, 380, 503, 219, 1150, 656, 514, 842, 944, 1590, 207, 273, 528]. Functions [1350, 953, 1074, 587, 1656, 1823, 663, 1709, 402]. Fundamental [1044, 223].

Funding [968, 156]. Further [1187]. Fusion [1618, 1180]. Future [730, 1212, 1042, 1514, 293, 129, 965, 1513, 1282, 1041]. Fuzzy [1691, 1815, 118, 1126]. Fuzzy-Joins [1126]. Fuzzy-Set [118].

G [503]. G-WHIZ [503]. Gaea [910]. Galax [1859]. Gamma [451, 552]. Gamma-Acyclic [451]. Garbage [1049, 1253, 1252]. Gateways [1294]. Gazing [872]. GeMBASE [1742]. General [1586, 397, 444, 563, 564, 58, 78]. Generalised [1365]. Generalization [423, 464]. Generalized [467, 218, 1779, 91, 1093, 1706, 698, 205, 61, 558, 6, 106, 1080, 1813, 472]. Generalizing [1052]. Generated [181]. Generating [1417]. Generation [1631, 1421, 1876, 283, 1866, 13, 929]. generations [680]. Generator [282]. Generic [1512, 1389, 1667, 1544, 1658, 1257, 1585]. Genetic [1444]. Genetic-Based [1444]. Genome [820, 1642, 1159]. Genomics [1722]. Genomics-Centric [1722]. Geo [1271]. Geo/Environmental [1271]. Geographical [1491]. Geometric [636, 685, 1742]. Geophysical [1162]. Georgraphic [876]. GERM [339]. Germany [1911, 1884, 1886]. Get [1753, 393]. GHOST [1396]. GIS [1201]. Given [1829]. Global [989, 718, 907, 1256, 1043, 1842, 1275, 1770, 522]. GlOSS [1052]. GMAP [1006]. GnatDb [1717]. Go [1390]. Goals [937]. going [1341]. Good [1473, 153]. Government [1606, 1603]. Governmental [1510]. Governors [971]. GRACE [550, 708]. Gradients [1571]. Graffiti [1643]. Grained [1751, 1800]. Gral [685]. Grammar [619]. Grammars [1712]. Granularities [1121]. Granularity [1396, 22, 809]. Graph [1008, 701, 451]. GraphDB [1000]. Graphical [1740, 691, 351]. Graphics [109, 265].

Graphs [571, 303, 1489, 1008, 1000, 1681]. Greece [1905]. Grid [1411, 1235, 1879, 1847]. Grid-Clustering [1411]. gRNA [1722]. Group [279, 1005, 1365]. Group-By [1005, 1365]. Grouping [1139, 714, 1813, 1828]. Groups [1170]. Groupware [1520]. Groupwise [1264]. Grown [1217]. GTE [1272]. Guaranteeing [879]. Guarantees [1305]. Guidance [66]. Guide [230, 351]. Guide-Line [230].

Hamming [1670, 936]. Handle [463].

Handling [854, 845, 104, 1372, 288]. Hands [1806]. Hands-Off [1806]. Happens [1472]. Hard [1120]. Hardware [36, 60, 300]. Harness [1199]. Hash [760, 853, 951, 1007, 498, 1289, 1388, 1365, 708, 748, 1561, 747, 678, 829, 401, 454, 746]. Hash-Based [760, 498, 747]. Hash-Join [1388, 829]. Hash-Partitioned [678]. Hashing [462, 1412, 553, 479, 269, 378, 193, 268, 478, 480, 770, 786, 402]. Hausdorff [1125]. Having [609]. hB [1092]. hB-tree [1092]. hcC [992]. hcC-tree [992]. Head [26]. Health [1169]. Healthcare [1738]. Hear [1382]. Heavy [1798]. her [408]. Heterogeneity [919, 353, 1354, 1294, 676]. Heterogeneous [1493, 1055, 470, 878, 1236, 1405, 1849, 1165, 975, 1563, 1089, 1292, 1406, 1494, 1748, 879, 1101, 841, 1376, 1242, 1349, 774]. Heterogenity [1046]. Heuristic [354]. Hidden [1675, 1552, 1538]. Hierarchic [421]. Hierarchical [1182, 1426, 1370, 1798, 1675, 1462, 1658, 17, 1239, 1398, 1797]. Hierarchically [1703]. Hierarchically-Clustered [1703]. Hierarchies [1003, 423, 1052, 1413, 738, 1496, 1189, 1019]. Hierarchy [1420]. High [1721, 1517, 1356, 1123, 1449, 1415, 1209, 552, 1781, 1412, 1411, 1487, 1085, 979, 1385, 1012, 1057, 1428, 867, 1277, 60, 1312, 1635, 1488, 584, 1234, 881, 427, 1153, 1298, 1280]. High-Concurrency [1057]. High-Dimensional [1123, 1411, 1312, 1488, 1298, 1234]. High-End [1209]. High-Level [427, 1153]. High-Performance [1517, 1428]. High-throughput [1635]. High-Volume [1781]. Highly [1817, 1380, 1584, 45, 1593, 1515, 855, 1722, 658, 807]. Highly-Accurate [1593]. Hilbert [1017]. Hints [789]. Hippocratic [1654]. Histogram [1377, 1679, 1790]. Histogram-Based [1377]. Histograms [1263, 931, 1761, 1305, 1404, 1827, 1450]. Historical [535, 834]. Histories [1088, 450]. History [337, 1761, 941, 603, 1363, 1320, 536]. Hit [528]. Hits [766]. Hitters [1798]. Hole [969]. Holistic [1782]. Hong [1910]. Hope [1101]. Horn [659, 661]. Hospital [94]. Hosted [1867]. Hot [1053, 372]. Huge [1314]. Human [81, 281, 1159]. Hybrid [758, 771, 708, 1713, 1249, 1375, 1595]. Hybrid-Range [771]. Hyper [1434]. Hyper-Programming [1434]. Hyperlarge [262]. Hypermedia [814]. Hyperqueries [1594]. Hypertext [1399, 734, 700]. Hypothetical [1461, 403].

I/O [1674, 216, 1135]. I/O-Conscious [1674]. I/O-Operations [216]. IBM [1358, 1719]. Iceberg [1307, 1799]. ICICLES [1457]. ICL [482]. ICOT [577]. IDA [98]. IDBD [359]. Idea [547]. Identifiers [769]. Identifying [1474]. Identity [1605]. II [1867, 48, 58]. Illuminating [1852]. ILOG [769]. Image [1548, 484, 1024, 856, 1479, 1158, 1561, 905, 585]. Imagery [1875]. Images [1620, 1030, 1128]. IML [328]. IML-Inscribed [328]. Immunization [1167]. Impact [989, 139, 140, 897]. Impacts [317]. Implement [1063, 1453]. Implementation [105, 339, 1427, 1444, 371, 644, 742, 1416, 501, 882, 1418, 927, 1728, 762, 43, 663, 208, 118, 1320, 280, 1159, 669, 166, 1136, 39, 379, 1176, 47, 539, 1107, 50, 403, 682]. Implementations [1543]. Implementing [147, 1395, 1522, 1859, 997, 946, 1222, 1408, 643, 656, 10, 1141, 881, 817]. Implications [1820, 1459, 1118, 513]. Implicit [792]. Imprecise [788, 870, 329]. Improve [807]. Improved [1017, 963, 940, 860]. Improvement [1187]. Improvements [502]. Improving [1838, 1299, 951, 1083, 1555, 1690, 1578, 1368, 1589]. Impure [371]. IMS [381]. In-Memory [1517]. Including [1005]. Inclusion [578, 601, 641, 715]. Incomplete [485, 1190, 888, 434, 331, 1179, 119, 684, 1312, 273]. Inconsistency [121]. Incorporating [85, 1651]. Incremental [1285, 1049, 991, 785, 851, 1459, 1309, 947, 1263, 1218, 1483, 597, 1567, 517, 1709, 519, 1131]. Independence [1407, 1265, 110, 1006, 1341]. Independent [1346, 601, 332, 310, 924, 307]. Indeterminacy [1481]. Index [1534, 1123, 1663, 1231, 1003, 1573, 497, 731, 35, 1058, 1553, 1385, 1574, 560, 1322, 1723, 615, 1488, 1872, 637, 1166, 992, 1467, 1677, 1257, 1019, 1793]. Index-Organized [1467]. Indexed [1664, 1782]. Indexes [1486, 1557, 892, 1164, 1396, 1662, 764, 1773]. Indexing [1836, 1311, 991, 1449, 1415, 1024, 734, 1433, 1858, 781, 668, 1676, 990, 1428, 1575, 664, 1368, 1126, 1811, 1369, 1624, 1476, 1581, 884]. India [1904, 1118, 1200]. Indicator [454]. Indices [885, 1391, 1673, 1098]. Industrial [1846, 1533, 1431, 1847, 1360, 1361, 929]. Industry [972, 1118, 1730]. Inequality [472]. Inexpensive [474]. Inference [600]. Inferring [1302]. Infinite [495]. Influence [1549]. Infological [125]. Informatics [1525, 1532, 1524]. Information [467, 1446, 1632, 1535, 176, 1646, 120, 1551, 1499, 704, 1352, 1523, 991, 128, 345, 1339, 1493, 456, 1525, 497, 46, 1190, 1220, 1238, 175, 336, 788, 888, 1037, 434, 167, 73, 183, 331, 94, 1155, 1160, 429, 634, 870, 316, 1043, 1165, 119, 1506, 415, 52, 1120, 173, 1759, 68, 920, 1359, 329, 546, 1406, 876, 1492, 1642, 905, 1510, 1748, 224, 1419, 714, 1210, 315, 209, 1756, 536, 1590, 366, 1719, 1199, 918, 679, 1629, 1866, 1730, 273, 1783, 1169, 1235, 1153, 1644, 617, 1051, 763, 682]. Information-Loss [704]. Information-Retrieval [905]. Information-Theoretic [679]. Information-Transport [209]. Informative [199]. Informed [623]. Informix [1612, 1112]. Informix-Online [1112]. Infrastructure [1525, 1722]. INGRES [404, 227, 1032]. Inheritance [1416, 1222]. Inhomogeneous [108]. Inlining [1649]. Innovation [1215]. Innovative [970]. Inputs [1724, 1567]. Inscribed [328]. Insertion [1385]. Inserts [1838]. INSITE [1505]. Installation [230]. Installations [1718]. Instant [1611]. Institutions [582]. Insurance [1274, 1169]. Integrate [1150]. Integrated [417, 469, 1156, 497, 632, 1845, 415, 567, 366, 263]. Integrates [1316]. Integrating [1438, 1193, 1269, 1405, 316, 1759, 1224, 944, 1465, 792, 327, 1039, 1880]. Integration [1745, 1632, 1014, 1499, 707, 481, 1238, 505, 975, 1506, 1627, 920, 1406, 362, 439, 1529, 1864, 1727, 1492, 1031, 1756, 1719, 307, 1799]. Integrator [1789]. Integrity [257, 641, 1481, 404, 294, 3, 957, 2, 1205, 895, 591, 377, 1187, 1226, 778, 220, 706, 844, 922, 529, 627, 1181, 256]. Intelligence [1538]. Intelligent [493, 66, 98, 1592, 1127]. Intensional [1383, 706]. Intensive [1421, 1422, 1160, 1536, 1458, 1625]. Interaction [412, 1505]. Interactions [180, 1089]. Interactive [225, 1559, 28, 359, 350, 109, 1362, 1069, 1429, 1577, 1749, 1589, 319, 201]. Interactiveness [1275]. InterBase [718, 773]. Interesting [1313]. Interface [1346, 101, 428, 648, 996, 503, 97, 179, 533, 385, 433, 693, 1111, 202, 558, 1495, 430, 264, 351]. Interfaces [727, 493, 311, 1789, 199, 643, 1270, 1417, 416, 1040, 391]. Interfacing [647]. Internal [149, 1608, 152, 57]. International [1883, 1900, 1896, 1909, 1907, 1895, 1910, 1901, 1887, 1898, 1891, 1902, 1908, 1911, 1885, 1906, 1905, 1881, 1882, 1897, 1892, 1890, 1894, 1903, 1889, 1886, 1884, 1899, 1888, 1893, 1904]. Internet [1598, 1509, 1741, 1236, 1786, 1735, 1594, 1213, 1283, 1874, 1731, 1460, 1199, 1687]. Internet-scale [1874]. Interoperability [848, 1293, 1163, 1275]. Interoperation [1181]. Interpolated [1455, 1870]. Interpretation [943]. Interpreter [656]. Interpreting [620, 1505]. Interrelational [338]. Intersystem [809]. Intertask [921]. Interval [904, 450, 396, 1139, 621, 1582]. Intervals [468, 1478]. Intra [1078, 1184]. Intra-operator [1078]. Intra-Transaction [1184]. Intranets [1199]. Introduction [145, 355, 247, 421]. intrusive [1823]. Intuitive [284]. Invalidation [1689]. Inverted [934]. Invertible [1074]. Investigation [1010]. Investment [1153]. Invited [355, 381, 1533, 337, 313, 317, 308, 312, 349, 327]. Invocation [1]. IR [1831, 1272]. IR-Style [1831]. Ireland [1900]. IrisNet [1874]. IRO [1342, 1177]. IRO-DB [1342, 1177]. IRRD [208]. Irregular [1496]. Irrelevant [579]. ISO [278]. ISO/TC97/SC5/WG3 [278]. Isolation [1702]. Issue [971]. Issues [143, 1207, 848, 1379, 1464, 1537, 1337, 1194, 1483, 1732, 119, 980, 897, 414, 821, 223, 77, 902, 320, 82, 1886, 476]. Italian [1605]. Italy [1909, 1890]. ITCIS [20]. Itemsets [1445]. Iterative [831, 543].

J2EE [1690]. Janeiro [1885]. Japan

[1883, 582]. Japanese [584]. Jasmin [477]. Jasmine [775]. Jasmine/C [775]. Java [1434]. Join [1330, 785, 303, 551, 498, 1667, 982, 1388, 1803, 1255, 1290, 1097, 1822, 1365, 222, 708, 748, 511, 747, 847, 1021, 1809, 1472, 298, 341, 860, 688, 678, 1567, 829, 757, 1247, 1183, 770, 703, 949, 1166, 1027, 437, 1783, 1019, 746, 1290]. Join-Order [982]. Joining [1724]. Joins [1291, 853, 1664, 1007, 854, 1801, 1289, 1588, 1255, 271, 1256, 1782, 1126, 1473, 1234, 1065, 846, 472, 1329, 1784]. Journals [1099]. June [1903]. Just [1432].

K-Nearest [1802]. Kaleidoscope [827]. Kansas [1533]. Kernel [1346, 122, 1465, 598]. Key [735, 963, 764]. Key-Value [764]. Keynote [1532]. Keyword [1744, 1857, 49, 1698, 1831]. KMS [533]. KNN [1581]. Knowledge [849, 493, 830, 783, 258, 903, 577, 820, 1155, 1383, 1423, 607, 534, 562, 922, 529, 466, 1528, 974, 523, 1540, 1338, 835, 955, 1341, 585, 681, 709]. Knowledge-Based [493, 258, 466, 585, 529, 974, 709]. Knowledge-Bases [562]. Knowledgeable [433]. KODA [1346]. Komagome [94]. Kong [1910]. KVL [764].

L [1069]. L/MRP [1069]. Laboratory [1037]. Lachesis [1818]. laguna [680]. Land [336]. Landscape [1118]. Langauges [107]. Language [726, 850, 943, 105, 775, 594, 1014, 928, 790, 827, 1262, 856, 692, 200, 1340, 247, 1702, 244, 250, 648, 1576, 946, 1293, 422, 99, 97, 51, 1418, 838, 556, 104, 805, 875, 659, 691, 1163, 458, 724, 1788, 655, 350, 568, 1221, 558, 155, 1611, 209, 923, 835, 1176, 756, 391, 532, 63]. Language-based [1702]. Languages [182, 318, 606, 383, 55, 945, 608, 507, 150, 223, 944, 1010, 78, 1061, 1161]. Lapses [942]. Large [673, 1016, 1032, 1909, 192, 1907, 1875, 1910, 849, 1901, 1094, 31, 1898, 28, 347, 171, 474, 1566, 971, 1550, 1201, 1902, 1337, 334, 1908, 662, 1911, 1832, 121, 1317, 217, 245, 22, 1906, 29, 1081, 36, 1116, 124, 1553, 1905, 83, 668, 1881, 1160, 1315, 1241, 1532, 85, 616, 1424, 1867, 1674, 1827, 1770, 335, 658, 1046, 10, 369, 54, 583, 9, 81, 1196, 21, 388, 1287, 1031, 1082, 1890, 123, 819, 1267, 1318, 1579, 1694, 25, 457, 47, 1142, 231, 1106, 1904, 249, 1886, 561, 1718, 934, 1883, 1613, 1900, 1896, 1895, 1887, 1891, 1885, 1893, 1423, 1882]. large [1897, 1154, 1892, 1894, 1889, 1884, 1899, 1888, 660]. Large-Scale [1875, 1160, 1674, 1046, 1287, 1718, 1613, 1423]. Latching [901, 765]. Late [1074]. Latency [1688, 1516]. Latency-Recency [1688]. LAURA [408]. Law [1171]. Layered [476]. Lazy [997, 1453, 1373, 1075]. LDAP [1599]. LDL [702, 532]. Learning [491, 1470, 1117, 1749, 1541]. Learns [815]. Least [88]. led [1749]. left [763]. left- [763]. Legacy [1275, 1243, 1435]. LegoDB [1746]. Length [1333, 1174]. LEO [1541]. Less [437]. Level [1415, 49, 1081, 796, 1417, 994, 1642, 11, 58, 881, 427, 110, 42, 1153, 1200]. Levels [9, 58]. Leveraging [1153]. lexicon [1154]. Lexicons [934]. LHAM [1320]. Libraries [1240, 1030]. Library [1543, 1626]. Lies [820]. Life [1044, 1629, 1274]. Light [1322, 1435]. light-weight [1435]. Lightweight [1702, 1460]. Like [220, 827, 1140, 756]. Limitations [308, 1393]. Line [373, 781, 1241, 1425, 230, 899, 451, 1704, 1679, 1325]. Lineage [1586]. Linear [1697, 1656, 697, 269, 378, 268, 786, 763]. Lingering [1464]. Link [1791]. Linking [1519]. List [1048]. Lists [862, 863]. Living [1502]. Lixto [1631, 1551]. Load [1182, 816, 795, 891, 829, 1183, 925, 1079, 1376, 1785]. Loading [1327, 1203, 1397, 1211, 1257, 1585, 985, 1048]. Local [1449, 962, 511, 1835, 522, 1769]. Local-alignment [1835]. Locally [926, 307]. Locating [1832]. Location [1862]. Location-Based [1862]. Locking [898, 1387, 495, 397, 197, 305, 901, 808, 277, 357, 1057, 509, 399, 963, 1809, 764, 765, 809, 443]. Lockprotocols [400]. Locks [22, 277]. Log [1842, 628, 1320, 1731]. Log-Based [628]. Log-Structured [1320]. Logging [1387, 945, 624, 216, 717]. Logic [849, 572, 834, 317, 762, 514, 944, 761, 874, 923, 532, 557, 319, 613, 570]. Logic-Based [532]. Logical [186, 865, 408, 530, 1648, 1186, 200, 174, 219, 12, 221, 438, 362, 213, 1225, 1700, 39, 419, 307, 652]. Long [802, 1604, 721, 1002]. Long-Running [802]. Long-Term [1604]. Look [1753]. Loop [1027, 1468]. Loops [762]. Loosely [613]. LoPiX [1627]. Loss [704]. Lotus [1520]. Low [743, 1297, 1012, 694, 1025]. Low-Cost [743, 1297, 1025]. LP [166]. LSD [686].

M [1259]. M-tree [1259]. MA [1881]. Machine [172, 490, 313, 552, 550, 73, 694, 1117]. Machines [186, 771, 289, 887, 549, 518, 770, 320]. Magic [587, 753]. Magnetic [662]. Mail [647]. Main [1077, 1356, 1701, 1557, 1255, 942, 979, 560, 787, 1369, 1225, 786, 900]. Main-Memory [1077, 1356, 1701, 1557, 942, 900]. Mainframe [567]. maintain [763]. Maintaining [840, 1373, 1699, 1686, 741]. Maintenance [1285, 257, 1612, 777, 851, 497, 1263, 1680, 1747, 1219, 895, 1704, 1483, 1450, 1709, 922, 1131, 1161]. Making [1715]. Manage [261, 1008, 1850]. Management [143, 467, 649, 865, 1032, 806, 1517, 1271, 1045, 339, 1512, 426, 1684, 233, 1379, 493, 1115, 909, 442, 538, 1660, 892, 554, 1637, 1568, 474, 496, 856, 1876, 707, 1725, 101, 1519, 824, 122, 227, 1157, 15, 353, 1351, 1422, 907, 1090, 1845, 1381, 796, 32, 198, 926, 1168, 948, 1215, 1879, 789, 1012, 556, 179, 65, 618, 563, 86, 634, 477, 560, 168, 1147, 24, 604, 663, 74, 1213, 1597, 369, 286, 1202, 1069, 61, 1616, 1602, 821, 165, 1460, 1642, 388, 787, 1100, 328, 224, 72, 142, 7, 567, 714, 96, 366, 1818, 1726, 819].

Management [537, 872, 1730, 278, 13, 912, 807, 82, 263, 557, 135, 476, 617, 1054, 240, 1622, 1886]. Manager [1356, 1830, 906, 1273, 979, 633, 721, 431, 1210, 1127, 1347, 1208, 1785, 719, 1279]. Managers [531, 13]. Managing [1855, 751, 937, 345, 919, 935, 1108, 261, 687, 910, 631, 1478, 1639, 1640, 1078, 540, 372, 1121, 841, 1199, 169]. Mangers [879]. Manipulating [726, 1455, 1221]. Manipulation [318, 383, 769, 436, 191, 691, 1627, 78, 63]. Manufacturing [370]. MANY [1046]. Map [1682]. MapInfo [1359]. Mapping [1875, 448, 1047, 820, 1448, 480, 1184, 1808]. Mappings [1807]. Markov [1241, 1679]. Markov-Chain [1241]. MARS [1776]. Masking [626]. Mass [1096, 49]. Massive [1648, 1474, 1334]. Master [1373]. Match [37]. Matching [1774, 1693, 1750, 1544, 1292, 1128]. Materialization [504, 1028]. Materialized [1285, 1486, 1232, 1343, 1214, 1809, 1322, 1323, 1131, 1230]. Matrix [192]. Matter [1419]. Matters [1643]. Max [1462]. Maximizing [1013, 1783]. Maximum [356]. MCC [576]. MDL [1706]. Measure [1814]. Measured [621]. Measurements [1286, 1391]. Measuring [757]. Mechanism [322, 410, 850, 214, 465, 461, 372]. Mechanisms [522]. Media [535, 1266, 1268]. Media360 [1522]. Mediate [681]. Mediation [1862]. Mediator [1506, 1180, 1742]. Medical [1271, 856, 1158, 1338]. Meet [937]. Meets [295]. Melodies [95]. Memex [1498]. Memories [163, 86]. Memory [1817, 1517, 1077, 1356, 1701, 1367, 937, 1557, 1266, 1370, 1725, 1007, 93, 907, 901, 92, 1255, 942, 979, 560, 24, 1472, 1658, 939, 1224, 81, 26, 829, 960, 787, 1369, 1225, 1095, 1146, 786, 900, 1254, 1793, 1580, 475]. Memory-Adaptive [960].

Memory-Contention [1007]. Memory-Resident [901, 1793, 1580]. MEMS [1816]. MEMS-based [1816]. Menu [391]. Menu-Based [391]. Merge [1667, 1388]. Merge-Join [1388]. Mergesort [1254, 1326]. Mergesorts [1054]. Merging [1236, 439, 1829]. MERLIN [67]. Message [1781, 327]. Messaging [1690]. Meta [1721, 608, 153]. Meta-Data [1721]. Meta-Translation [608]. Metadata [1512, 1144, 369, 1141, 804]. Meter [1352]. Method [1063, 831, 696, 1259, 683, 499, 174, 183, 620, 1482, 196, 708, 748, 365, 616, 852, 1002, 891, 764, 765, 940, 1723, 1320, 678, 717, 315, 779, 1582, 1825, 873, 1581]. Methodologic [230]. Methodology [408, 793, 260, 362, 868, 188, 70]. Methods [1064, 462, 1880, 812, 662, 914, 234, 987, 306, 669, 1217, 58, 1149, 1298, 1580]. Metric [1094, 1259, 1797]. Metrics [559, 1101, 780]. Metropolitan [94]. Mexico [1889]. Micro [1200, 162]. Micro-DBMS [162]. Micro-level [1200]. Microcomputer [169]. MICRONET [169, 379]. Microprocessor [368, 170]. Microsoft [1216, 1386, 1387, 1231, 1289]. Middle [1614]. Middleware [1508, 1705, 1734, 1354, 1294]. Migration [1241, 1102, 813]. Millennium [1379]. Mind [619]. MindReader [1300]. MineSet [1209]. Mini [10]. Mini-Computer [10]. minimization [1772]. Minimizing [1174, 216]. Mining [904, 1016, 1333, 1498, 1751, 1444, 1571, 1309, 1752, 1384, 1753, 1357, 1795, 1371, 1443, 1274, 1315, 1331, 1086, 1805, 1117, 1570, 1140, 1314, 1529, 987, 1864, 1731, 1806, 1287, 1699, 1611, 1082, 1191, 1332, 1080, 1376, 1767, 1527, 1169, 1235, 1445, 1868, 1765]. Mining-Based [1805]. MIRO [1342, 1438]. MIRO-Web [1342]. Mirror [1440]. Mirroring [959]. Mirrors [1678]. Mismatch [805]. Mixed [1776, 1778, 676]. MLS [869]. MMDBMS [1440]. Mobile [1720, 1845, 855, 969, 1739]. Mobilizing [1849]. MOBY [586]. Modal [422]. Mode [1778]. Model [913, 1346, 487, 417, 826, 194, 408, 31, 212, 494, 394, 827, 360, 424, 8, 801, 856, 1389, 802, 1148, 773, 797, 174, 914, 1576, 1073, 1177, 882, 502, 798, 1008, 646, 55, 503, 219, 124, 877, 1443, 196, 326, 1150, 53, 295, 38, 725, 563, 160, 436, 870, 894, 896, 1506, 415, 114, 564, 290, 738, 213, 862, 558, 787, 520, 864, 863, 593, 1062, 1184, 372, 421, 58, 312, 153, 523, 677, 590, 869, 1841, 835, 110, 42, 1687, 846, 774, 1870, 358, 256, 356, 14, 569, 41, 528]. Model-Based [1506]. Modeling [410, 791, 128, 790, 144, 683, 797, 1171, 1000, 768, 1029, 416, 1192, 328, 869, 342]. Modelling [255, 297, 345, 296, 619, 1033, 429, 458, 999, 1869, 672]. Models [116, 1063, 447, 318, 76, 946, 196, 56, 1658, 241, 17, 684, 1829, 5, 1419, 210, 682]. Modern [1390, 1145]. Modification [957]. Modified [1092]. Modular [1063, 986, 170, 525]. Modularization [410]. Module [1743]. Modules [926, 26]. MOLAP [1704]. Molecular [447, 1841]. MOMIS [1499]. Money [1352]. Monitor [404]. Monitoring [309, 1660, 1804, 1739, 220, 728, 729, 1672]. Monotonic [783]. Montréal [1887]. Morphing [1794]. Morphology [967, 490]. Movable [26]. Movable-Head [26]. Move [983]. Move-Around [983]. Movie [1025]. Moving [1802, 1477]. MR [520]. MRP [1069]. Mudular [657]. Multi [535, 857, 1840, 475, 1669, 1571, 1293, 1801, 1479, 1227, 736, 796, 948, 1400, 808, 1163, 1114, 1561, 847, 1303, 1079, 864, 1184, 11, 1713, 637, 1318, 949, 1167, 1796, 1841, 1466, 1364, 110, 42, 1783, 476, 555, 763]. Multi-Attribute [736]. Multi-Database [1163]. Multi-Dimensional [1669, 1571, 1293, 1400, 1114, 637, 1796, 1466, 1227, 1841]. Multi-Dimensionally [1840]. Multi-Feature [1479]. Multi-Join [949]. Multi-Joins [1801]. Multi-Layered [476]. Multi-Level [796, 11, 110, 42]. multi-linear [763]. Multi-Media [535]. Multi-node [808]. Multi-Paradigm [948, 1167]. Multi-Phase [1303]. Multi-Precision [1561]. Multi-Processor [475, 1184]. Multi-Resolution [864, 1318]. Multi-Resource [1079]. Multi-System [857, 555]. Multi-Threading [1364]. Multi-version [1713]. Multi-Way [847, 1783]. Multiaction [764]. Multiattribute [551]. Multiclass [937]. Multidatabase [1036, 423, 773, 902]. Multidatabases [848]. Multidimensional [1188, 1232, 686, 1258, 1708, 852, 480, 1811, 1366, 1469, 1592, 1217, 1189, 1323, 1257]. Multifractals [1171]. Multilevel [1186]. Multilingualism [1768]. Multimedia [1070, 456, 1288, 1630, 1133, 1270, 1069, 1287, 1267, 1239, 386, 617]. Multiple [1438, 237, 1697, 571, 1170, 101, 1392, 1081, 444, 1219, 695, 668, 990, 117, 399, 1842, 939, 26, 879, 657, 868, 30, 446, 88, 1300]. Multiple-Level [1081]. Multiple-Query [939]. Multiple-View [1219]. Multiprocessor [1557, 866, 498, 771, 845, 887, 747, 549, 688, 829, 770, 1054]. Multiprocessors [949]. Multipurpose [392]. Multirelations [507]. Multiscale [1827]. Multitiered [1689]. Multiuser [746]. Multivalued [203, 204]. Multiversion [898, 394, 1568, 1149]. Multiview [868]. Mumbai [1904]. Mutual [205, 177]. MV3R [1582]. MV3R-Tree [1582].

N [1769]. Name [160, 482]. Name-Tracing [482]. Nanotechnology [1758]. Nanotechnology-based [1758]. NAOS [986]. National [583, 67, 584, 1338]. Nationwide [252]. Native [1854]. Natural [200, 99, 97, 568, 155, 391]. Navigable [1652]. Navigating [1613, 1261]. Navigation [428]. Navigational [427].

Navigations [1019]. NCR [929]. ND [1811, 1293]. nD-SQL [1293]. ND-Tree [1811]. NDB [161]. Near [1094, 1241]. Near-Line [1241]. Nearest [1480, 1487, 1802, 1158, 1710, 1666]. NEATO [1138]. Necessarily [114]. Need [971, 146, 1147, 973]. Needed [19]. Needs [1001, 1851]. neg [1222]. Negation [1773]. Neighbor [1094, 1487, 1802, 1158, 1710, 1666]. Neighborhood [1250]. Neighbour [1480]. Nested [675, 1291, 1003, 605, 644, 500, 646, 782, 471, 628, 689, 717, 598, 1027, 1364, 446, 800]. Nested-Loop [1027]. Nests [605]. NetBook [523]. NetCube [1570]. Netherlands [1896, 20]. Nets [328]. Network [1836, 1379, 383, 1804, 1335, 1275, 511, 284, 108, 1826, 584, 1538, 169, 427, 1208, 358, 1452, 41]. Networked [1380, 1381]. Networking [1740, 1760]. Networks [1278, 975, 52, 7, 1528, 1249]. Neural [1335, 975]. NeuroRule [1086]. Neuroscience [1506]. News [1013]. Next [1379, 1876, 1601, 929, 680]. Next-Generation [1876, 929]. NexusScout [1862]. NF2 [558, 756]. NHS [1147]. NICNET [1200]. No [1382]. Node [1092, 993, 1310, 1780, 808]. Node-Selecting [1780]. Noise [1087]. Non [407, 1646, 1077, 475, 397, 836, 1667, 783, 499, 1823, 277, 357, 661, 1709, 223, 1811, 1797, 672]. Non-blocking [1667]. Non-Distributive [1709]. Non-Equijoin [836]. Non-First-Normal-Form [407]. Non-Horn [661]. Non-intrusive [1823]. Non-Monotonic [783]. Non-ordered [1811]. Non-Procedural [223]. Non-Relevant [1646]. Non-Two-Phase [397, 277, 357]. Non-Uniform [1077, 499]. Non-Vector [1797]. Nonlinear [34, 1823]. Nonnumeric [300]. Nonpoint [686]. Nonrecursive [542]. Normal [407, 272, 406, 782, 361, 114, 158, 613].

Normalization [145]. Normalize [284]. Normalized [646, 114]. Norms [1670, 1476]. Not-Necessarily-Normalized [114]. Note [129]. Notes [1520]. Nothing [866, 925]. Novel [1433, 1385, 765, 388, 1477]. November [1890]. Now-Relative [1311]. NT [717]. Null [304, 642, 203]. Number [250, 1072]. Numeric [1143]. Numerical [1130, 1452]. NY [1906].

O [1414, 1135]. O-Conscious [1674]. O-O [1414]. O-Operations [216]. O2 [1060, 724, 1252, 719]. OASIS [1835]. Object [486, 726, 850, 913, 1327, 1049, 1132, 1253, 967, 1014, 826, 791, 1291, 538, 1156, 1508, 767, 1262, 1003, 1568, 856, 1245, 986, 740, 824, 997, 1060, 1074, 914, 383, 1073, 1178, 1177, 996, 814, 450, 1037, 1497, 906, 769, 948, 895, 768, 1405, 1155, 755, 859, 954, 1011, 805, 650, 725, 965, 1478, 634, 894, 896, 425, 564, 778, 516, 1368, 1111, 1647, 1180, 1477, 933, 868, 1184, 595, 792, 992, 923, 813, 1061, 858, 719, 1456, 890, 617, 1019, 1039]. Object- [914, 1456]. Object-Bases [859]. Object-Centered [1155]. Object-Orientation [923]. Object-Oriented [486, 850, 1132, 1014, 1291, 767, 856, 1245, 986, 1074, 383, 1073, 1178, 1177, 996, 1037, 948, 895, 1405, 805, 650, 725, 965, 896, 564, 778, 516, 1368, 1111, 868, 595, 792, 1061, 858, 617, 1019, 1039, 726]. Object-Relational [1291, 1508, 1245, 1478, 1647]. Objectivity [1360]. Objects [649, 720, 447, 930, 1395, 1119, 1070, 917, 1148, 998, 687, 1288, 686, 1222, 733, 875, 1038, 713, 635, 728, 637, 915, 539, 1272, 527]. OBK [1721]. Obligations [1684]. Observation [1875, 559]. Observations [78]. Obtaining [394, 1179]. Occurrence [499]. October [1883, 1885, 1890]. ODBMSs [1375]. Ode [825]. OdeFS [996]. ODMG [1262]. ODMG-Compliant [1262]. Off [1806, 677]. Offering [1454]. Office [488, 456, 458, 457, 386, 483]. Office/Workstation [457]. Official [1141]. OFS [263]. O’GRAPE [1860]. OID [1047, 1368]. OIS [516]. OLAP [1461, 1386, 1293, 1497, 1609, 1708, 1463, 1496, 1705, 1137, 1592]. OLTP [1821]. On- [1013]. On-Demand [1596, 1239]. on-going [1341]. On-Line [1425, 899, 1704, 1679, 1325]. On-the-Fly [519]. Once [1734]. One [1251, 1421, 1753, 1547, 1708, 1382]. One-Pass [1251, 1547, 1708]. One-To-One [1421]. Online [1721, 1584, 1504, 1665, 1835, 1429, 1364, 1112]. Only [1753, 464]. Ontology [1606]. Ontology-based [1606]. OODB [1027, 985, 1048]. OODBMS [984]. OODBs [1389]. Open [702, 1645, 1537, 1862, 640, 1112, 42, 680]. Open-Ended [640, 42]. Opened [642]. Opening [139]. Operating [1617, 444, 633, 764]. Operation [298, 534, 195]. Operational [1648, 1091]. Operations [407, 159, 462, 988, 1097, 553, 418, 341, 515, 216]. Operator [1830, 551, 1795, 1140, 955, 1078]. Operators [573, 1410]. OPOSSUM [1090]. Opportunistic [1290]. Opportunities [1610, 1393]. Optical [662, 833]. Optimal [1310, 1411, 1305, 1708, 52, 10, 453, 18]. Optimally [1240]. Optimisation [1010]. Optimistic [794, 234, 395, 306]. Optimizable [31]. Optimization [799, 127, 1350, 715, 572, 588, 953, 1005, 1134, 1427, 707, 1407, 1403, 878, 1422, 1301, 1246, 1260, 1824, 978, 1066, 1656, 1823, 861, 645, 954, 343, 1800, 542, 961, 950, 1788, 887, 983, 471, 847, 1067, 658, 344, 1472, 1394, 16, 689, 757, 522, 703, 677, 949, 622, 402, 1813, 1349, 911, 1718]. Optimizations [1256]. Optimize [1302, 587]. Optimized [214, 1367, 1143, 1770, 745, 1825]. Optimizer [1542, 1206, 1177, 828, 656, 544, 952, 657, 1541]. Optimizers [1138]. Optimizing [1471, 1652, 473, 1683, 1245, 833, 1479, 1244, 533, 859, 1673, 1139, 613, 609]. OQL [726, 835]. OQL/T [835]. Oracle [1844, 1837, 1838, 1345, 1343, 1279, 1113, 1344, 1197, 1361]. Oracle7 [1198]. Oracle8i [1509, 1467, 1624]. Oracle9i [1725]. ORDBMS [1410, 1743]. Order [982, 1004, 518]. Ordered [1788, 215, 1811]. Ordering [1535, 1770, 761, 879, 1828]. Organization [89, 1218, 300, 1355, 918, 152, 88, 14, 936]. Organizations [1258]. Organized [1467]. Organizers [1535]. Organizing [347]. Orientation [923]. Oriented [486, 850, 1132, 1253, 1014, 826, 791, 1291, 767, 856, 1245, 986, 740, 824, 1074, 1714, 914, 383, 1073, 1178, 1177, 996, 1037, 903, 99, 948, 895, 1405, 805, 650, 725, 965, 107, 896, 564, 778, 516, 1368, 1111, 868, 595, 792, 992, 1061, 858, 817, 617, 1019, 1039, 726]. OrientStore [1854]. Orleans [232]. Orthogonal [1418]. Orthogonality [1139]. OSAM* [835]. OSAM*/T [835]. OSIRIS [1860]. Outerjoins [463]. Outliers [1315, 1383]. Output [1814, 1097, 48, 1783]. Overhead [1012]. Overlap [1864]. Overlapping [1405]. Overview [143, 1035, 1607, 550, 1530, 719].

P [1480, 1572]. P-Sphere [1480]. Page [1770, 341, 809, 1731]. Page-Fetches [341]. Pages [1490, 958]. Pairs [1572]. Panel [1512, 848, 1514, 84, 1431, 968, 820, 837, 83, 969, 970, 897, 1600, 876, 1602, 1099, 822, 1100, 1513, 1101, 872, 823, 1601]. Paper [355, 381, 337, 313, 317, 308, 312, 349, 133, 327, 959, 132]. Paradigm [948, 1331, 390, 1167]. Paradise [1022]. Parallel [1345, 1077, 1182, 1350, 1617, 883, 760, 854, 812, 1108, 550, 1246, 1357, 814, 1066, 988, 927, 772, 690, 748, 1197, 950, 852, 694, 847, 1078, 1196, 1183, 1211, 925, 1079, 574, 239, 786, 1191, 1234, 1112, 1466, 1634, 379, 1176, 1376, 846, 936]. Parallel-Join [1183]. Parallelism [978, 1078, 862, 1184, 1107, 900]. Parallelism-Communication [978]. Parallelizer [1303]. Parameters [262]. Parametric [1301, 1824, 1656, 1823, 861, 1404]. Paraphrasing [568]. Parity [743]. Part [1514, 1513]. Partial [37, 646, 1814, 269, 378]. Partial-Match [37]. Partially [623, 934]. Partition [845]. Partitioned [782, 1842, 1021, 678, 1048]. Partitioned-List [1048]. Partitioning [771, 518, 411]. Pass [1251, 1547, 1708, 1362]. Passenger [1720]. Passive [839]. Past [754]. Pasta [691]. Pasta-3 [691]. Path [1598, 1771, 1174, 998, 1178, 326, 1575, 1679, 248]. Paths [237, 701]. Pattern [1853, 1711, 1752, 1384, 1378, 1028, 1611, 1767]. Pattern-Based [1853]. Patterns [1333, 1779, 1313, 1765]. Paving [1819, 1540]. Pay [1268]. Pay-Per-View [1268]. PC [1042, 1376]. Peer [1745, 1736, 1595]. Peer-to-Peer [1745, 1736, 1595]. Percentile [695]. Perfect [454]. Performance [638, 1556, 1838, 1032, 1517, 287, 172, 1212, 1356, 818, 981, 442, 639, 69, 313, 552, 908, 742, 93, 1047, 662, 1083, 502, 559, 796, 988, 926, 945, 1418, 927, 124, 979, 645, 1012, 1286, 1391, 1428, 633, 1483, 1792, 544, 10, 60, 17, 891, 744, 1320, 621, 1454, 829, 1818, 312, 215, 443, 1589, 846, 1298, 890, 985, 1280, 900, 1054, 1294]. Performances [1389]. Performing [1723]. Periodic [1185]. Perpetual [1270]. Persistent [919, 596, 1008, 945, 1418, 1009, 1368, 1135]. Personal [1850, 432, 431]. Personalization [1507, 1733, 1504]. Personalized [653]. Perspective [417, 849, 1545, 778, 1520, 974, 774]. Pervasive [1637]. PESTO [1156]. Peterlee [64]. Phantom [395]. Phase [397, 236, 277, 357, 961, 399, 1303]. Philosophies [238]. Phrase [1774]. Physic [1721]. Physical [1648, 1407, 4, 1258, 1225, 832, 188, 58, 110, 1006, 324, 358]. Physics [1277]. PicoDBMS [1628]. PicoDMBS [1442]. Pictorial [311]. Picture [1023]. Pictures [798, 1098]. Piecewise [1656, 418]. PIER [1786]. Pilot [209]. Pilot-System [209]. Pipeline [1589]. Pipelined [853, 978]. Placement [814, 772, 1816]. Placing [1240]. Plain [1340]. Plan [1350, 1657, 955, 1200]. Plan-Operator [955]. Plan-Per-Tuple [1350]. Planning [952, 1200, 270]. Planning- [952]. Plans [1542, 1407]. Platform [1526, 1862, 1727, 1736, 1841, 1615]. Platforms [1112]. Play [1460]. Plots [1480]. Plug [1460]. Plug-and-Play [1460]. PM3 [1418]. Point [686, 528, 660]. Pointer [890]. Points [34, 1381, 1802, 446, 18]. Pol [42]. Policies [513]. Policy [1684, 509, 1713]. Polygen [774]. Polymer [1133]. Polyphony [95]. Pool [372, 1347]. POREL [108]. Portable [928]. Portal [1733, 1623, 1866]. Portals [430, 1613]. Portugal [208]. Possibilistic [712]. Possibly [1769]. POSTGRES [593, 614, 1453]. Postgres-R [1453]. Potter [1577]. Power [812, 691]. Powered [1791]. Powerful [594, 1497, 1028, 342]. PQ [1887]. Practical [1063, 849, 956, 1464, 971, 1645, 854, 144, 1104, 1008, 314, 295, 363, 1194, 365, 1141, 1425]. Practice [348, 957, 1724, 132, 131, 285, 133, 1354]. Practices [1523]. Practitioner [1102]. Pragmatic [405, 432, 259, 409]. Pre [1833, 509, 1425, 1496, 1813]. Pre-Aggregation [1833, 1425]. Pre-Analysis [509]. Pre-Computed [1496]. Pre-Grouping [1813]. Preaggregation [1814]. Precision [1561, 1454]. Precision-Performance [1454]. Predeclared [356]. Predicate [531, 244, 250, 983, 242]. Predicates [1134, 1255, 251, 1304]. Predicative [395]. Prediction [1555, 312]. Predictions [1241, 964]. Predictive [816, 1825]. Preface [1535]. Preference [1728]. Preferences [1001, 1668, 607]. Prefetch [1395]. Prefetching [1673, 1713]. Prefix [1426]. Preliminary [435, 567]. Preparation [1674]. Preparatory [515]. Presence [1087, 953, 1189, 1813]. Present [1042]. Presentation [392, 1621]. Preservation [1604]. Preserving [176, 1836, 429, 1004, 518]. Presumed [961]. Prevention [495, 1555]. PRIMA [629]. Primitives [1820]. Principle [1044, 879, 1605]. Principles [1240, 45, 1144, 899, 1757]. Priorities [840]. Priority [723, 789, 789]. Priority-Based [789]. Priority-Hints [789]. PRISMA [927, 900]. PRISMA/DB [927, 900]. Privacy [1836, 1876, 137, 138, 1699]. Privacy-Enhanced [1876]. Privacy-Preserving [1836]. Private [502]. prize [959]. Proactivity [1275]. Probabilistic [592, 1403, 1238, 788, 1696]. Problem [105, 1045, 237, 28, 1545, 1351, 1214, 191, 395, 1294]. Problems [348, 972, 805, 1804, 1120, 1141, 821, 367]. Problems/Challenges [972]. Procedural [1195, 223]. Procedure [932]. Procedures [31, 645, 1061]. Proceedings [1909, 1907, 1898, 1908, 1911, 1906, 1905, 1881, 1903, 1904, 1883, 1900, 1896, 1895, 1910, 1901, 1891, 1902, 1893, 1882, 1894, 1886, 1899, 1888]. Process [1521, 1537, 1555, 906, 1484, 1587, 536, 1716]. Process-Centered [906]. Processable [221]. Processes [1639, 1640]. Processing [467, 1276, 1299, 799, 373, 1559, 928, 1840, 667, 309, 1451, 571, 354, 1264, 101, 34, 423, 605, 889, 1470, 1781, 254, 1074, 1178, 1638, 1801, 464, 1650, 1297, 1076, 1778, 258, 736, 238, 795, 690, 697, 1703, 755, 1594, 506, 1780, 1878, 432, 1278, 510, 887, 1599, 68, 1651, 765, 546, 1600, 1826, 1425, 1477, 300, 1429, 574, 328, 330, 251, 463, 466, 1095, 239, 770, 584, 1018, 1166, 955, 1149, 780, 1828, 319, 1280, 570, 472, 240, 1581, 1769, 1329]. Processor [1390, 475, 1843, 85, 168, 1661, 226, 1184]. Processors [167, 170, 411]. Proclamation [877]. Proclamation-Based [877]. Product [1206, 1105]. Production [840, 777, 851, 883, 919, 947, 283, 974, 817]. Profiles [1688]. Profiling [1741]. Program [1535, 428, 103, 283, 98, 243, 248, 430, 62, 229, 42]. Programmable [1722]. Programming [850, 943, 775, 182, 849, 197, 945, 1418, 838, 805, 724, 1417, 842, 1010, 521, 78, 87, 1434]. Programs [543, 626, 587, 514, 761, 752]. Progress [95, 1109]. Progressive [394, 1667, 1569]. Project [1499, 252, 175]. Projecting [1777]. Projection [301, 786]. Projections [1697, 380, 872]. Projects [156]. PROLOG [483]. Promises [965, 1103]. Proof [318]. Propagation [1306, 1578, 1748]. Propel [1615]. Properties [811, 407, 578, 176, 1139, 1590]. Property [1032, 358]. PROQUEL [655]. ProTDB [1696]. Protect [1120]. Protected [807]. Protecting [365]. Protection [120, 55, 183]. Protein [1695]. Protocol [751, 1701, 397]. Protocols [394, 305, 277, 357, 1809, 809]. Prototype [1620, 629, 634, 714]. Prototyping [1722, 655, 1105]. Provide [706]. Providers [1602]. Providing [602, 976, 1155, 1102, 940, 1419, 912]. Proving [228]. Provisioning [1587]. Provisions [1684]. Proximity [1857, 1284, 1234, 1398]. Proxy [1558]. PRTV [47]. Pruning [1316]. Pseudo [886]. Pseudo-Ranked [886]. PSJ [610]. PSJ-Queries [610]. Psychological [82]. PTT [20]. Public [1720, 1316]. Publish [1750, 1503, 1635, 1865]. Publish/Subscribe [1750, 1503, 1865]. Published [1834]. Publishing [1712, 1508, 1776, 1447]. Purpors [444]. Purpose [78]. Push [1507]. Pushing [1856]. Putting [607].

Quadtrees [1125]. Qualified [1001].

Quality [1500, 1145, 1555, 1305, 1804, 1578, 1805, 1406, 973]. Quality-driven [1406]. Quantifiable [1331]. Quantification [1245]. Quantifiers [605]. Quantiles [1251, 1680]. Quantitative [1548, 400, 1298]. Quantities [45]. Quasi [718]. Queries [1471, 1856, 1461, 1490, 1071, 1652, 1560, 1771, 37, 572, 588, 1426, 1659, 1264, 1134, 1401, 1402, 473, 1302, 1683, 1591, 1245, 423, 605, 889, 470, 1403, 1307, 1074, 543, 788, 1812, 1593, 1547, 450, 1801, 464, 1650, 1824, 1479, 798, 1244, 238, 690, 1822, 1802, 1703, 506, 725, 1780, 1665, 542, 762, 691, 607, 508, 643, 1462, 924, 1139, 471, 847, 544, 1563, 1463, 706, 689, 860, 1567, 589, 1454, 1649, 1485, 1762, 1773, 574, 251, 1562, 770, 243, 248, 792, 1172, 1027, 1065, 622, 215, 1364, 1582, 1825, 1543, 1783, 319, 613, 610, 472]. Queries [672, 609, 1772]. Query [1439, 726, 1817, 1299, 105, 799, 1559, 1840, 1386, 898, 1156, 1295, 1206, 827, 1451, 571, 354, 1493, 953, 1005, 1262, 938, 1427, 856, 971, 707, 250, 1470, 1781, 878, 648, 499, 35, 1564, 1301, 1457, 1177, 1246, 1638, 1657, 1260, 422, 1297, 1076, 1397, 1130, 1778, 258, 736, 99, 1066, 608, 1656, 1823, 620, 861, 1377, 645, 556, 755, 954, 1594, 343, 1404, 1050, 1878, 418, 432, 691, 828, 656, 1788, 887, 983, 1270, 350, 1770, 568, 1067, 1661, 658, 344, 1394, 1138, 939, 1448, 952, 546, 1731, 757, 1826, 1477, 1183, 1010, 463, 466, 657]. Query [1611, 1095, 1146, 522, 1872, 1018, 677, 835, 622, 1061, 1176, 1161, 1813, 1589, 1149, 1242, 1349, 780, 1107, 1828, 610, 1765, 570, 240, 1769, 1]. Query-Answers [1377]. Query-By-Example [418, 1]. Query-driven [1872]. Query-Result [1183]. Query/Browser [1156]. Query/Update [422]. Querying [916, 1088, 1554, 1853, 1170, 1186, 1392, 1753, 1000, 1033, 1786, 855, 1300, 1165, 1575, 1296, 1561, 1328, 1495, 1393, 1565]. Quest [206]. Question [181].

Question-Answering [181]. Questions [244, 245]. Queues [919]. QUIET [1872]. Quintillabit [262]. QUIST [343]. Quorum [674, 751, 750]. Quotient [1707].

R [1838, 1311, 337, 797, 1310, 1858, 1256, 1017, 1453, 1057, 1810, 637, 1718]. R* [544]. R-Tree [1311, 1017]. R-Trees [1838, 1057, 1810, 1310, 1256]. R/3 [1718]. Railways [584]. Rainbow [1501]. RainForest [1317]. RAM [716]. Random [886, 1490, 23, 1389, 833, 545, 710, 401]. Randomized [982]. Randomizing [365]. Range [1426, 1591, 771, 1462, 963, 1233, 1121]. Range-Max [1462]. Range-Sum [1426, 1591]. Ranked [886, 1724, 1567]. ranking [1154]. Ranks [1236]. Rapid [1532, 655, 1736]. RasDaMan [1271, 1437]. Raster [1329]. Rate [1783]. Ratio [1331]. RDB [344]. RDB/V1 [344]. RDBMS [1619, 1858, 1475, 1359, 1112]. Re [1029, 1663]. Re-Engineering [1029]. RE-Tree [1663]. Reactive [986]. Read [464, 745, 1326]. Read-Ahead [1326]. Read-Only [464]. Reading [519, 958]. Real [638, 722, 1348, 909, 1533, 1321, 1058, 1845, 794, 1273, 1085, 737, 659, 1739, 822, 1579, 1629, 1672]. Real-Life [1629]. Real-Time [722, 1348, 909, 1321, 1058, 794, 1085, 737, 1739, 822, 638, 1273, 1579]. Realities [1103]. Reality [965, 1600]. Reason [1064]. Reasoning [1185, 524, 1023]. Rebound [171]. Recency [1688]. Recognition [451]. Recommender [1645]. Reconciliation [1500]. Reconciling [1873]. Reconstruction [402]. Record [49, 1723, 157]. Recording [1218]. Records [1604, 736]. Recovering [1220, 942]. Recovery [750, 31, 1387, 197, 1092, 1714, 749, 739, 1842, 1084, 809, 940, 628, 717, 288]. Recovery-Oriented [1714]. Recursion [503, 697]. Recursive [587, 690, 589, 1649, 574, 613]. Recursively [504]. Red [1612]. Redo [1084]. Reduced [326]. Reducing [674, 1295, 765, 1775]. Reduction [1449, 699]. Redundancy [1775, 59, 54, 414, 88]. Redundant [257, 1776, 756]. Reefs [506]. REFEREE [1645]. Reference [1253, 1291]. References [1325]. Referencial [641]. Referencing [1568]. Referential [294, 1205, 778, 803]. Refining [491]. Reflect [1001]. Reflections [361]. Reformulation [627]. Reformulations [1270]. Refresh [624]. Region [1258, 1233]. Register [86]. Regression [1669, 1233, 1767]. Regression-Based [1767]. Regular [1663, 1384, 1575, 701, 752]. Related [602]. Relating [150, 1038]. Relation [405, 500, 104, 612, 114, 202, 349]. Relational [435, 186, 1744, 1655, 1327, 105, 407, 176, 120, 172, 449, 484, 1386, 1746, 462, 1291, 279, 1815, 321, 1508, 715, 163, 1170, 1264, 1636, 496, 1245, 707, 481, 178, 644, 113, 35, 501, 550, 888, 687, 304, 1803, 685, 2, 90, 64, 111, 608, 1698, 1831, 1822, 331, 573, 1091, 104, 343, 1478, 107, 1163, 870, 203, 24, 471, 382, 568, 59, 512, 114, 10, 1502, 597, 803, 461, 298, 713, 534, 706, 292, 684, 545, 1111, 1647, 223, 864, 1184, 372, 195, 522, 11, 598, 703, 243, 1393, 1447, 1565, 1018, 881, 869]. Relational [106, 291, 169, 427, 115, 307, 47, 732, 557, 1456, 319, 483, 569, 240, 1816, 1633, 41, 800, 63, 1539]. Relations [322, 611, 1556, 373, 1395, 579, 380, 646, 782, 505, 1091, 755, 508, 1139, 43, 1827, 205, 241, 177, 874, 239, 598, 403, 1775]. Relationship [850, 192, 339, 8, 683, 556, 599, 439, 285]. Relationships [651, 1290, 1023, 1098]. Relative [1311, 1488]. Relaxed [1083]. Relevance [843]. Relevant [1646]. Reliable [309, 551, 743, 286, 897, 1224]. REMIT [568]. REMORA [175]. Remote [1201]. Removal [357, 1051]. Renaissance [95]. Reordering [1429, 1146]. Reorganization [34, 597, 517, 899, 342, 1325]. Repeating [1363, 754]. Replacement [1173, 1012, 1713]. Replica [1373]. Replicas [1205]. Replicated [751, 865, 489, 993, 1083, 867, 1454, 1373, 1452]. Replication [654, 639, 1157, 993, 1453]. Replication-Transparent [654]. Report [435, 279, 95, 232, 278, 1718]. Reports [1593]. Repositories [1625, 893, 1762, 1686]. Repository [1436, 1721, 1035, 1216, 1566]. Representation [636, 835, 527, 660]. Representations [5]. Representative [1474]. Representing [1122, 331]. Requests [181, 1153]. Requirement [1029, 210]. Requirement-Based [1029]. Requirements [1817, 1157, 974]. Research [1514, 1533, 84, 1501, 167, 577, 968, 1155, 969, 576, 970, 156, 72, 1159, 1513, 1524, 966, 229, 1354, 730]. ResearchIndex [1645]. Reservation [584]. Resident [722, 1251, 901, 1793, 1580]. Resilient [865, 1764]. Resolution [864, 1318]. Resolving [804]. Resource [310, 723, 1399, 1268, 1288, 688, 1079, 879, 1700]. Resources [1648, 1491, 1246, 52]. Response [937, 1085, 26]. Responses [679]. Responsive [1007]. Restart [31, 940]. Restarting [495]. Restructuring [782, 9, 102, 18]. Result [1652, 1404, 1673, 204, 1183]. Result-size [1404]. Results [1207, 1397, 38, 511, 177, 48]. Rethinking [1441]. Retrieval [192, 488, 1437, 484, 991, 1620, 1070, 1636, 1036, 1190, 181, 167, 534, 905, 366, 1023, 1098, 249, 1468, 1039, 527]. Retrieval-By-Unification [534]. Retrievals [833]. Reuse [1496, 1031, 1153]. Reverse [1710]. Review [164]. Reviewers [1535]. Revisited [778, 1048]. Revolution [1760, 1601]. Revolutionize [474]. Rewriting [1800]. Right [853, 763]. Right- [763]. Right-Deep [853]. Rights [1873]. Rio [1885]. RISC [1441]. RISC-Style [1441]. RoadRunner [1550]. ROBOT [99]. Robust [1833, 748, 891, 1818, 779].

Robustness [475]. Role

[116, 212, 697, 1181]. Roles [913]. ROLEX [1652]. Rollback [31]. Rollback/Restart/Recovery [31]. Rollups [1592]. Roma [1909]. Rotating [24]. Routing [1787, 555]. RP* [1004]. RQA [589]. RQA/FQI [589]. RRXF [1775]. RTMonitor [1739]. Rubicon [1729]. Rufus [918]. Rule [840, 1015, 818, 824, 533, 875, 1303, 922, 761, 1699, 819, 1376, 763]. Rule-Based [875]. Rule-Data [533]. Rule-Driven [1303]. Rules [1016, 777, 851, 883, 919, 947, 15, 998, 1143, 1081, 505, 1331, 656, 1140, 1314, 1313, 1082, 881, 1080, 1142, 817]. Run [1422]. Run-time [1422]. Running [802, 926]. Runs [695].

S [1865]. S-ToPSS [1865]. Sabre [326]. Safe [716, 1248, 509, 803]. Safely [1325]. Samples [1457]. Sampling [886, 1685, 1593, 1072, 1675, 1833, 545, 710, 1142]. Sampling-Based [1072]. Santiago [1901]. SAP [1718]. SAR [1910]. SASH [1790]. Satisfaction [434]. Satisfiability [599]. Satisfying [250, 380]. Save [1620]. SC5 [278]. Scalability [1517, 908]. Scalable [1330, 1380, 1509, 1711, 1504, 1004, 1475, 1570, 1089, 1485, 1191, 1332, 1112, 1626]. Scalablity [1198]. Scale [1598, 1875, 1566, 1160, 1867, 1674, 1046, 1287, 123, 1718, 1613, 1423, 1874]. Scaling [1087, 1442, 1584]. Scan [90]. Scan-Driven [90]. Scans [1723]. Scheduling [638, 722, 723, 1830, 121, 1246, 1268, 1288, 1784, 341, 974, 1700, 1589]. Schema [1863, 984, 806, 147, 491, 212, 237, 148, 211, 692, 101, 1693, 1060, 175, 1090, 246, 805, 650, 1544, 603, 54, 1854, 920, 1448, 1292, 16, 102, 439, 315, 419, 57, 40, 88]. Schema-driven [1863]. Schemas [715, 793, 149, 1844, 1808]. SchemaSQL [1163, 1408]. Schemata [599, 420]. Scheme [993, 772, 357, 668, 643, 190, 1624, 1767].

Schemes [601, 449, 303, 1003, 448, 184, 926, 794, 89, 517, 398, 302]. Schems [451]. Science [1115, 1215, 1133]. Sciences [1523]. Scientific [910, 1609, 1155, 1133, 1099, 445, 1694, 912, 1903, 911, 1041]. Scopes [1491]. Scotland [1907]. Scrap [1243]. Scream [1382]. SDC [748]. Search [673, 1087, 1032, 192, 1646, 1857, 23, 1094, 1259, 1763, 1269, 1752, 1412, 1284, 735, 1789, 1093, 1698, 1831, 1675, 1158, 191, 85, 828, 950, 168, 1673, 1114, 43, 1674, 1770, 1283, 1312, 1562, 1267, 215, 452, 1666, 1687, 1298, 1626, 1791]. Searches [1480, 1835]. Searching [1744, 189, 1751, 1695, 616, 1272, 934]. Seat [584]. second [1904]. Secondary [1695, 1780, 223, 5]. Secure [1683, 122, 1321, 977, 1717]. Secured [1653, 1873]. Security [1684, 345, 137, 227, 138, 976, 566]. Segmentations [1314]. Segmented [853, 1673]. Segments [781]. Select [551]. Select-Join [551]. Selecting [1780]. Selection [1486, 1232, 1231, 1401, 1402, 1003, 1545, 1648, 35, 982, 1178, 1657, 1675, 1323, 14]. Selective [1446, 497]. Selectivity [1598, 1071, 1400, 1335, 1679, 658, 1265]. Self [1441, 606, 1546, 1457, 111, 1219, 1679, 1790, 1644, 1736]. Self-Adaptive [1790]. Self-Contained [606]. Self-Descriptive [111]. Self-Maintenance [1219]. SELF-SERV [1736]. Self-similarity [1546]. Self-Tuning [1441, 1457, 1679, 1644]. Sell [1616]. Sell-side [1616]. Semantic [257, 791, 848, 919, 360, 424, 705, 1427, 1763, 1173, 244, 798, 2, 55, 1378, 556, 506, 343, 377, 1871, 975, 119, 512, 1046, 1495, 7, 902, 153, 677, 804, 1540, 256, 206, 1877, 1865]. Semantics [485, 1064, 1386, 318, 297, 1026, 1302, 1877, 812, 460, 383, 433, 507, 1707, 338, 382, 590, 420, 265, 660]. Semantics-Based [1064]. Semantics-Knowledgeable [433]. Semi [1751, 918, 1068, 472, 1613]. Semi-Joins [472]. Semi-Structured [1751, 918, 1068, 1613]. Semistructured [1285, 1853, 1438, 1573, 1392, 1260]. Sensing [1201, 1874]. Sensitive [1120, 1705]. Sensor [1752]. Separability [324, 358]. Sept [1902]. September [1909, 1907, 1895, 1901, 1898, 1908, 1911, 1881, 1894, 1889, 1904, 1886, 1884, 1888]. Sequence [484, 1136, 1128, 1476]. Sequences [1695, 1553, 1800, 1296, 1835, 1162]. Sequential [214, 1384, 401, 1611]. Serial [1701, 931]. Serializability [718, 396, 994, 879]. Series [1087, 1669, 1157, 1164, 1474, 1371]. SERV [1736]. Serve [583]. Server [1049, 1509, 1279, 1387, 1266, 1231, 740, 1022, 908, 630, 907, 962, 1357, 1289, 1388, 444, 1197, 1359, 1025, 1460, 1204, 1135, 1239, 1375, 1345, 1109]. Server-Directed [1135]. Servers [1653, 1704, 1792, 1754]. ServerWare [1206]. Service [1145, 1587, 1602, 1494, 1734]. ServiceGlobe [1735]. Services [1745, 1863, 1875, 1852, 1846, 1615, 1607, 1157, 1845, 1879, 1735, 1690, 1874, 1731, 1611, 1736, 1608, 1601, 1644, 1452]. Session [1846, 139, 1847]. Set [71, 759, 1377, 1790, 118, 1473, 372, 958, 817]. Set-Oriented [817]. Set-Valued [1377]. Sets [1291, 1474, 990, 1196, 862, 1549]. seventeenth [1898]. seventh [1909, 1888]. SGML [1210]. Shadowing [666]. Shapes [1088]. Shared [475, 1557, 866, 197, 1246, 22, 124, 277, 1618, 809, 829, 925, 1784]. Shared-Disk [1618]. Shared-Memory [1557, 475]. Shared-Nothing [866]. Sharing [943, 766, 808, 425, 555]. Shedding [1785]. Shift [86]. Shooting [1665]. Show [265]. Shrinking [1796]. Shrinking-Based [1796]. sic [1882]. Side [1852, 1616]. Sideway [1647]. Signature [696, 499, 936]. Signatures [1261]. Similarity [1087, 1299, 1646, 1559, 1560, 1259, 1752, 1412, 1561, 1562, 1267, 1098, 1298, 1546].

Similarity-Search [1298]. Simple [594, 394, 701, 765, 545, 915]. Simplification [181]. Simplified [104]. Simplify [1292, 416]. Simulation [125, 1389, 124, 1773]. Singapore [1891]. Single [160, 378, 1362, 239]. Single-Assignment [239]. Single-File [378]. SIRIUS [252, 353]. SIRIUS-DELTA [353]. SIT [1629]. SIT-IN [1629]. Site [1421, 1248, 1351]. Sites [1550, 1623, 1558, 1536, 1458, 1422, 1625]. Situation [729]. Sixth [1887]. Size [818, 380, 708, 1708, 698, 1121, 1404]. Sketches [1474]. Skew [854, 845, 748, 846]. Skewed [766, 885, 1171, 658]. Ski [433]. Sky [1665]. Skyline [1665, 1569]. Sliced [452]. Slicing [1812]. Sliding [1801]. Small [1322, 1717, 249, 1868]. Small-Footprint [1717]. Smart [1729, 1715]. Smartcard [1442]. SMOOTH [1630]. Snapshot [624, 1129]. Snapshots [253, 322]. Social [1510]. Society [1540]. Socratic [1420]. SOFIS [96]. Software [550, 1118, 893, 210, 807, 135, 136, 340, 70, 134]. Solution [1350, 1433, 1602]. Solutions [1614, 1732, 1518, 367]. Solver [105]. Solving [28, 805, 395]. Some [217, 245, 511, 980, 290, 292, 81, 177, 504, 367]. Sometimes [1353]. Sophisticate [145, 355]. Sophisticated [349]. Sort [171, 1667, 90, 396, 694, 549, 452]. Sort-based [1667]. Sorted [695]. Sorter [171]. Sorting [301, 960, 714]. Source [608, 1165, 774]. Source-to-Source [608]. Sources [1438, 1059, 1493, 1832, 1236, 1244, 1150, 1165, 1807, 1563, 1243, 1419, 1242, 1783, 1435]. Space [1620, 33, 1246, 1480, 1052, 636, 1382, 1505]. Space-Shared [1246]. Spaces [1094, 1449, 1259, 1487, 950, 1811, 1488, 1298, 1797]. Spain [1898]. SPARC [117, 150]. SPARCOM [192]. Sparse [192, 1228]. Spatial [1619, 799, 1330, 1071, 989, 648, 686, 198, 988, 1097, 1256, 1827, 564, 1359, 987, 1328, 1826, 1562, 871, 669, 779, 1121, 1318, 1023, 1098, 1235, 1329, 1580]. Spatially [1021]. SpatialWare [1359]. Spatio [1632, 1437, 1629, 1582, 1825]. Spatio-Temporal [1632, 1437, 1629, 1582, 1825]. Specialized [167]. Specific [1269, 1648, 1818]. Specification [225, 375, 428, 384, 460, 184, 882, 422, 957, 185, 440, 48, 595, 207, 1105]. Specification-Based [375]. Specifications [376, 3, 164, 333]. Specified [934]. Specifying [921, 946]. Speculative [1056]. Speed [1503]. Speeding [1797]. Speeding-up [1797]. Sphere [1480]. SPIRIT [1384]. Split [190, 640]. Split-Transactions [640]. Splitting [1310, 246, 1258, 1233]. Sponsors [1535]. Spreading [748]. Spreadsheets [1869]. SPRINT [1191]. SQL [1387, 1486, 1295, 1206, 337, 1231, 1820, 1193, 1725, 1269, 1293, 1289, 1388, 1728, 1800, 1408, 1195, 1140, 860, 558, 1204, 1334, 756, 1456, 1868, 1869, 609]. SQL-like [1140, 756]. SQL-Type [558]. SQL/Data [337]. SQL3 [930, 1416]. Staging [1596, 1107]. Staircase [1803]. Stand [1276]. Stand-alone [1276]. Standardization [279, 79]. Standards [1875, 293, 1755, 1141]. Standards-Based [1875]. Stanford [6]. Star [473, 1703, 872, 1799]. Star-Cubing [1799]. Starburst [785, 696, 721, 817]. Stars [1665]. Starting [1353]. State [1537, 821, 1167, 1041]. State-of-the-Art [821]. State-Wide [1167]. StateMachine [1743]. Statement [139, 1800]. Statistical [790, 347, 1124, 1148, 334, 1144, 368, 413, 365, 1139, 711, 1120, 369, 441, 1137, 367, 1903, 231, 1235, 1672]. Statistics [1839, 1117, 1864, 1141, 213]. StatMiner [1864]. StatStream [1672]. Status [279, 293, 278]. STDL [928]. Steps [1803]. STING [1235]. Stochastic [1334]. Stockholm [1892, 1903]. Storage [674, 1554, 1356, 1746, 1096, 1420, 360, 1636, 474, 1240, 1433, 1776, 630, 1758, 49, 743, 646, 1794, 1858, 979, 1286, 1780, 1002, 286, 1854, 190, 1025, 1602, 787, 5, 1287, 1355, 1818, 1189, 614, 1239, 1816, 1775]. Store [866, 510]. Stored [48, 622]. Stores [936]. Storing [95]. Story [1341]. Strategies [496, 254, 968, 808, 1011, 1578, 950, 963, 612, 589, 925, 1787, 288, 50, 890, 1458, 1326]. Strategy [1077, 1514, 760, 470, 771, 828, 1258, 344, 1069, 678, 589, 574, 1513]. Stratifiable [1187]. Stream [1830, 1878, 1806, 1494, 1208, 1785]. Streaming [1659, 1843, 1764, 1783]. Streams [1766, 1660, 1669, 1670, 1798, 1753, 1547, 1801, 1710, 1671, 862, 1767, 1868, 1672, 1784]. String [1588, 1574]. Striping [743, 1053]. Striptease [1850]. Strongly [850]. Structural [1664, 1649]. Structurally [1649]. Structure [105, 1123, 1122, 1663, 731, 500, 1695, 1662, 54, 118, 1322, 165, 1328, 1681, 453, 1488, 992, 152]. Structured [1751, 1148, 1416, 1320, 284, 1210, 918, 1068, 409, 1039, 1613]. Structured-Text [1039]. Structures [865, 484, 360, 45, 238, 221, 1574, 1091, 560, 1004, 803, 315, 366, 48, 1332, 291, 1841, 756, 807, 1677, 110, 1257, 1793, 63]. Structuring [190, 446]. Studies [818, 230]. Study [1032, 172, 639, 1339, 1614, 740, 29, 238, 645, 560, 549, 679, 1298, 890, 985, 1054]. Style [1441, 1831]. Sub [242]. Sub-Predicate [242]. SUBJECT [347]. Subqueries [605]. Subscribe [1750, 1503, 1635, 1865]. Subset [1205, 1255]. Substitutability [1416]. Substitution [735]. Substring [1400]. Subsystem [3, 957, 91]. Subtype [157]. Success [1464]. Sum [1426, 1591]. Summaries [1547]. Summarization [1820, 1706, 1871]. Summarize [1680, 1707]. Summarizing [1827]. Summary [1220]. Sundgren [132]. Super [748]. Superimposed [668]. Superjoin [548]. SuperPages [1272]. SuperSQL [1621].

Supervised [1631]. Support

[1439, 857, 1652, 1606, 1110, 360, 1614, 126, 422, 1720, 1795, 1034, 733, 1091, 755, 117, 458, 173, 109, 635, 738, 1369, 700, 745, 523, 912, 207, 42, 585, 1445]. Supporting [575, 984, 172, 1185, 424, 1092, 629, 1822, 1385, 645, 1810, 663, 1750, 1195, 1567, 863, 868, 1611, 598, 1167, 1543, 1019]. Surf [1498]. Surface [1742]. Surfing [1547]. Surprising [1333]. Surrogate [389]. Surrogates [371]. Survey [129, 74, 72, 188]. Survives [1067]. SVP [862]. Sweden [1903]. Sweeping [1330]. Sweeping-Based [1330]. Swiss [1274]. Switzerland [1902]. Swizzling [890]. Sybase [1542]. Symmetric [1248, 949]. Synchronization [276]. Synergistic [346]. Synopses [1681]. Syntax [158]. Synthesis [112]. Synthesizing [776]. Synthetic [113]. System [186, 467, 1733, 840, 1032, 352, 857, 299, 218, 287, 1857, 1348, 1271, 339, 492, 818, 848, 233, 1115, 1620, 1059, 1096, 321, 538, 1617, 347, 424, 696, 1441, 473, 456, 1036, 986, 404, 1209, 481, 101, 197, 45, 423, 1776, 1693, 46, 122, 1157, 66, 997, 1060, 353, 175, 626, 550, 126, 1177, 996, 1144, 422, 1037, 1720, 901, 685, 2, 90, 444, 1858, 99, 36, 1501, 1747, 926, 608, 1738, 94, 645, 556, 1155, 647, 343, 618, 1050, 1044, 1630, 633, 582, 1408, 85, 1133, 1147, 338, 663, 1750, 1842, 1084, 568, 1506, 1627, 1854, 118, 1202]. System [329, 583, 1616, 1496, 1635, 1577, 1465, 283, 224, 1419, 902, 1184, 372, 1210, 209, 1749, 96, 366, 6, 584, 1136, 312, 1199, 918, 1629, 614, 207, 25, 169, 320, 187, 263, 386, 557, 393, 1717, 1169, 201, 392, 617, 235, 1054, 1039, 240, 555, 1622, 1633, 337, 1865, 1204]. Systematic [387, 1805, 333]. Systems [602, 1439, 143, 1276, 1049, 192, 1015, 1212, 1077, 274, 586, 275, 1042, 426, 387, 381, 1701, 468, 531, 1182, 989, 909, 442, 1420, 495, 588, 1557, 76, 69, 1055, 474, 496, 1193, 596, 1645, 497, 810, 1702, 740, 227, 184, 15, 336, 962, 261, 305, 1832, 266, 1321, 814, 1308, 181, 368, 185, 795, 1093, 1418, 772, 183, 1273, 577, 1879, 179, 1160, 1668, 965, 1768, 75, 1163, 560, 168, 316, 1043, 852, 1882, 74, 661, 425, 415, 1213, 1658, 60, 1078, 286, 68, 346, 17, 765, 897, 242, 414, 61, 292, 1406, 455, 1368, 876]. Systems [157, 165, 1180, 280, 925, 1079, 77, 1748, 216, 328, 567, 209, 306, 11, 779, 819, 964, 745, 524, 841, 79, 1103, 1023, 106, 278, 807, 1903, 1376, 327, 858, 288, 774, 1456, 136, 476, 1280, 483, 356, 1595, 1769, 800, 682, 676].

T [835]. Table [1410, 268, 1467, 454]. Tables [1819, 1840, 1703, 1475, 1120, 1126, 1]. Tabular [1816]. Tagging [774]. Talk [1533]. Taming [1638]. Tandem [1206]. Target [419]. Task [279]. Tavant [1616]. TAXIS [290]. Taxonomy [1261, 56, 846]. TC97 [278]. TDMS [238]. Teach [1803]. Teams [1289, 1365]. Technical [143]. Technique [334, 587, 1794, 1835, 1649, 1811, 342, 884]. Techniques [1327, 192, 1442, 1560, 1880, 1427, 1174, 1047, 1800, 1788, 621, 1010, 188, 669, 1757, 1332, 539, 1149, 1585, 1169, 1272]. Technologies [1737, 1431, 139, 1738, 837, 140, 1739, 1277, 1641, 1754, 388]. Technology [1507, 1035, 1096, 141, 129, 73, 1639, 1640, 1623, 1109, 897, 1529, 1510, 1530, 1848, 1282, 1167, 1644, 1354, 1294]. Telcordia [1500]. Telecom [1085]. Telecommunication [857, 1348]. Telidon [281]. Temporal [1632, 1437, 1185, 1151, 1152, 376, 1333, 935, 651, 856, 1481, 889, 731, 834, 888, 1812, 1583, 1482, 733, 440, 887, 1021, 603, 1463, 1368, 922, 421, 703, 537, 1162, 1629, 1166, 835, 1582, 1825, 1767, 732, 652]. Temporal-Hierarchic [421]. tenth [1891]. Terabit [92]. Terabyte [1203]. TeraBytes [1638]. Teradata [1110, 1617]. Term [1604]. Termination [1409]. Terms [934]. Tertiary [1240, 93, 1286, 1095, 1146]. Test [64]. Testing [1645, 1334, 1280]. Tests [242]. Text [991, 1261, 1697, 662, 619, 167, 1030, 1283, 540, 328, 1039, 884]. Textual [664]. Their [78, 379, 420, 326]. them [491]. Theoretic [118, 679, 944]. Theoretical [144, 314]. Theory [675, 145, 318, 592, 71, 144, 305, 957, 396, 295, 132, 131, 285, 133]. There [1101]. Thinking [1644]. Third [1883, 1905]. Thirteenth [1894]. THM [421]. Those [1430]. Thought [1455]. Thousands [1672]. Thrashing [891]. Threading [1364]. Threads [1009]. Three [147, 667, 740, 9]. Throughput [743, 1085, 867, 1635]. Tier [1517, 1614]. TIMBER [349]. Time [602, 722, 1087, 1390, 1348, 937, 909, 1669, 481, 1157, 731, 1164, 1246, 1321, 1480, 1058, 794, 1085, 620, 1474, 504, 1371, 580, 1676, 429, 737, 1739, 1296, 621, 26, 822, 1166, 1476, 1672, 638, 1422, 1273, 1579]. Time- [1246]. Time-Related [602]. Time-Series [1087, 1669]. Timely [863]. Timestamp [675, 275, 399, 1582]. Timestamp-Based [675, 275]. Timestamping [1583]. Timestamps [468]. Tioga [912]. tm [1209]. Together [1030]. Tokyo [1883]. Tolerance [807]. Tool [125, 1500, 1231, 359, 1740, 438, 268, 1570, 1505, 1634, 1861, 1006, 525]. Toolkit [693, 1138]. Tools [727, 492, 130, 1880, 69, 314, 260, 1805, 1102, 292, 230, 224]. Top [1401, 1402, 1403, 1090, 1822, 994, 881, 1799, 1769]. Top- [1401, 1402]. Top-Down [994, 1799]. Top-k [1822]. Top-N [1769]. TOPAZ [1303]. Topological [396, 1827]. Topology [590]. TOPSS [1750, 1865]. Toronto [1865]. Total [94]. Toto [1533]. TPC [1207]. TPC-D [1207]. TPR* [1825]. TPR*-Tree [1825]. Tracing [1586, 482]. Tracking [1036, 1167]. Trade [677]. Trade-off [677]. Tradeoff [978, 1454]. Tradeoffs [770]. Traffic [1804, 1208]. Trails [1498]. Trajectories [1477]. Transaction [1063, 928, 667, 1701, 309, 233, 801, 773, 993, 308, 957, 29, 796, 795, 1583, 465, 1044, 633, 1278, 477, 765, 809, 738, 1460, 1062, 1184, 899, 745, 476, 1280, 356].

Transaction-Based [899]. Transactional [1253, 802]. Transactions [638, 722, 435, 1064, 857, 675, 487, 228, 877, 739, 316, 1842, 1226, 764, 628, 364, 640, 776, 717, 315, 1281, 288, 446]. Transducer [1661]. Transducer-Based [1661]. Transfer [809]. Transformation [1059, 793, 100, 196, 59, 1247, 1562, 1813, 610]. Transformation-Based [1247]. Transformational [323]. Transformations [176, 1586, 982, 1405, 243, 763]. Transforming [839]. Transitioning [1533]. Transitive [611, 758, 760, 759, 927, 573, 671, 698, 612]. Translating [543, 1692, 609]. Translation [1087, 1712, 572, 1493, 608, 461, 920, 1292, 291, 1861]. Translator [580]. Transparent [654]. Transport [1720, 209]. Transposed [189, 215, 526]. Transposition [561]. Traversal [1256]. Tree [751, 1311, 1711, 1779, 885, 1317, 238, 1780, 43, 764, 689, 1811, 1465, 1316, 779, 637, 1582, 1825, 1123, 1259, 1092, 686, 1482, 1017, 1488, 992, 1663]. Trees [1838, 886, 23, 1652, 853, 951, 605, 1143, 1480, 735, 1093, 326, 1057, 191, 1810, 1114, 325, 1002, 515, 1233, 710, 670, 1217, 995, 1121, 1310, 1256, 453]. TreeScape [1496]. Trends [1537, 1474]. Tribeca [1208]. Trie [402]. Trigger [1409]. Triggered [737]. Triggers [1464, 1193, 825]. Troubleshooting [1757]. tu [1539]. Tuning [981, 1441, 1457, 845, 708, 1679, 1872, 880, 1757, 1644]. Tuple [1350, 368, 1787]. Tuples [250]. Turmoil [1119]. TURTLE [1132]. Tutorial [1751, 197, 1752, 1753, 38, 1878, 1754, 1755, 1756, 1757]. Twelfth [1893]. Twenty [1909, 1907, 1910, 1906, 1905, 1904]. Twenty-Eighth [1910]. Twenty-fifth [1907]. Twenty-fourth [1906]. twenty-second [1904]. Twenty-seventh [1909]. Twenty-third [1905]. Twig [1683, 1782]. Two [1427, 397, 749, 236, 277, 357, 961, 399, 994, 230]. Two-Phase [236, 399]. Type [488, 359, 558, 1068].

Typed [850]. Types [811, 447, 1438, 154, 1415, 1430, 1416, 663, 1223].

U.S. [27]. UB [1465]. UB-Tree [1465]. Ubiquitous [1028, 1538]. UDL [247]. Ugly [1473]. UK [1907]. Ultra [1515]. Uncertain [870]. Uncooperative [1769]. Undecomposable [177]. Understanding [212, 1555, 513]. Undo [216]. Undo-Logging [216]. UNIBASE [469]. Unification [534]. Unified [322, 8, 247, 605, 1144, 103, 1443, 478]. Uniform [1077, 889, 824, 499]. Unifying [565]. Union [195]. Universal [405, 1279, 1427, 1245, 1407, 643, 202, 1358, 1349]. Universality [931]. Universe [1680]. Unix [1110]. Unlimited [45]. Unloading [1203]. Unnesting [860]. Unrolling [1409]. Unsolved [821]. Unstructured [1161]. Untrusted [1653]. Updatability [178]. Update [485, 705, 276, 1591, 422, 945, 580, 1578, 1226, 461, 513, 966, 420]. Updated [1549]. Updates [1175, 214, 579, 321, 997, 1802, 784, 1662, 875, 643, 1810, 924, 1723, 1372, 536, 30, 1020, 672, 959]. Updating [916, 579, 489, 1325]. USA [1895, 1906, 1881]. Usable [391]. Usage [213]. Use [154, 1266, 982, 1497, 4, 1194, 920, 1848]. Used [1030]. Useful [396, 970]. User [1542, 225, 493, 1350, 1134, 141, 1415, 648, 311, 1001, 99, 64, 1410, 199, 146, 179, 83, 618, 663, 664, 416, 1469, 1267, 1040, 106, 457, 681, 351]. User-Adaptable [618, 1267]. User-Adaptive [1469]. User-Defined [1350, 1415, 1410, 664, 1134]. User-Optimizer [1542]. User-Oriented [99]. Users [972, 1505, 1625]. Uses [1620]. usind [760]. Using [1439, 186, 1064, 1299, 857, 1253, 1071, 257, 750, 1688, 28, 1261, 1333, 1451, 1636, 551, 853, 938, 1685, 1648, 1670, 1340, 1489, 797, 1171, 80, 1238, 1143, 1205, 845, 1256, 1474, 1150, 755, 1739, 1165, 975, 1226, 1834, 1292, 1723, 515, 706, 678, 1864, 1731, 1485, 787, 879, 7, 372, 1488, 1562, 536, 1135, 312, 1167, 590, 1098, 1166, 1172, 807, 1242, 1398, 1445, 1456, 482, 358, 1272, 1869, 1536, 14, 1542, 1646, 571, 1645, 1125, 1720, 1858, 1017, 1621, 1154, 1196, 1435, 1624, 934]. Utility [742]. Utilization [175, 688, 190]. Utilizations [1013].

V [1002]. V-Trees [1002]. V1 [344]. Vague [788]. Validation [1628, 128, 544, 529, 627]. Value [1056, 1519, 1416, 1388, 764, 1647, 1681, 1265]. Value-cognizant [1056]. Valued [1377, 88]. Values [1593, 304, 642, 1072, 203, 658]. Vancouver [1899]. Variables [756]. Variant [1091]. Various [1211]. Vector [1052, 1002, 1797]. Vector-Space [1052]. Vehicle [64]. Verification [225, 128, 440]. Versatile [1006]. Version [565, 630, 563, 378, 399, 249, 1713]. Versioned [720]. Versioning [1225]. Versions [767, 481, 1704, 650, 1226, 1597, 1372, 1192, 915]. Versus [113, 330, 1281]. Vertical [1241]. Very [1883, 673, 1900, 1896, 1909, 1907, 1895, 1910, 1901, 1887, 1898, 971, 1891, 1902, 1337, 1908, 1911, 1885, 217, 245, 1906, 36, 1116, 1905, 83, 668, 1881, 1424, 1867, 1882, 1897, 54, 583, 81, 388, 1892, 1890, 1318, 1579, 1894, 25, 457, 1889, 1106, 1904, 1886, 561, 1884, 1899, 1888, 1893, 1046]. via [943, 1490, 1697, 1412, 1639, 1640, 770]. Viable [1644]. Viator [1740]. Video [1033]. Videotex [281]. Vienna [966]. View [1232, 426, 1652, 1689, 851, 8, 1545, 1268, 1740, 1219, 83, 580, 117, 461, 513, 362, 1323, 1161, 756, 135, 136, 1230, 528, 680]. Viewers [932]. Viewpoint [134]. Views [1285, 1374, 1436, 322, 1486, 225, 1343, 405, 1566, 178, 1839, 464, 1214, 1008, 504, 1809, 653, 1485, 868, 1565, 291, 1172, 1131, 41]. VIMSYS [798]. VIPAS [1791]. Virtual [1620, 193, 1579, 1791]. Virtues [308]. Vision [1511]. Vista [1612]. Visual [1132, 1551, 1853, 1262, 503, 1270, 1105, 1633]. Visualization [917, 1209, 1090, 912]. visXcerpt [1853]. VLDB [1900, 1535, 1895, 1910, 1514, 1617, 1902, 581, 1511, 1908, 1911, 1893, 1046, 1513, 1894, 1899]. VLDB-2020 [1511]. VLDB’98 [1358]. VLDBs [1110, 1362]. VLKDBs [783]. VLSI [467, 486, 298, 452]. VLSI/CAD [486]. VM [46]. VM/370 [46]. VMLDB [1046]. Voice [970]. Volume [1781, 1385, 1870]. VOODB [1389]. Voting [625]. vs [1480, 1215, 265]. VSAM [33]. VXMLR [1633].

W3QS [1050]. W4F [1435]. Walking [1579]. Walks [1490]. Want [1430]. Warehouse [1586, 1324, 1113, 1385, 1704, 1849, 1483, 1597, 1202, 1211, 1127, 1531, 1229]. Warehouses [1691, 1433, 1337, 1413, 1424, 1528, 1466, 1634]. Warehousing [1610, 1612, 1431, 1309, 1076, 1214, 1219, 1218, 1484, 1587, 1129, 1322, 1230]. WARLOCK [1634]. Warping [1676]. Watch [1803]. WATCHMAN [1127]. Watermarking [1655]. waters [680]. WaveCluster [1318]. Wavelet [1450]. Wavelet-Based [1450]. Wavelets [1451, 1547]. Way [1819, 1453, 847, 1540, 1783]. Weave [1237, 1536]. Weaving [1556]. Web [1745, 1863, 1733, 1632, 1237, 1490, 1875, 1551, 1438, 1688, 1852, 1846, 1421, 1880, 1459, 1550, 1614, 1877, 1470, 1546, 1491, 1607, 1504, 1342, 1564, 1351, 1422, 1789, 1675, 1050, 1423, 1578, 1792, 1673, 1623, 1732, 1867, 1674, 1791, 1558, 1641, 1625, 1731, 1503, 1692, 1552, 1762, 1435, 1505, 1736, 1608, 1866, 1540, 1861, 1601, 1536, 1458]. Web-Accessible [1564]. Web-site [1351]. Web-sites [1422, 1625]. WebFilter [1635]. WebService [1860]. Weight [1322, 1435]. Weird [1430]. Welfare [583]. Well [420]. Well-Behaved [420]. We’re [1533]. West [1884, 1886]. West-Berlin [1886]. WG3 [278]. Wheel [1577]. Where [1390]. White [1432]. WHIZ [503]. Who [1851, 1040]. Wide [1278, 1167, 1050]. Wide-Area [1278]. will [474]. Window [1801, 1784]. Windows [554]. Wireless [1643]. WISE [1789]. WISE-Integrator [1789]. Wishful [1644]. within [1619, 1155, 117]. Without [754, 1265, 1154]. Work [1523]. Workfile [1054]. Workflow [1522, 1610, 1639, 1640, 1100, 1167, 1622]. Workflow-Implementing [1522]. Workflows [1700]. Workload [937, 1820]. Workloads [981, 1790, 939]. Workshop [232]. Workspace [502]. Workspaces [1855]. Workstation [740, 567, 457]. Workstation-Mainframe [567]. Workstation-Server [740]. Workstations [681]. World [642, 1845, 932, 1050]. World-Wide [1050]. WorlInfo [1632]. Wrap [1243]. Wrapper [1631, 1243]. wrappers [1435]. Write [717, 745, 807]. Write-Ahead [717]. Writesets [356].

X [1123, 1199]. X-tree [1123]. Xcerpt [1853]. XISS [1858]. XISS/R [1858]. XMark [1726]. XML [1436, 1745, 1855, 1598, 1446, 1774, 1856, 1857, 1853, 1746, 1771, 1508, 1711, 1636, 1775, 1664, 1566, 1851, 1763, 1430, 1776, 1781, 1564, 1812, 1778, 1858, 1782, 1780, 1575, 1679, 1661, 1502, 1563, 1597, 1777, 1599, 1627, 1394, 1854, 1625, 1844, 1696, 1649, 1635, 1681, 1773, 1726, 1719, 1393, 1447, 1565, 1624, 1539, 1765, 1682, 1633]. XML-based [1635]. XML-Relational [1633]. XPath [1650, 1772]. XPathLearner [1679]. XPERANTO [1508]. XPRS [665]. XPS [1203, 1112]. XQRL [1843]. Xquec [1856]. XQuery [1779, 1843, 1859]. XSEarch [1763]. XSL [1651]. XXL [1543].

Year [1464, 581, 964, 872]. Yellow [1731]. York [1906].

Zero [704, 1670, 1849, 1516]. Zero-Delay [1849]. ZOO [1168]. Zurich [1902].

References

Zloof:1975:QEI

[1] Moshé M. Zloof. Query-by-example: the invocation and definition of tables and forms. In Kerr [1881], pages 1– 24. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                              /www.vldb.org/dblp/db/conf/vldb/

Zloof75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Hammer:1975:SIR

[2] Michael Hammer and Dennis McLeod. Semantic integrity in a relational data base system. In Kerr [1881], pages 25– 47. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                              /www.vldb.org/dblp/db/conf/vldb/
                                              HammerM75.html. ACM SIGMOD v. 1,

no. 1, September 1975.

Eswaran:1975:FSS

[3] Kapali P. Eswaran and Donald D. Chamberlin. Functional specifications of subsystem for database integrity. In Kerr [1881], pages 48–68. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                              /www.vldb.org/dblp/db/conf/vldb/
                                              EswaranC75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Hoffer:1975:UCA

[4] Jeffrey A. Hoffer and Dennis G. Severance. The use of cluster analysis in physical data base design. In Kerr [1881], pages 69–86. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/HofferS75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Reiter:1975:DMS

[5] Allen Reiter. Data models for secondary storage representations. In Kerr [1881], pages 87–119. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Reiter75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Schroeder:1975:SGD

[6] J. R. Schroeder, W. C. Kiefer, Richard L. Guertin, and W. J. Berman. Stanford’s generalized database system. In Kerr [1881], pages 120–143. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SchroederKGB75.html. ACM SIG-

MOD v. 1, no. 1, September 1975.

Roussopoulos:1975:USN

[7] Nick Roussopoulos and John Mylopoulos. Using semantic networks for database management. In Kerr [1881], pages 144–172. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
RoussopoulosM75.html. ACM SIG-

MOD v. 1, no. 1, September 1975.

Chen:1975:ERM

[8] Peter P. Chen. The enity-relationship model: Toward a unified view of data. In Kerr [1881], page 173. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Chen75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Navathe:1975:RLD

[9] Shamkant B. Navathe and James P. Fry. Restructuring for large data bases: Three levels of abstraction. In Kerr [1881], page 174. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
NavatheF75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Manacher:1975:FIL

[10] Glenn K. Manacher. On the feasibility of implementing a large relational data base with optimal performance on a mini-computer. In Kerr [1881], pages 175–201. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
Manacher75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Schmid:1975:MLA

[11] Hans Albrecht Schmid and Philip A. Bernstein. A multi-level architecture for relational data base systems. In Kerr [1881], pages 202–226. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/SchmidB75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Hubbard:1975:ALF

[12] George U. Hubbard and Norman Raver. Automating logical file design. In Kerr [1881], pages 227–253.

ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HubbardR75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Stemple:1975:DBM

[13] David W. Stemple. A data base management facility for automatic generation of data base managers. In Kerr [1881], page 254. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Stemple75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Yao:1975:SFO

[14] S. Bing Yao and Alan G. Merten. Selection of file organization using an analytic model. In Kerr [1881], pages 255–267. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/YaoM75.html. ACM

SIGMOD v. 1, no. 1, September 1975.

Fernandez:1975:DEA

[15] Eduardo B. Fernández, Rita C. Summers, and Tomás Lang. Definition and evaluation of access rules in data management systems. In Kerr [1881], pages 268–285. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
FernandezSL75.html. ACM SIGMOD

v. 1, no. 1, September 1975.

Mitoma:1975:ADB

[16] Michael F. Mitoma and Keki B. Irani. Automatic data base schema design and optimization. In Kerr [1881], pages 286–321. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

                                        US$15.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/MitomaI75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Miyamoto:1975:HPA

[17] Isao Miyamoto. Hierarchical performance analysis models for data base systems. In Kerr [1881], pages 322–352. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Miyamoto75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Winslow:1975:OCD

[18] Leon E. Winslow and J. C. Lee. Optimal choice of data restructuring points. In Kerr [1881], pages 353–363. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        WinslowL75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Comba:1975:NDC

[19] Paul G. Comba. Needed: Distributed control. In Kerr [1881], pages 364–375. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Comba75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Karsner:1975:DEN

[20] G. R. Karsner. Database experience in the netherlands PTT (ITCIS). In Kerr [1881], pages 376–404. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

                                        US$15.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Karsner75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Peebles:1975:CAL

[21] Richards Peebles and Eric G. Manning. A computer architecture for large (distributed) data bases. In Kerr [1881], pages 405–427. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
PeeblesM75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Gray:1975:GLL

[22] Jim Gray, Raymond A. Lorie, Gianfranco R. Putzolu, and Irving L. Traiger. Granularity of locks in a large shared data base. In Kerr [1881], pages 428–451. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/GrayLPT75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Bayer:1975:EST

[23] Rudolf Bayer and J. K. Metzger. On the encipherment of search trees and random access files. In Kerr [1881], page 452. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/BayerM75.html. ACM

SIGMOD v. 1, no. 1, September 1975.

Lin:1975:DRA

[24] Chyuan Shiun Lin, Diane C. P. Smith, and John Miles Smith. The design of a rotating associative array memory for a relational database management application. In Kerr [1881], pages 453–455. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

LinSS75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Su:1975:CCS

[25] Stanley Y. W. Su and G. Jack Lipovski. CASSM: a cellular system for very large data bases. In Kerr [1881], pages 456–472. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/

SuL75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Omahen:1975:ERT

[26] Kenneth Omahen. Estimating response time for auxiliary memory configurations with multiple movable-head disk modules. In Kerr [1881], pages 473–495. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Omahen75.html. ACM

SIGMOD v. 1, no. 1, September 1975.

Simonson:1975:DUB

[27] Walter E. Simonson and William T. Alsbrooks. A DBMS for the U.S. Bureau of the Census. In Kerr [1881], pages 496–498. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        SimonsonA75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Carlson:1975:ULD

[28] Eric D. Carlson. Using large data bases for interactive problem solving. In Kerr [1881], pages 499–501. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Carlson75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Hackathorn:1975:ATA

[29] Richard D. Hackathorn. Analyzing transaction activity to a large data base: An empirical study. In Kerr [1881], pages 502–504. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
Hackathorn75.html. ACM SIGMOD

v. 1, no. 1, September 1975.

Sekino:1975:MCU

[30] Lucia C. Sekino. Multiple concurrent updates. In Kerr [1881], pages 505–507. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Sekino75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Browne:1975:OMA

[31] James C. Browne and G. L. Lasseter. An optimizable model for application of rollback/restart/recovery procedures for large data bases. In Kerr [1881], pages 508–510. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/BrowneL75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Heacox:1975:EDD

[32] H. C. Heacox, E. S. Cosloy, and J. B. Cohen. An experiment in dedicated data management. In Kerr [1881], pages 511–513. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        HeacoxCC75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Chin:1975:AVF

[33] Y. H. Chin. Analysis of VSAM’s free-space behavior. In Kerr [1881], pages 514–515. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Chin75.html. ACM

SIGMOD v. 1, no. 1, September 1975.

Das:1975:RPF

[34] K. Sundar Das, Toby J. Teorey, and S. Bing Yao. Reorganization points for file designs with nonlinear processing costs. In Kerr [1881], pages 516–518. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

DasTY75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Farley:1975:QEI

[35] Gilles Farley and Stewart A. Schuster. Query execution and index selection for relational data bases. In Kerr [1881], page 519. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

                                        US$15.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/FarleyS75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Healey:1975:HSA

[36] Robert Healey and Bradford Heckman. Hardware and system architecture for a very large database. In Kerr [1881], pages 520–522. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

US$15.00. URL http://www.vldb.
org/dblp/db/conf/vldb/HealeyH75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Burkhard:1975:PMQ

[37] Walter A. Burkhard. Partialmatch queries and file designs. In Kerr [1881], pages 523–525. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
Burkhard75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Kerr:1975:RRA

[38] Douglas S. Kerr. Recent results on the attribute based data model — A tutorial. In Proceedings of the International Conference on Very Large Data Bases, Framingham, MA, USA, September 22–24, 1975 [1881], pages 526–528. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Kerr75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Shoshani:1975:ILD

[39] Arie Shoshani and K. Brandon. On the implementation of a logical data base converter. In Kerr [1881], pages 529–531. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
ShoshaniB75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Tsichritzis:1975:FCS

[40] Dennis Tsichritzis. Features of a conceptual schema. In Kerr [1881], pages 532–534. ISSN 0278-2596.

LCCN QA76.9.D3 I55 1975. US$15.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Tsichritzis75.html.

ACM SIGMOD v. 1, no. 1, September 1975.

Zimmerman:1975:DVD

[41] Klaus Zimmerman. Different views of a data base: Coexistence between network model and relational model. In Kerr [1881], pages 535–537. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Zimmerman75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Tsubaki:1975:MLD

[42] Masaaki Tsubaki. Multi-level data model in DPLS — database, dynamic program control & open-ended pol support. In Kerr [1881], pages 538–539. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Tsubaki75.html. ACM SIGMOD v. 1,

no. 1, September 1975.

Lien:1975:BST

[43] Y. Edmund Lien, Carl E. Taylor, James R. Driscoll, and Mark L. Reynolds. Binary search tree complex — towards the implementation of relations. In Kerr [1881], pages 540–542. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LienTDR75.html. ACM SIGMOD v. 1,

no. 1, September 1975.

Chu:1975:FDD

[44] Wesley W. Chu and E. Nahouraii. File directory design considerations for distributed data bases. In Kerr [1881], pages 543–545. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/ChuN75.html. ACM

SIGMOD v. 1, no. 1, September 1975.

Dathe:1975:DPD

[45] G. Dathe, K.-H. Dreckmann, and G. Henselmann. Design principles of a data base system for unlimited quantities of data with highly complex structures. In Kerr [1881], pages 546–548. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DatheDH75.html. ACM SIGMOD v. 1,

no. 1, September 1975.

Donovan:1975:EVB

[46] John J. Donovan, Rafael Fessel, Stuart G. Greenberg, and Loius M. Gutentag. An experimental VM/ 370 based information system. In Kerr [1881], pages 549–553. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
DonovanFGG75.html. ACM SIGMOD

v. 1, no. 1, September 1975.

Todd:1975:PEI

[47] Stephen Todd. PRTV: An efficient implementation for large relational data bases. In Kerr [1881], pages 554–556. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Todd75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Senko:1975:SSD

[48] Michael E. Senko. Specification of stored data structures and desired output results in DIAM II with FORAL. In Kerr [1881], pages 557–571. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/

Senko75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Gey:1975:KAM

[49] Fredric C. Gey and Marilyn M. Mantei. Keyword access to a mass storage device at the record level. In Kerr [1881], pages 572–588. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/

GeyM75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Weldon:1975:ISC

[50] Jay-Louise Weldon. Implementation strategies for the census data base. In Kerr [1881], pages 589–590. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975.

                                        US$15.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Weldon75.

html. ACM SIGMOD v. 1, no. 1, September 1975.

Hill:1975:DAL

[51] Jeffrey M. Hill. Datalanguage: The access language on the datacomputer. In Kerr [1881], pages 591–592. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/

Hill75.html. ACM SIGMOD v. 1, no. 1, September 1975.

Mahmoud:1975:OAR

[52] Samy A. Mahmoud and J. Spruce Riordon. Optimal allocation of resources in distributed information networks. In Kerr [1881], page ?? ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MahmoudR75.html. ACM SIGMOD v.

1, no. 1, September 1975.

Kent:1976:NCC

[53] William Kent. New criteria for the conceptual model. In Lockemann and Neuhold [1882], pages 1– 12. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

FL60.00. URL http://www.vldb.org/
dblp/db/conf/vldb/Kent76.html.

Meltzer:1976:SRC

[54] Herbert S. Meltzer. Structure and redundancy in the conceptual schema in the administration of very large data bases. In Lockemann and Neuhold [1882], pages 13–25. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Meltzer76.html.

Hartson:1976:SMD

[55] H. Rex Hartson and David K. Hsiao. A semantic model for data base protection languages. In Lockemann and Neuhold [1882], pages 27–42. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/HartsonH76.html.

Kerschberg:1976:TDM

[56] Larry Kerschberg, Anthony C. Klug, and Dennis Tsichritzis. A taxonomy of data models. In Lockemann and Neuhold [1882], pages 43–64. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56

1976.  FL60.00.  URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        KerschbergKT76.html.

Tompa:1976:CEI

[57] Frank Wm. Tompa. Choosing an efficient internal schema. In Lockemann and Neuhold [1882], pages 65– 77. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

                                        FL60.00. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Tompa76.html.

Senko:1976:DIL

[58] Michael E. Senko and Edward B. Altman. DIAM II and levels of abstraction — the physical device level: a general model for access methods. In Lockemann and Neuhold [1882], pages 79–94. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

                                        FL60.00. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SenkoA76.html.

Lozinskii:1976:RTR

[59] Eliezer L. Lozinskii. Relational transformation and a redundancy in relational data base. In Lockemann and Neuhold [1882], pages 95–102. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Lozinskii76.html.

McGregor:1976:HPH

[60] D. R. McGregor, R. G. Thomson, and W. N. Dawson. High performance hardware for database systems. In Lockemann and Neuhold [1882], pages 103–116. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
McGregorTD76.html.

Nahouraii:1976:ADC

[61] E. Nahouraii, L. O. Brooks, and Alfonso F. Cardenas. An approach to data communication between different generalized data base management systems. In Lockemann and Neuhold [1882], pages 117–142. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/NahouraiiBC76.html.

Su:1976:APC

[62] Stanley Y. W. Su. Application program conversion due to data base changes. In Lockemann and Neuhold [1882], pages 143–157. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

FL60.00. URL http://www.vldb.org/
dblp/db/conf/vldb/Su76.html.

terBekke:1976:DML

[63] J. H. ter Bekke. A data manipulation language for relational data structures. In Lockemann and Neuhold [1882], pages 159–168. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bekke76.html.

Hitchcock:1976:UEP

[64] Peter Hitchcock. User extensions to the Peterlee Relational Test Vehicle. In Lockemann and Neuhold [1882], pages 169–180. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hitchcock76.html.

Kellogg:1976:DCD

[65] Charles Kellogg, Philip Klahr, and Larry Travis. A deductive capability for data management. In Lockemann and Neuhold [1882], pages 181–196. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56

1976.  FL60.00.  URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        KelloggKT76.html.

Erbe:1976:EGC

[66] R. Erbe and Georg Walch. An enduser guidance component for an intelligent data base system. In Lockemann and Neuhold [1882], pages 197– 210. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

                                        FL60.00. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ErbeW76.html.

Noerr:1976:MDN

[67] P. L. Noerr and M. C. White. MERLIN — design of a national bibliographic database. In Lockemann and Neuhold [1882], pages 211– 222. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976.

                                        FL60.00. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/NoerrW76.html.

Merten:1977:IPS

[68] Alan G. Merten. Information processing systems in a database environment. In ACM [1883], page 2. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

                                        1977. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Merten77.html.

Chen:1977:DPT

[69] Peter P. Chen and S. Bing Yao. Design and performance tools for data base systems. In ACM [1883], pages 3– 15. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/ChenY77.html.

Yeh:1977:TDM

[70] Raymond T. Yeh and Jerry W. Baker. Toward a design methodology for DBMS: a software engineering approach. In ACM [1883], pages 16– 27. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/YehB77.html.

Childs:1977:EST

[71] D. L. Childs. Extended set theory. In ACM [1883], pages 28– 46. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/Childs77.html.

Rothnie:1977:SRD

[72] James B. Rothnie and Nathan Goodman. A survey of research and development in distributed database management. In ACM [1883], pages 48–62. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RothnieG77.html.

Hsiao:1977:DMA

[73] David K. Hsiao and Stuart E. Madnick. Database machine architecture in the context of information technology evolution. In ACM [1883], pages 63– 84. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/HsiaoM77.html.

Lowenthal:1977:SAD

[74] Eugene I. Lowenthal. A survey: The application of data base management computers in distributed systems. In ACM [1883], pages 85–92. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Lowenthal77.html.

Kunii:1977:DCD

[75] Tosiyasu L. Kunii and Hideko S. Kunii. Design criteria for distributed database systems. In ACM [1883], pages 93– 104. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

                                        1977. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KuniiK77.html.

Chandy:1977:MDS

[76] K. Mani Chandy. Models of distributed systems. In ACM [1883], pages 105– 120. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

                                        1977. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Chandy77.html.

Ramamoorthy:1977:AID

[77] C. V. Ramamoorthy, Gary S. Ho, T. Kirshnarao, and Benjamin W. Wah. Architectural issues in distributed data base systems. In ACM [1883], pages 121–126. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        RamamoorthyHKW77.html.

Stonebraker:1977:ODM

[78] Michael Stonebraker and Lawrence A. Rowe. Observations on data manipulation languages and their embedding in general purpose programming languages. In ACM [1883], pages 128–143. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
StonebrakerR77.html.

Sibley:1977:SDS

[79] Edgar H. Sibley. Standardization and database systems. In ACM [1883], pages 144–155. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sibley77.html.

Fields:1977:UNC

[80] Craig Fields and Nicholas Negroponte. Using new clues to find data. In ACM [1883], pages 156–158. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FieldsN77.html.

Nickerson:1977:SCH

[81] R. S. Nickerson. Some comments on human archival memory as a very large data base. In ACM [1883], pages 159–168. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Nickerson77.html.

Thomas:1977:PID

[82] John C. Thomas. Psychological issues in data base management. In ACM [1883], pages 169–186. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Thomas77.html.

Kamijo:1977:PUV

[83] Fumihiko Kamijo, Alois Strnad, D. S. Appleton, and Carl William. Panel: User’s view of very large database applications. In ACM [1883], pages 188–191. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KamijoSAW77.html.

Champine:1977:PDD

[84] George Champine, Craig Fields, Koichi Furukawa, David Jefferson, Fritz Krückeberg, Giora Pelled, Charles Purcell, John Rundell, and Tadahiko Takahama. Panel: Directions in database research. In ACM [1883], pages 194–202. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChampineFFJKPPRT77.html.

Lang:1977:AEL

[85] T. Lang, E. Nahouraii, K. Kasuga, and Eduardo B. Fernández. An architectural extension for a large database system incorporating a processor for disk search. In ACM [1883], pages 204– 210. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

                                        I55 1977. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/LangNKF77.
                                        html.

Kluge:1977:DFM

[86] Werner E. Kluge. Data file management in shift register memories. In ACM [1883], pages 211–212. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

URL http://www.vldb.org/dblp/
db/conf/vldb/Kluge77.html.

Su:1977:APC

[87] Stanley Y. W. Su. Associative programming in CASSM and its applications. In ACM [1883], pages 213– 228. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

I55 1977. URL http://www.vldb.
org/dblp/db/conf/vldb/Su77.html.

Yamamoto:1977:DBM

[88] Sumiyasu Yamamoto, Shinsei Tazawa, Kazuhiko Ushio, and Hideto Ikeda. Design of a balanced multiple valued file organization schema with the least redundancy. In ACM [1883], page 230. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/YamamotoTUI77.
html.

Ikeda:1977:ECF

[89] Hideto Ikeda. Evaluation of combinatorial file organization schemes. In ACM [1883], pages 231–235. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ikeda77.html.

Harder:1977:SDS

[90] Theo Härder. A scan-driven sort facility for a relational database system. In ACM [1883], pages 236– 244. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/Harder77.html.

Hakozaki:1977:CDG

[91] Katsuya Hakozaki, Takenori Makino, Masayuki Mizuma, Mamoru Umemura, and Shigeki Hiyoshi. A conceptual design of a generalized database subsystem. In ACM [1883], pages 246–253. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HakozakiMMUH77.html.

Heard:1977:DAT

[92] H. G. Heard. A direct access terabit archival memory. In ACM [1883], pages 254–258. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Heard77.html.

Eastlake:1977:TMA

[93] Donald E. Eastlake III. Tertiary memory access and performance in the datacomputer. In ACM [1883], pages 259–268. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Eastlake77.html.

Itoh:1977:THI

[94] Ichiji Itoh, Shigekoto Kaihara, Tadao Saito, and Mituhiro Sato. Total hospital information system of komagome metropolitan hospital. In ACM [1883], pages 270–276. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ItohKSS77.html.

Lincoln:1977:EDS

[95] Harry B. Lincoln. Encoding, decoding and storing melodies for a data base of renaissance polyphony: a progress report. In ACM [1883], pages 277– 282. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

I55 1977. URL http://www.vldb.
org/dblp/db/conf/vldb/Lincoln77.
html.

Scheber:1977:ASD

[96] Anton Scheber. Architecture of the SOFIS data base management system. In ACM [1883], pages 283– 290. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

I55 1977. URL http://www.vldb.
org/dblp/db/conf/vldb/Scheber77.
html.

Hendrix:1977:DNL

[97] Gary G. Hendrix, Earl D. Sacerdoti, Daniel Sagalowicz, and Jonathan Slocum. Developing a natural language interface to complex data. In ACM [1883], page 292. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HendrixSSS77.html.

Sagalowicz:1977:IID

[98] Daniel Sagalowicz. IDA: An intelligent data access program. In ACM [1883], pages 293–302. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sagalowicz77.html.

Harris:1977:UOD

[99] Larry R. Harris. User-oriented data base query with the ROBOT natural language query system. In ACM [1883], pages 303–312. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Harris77.html.

Falkenberg:1977:CTD

[100] Eckhard D. Falkenberg. Coexistence and transformation of data. In ACM [1883], pages 314–317. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Falkenberg77.html.

Dale:1977:PIM

[101] Alfred G. Dale and C. V. Yurkanan. A processing interface for multiple external schema access to a data base management system. In ACM [1883], pages 318–325. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DaleY77.html.

Navathe:1977:SAD

[102] Shamkant B. Navathe. Schema analysis for database restructuring. In ACM [1883], page 326. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Navathe77.html.

Housel:1977:UAP

[103] Barron C. Housel. A unified approach to program and data conversion. In ACM [1883], pages 327– 336. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/Housel77.html.

Kambayashi:1977:RDL

[104] Yahiko Kambayashi, Katsumi Tanaka, and Shuzo Yajima. A relational data language with simplified binary relation handling capability. In ACM [1883], pages 338–350. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KambayashiTY77.html.

Antonacci:1977:SIR

[105] F. Antonacci, P. Dell’Orco, V. N. Spadavecchia, and A. Turtur. Structure and implementation of a relational query language for the problem solver. In ACM [1883], pages 351–358. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AntonacciDST77.html.

Spath:1977:GEU

[106] C. R. Spath and L. S. Schneider. A generalized end-user facility architecture for relational database systems. In ACM [1883], pages 359– 369. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/SpathS77.html.

Lacroix:1977:DOR

[107] Michel Lacroix and Alain Pirotte. Domain-oriented relational langauges. In ACM [1883], pages 370–378. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/LacroixP77.html.

Neuhold:1977:PDD

[108] Erich J. Neuhold and Horst Biller. POREL: a distributed data base on an inhomogeneous computer network. In ACM [1883], pages 380–395. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/NeuholdB77.html.

Meder:1977:DBS

[109] H. G. Meder and Frank P. Palermo. Data base support and interactive graphics. In ACM [1883], pages 396– 402. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

                                        1977. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MederP77.html.

Toh:1977:MLS

[110] Tetsuo Toh, Seiichi Kawazu, and Kenji Suzuki. Multi-level structures of the DBTG data model for an achievement of physical data independence. In ACM [1883], pages 403– 414. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

                                        1977. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/TohKS77.html.

Hotaka:1977:SDR

[111] Ryosuke Hotaka and Masaaki Tsubaki. Self-descriptive relational data base. In ACM [1883], pages 415– 426. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

                                        I55 1977. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/HotakaT77.
                                        html.

Vetter:1977:DBD

[112] M. Vetter. Data base design by applied data synthesis. In ACM [1883], pages 428–440. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

I55 1977. URL http://www.vldb.
org/dblp/db/conf/vldb/Vetter77.
html.

Fagin:1977:DVS

[113] Ronald Fagin. The decomposition versus synthetic approach to relational database design. In ACM [1883], pages 441–446. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Fagin77.html.

Makinouchi:1977:CNF

[114] Akifumi Makinouchi. A consideration on normal form of not-necessarily-normalized relation in the relational data model. In ACM [1883], pages 447– 453. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/Makinouchi77.
html.

Tanaka:1977:DCR

[115] Yuzuru Tanaka and Takao Tsuda. Decomposition and composition of a relational database. In ACM [1883], pages 454–462. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

I55 1977. URL http://www.vldb.
org/dblp/db/conf/vldb/TanakaT77.
html.

Bachman:1977:RCD

[116] Charles W. Bachman and Manilal Daya. The role concept in data models.

In ACM [1883], pages 464–476. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

URL http://www.vldb.org/dblp/
db/conf/vldb/BachmanD77.html.

Klug:1977:MVS

[117] Anthony C. Klug and Dennis Tsichritzis. Multiple view support within the ANSI/SPARC framework. In ACM [1883], pages 477–488. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KlugT77.html.

Mizumoto:1977:IFS

[118] Masaharu Mizumoto, Motohide Umano, and Kokichi Tanaka. Implementation of a fuzzy-set theoretic data structure system (abstract). In ACM [1883], page 490. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MizumotoT77.html.

Lipski:1977:SIC

[119] Witold Lipski, Jr. On semantic issues connected with incomplete information data bases (abstract). In ACM [1883], page 491. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Lipski77.html.

Bancilhon:1977:PIR

[120] François Bancilhon and Nicolas Spyratos. Protection of information in relational data bases. In ACM [1883], pages 494– 500. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/BancilhonS77.
html.

Gardarin:1977:SAA

[121] Georges Gardarin and P. Lebeux. Scheduling algorithms for avoiding inconsistency in large databases. In ACM [1883], pages 501–506. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GardarinL77.html.

Downs:1977:KDS

[122] Deborah Downs and Gerald J. Popek. A kernel design for a secure data base management system. In ACM [1883], pages 507–514. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DownsP77.html.

Scott:1977:ALS

[123] George M. Scott. Auditing large scale data bases. In ACM [1883], pages 515– 522. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55

1977. URL http://www.vldb.org/
dblp/db/conf/vldb/Scott77.html.

Hulten:1977:SMP

[124] Christer Hulten and Lars Söderlund. A simulation model for performance analysis of large shared data bases. In ACM [1883], pages 524–532. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HultenS77.html.

Berild:1977:CTD

[125] Stig Berild and Sam Nachmens. CS4: a tool for database design by infological simulation (abstract). In ACM [1883], page 533. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BerildN77.html.

Gambino:1977:DBD

[126] Thomas J. Gambino and Rob Gerritsen. A data base design decision support system. In ACM [1883], pages 534–544. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GambinoG77.html.

Berelian:1977:EO

[127] Elias Berelian and Keki B. Irani. Evaluation and optimization. In ACM [1883], pages 545–555. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BerelianI77.html.

Bubenko:1977:VVA

[128] Janis A. Bubenko, Jr. Validation and verification aspects of information modeling. In ACM [1883], pages 556– 566. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3

                                        I55 1977. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Bubenko77.
                                        html.

Housel:1977:NSF

[129] Barron C. Housel and Alan G. Merten. A note on the survey: “Future of Database Technology”. In ACM [1883], pages 568–570. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HouselM77.html.

Bubenko:1978:DBD

[130] Janis A. Bubenko, Jr. and S. Bing Yao. Data base design tools. In Yao [1884], page 2. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BubenkoY78.html.

Sundgren:1978:DBD

[131] Bo Sundgren. Data base design in theory and practice. In Yao [1884], pages 3–16. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Sundgren78.html. See comments [132, 133].

Smith:1978:CPS

[132] John Miles Smith. Comments on the paper “Data Base Design in Theory and Practice” by Bo Sundgren. In Yao [1884], pages 17– 18. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Smith78.html. See [131, 133].

Theys:1978:CPS

[133] M. Theys. Comments on the paper “Data Base Design in Theory and Practice”. In Yao [1884], pages 19– 20. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Theys78.html. See [132, 131].

Yeh:1978:DBS

[134] Raymond Y. Yeh. Data base and software engineering — A design viewpoint. In Yao [1884], page 22. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        Yeh78.html.

Wasserman:1978:SEV

[135] Anthony I. Wasserman. A software engineering view of data base management. In Yao [1884], pages 23– 35. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Wasserman78.html.

Weber:1978:SEV

[136] Herbert Weber. A software engineering view of data base systems. In Yao [1884], pages 36–51. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Weber78.html.

Davida:1978:SP

[137] George I. Davida. Security and privacy. In Yao [1884], page 54. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Davida78.
                                        html.

Hsiao:1978:PSD

[138] David K. Hsiao, Douglas S. Kerr, and Stuart E. Madnick. Privacy and security of data communications and data bases. In Yao [1884], pages 55–67. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/HsiaoKM78.
                                        html.

Hsiao:1978:OSS

[139] David K. Hsiao. An opening statement for the session in impact of new technologies. In Yao [1884], page 70. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hsiao78.html.

Kluge:1978:INT

[140] Werner E. Kluge. Impact of new technologies. In Yao [1884], page 71. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
Kluge78.html.

Chen:1978:CTD

[141] Tien Chi Chen. Computer technology and the database user. In Yao [1884], pages 72–86. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Chen78.html.

Rothnie:1978:DDM

[142] James B. Rothnie. Distributed database management. In Yao [1884], page 88. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Rothnie78.html.

Adiba:1978:IDD

[143] Michel E. Adiba, Jean-Claude Chupin, Robert Demolombe, Georges Gardarin, and Jean Le Bihan. Issues in distributed data base management systems: A technical overview. In Yao [1884], pages 89–110. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/AdibaCDGB78.html.

Delobel:1978:DBT

[144] Claude Delobel. Data base theory and modeling — theoretical and practical aspects. In Yao [1884], page 112. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Delobel78.
                                        html.

Beeri:1978:SID

[145] Catriel Beeri, Philip A. Bernstein, and Nathan Goodman. A sophisticate’s introduction to database normalization theory. In Yao [1884], pages 113–124. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/BeeriBG78.
                                        html.

Johnson:1978:WEU

[146] F. E. Johnson. What end-user facilities do we need? In Yao [1884], page 126. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Johnson78.html.

Berg:1978:ITS

[147] John L. Berg. Implementing a three schema DBMS framework. In Yao [1884], pages 128–129. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Berg78.html.

Clemons:1978:ESC

[148] Eric K. Clemons. The external schema and CODASYL. In Yao [1884], page 130. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Clemons78.html.

Gerritsen:1978:CIS

[149] Rob Gerritsen. Conceptual and internal schemas in CODASYL. In Yao [1884], page 131. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Gerritsen78.html.

Manola:1978:RCDa

[150] Frank Manola. On relating the CO-DASYL database languages and the ANSI/SPARC framework. In Yao [1884], page 132. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Manola78.html.

Smith:1978:CCF

[151] Diane C. P. Smith. Conversion and the CODASYL framework. In Yao [1884], pages 133–134. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Smith78a.
html.

Stocker:1978:EOI

[152] Peter M. Stocker. Efficient organization of internal DBMS structure. In Yao [1884], page 136. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Stocker78.
html.

Sharman:1978:WSS

[153] G. Sharman. What is a “good” semantic or meta model? In Yao [1884], page 138. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sharman78.html.

Brodie:1978:WUA

[154] Michael L. Brodie and Joachim W. Schmidt. What is the use of abstract data types? In Yao [1884], pages 140– 141. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BrodieS78.html.

Raben:1978:NLD

[155] Joseph Raben. Natural language data bases. In Yao [1884], page 144. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Raben78.html.

Manola:1978:DBR

[156] Frank Manola. Data base research projects and funding. In Yao [1884], page 146. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Manola78a.html.

Palmer:1978:RSF

[157] I. R. Palmer. Record subtype facilities in database systems. In Yao [1884], pages 148–155. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Palmer78.
                                        html.

Smith:1978:NFA

[158] John Miles Smith. A normal form for abstract syntax. In Yao [1884], pages 156–162. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Smith78b.html.

Bradley:1978:ODB

[159] J. Bradley. Operations data bases. In Yao [1884], pages 164–176. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Bardley78.
html.

Kleefstra:1978:DBD

[160] Willem J. Kleefstra. Data base description with a single name category data model. In Yao [1884], pages 177– 185. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Kleefstra78.html.

Sharman:1978:DDF

[161] G. C. H. Sharman and N. Winterbottom. The data dictionary facilities of NDB. In Yao [1884], pages 186– 197. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SharmanW78.html.

Ting:1978:MDD

[162] P. D. Ting and Dennis Tsichritzis. A micro-DBMS for a distributed data base. In Yao [1884], pages 200– 206. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TingT78.html.

Chang:1978:BMR

[163] H. Chang. On bubble memories and relational data base. In Yao [1884], pages 207–229. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Chang78.html.

Manola:1978:RCDb

[164] Frank Manola. A review of the 1978 CODASYL database specifications. In Yao [1884], pages 232–242. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Manola78b.
                                        html.

Paolini:1978:ASD

[165] Paolo Paolini. An alternative structure for data base management systems. In Yao [1884], pages 243–254. ISBN ???? LCCN QA76.9.D3 I55 1978.

                                        US$25.00. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Paolini78.
                                        html.

Senko:1978:FLD

[166] Michael E. Senko. FORAL LP: Design and implementation. In Yao [1884], pages 255–267. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Senko78.html.

Hollaar:1978:CRS

[167] Lee A. Hollaar and David C. Roberts. Current research into specialized processors for text information retrieval. In Yao [1884], pages 270–279. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/HollaarR78.html.

Leilich:1978:SPD

[168] Hans-Otto Leilich, Günther Stiege, and Hans Christoph Zeidler. A search processor for data base management systems. In Yao [1884], pages 280– 287. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LeilichSZ78.html.

Su:1978:MMN

[169] Stanley Y. W. Su, Stefan Lupkiewicz, Chang jung Lee, Der Her Lo, and Keith L. Doty. MICRONET: a microcomputer network system for managing distributed relational databases. In Yao [1884], pages 288–298. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/SuLLLD78.
html.

Ozkarahan:1978:MBM

[170] Esen A. Ozkarahan and Kemal Oflazer. Microprocessor based modular database processors. In Yao [1884], pages 300–311. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/OzkarahanO78.html.

Chen:1978:RSE

[171] T. C. Chen, Vincent Y. Lum, and C. Tung. The rebound sorter: An efficient sort engine for large files. In Yao [1884], pages 312–318. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/ChenLT78.
html.

Banerjee:1978:PSD

[172] Jayanta Banerjee and David K. Hsiao. Performance study of a database machine in supporting relational databases. In Yao [1884], pages 319–329. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BanerjeeH78.html.

Mason:1978:DAS

[173] P. J. Mason. A DBMS architecture to support information analysis (abstract). In Yao [1884], page 332. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Mason78.html.

Flory:1978:MML

[174] André Flory and Jacques Kouloumdjian. A model and a method for logical data base design. In Yao [1884], pages 333–341. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        FloryK78.html.

Foucaut:1978:CDI

[175] O. Foucaut and Colette Rolland. Concepts for design of an information system conceptual schema and its utilization in the REMORA project. In Yao [1884], pages 342–350. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/FoucautR78.html.

Arora:1978:IPP

[176] Adarsh K. Arora and C. Robert Carlson. The information preserving properties of relational database transformations. In Yao [1884], pages 352– 359. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AroraC78.html.

Nicolas:1978:MDS

[177] Jean-Marie Nicolas. Mutual dependencies and some results on undecomposable relations. In Yao [1884], pages 360–367. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Nicolas78.html.

Dayal:1978:URV

[178] Umeshwar Dayal and Philip A. Bernstein. On the updatability of relational views. In Yao [1884], pages 368– 377. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DayalB78.html.

Kameny:1978:EEU

[179] I. Kameny, J. Weiner, M. Crilley, J. Burger, R. Gates, and David Brill. EUFID: The end user friendly interface to data management systems. In Yao [1884], pages 380– 391. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KamenyWCBGB78.html.

Lee:1978:CAD

[180] Ronald M. Lee. Conversational aspects of database interactions. In Yao [1884], pages 392–399. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Lee78.html.

Grishman:1978:SRR

[181] Ralph Grishman. The simplification of retrieval requests generated by question-answering systems. In Yao [1884], pages 400–406. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Grishman78.html.

Atkinson:1978:PLD

[182] Malcolm P. Atkinson. Programming languages and databases. In Yao [1884], pages 408–419. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Atkinson78.html.

Huber:1978:IPM

[183] Rainer Hüber and Peter C. Lockemann. Information protection by method base systems. In Yao [1884], pages 420– 426. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HuberL78.html.

Ehrig:1978:ASS

[184] Hartmut Ehrig, Hans-Jörg Kreowski, and Herbert Weber. Algebraic specification schemes for data base systems. In Yao [1884], pages 427–440. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/EhrigKW78.
html.

Hebalkar:1978:ASD

[185] Prakash G. Hebalkar. Application specification for distributed data base systems. In Yao [1884], pages 442– 449. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hebalkar78.html.

Adiba:1978:DDB

[186] Michel E. Adiba, Jean-Yves Caleca, and Christian Euzet. A distributed data base system using logical relational machines. In Yao [1884], pages 450–461. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AdibaCE78.html.

Toth:1978:ASA

[187] K. C. Toth, Samy A. Mahmoud, J. Spruce Riordon, and O. Sherif. The ADD system: An architecture for distributed databases. In Yao [1884], pages 462–471. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/TothMRS78.
html.

Schkolnick:1978:SPD

[188] Mario Schkolnick. A survey of physical database design methodology and techniques. In Yao [1884], pages 474– 487. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Schkolnick78.html.

Batory:1978:STF

[189] Don S. Batory. On searching transposed files (abstract). In Yao [1884], page 488. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

URL http://www.vldb.org/dblp/
db/conf/vldb/Batory78.html.

Nakamura:1978:ASU

[190] Toshiyuki Nakamura and Tetsuo Mizoguchi. An analysis of storage utilization factor in block split data structuring scheme. In Yao [1884], pages 489– 495. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
NakamuraM78.html.

Kung:1978:CDM

[191] H. T. Kung and Philip L. Lehman. A concurrent database manipulation problem: Binary search trees (abstract). In Yao [1884], page 498.

ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
KungL78.html.

Ashany:1978:ASM

[192] Ron Ashany. Application of sparse matrix techniques to search, retrieval, classification and relationship analysis in large data base systems — SPAR-COM. In Yao [1884], pages 499– 516. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ashany78.html.

Litwin:1978:VHD

[193] Witold Litwin. Virtual hashing: a dynamically changing hashing. In Yao [1884], pages 517–523. ISBN ???? LCCN QA76.9.D3 I55 1978.

US$25.00. URL http://www.vldb.
org/dblp/db/conf/vldb/Litwin78.
html.

Borkin:1978:DME

[194] Sheldon A. Borkin. Data model equivalence. In Yao [1884], pages 526– 534. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Borkin78.html.

Sagiv:1978:EAR

[195] Yehoshua Sagiv and Mihalis Yannakakis. Equivalence among relational expressions with the union and difference operation. In Yao [1884], pages 535–548. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SagivY78.html.

Kalinichenko:1978:DMT

[196] Leonid A. Kalinichenko. Data models transformation method based on axiomatic data model extension. In Yao [1884], pages 549–555. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00. URL http://

www.vldb.org/dblp/db/conf/vldb/
Kalinichenko78.html.

Date:1979:LRS

[197] C. J. Date. Locking and recovery in a shared database system: An application programming tutorial. In Furtado and Morgan [1885], pages 1–15. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/Date79.html.

Herot:1979:SMD

[198] Christopher F. Herot. Spatial management of data (abstract). In Furtado and Morgan [1885], page 16. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Herot79.html.

Janas:1979:TMI

[199] Jürgen M. Janas. Towards more informative user interfaces. In Furtado and Morgan [1885], pages 17–23. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Janas79.html.

Dahl:1979:LDD

[200] Verónica Dahl. Logical design of deductive natural language consultable data bases. In Furtado and Morgan [1885], pages 24–31. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/Dahl79.html.

Wilson:1979:SIE

[201] Gerald A. Wilson and Sandra B. Salazar. A system for interactive error detection. In Furtado and Morgan [1885], pages 32–51. LCCN QA76.9.D3

                                        I55 1979. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/WilsonS79.
                                        html.

Osborn:1979:TUR

[202] Sylvia L. Osborn. Towards a universal relation interface. In Furtado and Morgan [1885], pages 52–60. LCCN QA76.9.D3 I55 1979. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Osborn79.html.

Lien:1979:MDN

[203] Y. Edmund Lien. Multivalued dependencies with null values in relational data bases. In Furtado and Morgan [1885], pages 61–66. LCCN QA76.9.D3 I55 1979. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Lien79.html.

Parker:1979:AAN

[204] D. Stott Parker and Claude Delobel. Algorithmic applications for a new result on multivalued dependencies. In Furtado and Morgan [1885], pages 67–74. LCCN QA76.9.D3

                                        I55 1979. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/ParkerD79.
                                        html.

Mendelzon:1979:GMD

[205] Alberto O. Mendelzon and David Maier. Generalized mutual dependencies and the decomposition of database relations. In Furtado and Morgan [1885], pages 75–82. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MendelzonM79.html.

dosSantos:1979:QCS

[206] Clesio Saraiva dos Santos. The quest for comprehensive semantic formalisms. In Furtado and Morgan [1885], pages 83–84. LCCN QA76.9.D3

I55 1979. URL http://www.vldb.
org/dblp/db/conf/vldb/Santos79.
html.

Studer:1979:FSD

[207] Rudi Studer. Functional specification of a decision support system. In Furtado and Morgan [1885], pages 85–93. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/Studer79.html.

Lucas:1979:IID

[208] Ana Lucas. Implementation of the IRRD data base in portugal. In Furtado and Morgan [1885], pages 94– 100. LCCN QA76.9.D3 I55 1979.

URL http://www.vldb.org/dblp/
db/conf/vldb/Lucas79.html.

Sandewall:1979:DLP

[209] Erik Sandewall. A description language and pilot-system executive for information-transport systems. In Furtado and Morgan [1885], pages 101– 110. LCCN QA76.9.D3 I55 1979.

URL http://www.vldb.org/dblp/
db/conf/vldb/Sandewall79.html.

Solvberg:1979:SRD

[210] Arne Sølvberg. Software requirement definition and data models. In Furtado and Morgan [1885], pages 111– 118. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Solvberg79.html.

Clemons:1979:ESF

[211] Eric K. Clemons. An external schema facility for CODASYL 1978. In Furtado and Morgan [1885], pages 119– 128. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Clemons79.html.

Bubenko:1979:RUM

[212] Janis A. Bubenko, Jr. On the role of ‘understanding models’ in conceptual schema design. In Furtado and Morgan [1885], pages 129–139. LCCN QA76.9.D3 I55 1979. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Bubenko79.html.

Oren:1979:SUC

[213] Ole Oren and Frode Aschim. Statistics for the usage of a conceptual data model as a basis for logical data base design. In Furtado and Morgan [1885], pages 140–145. LCCN QA76.9.D3 I55

                                        1979. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/OrenA79.html.

Arditi:1979:OBM

[214] Joel Arditi. An optimized backout mechanism for sequential updates. In Furtado and Morgan [1885], pages 147–154. LCCN QA76.9.D3 I55

                                        1979. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Arditi79.html.

Svensson:1979:SPC

[215] Per Svensson. On search performance for conjunctive queries in compressed, fully transposed ordered files. In Furtado and Morgan [1885], pages 155– 163. LCCN QA76.9.D3 I55 1979.

URL http://www.vldb.org/dblp/
db/conf/vldb/Svensson79.html.

Reuter:1979:MOU

[216] Andreas Reuter. Minimizing the I/O-operations for undo-logging in database systems. In Furtado and Morgan [1885], pages 164–172. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Reuter79.html.

Ghosh:1979:SVL

[217] Sakti P. Ghosh, Alfonso F. Cardenas, Ignacio Mijares, and Gio Wiederhold. Some very large data bases in developing countries. In Furtado and Morgan [1885], pages 173–182. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GhoshCMW79.html.

Bach:1979:ASG

[218] Maurice J. Bach, Nancy H. Coguen, and Michael M. Kaplan. The ADAPT system: a generalized approach towards data conversion. In Furtado and Morgan [1885], pages 183–193. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BachCK79.html.

Housel:1979:FDM

[219] Barron C. Housel, Vance E. Waddle, and S. Bing Yao. The functional dependency model for logical database design. In Furtado and Morgan [1885], pages 194–208. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HouselWY79.html.

Melo:1979:MIC

[220] Rubens N. Melo. Monitoring integrity constraints in a CODASYL-like DBMS. In Furtado and Morgan [1885], pages 209–218. LCCN QA76.9.D3 I55

                                        1979. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Melo79.html.

Irani:1979:DDP

[221] Keki B. Irani, Subir Purkayastha, and Toby J. Teorey. A designer for DBMSprocessable logical database structures. In Furtado and Morgan [1885], pages 219–231. LCCN QA76.9.D3

                                        I55 1979. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/IraniPT79.
                                        html.

Kent:1979:EJ

[222] William Kent. The entity join. In Furtado and Morgan [1885], pages 232–238. LCCN QA76.9.D3 I55

                                        1979. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Kent79.html.

Pirotte:1979:FSI

[223] Alain Pirotte. Fundamental and secondary issues in the design of nonprocedural relational languages. In Furtado and Morgan [1885], pages 239– 250. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Pirotte79.html.

Rolland:1979:TIS

[224] Colette Rolland, S. Leifert, and Christian Richard. Tools for information system dynamics management. In Furtado and Morgan [1885], pages 251– 261. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/RollandLR79.html.

Baldissera:1979:ISF

[225] C. Baldissera, Stefano Ceri, Giuseppe Pelagatti, and Giampio Bracchi. Interactive specification and formal verification of user’s views in data bases design. In Furtado and Morgan [1885], pages 262–272. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BaldisseraCPB79.html.

Luque:1979:DCP

[226] Emilio Luque, José J. Ruz, Ana Ripoll, and Alfredo Bautista. Database concurrent processor. In Furtado and Morgan [1885], pages 273–279. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LuqueRRB79.html.

Downs:1979:DBM

[227] Deborah Downs and Gerald J. Popek. Data base management systems security and INGRES. In Furtado and Morgan [1885], pages 280–290. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DownsP79.html.

Gardarin:1979:PCD

[228] Georges Gardarin and Michel A. Melkanoff. Proving consistency of database transactions. In Furtado and Morgan [1885], pages 291–298. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GardarinM79.html.

Taylor:1979:DPC

[229] Robert W. Taylor, James P. Fry, Ben Shneiderman, Diane C. P. Smith, and Stanley Y. W. Su. Database program conversion: a framework for research. In Furtado and Morgan [1885], pages 299–312. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/TaylorFSSS79.
html.

Quintella:1979:MGL

[230] H. M. Quintella, G. A. de Gusmão, M. R. de Oliveira, L. T. de Albuquerque, and F. R. Leão. A methodologic guide-line for the installation of data base development tools and data administration tools: Two case studies from Brazil. In Furtado and Morgan [1885], pages 313–318. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
QuintellaGOAL79.html.

Turner:1979:DLS

[231] M. J. Turner, R. Hammond, and P. Cotton. A DBMS for large statistical databases. In Furtado and Morgan [1885], pages 319–327. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TurnerHC79.html.

Lum:1979:NOD

[232] Vincent Y. Lum, Sakti P. Ghosh, Mario Schkolnick, R. W. Taylor, D. Jefferson, Stanley Y. W. Su, James P. Fry, Toby J. Teorey, B. Yao, D. S. Rund, B. Kahn, Shamkant B. Navathe, D. Smith, L. Aguilar, W. J. Barr, and P. E. Jones. 1978 new orleans data base design workshop report. In Furtado and Morgan [1885], pages 328–339. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LumGSTJSFTYRKNSABJ79.html.

Breitwieser:1979:TCM

[233] H. Breitwieser and U. Kersten. Transaction and catalog management of the distributed file management system DISCO. In Furtado and Morgan [1885], pages 340–350. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BreitwieserK79.html.

Kung:1979:OMC

[234] H. T. Kung and John T. Robinson. On optimistic methods for concurrency control. In Furtado and Morgan [1885], pages 351–?? LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/KungR79.html.

Wood:1979:DAD

[235] C. Wood and Eduardo B. Fernández. Decentralized authorization in a database system. In Furtado and Morgan [1885], pages 352–359. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WoodF79.html.

Kawazu:1979:TPD

[236] Seiichi Kawazu, Susumu Minami, Kenji Itoh, and Katsuni Teranaka. Twophase deadlock detection algorithm in distributed databases. In Furtado and Morgan [1885], pages 360–367. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KawazuMIT79.html.

Buneman:1979:PMP

[237] Peter Buneman. The problem of multiple paths in a database schema. In Furtado and Morgan [1885], pages 368– 372. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Buneman79.html.

Hardgrave:1979:APB

[238] W. Terry Hardgrave. Ambiguity in processing Boolean queries on TDMS tree structures: a study of four different philosophies. In Furtado and Morgan [1885], pages 373–397. LCCN QA76.9.D3 I55 1979. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Hardgrave79.html.

Schmidt:1979:PPR

[239] Joachim W. Schmidt. Parallel processing of relations: a single-assignment approach. In Furtado and Morgan [1885], pages 398–408. LCCN QA76.9.D3

                                        I55 1979. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Schmidt79.
                                        html.

Youssefi:1979:QPR

[240] Karel Youssefi and Eugene Wong. Query processing in a relational database management system. In Furtado and Morgan [1885], pages 409– 417. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/YoussefiW79.html.

Merrett:1979:DMR

[241] T. H. Merrett and Ekow J. Otoo. Distribution models of relations. In Furtado and Morgan [1885], pages 418– 425. LCCN QA76.9.D3 I55 1979.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/MerrettO79.html.

Munz:1979:ASP

[242] Rudolf Munz, H.-J. Schneider, and Frank Steyer. Application of subpredicate tests in database systems.

In Furtado and Morgan [1885], pages 426–435. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/MunzSS79.html.

Setzer:1979:PDT

[243] V. W. Setzer. Program development by transformations applied to relational database queries. In Furtado and Morgan [1885], pages 436–443. LCCN QA76.9.D3 I55

1979. URL http://www.vldb.org/
dblp/db/conf/vldb/Setzer79.html.

Demolombe:1979:SCQ

[244] Robert Demolombe. Semantic checking of questions expressed in predicate calculus language. In Furtado and Morgan [1885], pages 444–450. LCCN QA76.9.D3 I55 1979. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Demolombe79.html.

Gotlieb:1980:SLQ

[245] C. C. Gotlieb. Some large questions about very large data bases. In Canadian Information Processing Society [1887], pages 3–7. LCCN QA76.9.D3

.I559 1980. URL http://www.vldb.
org/dblp/db/conf/vldb/Gotlieb80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Kent:1980:SCS

[246] William Kent. Splitting the conceptual schema. In Canadian Information Processing Society [1887], pages 10–14. LCCN QA76.9.D3 .I559 1980.

URL http://www.vldb.org/dblp/
db/conf/vldb/Kent80.html. Spon-

sored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Date:1980:IUD

[247] C. J. Date. An introduction to the unified database language (UDL). In Canadian Information Processing Society [1887], pages 15–32. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Date80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Shneiderman:1980:PEC

[248] Ben Shneiderman and Glenn Thomas. Path expressions for complex queries and automatic database program conversion. In Canadian Information Processing Society [1887], pages 33– 44. LCCN QA76.9.D3 .I559 1980.

URL http://www.vldb.org/dblp/
db/conf/vldb/ShneidermanT80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Walker:1980:RSV

[249] Adrian Walker. On retrieval from a small version of a large data base. In Canadian Information Processing Society [1887], pages 47–54. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Walker80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Demolombe:1980:ENT

[250] Robert Demolombe. Estimation of the number of tuples satisfying a query expressed in predicate calculus language. In Canadian Information Processing Society [1887], pages 55–63. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Demolombe80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Rosenkrantz:1980:PCP

[251] Daniel J. Rosenkrantz and Harry B. Hunt III. Processing conjunctive predicates and queries. In Canadian Information Processing Society [1887], pages 64–72. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RosenkrantzH80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Bihan:1980:SFN

[252] Jean Le Bihan, Christian Esculier, Gerard Le Lann, Witold Litwin, Georges Gardarin, S. Sedillort, and L. Treille. SIRIUS: a French nationwide project on distributed data bases. In Canadian Information Processing Society [1887], pages 75–85. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BihanELLGST80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Adiba:1980:DS

[253] Michel E. Adiba and Bruce G. Lindsay. Database snapshots. In Canadian Information Processing Society [1887], pages 86–91. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/AdibaL80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Epstein:1980:ADD

[254] Robert S. Epstein and Michael Stonebraker. Analysis of distributed data base processing strategies. In Canadian Information Processing Society [1887], pages 92–101. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

EpsteinS80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Brodie:1980:DAD

[255] Michael L. Brodie. Data abstraction, databases and conceptual modelling. In Canadian Information Processing Society [1887], pages 105–108. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Brodie80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Wilson:1980:CMS

[256] Gerald A. Wilson. A conceptual model for semantic integrity checking. In Canadian Information Processing Society [1887], pages 111–125. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Wilson80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Bernstein:1980:FMS

[257] Philip A. Bernstein, Barbara T. Blaustein, and Edmund M. Clarke. Fast maintenance of semantic integrity assertions using redundant aggregate data. In Canadian Information Processing Society [1887], pages 126–136. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BernsteinBC80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Hammer:1980:KBQ

[258] Michael Hammer and Stanley B. Zdonik. Knowledge-based query processing. In Canadian Information Processing Society [1887], pages 137–147. LCCN QA76.9.D3

                                        .I559 1980. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/HammerZ80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Spewak:1980:PAD

[259] Steven H. Spewak. A pragmatic approach to database design. In Canadian Information Processing Society [1887], pages 151–152. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Spewak80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Jefferson:1980:DAD

[260] David K. Jefferson. The development and application of data base design tools and methodology. In Canadian Information Processing Society [1887], pages 153–154. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Jefferson80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Freilich:1980:MSM

[261] Lenny Freilich and Paul J. Sadowski. Managing systems that manage data. In Canadian Information Processing Society [1887], page 155. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FreilichS80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Post:1980:QPH

[262] Jonathan V. Post. Quintillabit: Parameters of a hyperlarge database. In Canadian Information Processing Society [1887], pages 156–158. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Post80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Tsichritzis:1980:OIF

[263] Dennis Tsichritzis. OFS: An integrated form management system. In Canadian Information Processing Society [1887], pages 161–166. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Tsichritzis80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Studer:1980:DID

[264] Rudi Studer. A dialogue interface for data base applications. In Canadian Information Processing Society [1887], pages 167–182. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Studer80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Wilson:1980:SVG

[265] Gerald A. Wilson and Christopher F. Herot. Semantics vs. graphics — to show or not to show. In Canadian Information Processing Society [1887], pages 183–197. LCCN QA76.9.D3

                                        .I559 1980. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/WilsonH80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Gardarin:1980:DDS

[266] Georges Gardarin, Nathan Goodman, Bruce G. Lindsay, Rudolf Munz, and James B. Rothnie. Distributed database systems. In Canadian Information Processing Society [1887], pages 201–202. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GardarinGLMR80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Eggers:1980:EAC

[267] Susan J. Eggers and Arie Shoshani. Efficient access of compressed data. In Canadian Information Processing Society [1887], pages 205–211. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

EggersS80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Litwin:1980:LHN

[268] Witold Litwin. Linear hashing: a new tool for file and table addressing. In Canadian Information Processing Society [1887], pages 212–223. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Litwin80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Larson:1980:LHP

[269] Per-Åke Larson. Linear hashing with partial expansions. In Canadian Information Processing Society [1887], pages 224–232. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Larson80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Walker:1980:DP

[270] W. A. Walker. Data planning. In Canadian Information Processing Society [1887], page 235. LCCN QA76.9.D3

                                        .I559 1980. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Walker80a.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Honeyman:1980:EJ

[271] Peter Honeyman. Extension joins. In Canadian Information Processing Society [1887], pages 239–244. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Honeyman80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Bernstein:1980:WDB

[272] Philip A. Bernstein and Nathan Goodman. What does Boyce-Codd Normal Form do? In Canadian Information Processing Society [1887], pages 245–259. LCCN QA76.9.D3 .I559

1980. URL http://www.vldb.org/
dblp/db/conf/vldb/BernsteinG80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Vassiliou:1980:FDI

[273] Yannis Vassiliou. Functional dependencies and incomplete information. In Canadian Information Processing Society [1887], pages 260–269. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Vassiliou80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Bayer:1980:DCC

[274] Rudolf Bayer, Klaus Elhardt, Hans Heller, and Angelika Reiser. Distributed concurrency control in database systems. In Canadian Information Processing Society [1887], pages 275– 284. LCCN QA76.9.D3 .I559 1980.

URL http://www.vldb.org/dblp/

                                        db/conf/vldb/BayerEHR80.html. Spon-

sored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Bernstein:1980:TBA

[275] Philip A. Bernstein and Nathan Goodman. Timestamp-based algorithms for concurrency control in distributed database systems. In Canadian Information Processing Society [1887], pages 285–300. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BernsteinG80a.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Cheng:1980:USD

[276] Wing Kai Cheng and Geneva G. Belford. Update synchronization in distributed databases. In Canadian Information Processing Society [1887], pages 301–308. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Cheng80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Kedem:1980:NTP

[277] Zvi M. Kedem and Abraham Silberschatz. Non-two-phase locking protocols with shared and exclusive locks. In Canadian Information Processing Society [1887], pages 309–317. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

KedemS80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Steel:1980:SRI

[278] Thomas B. Steel, Jr. Status report on ISO/TC97/SC5/WG3 — data base management systems. In Canadian Information Processing Society [1887], pages 321–325. LCCN QA76.9.D3 .I559 1980. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Steel80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Brodie:1980:SRA

[279] Michael L. Brodie. Standardization and the relational approach to databases: An ANSI task group status report. In Canadian Information Processing Society [1887], pages 326–327. LCCN QA76.9.D3

.I559 1980. URL http://www.vldb.
org/dblp/db/conf/vldb/Brodie80a.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Peeters:1980:EEI

[280] Emile Peeters. EEV evaluation and implementation of database systems. In Canadian Information Processing Society [1887], pages 328–329. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Peeters80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Phillips:1980:THF

[281] Dorothy A. Phillips. Telidon and the human factors of Videotex data bases. In Canadian Information Processing Society [1887], pages 330–331. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Phillips80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Tamir:1980:DDB

[282] M. Tamir, T. Missinai, Joel Arditi, E. Raban, S. Sole, S. Timor, and E. Zukovsky. DB1 — A DBMS-based application generator. In Canadian Information Processing Society [1887], pages 335–342. LCCN QA76.9.D3 .I559

                                        1980. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/TamirMARSTZ80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Risch:1980:PPG

[283] Tore Risch. Production program generation in a flexible data dictionary system. In Canadian Information Processing Society [1887], pages 343– 349. LCCN QA76.9.D3 .I559 1980.

URL http://www.vldb.org/dblp/
db/conf/vldb/Risch80.html. Spon-

sored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Navathe:1980:IAN

[284] Shamkant B. Navathe. An intuitive approach to normalize network structured data. In Canadian Information Processing Society [1887], pages 350–358. LCCN QA76.9.D3

.I559 1980. URL http://www.vldb.
org/dblp/db/conf/vldb/Navathe80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Swartwout:1980:DTR

[285] D. E. Swartwout. Database theory and its relationship to practice. In Canadian Information Processing Society [1887], page 361. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Swartwout80.html. Sponsored by

IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Menasce:1980:DRS

[286] Daniel A. Menascé and Oscar E. Landes. On the design of a reliable storage component for distributed database management systems. In Canadian Information Processing Society [1887], pages 365–375. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

MenasceL80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Badal:1980:AEC

[287] Dushan Z. Badal. The analysis of the effects of concurrency control on distributed database system performance. In Canadian Information Processing Society [1887], pages 376–383. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Badal80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Walter:1980:SHT

[288] Bernd Walter. Strategies for handling transactions in distributed data base systems during recovery. In Canadian Information Processing Society [1887], pages 384–389. LCCN QA76.9.D3 .I559

1980. URL http://www.vldb.org/
dblp/db/conf/vldb/Walter80.html.

Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Hawthorn:1980:DM

[289] Paula B. Hawthorn. Database machines. In Canadian Information Processing Society [1887], pages 393– 395. LCCN QA76.9.D3 .I559 1980.

URL http://www.vldb.org/dblp/
db/conf/vldb/Hawthorn80.html. Spon-

sored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Mylopoulos:1980:SFT

[290] John Mylopoulos and Harry K. T. Wong. Some features of the TAXIS data model. In Canadian Information Processing Society [1887], pages 399–410. LCCN QA76.9.D3 .I559

1980. URL http://www.vldb.org/
dblp/db/conf/vldb/MylopoulosW80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Spyratos:1980:TSR

[291] Nicolas Spyratos. Translation structures of relational views. In Canadian Information Processing Society [1887], pages 411–416. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Spyratos80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Nambiar:1980:SAT

[292] K. K. Nambiar. Some analytic tools for the design of relational database systems. In Canadian Information Processing Society [1887], pages 417–428. LCCN QA76.9.D3

                                        .I559 1980. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Nambiar80.

html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Deutsch:1980:DSC

[293] Donald R. Deutsch and Eric K. Clemons. DBMS standards: Current status and future directions. In Canadian Information Processing Society [1887], pages 431–433. LCCN QA76.9.D3 .I559 1980. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

DeutschC80.html. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Date:1981:RI

[294] C. J. Date. Referential integrity. In Zaniolo and Delobel [1888], pages 2– 12. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Date81.html. IEEE

catalog no. 81CH1701-2.

Kent:1981:DMT

[295] William Kent. Data model theory meets a practical application. In Zaniolo and Delobel [1888], pages 13– 22. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Kent81.html. IEEE

catalog no. 81CH1701-2.

DeAntonellis:1981:MED

[296] Valeria De Antonellis and Bruna Zonta. Modelling events a data base application design. In Zaniolo and Delobel [1888], pages 23–31. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AntonellisZ81.html. IEEE catalog

no. 81CH1701-2.

Brodie:1981:MBS

[297] Michael L. Brodie. On modelling behavioural semantics of databases. In Zaniolo and Delobel [1888], pages 32– 42. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Brodie81.html. IEEE

catalog no. 81CH1701-2.

Menon:1981:DAR

[298] M. J. Menon and David K. Hsiao. Design and analysis of a relational join operation for VLSI. In Zaniolo and Delobel [1888], pages 44–55. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

MenonH81.html. IEEE catalog no. 81CH1701-2.

Armisen:1981:CBE

[299] J. P. Armisen and J. Y. Caleca. A commercial back-end data base system. In Zaniolo and Delobel [1888], pages 56–65. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ArmisenC81.html. IEEE catalog no.

81CH1701-2.

Pramanik:1981:HON

[300] Sakti Pramanik. Hardware organization for nonnumeric processing. In Zaniolo and Delobel [1888], pages 66– 75. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Pramanik81.html. IEEE

catalog no. 81CH1701-2.

Muraszkiewicz:1981:CSP

[301] M. Muraszkiewicz. Concepts of sorting and projection in a cellular array. In Zaniolo and Delobel [1888], pages 76–80. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Muraszkiewicz81.html. IEEE catalog

no. 81CH1701-2.

Yannakakis:1981:AAD

[302] Mihalis Yannakakis. Algorithms for acyclic database schemes. In Zaniolo and Delobel [1888], pages 82–94. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Yannakakis81.html. IEEE catalog no.

81CH1701-2.

Chase:1981:JGA

[303] Karen Chase. Join graphs and acyclic database schemes. In Zaniolo and Delobel [1888], pages 95–100. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Chase81.html. IEEE catalog no. 81CH1701-2.

Goldstein:1981:CNV

[304] Billie S. Goldstein. Constraints on null values in relational databases. In Zaniolo and Delobel [1888], pages 101– 110. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Goldstein81.html. IEEE

catalog no. 81CH1701-2.

Fussell:1981:TCL

[305] Donald S. Fussell, Zvi M. Kedem, and Abraham Silberschatz. A theory of correct locking protocols for database systems. In Zaniolo and Delobel [1888], pages 112–124. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FussellKS81.html. IEEE catalog no.

81CH1701-2.

Schlageter:1981:OMC

[306] Gunter Schlageter. Optimistic methods for concurrency control in distributed database systems. In Zaniolo and Delobel [1888], pages 125–130. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Schlageter81.html. IEEE catalog no.

81CH1701-2.

Tanaka:1981:LIL

[307] Katsumi Tanaka and Yahiko Kambayashi. Logical integration of locally independent relational databases into a distributed database. In Zaniolo and Delobel [1888], pages 131–141. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TanakamK81.html. IEEE catalog no.

81CH1701-2.

Gray:1981:TCV

[308] Jim Gray. The transaction concept: Virtues and limitations (invited paper). In Zaniolo and Delobel [1888], pages 144–154. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Gray81.html. IEEE catalog no. 81CH1701-2.

Borr:1981:TME

[309] Andrea J. Borr. Transaction monitoring in ENCOMPASS: Reliable distributed transaction processing. In Zaniolo and Delobel [1888], pages 155– 165. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Borr81.html. IEEE

catalog no. 81CH1701-2.

Beeri:1981:RCI

[310] Catriel Beeri and Ron Obermarck. A resource class independent deadlock detection algorithm. In Zaniolo and Delobel [1888], pages 166–178. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BeeriO81.html. IEEE catalog no. 81CH1701-2.

Fields:1981:UIP

[311] C. Fields. User interfaces for pictorial data bases. In Zaniolo and Delobel [1888], page 180. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.
org/dblp/db/conf/vldb/Fields81.

html. IEEE catalog no. 81CH1701-2.

Sevcik:1981:DBS

[312] Kenneth C. Sevcik. Data base system performance prediction using an analytical model (invited paper). In Zaniolo and Delobel [1888], pages 182–198. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/Sevcik81.html.

IEEE catalog no. 81CH1701-2.

DeWitt:1981:PED

[313] David J. DeWitt and Paula B. Hawthorn. A performance evaluation of data base machine architectures (invited paper). In Zaniolo and Delobel [1888], pages 199–214. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

DeWittH81.html. IEEE catalog no. 81CH1701-2.

Hainaut:1981:TPT

[314] Jean-Luc Hainaut. Theoretical and practical tools for data base design. In Zaniolo and Delobel [1888], pages 216–224. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.
org/dblp/db/conf/vldb/Hainaut81.

html. IEEE catalog no. 81CH1701-2.

Sakai:1981:MDI

[315] Hirotaka Sakai. A method for defining information structures and transactions in conceptual schema design. In Zaniolo and Delobel [1888], pages 225–234. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Sakai81.html. IEEE

catalog no. 81CH1701-2.

Leonard:1981:ISD

[316] Michel Léonard and B. T. Luong. Information systems design approach integrating data and transactions. In Zaniolo and Delobel [1888], pages 235–246. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LeonardL81.html. IEEE catalog no.

81CH1701-2.

Gallaire:1981:ILD

[317] Hervé Gallaire. Impacts of logic and databases (invited paper). In Zaniolo and Delobel [1888], pages 248– 259. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Gallaire81.html. IEEE

catalog no. 81CH1701-2.

Borgida:1981:DMD

[318] Alexander Borgida and Harry K. T. Wong. Data models and data manipulation languages: Complementary semantics and proof theory. In Zaniolo and Delobel [1888], pages 260–271. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BorgidaW81.html. IEEE catalog no.

81CH1701-2.

Warren:1981:EPI

[319] David H. D. Warren. Efficient processing of interactive relational data base queries expressed in logic. In Zaniolo and Delobel [1888], pages 272–281. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/Warren81.html.

IEEE catalog no. 81CH1701-2.

Su:1981:DBM

[320] Stanley Y. W. Su. Data base machines: System evaluation and conversion issues. In Zaniolo and Delobel [1888], page 284. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.
org/dblp/db/conf/vldb/Su81.html.

IEEE catalog no. 81CH1701-2.

Cammarata:1981:DUR

[321] Stephanie J. Cammarata. Deferring updates in a relational data base system. In Zaniolo and Delobel [1888], pages 286–292. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Cammarate81.html. IEEE catalog no.

81CH1701-2.

Adiba:1981:DRU

[322] Michel E. Adiba. Derived relations: a unified mechanism for views, snapshots, and distributed data. In Zaniolo and Delobel [1888], pages 293– 305. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/Adiba81.html. IEEE

catalog no. 81CH1701-2.

Koenig:1981:TFA

[323] Shaye Koenig and Robert Paige. A transformational framework for the automatic control of derived data. In Zaniolo and Delobel [1888], pages 306–318. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.
org/dblp/db/conf/vldb/KoenigP81.

html. IEEE catalog no. 81CH1701-2.

Whang:1981:SAP

[324] Kyu-Young Whang, Gio Wiederhold, and Daniel Sagalowicz. Separability — an approach to physical data base design. In Zaniolo and Delobel [1888], pages 320–332. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.
org/dblp/db/conf/vldb/WhangWS81.

html. IEEE catalog no. 81CH1701-2.

Lomet:1981:DBT

[325] David B. Lomet. Digital B-trees. In Zaniolo and Delobel [1888], pages 333– 344. LCCN QA76.9.D3 I559 1981.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Lomet81.html. IEEE

catalog no. 81CH1701-2.

Karlsson:1981:RCT

[326] Kjell Karlsson. Reduced cover-trees and their application in the Sabre access path model. In Zaniolo and Delobel [1888], pages 345–353. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Karlsson81.html. IEEE catalog no.

81CH1701-2.

Tsichritzis:1981:IDB

[327] Dennis Tsichritzis. Integrating data base and message systems (invited paper). In Zaniolo and Delobel [1888], pages 356–362. LCCN QA76.9.D3 I559

                                        1981. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Tsichritzis81.

html. IEEE catalog no. 81CH1701-2.

Richter:1981:IIN

[328] Gernot Richter. IML-inscribed nets for modeling text processing and data (base) management systems. In Zaniolo and Delobel [1888], pages 363–375. LCCN QA76.9.D3 I559

                                        1981.  URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Richter81.

html. IEEE catalog no. 81CH1701-2.

Montgomery:1981:AIS

[329] Christine A. Montgomery and Enrique H. Ruspini. The active information system: a data-driven system for the analysis of imprecise data. In Zaniolo and Delobel [1888], pages 376–384. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/MontgomeryR81.

html. IEEE catalog no. 81CH1701-2.

Ries:1981:DDP

[330] Daniel R. Ries. Distributed data processing versus distributed data base processing. In Zaniolo and Delobel [1888], page 386. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Ries81.html. IEEE catalog no. 81CH1701-2.

Imielinski:1981:RII

[331] Tomasz Imielinski and Witold Lipski, Jr. On representing incomplete information in a relational data base. In Zaniolo and Delobel [1888], pages 388–397. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/ImielinskiL81.

html. IEEE catalog no. 81CH1701-2.

Bancilhon:1981:ICD

[332] François Bancilhon and Nicolas Spyratos. Independent components of databases. In Zaniolo and Delobel [1888], pages 398–408. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/BancilhonS81.

html. IEEE catalog no. 81CH1701-2.

Veloso:1981:SDC

[333] Paulo A. S. Veloso, José Mauro V. de Castilho, and Antonio L. Furtado. Systematic derivation of complementary specifications. In Zaniolo and Delobel [1888], pages 409–421. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
VelosoCF81.html. IEEE catalog no.

81CH1701-2.

Eggers:1981:CTL

[334] Susan J. Eggers, Frank Olken, and Arie Shoshani. A compression technique for large statistical data-bases. In Zaniolo and Delobel [1888], pages 424–434. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        EggersOS81.html. IEEE catalog no.

81CH1701-2.

Lynch:1981:ADC

[335] Clifford A. Lynch and E. B. Brownrigg. Application of data compression to a large bibliographic data base. In Zaniolo and Delobel [1888], pages 435–447. LCCN QA76.9.D3 I559

                                        1981. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/LynchB81.html.

IEEE catalog no. 81CH1701-2.

Frank:1981:ADL

[336] Andrew U. Frank. Application of DBMS to land information systems. In Zaniolo and Delobel [1888], pages 448–453. LCCN QA76.9.D3 I559 1981.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Frank81.html. IEEE

catalog no. 81CH1701-2.

Chamberlin:1981:HSR

[337] Donald D. Chamberlin, A. M. Gilbert, and Robert A. Yost. A history of System R and SQL/Data System (invited paper). In Zaniolo and Delobel [1888], pages 456–464. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChamberlinGY81.html. IEEE catalog

no. 81CH1701-2.

Lien:1981:DDS

[338] Y. Edmund Lien, Jonathan E. Shopiro, and Shalom Tsur. DSIS — A database system with interrelational semantics. In Zaniolo and Delobel [1888], pages 465–477. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/LienST81.html.

IEEE catalog no. 81CH1701-2.

Benneworth:1981:IGE

[339] R. L. Benneworth, C. D. Bishop, C. J. M. Turnbull, W. D. Holman, and F. M. Monette. The implementation of GERM, an entityrelationship data base management system. In Zaniolo and Delobel [1888], pages 478–484. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BenneworthBTHM81.html. IEEE cata-

log no. 81CH1701-2.

Weber:1981:DBS

[340] Herbert Weber. Data bases for software development. In Zaniolo and Delobel [1888], page 486. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Weber81.html. IEEE catalog no. 81CH1701-2.

Merrett:1981:SPF

[341] T. H. Merrett, Yahiko Kambayashi, and H. Yasuura. Scheduling of pagefetches in join operations. In Zaniolo and Delobel [1888], pages 488– 498. LCCN QA76.9.D3 I559 1981.

URL http://www.vldb.org/dblp/
db/conf/vldb/MerrettKY81.html. IEEE

catalog no. 81CH1701-2.

Soderlund:1981:CDB

[342] Lars Söderlund. Concurrent data base reorganization — assessment of a powerful technique through modeling. In Zaniolo and Delobel [1888], pages 499–509. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Soderlund81.html. IEEE catalog no.

81CH1701-2.

King:1981:QSS

[343] Jonathan J. King. QUIST: a system for semantic query optimization in relational databases. In Zaniolo and Delobel [1888], pages 510–517. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

King81.html. IEEE catalog no. 81CH1701-2.

Makinouchi:1981:OSQ

[344] Akifumi Makinouchi, Masayoshi Tezuka, Hajime Kitakami, and S. Adachi. The optimization strategy for query evaluation in RDB/V1. In Zaniolo and Delobel [1888], pages 518–529. LCCN QA76.9.D3 I559 1981. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        MakinouchiTKA81.html. IEEE catalog

no. 81CH1701-2.

Bussolati:1981:DAM

[345] U. Bussolati and Giancarlo Martella. A database approach to modelling and managing security information. In Zaniolo and Delobel [1888], pages 532–542. LCCN QA76.9.D3 I559

                                        1981. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BussolatiM81.

html. IEEE catalog no. 81CH1701-2.

Minsky:1981:SAD

[346] Naftaly H. Minsky. Synergistic authorization in database systems. In Zaniolo and Delobel [1888], pages 543–552. LCCN QA76.9.D3 I559

1981. URL http://www.vldb.org/
dblp/db/conf/vldb/Minsky81.html.

IEEE catalog no. 81CH1701-2.

Chan:1981:SDD

[347] Paul Chan and Arie Shoshani. SUB-JECT: a directory driven system for organizing and accessing large statistical databases. In Zaniolo and Delobel [1888], pages 553–563. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

ChanS81.html. IEEE catalog no. 81CH1701-2.

Curtice:1981:DDA

[348] Robert M. Curtice. Data dictionaries: An assessment of current practice and problems. In Zaniolo and Delobel [1888], pages 564–570. LCCN QA76.9.D3 I559 1981. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Curtice81.html. IEEE catalog no. 81CH1701-2.

Stonebraker:1982:TSR

[349] Michael Stonebraker and Joseph Kalash. TIMBER: a sophisticated relation browser (invited paper). In VLDB Endowment [1889], pages 1– 10. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
StonebrakerK82.html.

Lochovsky:1982:IQL

[350] Frederick H. Lochovsky and Dennis Tsichritzis. An interactive query language for external data bases. In VLDB Endowment [1889], pages 11– 21. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LochovskyT82.html.

Wong:1982:GGU

[351] Harry K. T. Wong and Ivy Kuo. GUIDE: Graphical user interface for database exploration. In VLDB Endowment [1889], pages 22–32. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/WongK82.html.

Andler:1982:SDD

[352] S. Andler, I. Ding, Kapali P. Eswaran, C. Hauser, Won Kim, James W. Mehl, and R. Williams. System D: a distributed system for availability. In VLDB Endowment [1889], pages 33– 44. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AndlerDEHKMW82.html.

Ferrier:1982:HDD

[353] Arlette Ferrier and Christine Stangret. Heterogeneity in the distributed database management system SIRIUS-DELTA. In VLDB Endowment [1889], pages 45–53. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

URL http://www.vldb.org/dblp/
db/conf/vldb/FerrierS82.html.

Chang:1982:HAD

[354] Jo-Mei Chang. A heuristic approach to distributed query processing. In VLDB Endowment [1889], pages 54–61. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/Chang82.html.

Bernstein:1982:SID

[355] Philip A. Bernstein and Nathan Goodman. A sophisticate’s introduction to distributed concurrency control (invited paper). In VLDB Endowment [1889], pages 62–76. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/BernsteinG82.
html.

Xu:1982:FMM

[356] J. Xu. A formal model for maximum concurrency in transaction systems with predeclared writesets. In VLDB Endowment [1889], pages 77–90. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/Xu82.html.

Kedem:1982:EDR

[357] Zvi M. Kedem, C. Mohan, and Abraham Silberschatz. An efficient deadlock removal scheme for non-two-phase locking protocols. In VLDB Endowment [1889], pages 91–97. ISBN 0-934613-14-1. LCCN QA76.9.D3

I559 1982. URL http://www.vldb.
org/dblp/db/conf/vldb/KedemMS82.
html.

Whang:1982:PDN

[358] Kyu-Young Whang, Gio Wiederhold, and Daniel Sagalowicz. Physical design of network model databases using the property of separability. In VLDB Endowment [1889], pages 98– 107. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WhangWS82.html.

Dahl:1982:IID

[359] Roland Dahl and Janis A. Bubenko, Jr. IDBD — an interactive design tool for CODASYL-DBTG-type data bases. In VLDB Endowment [1889], pages 108–121. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

                                        1982. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/DahlB82.html.

Chan:1982:SAS

[360] Arvola Chan, Sy Danberg, Stephen Fox, Wen-Te K. Lin, Anil Nori, and Daniel R. Ries. Storage and access structures to support a semantic data model. In VLDB Endowment [1889], pages 122–130. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/ChanDFLNR82.html.

LeDoux:1982:RBC

[361] Carol Helfgott LeDoux and D. Stott Parker. Reflections on Boyce-Codd Normal Form. In VLDB Endowment [1889], pages 131–141. ISBN 0-934613-14-1. LCCN QA76.9.D3

                                        I559 1982. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/LeDouxP82.
                                        html.

Navathe:1982:MVI

[362] Shamkant B. Navathe and Suresh G. Gadgil. A methodology for view integration in logical database design. In VLDB Endowment [1889], pages 142–164. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NavatheG82.html.

Kent:1982:CPD

[363] William Kent. Choices in practical data design. In VLDB Endowment [1889], pages 165–180. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/Kent82.html.

Neumann:1982:CTC

[364] Thomas Neumann and Christoph Hornung. Consistency and transactions in CAD database. In VLDB Endowment [1889], pages 181–188. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

URL http://www.vldb.org/dblp/
db/conf/vldb/NeumannH82.html.

Leiss:1982:RPM

[365] Ernst L. Leiss. Randomizing, A practical method for protecting statistical databases against compromise. In VLDB Endowment [1889], pages 189– 196. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Leiss82.html.

Schek:1982:DSI

[366] Hans-Jörg Schek and Peter Pistor. Data structures for an integrated data base management and information retrieval system. In VLDB Endowment [1889], pages 197–207. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/SchekP82.html.

Shoshani:1982:SDC

[367] Arie Shoshani. Statistical databases: Characteristics, problems, and some solutions. In VLDB Endowment [1889], pages 208–222. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Shoshani82.html.

Hawthorn:1982:MAT

[368] Paula B. Hawthorn. Microprocessor assisted tuple access, decompression and assembly for statistical database systems. In VLDB Endowment [1889], pages 223–233. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Hawthorn82.html.

McCarthy:1982:MML

[369] John L. McCarthy. Metadata management for large statistical databases. In VLDB Endowment [1889], pages 234–243. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        McCarthy82.html.

Navathe:1982:DCA

[370] Shamkant B. Navathe. Databases for computer aided design and manufacturing. In VLDB Endowment [1889], page 244. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Navathe82.html.

Deen:1982:IIS

[371] S. Misbah Deen. An implementation of impure surrogates. In VLDB Endowment [1889], pages 245–256. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

                                        1982. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Deen82.html.

Sacco:1982:MMB

[372] Giovanni Maria Sacco and Mario Schkolnick. A mechanism for managing the buffer pool in a relational database system using the hot set model. In VLDB Endowment [1889], pages 257–262. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/SaccoS82.html.

Bancilhon:1982:LPC

[373] François Bancilhon, Philippe Richard, and Michel Scholl. On line processing of compacted relations. In VLDB Endowment [1889], pages 263–269. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/BancilhonRS82.
html.

Fernandez-Villasenor:1982:DAD

[374] Yolanda Fernández-Villaseñor. Database applications in developing countries. In VLDB Endowment [1889], pages 270–272. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Fernandez-Villasenor82.html.

Balzer:1982:SBC

[375] Robert Balzer, David Dyer, M. Fehling, and S. Saunders. Specificationbased computing environments. In VLDB Endowment [1889], pages 273– 279. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BalzerDFS82.html.

Castilho:1982:TFD

[376] José Mauro V. de Castilho, Marco A. Casanova, and Antonio L. Furtado. A temporal framework for database specifications. In VLDB Endowment [1889], pages 280–291. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/

                                        dblp/db/conf/vldb/CastilhoCF82.
                                        html.

Lafue:1982:SID

[377] Gilles M. E. Lafue. Semantic integrity dependencies and delayed integrity checking. In VLDB Endowment [1889], pages 292–299. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

                                        1982. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Lafue82.html.

Larson:1982:SFV

[378] Per-Åke Larson. A single-file version of linear hashing with partial expansions. In VLDB Endowment [1889], pages 300–309. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

                                        1982. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Larson82.html.

Su:1982:PAT

[379] Stanley Y. W. Su and Krishna P. Mikkilineni. Parallel algorithms and their implementation in MICRONET. In VLDB Endowment [1889], pages 310–324. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SuM82.html.

Gelenbe:1982:SPR

[380] Erol Gelenbe and Danièle Gardy. The size of projections of relations satisfying a functional dependency. In VLDB Endowment [1889], pages 325– 333. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GelenbeG82.html.

Bjorner:1982:FDS

[381] Dines Bjørner and Hans Henrik Løvengreen. Formalization of database systems — and a formal definition of IMS (invited paper). In VLDB Endowment [1889], pages 334– 347. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BjornerL82.html.

Louis:1982:DDS

[382] Georges Louis and Alain Pirotte. A denotational definition of the semantics of DRC, A domain relational calculus. In VLDB Endowment [1889], pages 348–356. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/LouisP82.html.

Gangopadhyay:1982:SND

[383] Dipayan Gangopadhyay, Umeshwar Dayal, and James C. Browne. Semantics of network data manipulation languages: An object-oriented approach. In VLDB Endowment [1889], pages 357–369. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GangopadhyayDB82.html.

Dosch:1982:ASD

[384] Walter Dosch, Gianfranco Mascari, and Martin Wirsing. On the algebraic specification of databases. In VLDB Endowment [1889], pages 370– 385. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DoschMW82.html.

Kerschberg:1982:DDD

[385] Larry Kerschberg. The database/data dictionary interface. In VLDB Endowment [1889], pages 386–387. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559

1982. URL http://www.vldb.org/
dblp/db/conf/vldb/Kerschberg82.
html.

Tsichritzis:1983:MOF

[386] Dennis Tsichritzis, Stavros Christodoulakis, P. Economopoulos, Christos Faloutsos, A. Lee, D. Lee, J. Vandenbroeck, and Carson C. Woo. A multimedia office filing system. In Schkolnick and Thanos [1890], pages 2– 7. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TsichritzisCEFLLVW83.html. This

conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Bitton:1983:BDS

[387] Dina Bitton, David J. DeWitt, and Carolyn Turbyfill. Benchmarking database systems A systematic approach. In Schkolnick and Thanos [1890], pages 8–19. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BittonDT83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Penny:1983:ANT

[388] Philippe Penny and Michael Picard. Application of novel technologies to the management of a very large date base. In Schkolnick and Thanos [1890], pages 20–29. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

PennyP83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Meier:1983:SCE

[389] Andreas Meier and Raymond A. Lorie. A surrogate concept for engineering databases. In Schkolnick and Thanos [1890], pages 30–32. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/MeierL83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Morgenstern:1983:ADP

[390] Matthew Morgenstern. Active databases as a paradigm for enhanced computing environments. In Schkolnick and Thanos [1890], pages 34–42. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/Morgenstern83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Thompson:1983:BUM

[391] Craig W. Thompson, Kenneth M. Ross, Harry R. Tennant, and Richard M. Saenz. Building usable menubased natural language interfaces to databases. In Schkolnick and Thanos [1890], pages 43–55. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/ThompsonRTS83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Wilson:1983:MPS

[392] Gerald A. Wilson, Eric A. Domeshek, Ellen L. Drascher, and Jeffrey S. Dean. The multipurpose presentation system. In Schkolnick and Thanos [1890], pages 56–69. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

WilsonDDD83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Vassiliou:1983:HDE

[393] Yannis Vassiliou, James Clifford, and Matthias Jarke. How does an expert system get its data? (extended abstract). In Schkolnick and Thanos [1890], pages 70–72. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/VassiliouCJ83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Buckley:1983:OPP

[394] Gael N. Buckley and Abraham Silberschatz. Obtaining progressive protocols for a simple multiversion database model. In Schkolnick and Thanos [1890], pages 74– 80. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BuckleyS83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Reimer:1983:SPP

[395] Manuel Reimer. Solving the phantom problem by predicative optimistic concurrency control. In Schkolnick and Thanos [1890], pages 81–88. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/Reimer83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Ibaraki:1983:DIT

[396] Toshihide Ibaraki, Tiko Kameda, and Toshimi Minoura. Disjointinterval topological sort: a useful concept in serializability theory (extended abstract). In Schkolnick and Thanos [1890], pages 89– 91. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

IbarakiKM83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Dasgupta:1983:NTP

[397] Partha Dasgupta and Zvi M. Kedem. A non-two-phase locking protocol for concurrency control in general databases. In Schkolnick and Thanos [1890], pages 92–94. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

DasguptaK83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Peinl:1983:ECD

[398] Peter Peinl and Andreas Reuter. Empirical comparison of database concurrency schemes. In Schkolnick and Thanos [1890], pages 97–108. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/PeinlR83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Lin:1983:BTM

[399] Wen-Te K. Lin and Jerry Nolte. Basic timestamp, multiple version timestamp, and two-phase locking. In Schkolnick and Thanos [1890], pages 109–119. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LinN83.html. This conference is sponsored by VLDB Endowment and cosponsored by IFIP et al.

Kiessling:1983:QCL

[400] Werner Kießling and G. Landherr. A quantitative comparison of lockprotocols for centralized databases. In Schkolnick and Thanos [1890], pages 120–130. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KiesslingL83.html. This conference

is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Orenstein:1983:DHF

[401] Jack A. Orenstein. A dynamic hash file for random and sequential accessing. In Schkolnick and Thanos [1890], pages 132–141. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Orenstein83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Torenvliet:1983:ROT

[402] Leen Torenvliet and Peter van Emde Boas. The reconstruction and optimization of trie hashing functions. In Schkolnick and Thanos [1890], pages 142–156. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TorenvlietB83.html. This conference

is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Woodfill:1983:IHR

[403] John Woodfill and Michael Stonebraker. An implementation of hypothetical relations. In Schkolnick and Thanos [1890], pages 157–166. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

WoodfillS83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Cremers:1983:AIM

[404] Armin B. Cremers and G. Doman. AIM — an integrity monitor for the database system INGRES. In Schkolnick and Thanos [1890], pages 167– 170. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

CremersD83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Biskup:1983:URV

[405] Joachim Biskup and Hans Hermann Brüggemann. Universal relation views: a pragmatic approach. In Schkolnick and Thanos [1890], pages 172–185. ISBN 0-934613-15-X. LCCN QA 76.9

D3 I61 1983. URL http://www.vldb.
org/dblp/db/conf/vldb/BiskupB83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Grahne:1983:DDF

[406] Gösta Grahne and Kari-Jouko Räihä. Database decomposition into fourth normal form. In Schkolnick and Thanos [1890], pages 186–196. ISBN 0-934613-15-X. LCCN QA 76.9 D3

I61 1983. URL http://www.vldb.
org/dblp/db/conf/vldb/GrahneR83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Arisawa:1983:OPN

[407] Hiroshi Arisawa, Kunihiko Moriya, and Takao Miura. Operations and the properties on non-first-normal-form relational databases. In Schkolnick and Thanos [1890], pages 197– 204. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

ArisawaMM83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Brown:1983:LFD

[408] Robert Brown and D. Stott Parker. LAURA: a formal data model and her logical design methodology. In Schkolnick and Thanos [1890], pages 206– 218. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BrownP83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Tucherman:1983:PAS

[409] Luiz Tucherman, Antonio L. Furtado, and Marco A. Casanova. A pragmatic approach to structured database design. In Schkolnick and Thanos [1890], pages 219–231. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/TuchermanFC83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Albano:1983:MMC

[410] Antonio Albano, Maurizio Capaccioli, M. E. Occiuto, and Renzo Orsini. A modularization mechanism for conceptual modeling. In Schkolnick and Thanos [1890], pages 232– 240. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

AlbanoCOO83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Sacca:1983:DPC

[411] Domenico Saccà and Gio Wiederhold. Database partitioning in a cluster of processors. In Schkolnick and Thanos [1890], pages 242–247. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/SaccaW83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Yu:1983:FAD

[412] Clement T. Yu, M. K. Siu, K. Lam, and C. H. Chen. File allocation in distributed databases with interaction between files. In Schkolnick and Thanos [1890], pages 248– 259. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

YuSLC83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Lefons:1983:AAS

[413] Ezio Lefons, Alberto Silvestri, and Filippo Tangorra. An analytic approach to statistical databases. In Schkolnick and Thanos [1890], pages 260– 274. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LefonsST83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Muro:1983:FRI

[414] Shojiro Muro, Toshihide Ibaraki, Hidehiro Miyajima, and Toshiharu Hasegawa. File redundancy issues in distributed database systems. In Schkolnick and Thanos [1890], pages 275–277. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

MuroIMH83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Macleod:1983:MII

[415] Ian A. Macleod. A model for integrated information systems. In Schkolnick and Thanos [1890], pages 280–289. ISBN 0-934613-15-X. LCCN QA 76.9

D3 I61 1983. URL http://www.vldb.
org/dblp/db/conf/vldb/Macleod83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Pilote:1983:DMA

[416] Michel Pilote. A data modeling approach to simplify the design of user interfaces. In Schkolnick and Thanos [1890], pages 290–301. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/Pilote83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Beech:1983:IDM

[417] David Beech and J. Samuel Feldman. The integrated data model: a database perspective. In Schkolnick and Thanos [1890], pages 302– 304. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BeechF83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Krishnamurthy:1983:QEO

[418] Ravi Krishnamurthy, Stephen P. Morgan, and Moshé M. Zloof. Query-by-example: Operations on piecewise continuous data (extended abstract). In Schkolnick and Thanos [1890], pages 305–308. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KrishnamurthyMZ83.html. This con-

ference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Stocker:1983:TLS

[419] Peter M. Stocker and R. Cantié. A target logical schema: The ACS. In Schkolnick and Thanos [1890], pages 309–310. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

StockerC83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Verroust:1983:CWB

[420] Anne Verroust. Characterization of well-behaved database schemata and their update semantics. In Schkolnick and Thanos [1890], pages 312– 321. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Verroust83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Schiel:1983:AIT

[421] Ulrich Schiel. An abstract introduction to the temporal-hierarchic data model (THM). In Schkolnick and Thanos [1890], pages 322–330. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/Schiel83.html.

This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Golshani:1983:MSA

[422] Forouzan Golshani, T. S. E. Maibaum, and M. R. Sadler. A modal system of algebras for database specification and query/update language support. In Schkolnick and Thanos [1890], pages 331–339. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GolshaniMS83.html. This conference

is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Dayal:1983:PQG

[423] Umeshwar Dayal. Processing queries over generalization hierarchies in a multidatabase system. In Schkolnick and Thanos [1890], pages 342– 353. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Dayal83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Chan:1983:SSD

[424] Arvola Chan, Umeshwar Dayal, Stephen Fox, and Daniel R. Ries. Supporting a semantic data model in a distributed database system. In Schkolnick and Thanos [1890], pages 354–363. ISBN 0-934613-15-X. LCCN QA 76.9

D3 I61 1983. URL http://www.vldb.
org/dblp/db/conf/vldb/ChanDFR83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Lyngbaek:1983:AOS

[425] Peter Lyngbæk and Dennis McLeod. An approach to object sharing in distributed datbase systems. In Schkolnick and Thanos [1890], pages 364– 375. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LyngbaekM83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Bertino:1983:VMD

[426] Elisa Bertino, Laura M. Haas, and Bruce G. Lindsay. View management in distributed data base systems. In Schkolnick and Thanos [1890], pages 376–378. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

BertinoHL83.html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Subieta:1983:HLN

[427] Kazimierz Subieta. High-level navigational facilities for network and relational databases. In Schkolnick and Thanos [1890], pages 380–386. ISBN 0-934613-15-X. LCCN QA 76.9 D3

I61 1983. URL http://www.vldb.
org/dblp/db/conf/vldb/Subieta83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Demo:1983:PAC

[428] G. Barbara Demo. Program analysis for conversion from a navigation to a specification database interface. In Schkolnick and Thanos [1890], pages 387–398. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Demo83.html. This conference is sponsored by VLDB Endowment and cosponsored by IFIP et al.

Klopprogge:1983:MIP

[429] Manfred R. Klopprogge and Peter C. Lockemann. Modelling information preserving databases: Consequences of the concept of time. In Schkolnick and Thanos [1890], pages 399–416. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61

1983. URL http://www.vldb.org/
dblp/db/conf/vldb/KlopproggeL83.

html. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Stonebraker:1984:DPN

[430] Michael Stonebraker and Lawrence A. Rowe. Database portals: a new application program interface. In Dayal et al. [1891], pages 3–13. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/StonebrakerR84.
html.

Lyngbaek:1984:PDM

[431] Peter Lyngbæk and Dennis McLeod. A personal data manager. In Dayal et al. [1891], pages 14–25. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984.

URL http://www.vldb.org/dblp/
db/conf/vldb/LyngbaekM84.html.

Krishnamurthy:1984:QPP

[432] Ravi Krishnamurthy and Stephen P. Morgan. Query processing on personal computers: a pragmatic approach (extended abstract). In Dayal et al. [1891], pages 26– 29. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KrishnamurthyM84.html.

King:1984:SSK

[433] Roger King and Stephen Melville. Ski: a semantics-knowledgeable interface. In Dayal et al. [1891], pages 30–33. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/KingM84.html.

Grahne:1984:DSD

[434] Gösta Grahne. Dependency satisfaction in databases with incomplete information. In Dayal et al. [1891], pages 37–45. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Grahne84.html.

Abiteboul:1984:TRD

[435] Serge Abiteboul and Victor Vianu. Transactions in relational databases (preliminary report). In Dayal et al. [1891], pages 46–56. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/AbiteboulV84.html.

Kobayashi:1984:DMM

[436] Isamu Kobayashi. A data manipulation model: An extension of the alpha expression. In Dayal et al. [1891], pages 57–62. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Kobayashi84.html.

Tan:1984:LCC

[437] K. P. Tan. A less costly constraints checking for join dependency. In Dayal et al. [1891], pages 63–68. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Tan84.html.

Joo:1984:DDA

[438] Tan Tahn Joo, Tan Kah Poh, and Goh Ah Moi. Datadict — A data analysis and logical database design tool. In Dayal et al. [1891], pages 71–77. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/JooPM84.html.

Navathe:1984:RMS

[439] Shamkant B. Navathe, T. Sashidhar, and Ramez Elmasri. Relationship merging in schema integration. In Dayal et al. [1891], pages 78– 90. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NavatheSE84.html.

Kung:1984:TFD

[440] David Chenho Kung. A temporal framework for database specification and verification. In Dayal et al. [1891], pages 91–99. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/Kung84.html.

Ozsoyoglu:1984:SD

[441] Gultekin Özsoyoglu and Z. Meral Özsoyoglu. Statistical databases. In Dayal et al. [1891], page 103. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/OzsoyogluO84.
html.

Carey:1984:PCC

[442] Michael J. Carey and Michael Stonebraker. The performance of concurrency control algorithms for database management systems. In Dayal et al. [1891], pages 107–118. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/CareyS84.html.

Tay:1984:CPL

[443] Y. C. Tay and Rajan Suri. Choice and performance in locking for databases. In Dayal et al. [1891], pages 119– 128. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TayS84.html.

Harder:1984:EMS

[444] Theo Härder and Peter Peinl. Evaluating multiple server DBMS in general purpors operating system environments. In Dayal et al. [1891], pages 129–140. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HarderP84.html.

Shoshani:1984:CSD

[445] Arie Shoshani, Frank Olken, and Harry K. T. Wong. Characteristics of scientific databases. In Dayal et al. [1891], pages 147–160. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ShoshaniOW84.
                                        html.

Walter:1984:NTM

[446] Bernd Walter. Nested transactions with multiple commit points: An approach to the structuring of advanced database applications. In Dayal et al. [1891], pages 161–171. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Walter84.html.

Batory:1984:MOA

[447] Don S. Batory and Alejandro P. Buchmann. Molecular objects, abstract data types, and data models: a framework. In Dayal et al. [1891], pages 172–184. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BatoryB84.html.

DAtri:1984:EMD

[448] Alessandro D’Atri and Domenico Saccà. Equivalence and mapping of database schemes. In Dayal et al. [1891], pages 187–195. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/DAtriS84.html.

Beeri:1984:CAD

[449] Catriel Beeri and Michael Kifer. Comprehensive approach to the design of relational database schemes. In Dayal et al. [1891], pages 196–207. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/BeeriK84.html.

Ginsburg:1984:IQO

[450] Seymour Ginsburg and Katsumi Tanaka. Interval queries on object histories: Extended abstract. In Dayal et al. [1891], pages 208– 217. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GinsburgT84.html.

Zhu:1984:LGG

[451] Yun zhou Zhu. Line graph of gammaacyclic database schems and its recognition algorithm. In Dayal et al. [1891], pages 218–221. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/Zhu84.html.

Tanaka:1984:BSV

[452] Yuzuru Tanaka. Bit-sliced VLSI algorithm for search and sort. In Dayal et al. [1891], pages 225–234. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/Tanaka84.html.

Prabhakar:1984:TOD

[453] T. V. Prabhakar and H. V. Sahasrabuddhe. Towards an optimal data-structure: CB-trees. In Dayal et al. [1891], pages 235–244. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/PrabhakarS84.
html.

Yang:1984:DPH

[454] Wei-Pang Yang and M. W. Du. A dynamic perfect hash function defined by an extended hash indicator table. In Dayal et al. [1891], pages 245–254. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/YangD84.html.

Navathe:1984:DES

[455] Shamkant B. Navathe and Reind P. van de Riet. Databases for expert systems. In Dayal et al. [1891], page 257. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NavatheR84.html.

Christodoulakis:1984:DMI

[456] Stavros Christodoulakis, J. Vanderbroek, J. Li, T. Li, S. Wan, Y. Wang, M. Papa, and Elisa Bertino. Development of a multimedia information system for an office environment. In Dayal et al. [1891], pages 261– 271. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChristodoulakisVLLWWPB84.html.

Tagg:1984:EUA

[457] Roger M. Tagg. End user access to very large databases in an automated office/workstation environment. In Dayal et al. [1891], pages 272–279. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Tagg84.html.

Lamersdorf:1984:LSO

[458] Winfried Lamersdorf, G. Müller, and Joachim W. Schmidt. Language support for office modelling. In Dayal et al. [1891], pages 280– 288. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LamersdorfMS84.html.

Morgenstern:1984:CED

[459] Matthew Morgenstern. Constraint equations: Declarative expression of constraints with automatic enforcement. In Dayal et al. [1891], pages 291–300. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Morgenstern84.html.

Ehrich:1984:SSE

[460] Hans-Dieter Ehrich, Udo W. Lipeck, and Martin Gogolla. Specification, semantics, and enforcement of dynamic database constraints. In Dayal et al. [1891], pages 301–308. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984.

URL http://www.vldb.org/dblp/
db/conf/vldb/EhrichLG84.html.

Masunaga:1984:RDV

[461] Yoshifumi Masunaga. A relational database view update translation mechanism. In Dayal et al. [1891], pages 309–320. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984.

URL http://www.vldb.org/dblp/
db/conf/vldb/Masunaga84.html.

Bratbergsengen:1984:HMR

[462] Kjell Bratbergsengen. Hashing methods and relational algebra operations. In Dayal et al. [1891], pages 323– 333. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bratbergsengen84.html.

Rosenthal:1984:EAF

[463] Arnon Rosenthal and David S. Reiner. Extending the algebraic framework of query processing to handle outerjoins. In Dayal et al. [1891], pages 334– 343. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        RosenthalR84.html.

Goldhirsch:1984:PRO

[464] David Goldhirsch and Laura Yedwab. Processing read-only queries over views with generalization. In Dayal et al. [1891], pages 344–348. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/GoldhirschY84.
                                        html.

Kim:1984:TME

[465] Won Kim, Raymond A. Lorie, Dan McNabb, and Wil Plouffe. A transaction mechanism for engineering design databases. In Dayal et al. [1891], pages 355–362. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KimLMP84.html.

Rosenthal:1984:EKB

[466] Arnon Rosenthal, Sandra Heiler, and Frank Manola. An example of knowledge-based query processing in a CAD/CAM DBMS. In Dayal et al. [1891], pages 363–370. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/RosenthalHM84.
                                        html.

Adiba:1984:IPC

[467] Michel E. Adiba and Gia Toan Nguyen. Information processing for CAD/VLSI on a generalized data management system. In Dayal et al. [1891], pages 371–374. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AdibaN84.html.

Boksenbaum:1984:CIT

[468] Claude Boksenbaum, Michèle Cart, Jean Ferrié, and Jean-François Pons. Certification by intervals of timestamps in distributed database systems. In Dayal et al. [1891], pages 377–387. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BoksenbaumCFP84.html.

Brzezinski:1984:UIA

[469] Z. Brzezinski, Janusz R. Getta, J. Rybnik, and W. Stepniewski. UNIBASE — an integrated access to databases. In Dayal et al. [1891], pages 388– 396. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BrzezinskiGRS84.html.

Deen:1984:SDC

[470] S. Misbah Deen, R. R. Amin, and M. C. Taylor. A strategy for decomposing complex queries in a heterogeneous DDB. In Dayal et al. [1891], pages 397–400. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/DeenAT84.html.

Lohman:1984:ONQ

[471] Guy M. Lohman, Dean Daniels, Laura M. Haas, Ruth Kistler, and Patricia G. Selinger. Optimization of nested queries in a distributed relational database. In Dayal et al. [1891], pages 403–415. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/LohmanDHKS84.
                                        html.

Yoshikawa:1984:PIQ

[472] Masatoshi Yoshikawa and Yahiko Kambayashi. Processing inequality queries based on generalized semi-joins. In Dayal et al. [1891], pages 416– 428. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        YoshikawaK84.html.

Chen:1984:OSQ

[473] Arbee L. P. Chen and Victor O. K. Li. Optimizing star queries in a distributed database system. In Dayal et al. [1891], pages 429–438. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ChenL84.html.

Childs:1984:ILC

[474] David L. Childs. Inexpensive large capacity storage will revolutionize the design of database management systems. In Dayal et al. [1891], page 441. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Childs84.html.

Borr:1984:RCD

[475] Andrea J. Borr. Robustness to crash in a distributed database: a non shared-memory multi-processor approach. In Dayal et al. [1891], pages 445–453. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Borr84.html.

Weikum:1984:AIT

[476] Gerhard Weikum and Hans-Jörg Schek. Architectural issues of transaction management in multi-layered systems. In Dayal et al. [1891], pages 454– 465. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WeikumS84.html.

Lai:1984:DTM

[477] Ming-Yee Lai and W. Kevin Wilkinson. Distributed transaction management in jasmin. In Dayal et al. [1891], pages 466–470. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/LaiW84.html.

Mullin:1984:UDH

[478] James K. Mullin. Unified dynamic hashing. In Dayal et al. [1891], pages 473–480. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Mullin84.html.

Kjellberg:1984:CH

[479] Peter Kjellberg and Torben U. Zahle. Cascade hashing. In Dayal et al. [1891], pages 481–492. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/KjellbergZ84.
html.

Otoo:1984:MFD

[480] Ekow J. Otoo. A mapping function for the directory of a multidimensional extendible hashing. In Dayal et al. [1891], pages 493–506. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Otoo84.html.

Dadam:1984:ITV

[481] Peter Dadam, Vincent Y. Lum, and H.-D. Werner. Integration of time versions into a relational database system. In Dayal et al. [1891], pages 509–522. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DadamLW84.html.

Ward:1984:NTU

[482] A. G. Ward. Name-tracing using the ICL content addressable filestore. In Dayal et al. [1891], pages 523– 528. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Ward84.html.

Woehl:1984:ACO

[483] K. Woehl. Automatic classification of office documents by coupling relational data bases and PROLOG expert systems. In Dayal et al. [1891], pages 529–532. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

                                        1984. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Woehl84.html.

Benn:1984:RRS

[484] Wolfgang Benn and Bernd Radig. Retrieval of relational structures for image sequence analysis. In Dayal et al. [1891], pages 533–536. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559

1984. URL http://www.vldb.org/
dblp/db/conf/vldb/BennR84.html.

Abiteboul:1985:USI

[485] Serge Abiteboul and Gösta Grahne. Update semantics for incomplete databases. In Pirotte and Vassiliou [1892], pages 1–12. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/AbiteboulG85.
html.

Afsarmanesh:1985:EOO

[486] Hamideh Afsarmanesh, Dennis McLeod, David Knapp, and Alice Parker. An extensible object-oriented approach to databases for VLSI/CAD. In Pirotte and Vassiliou [1892], pages 13–24. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AfsarmaneshMKP85.html.

Bancilhon:1985:MCT

[487] François Bancilhon, Won Kim, and Henry F. Korth. A model of CAD transactions. In Pirotte and Vassiliou [1892], pages 25–33. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/BancilhonKK85.
html.

Barbic:1985:TCO

[488] Federico Barbic and Fausto Rabitti. The type concept in office document retrieval. In Pirotte and Vassiliou [1892], pages 34–48. ISBN 0-934613-17-6. LCCN QA 76.9 D3

I61 1985. URL http://www.vldb.

                                        org/dblp/db/conf/vldb/BarbicR85.
                                        html.

Blaustein:1985:URD

[489] Barbara T. Blaustein and Charles W. Kaufman. Updating replicated data during communications failures. In Pirotte and Vassiliou [1892], pages 49–58. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BlausteinK85.html.

Boral:1985:DMM

[490] Haran Boral and Steve Redfield. Database machine morphology. In Pirotte and Vassiliou [1892], pages 59–71. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BoralR85.html.

Borgida:1985:AED

[491] Alexander Borgida and Keith E. Williamson. Accommodating exceptions in databases, and refining the schema by learning from them. In Pirotte and Vassiliou [1892], pages 72–81. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BorgidaW85.html.

Bouzeghoub:1985:DDT

[492] Mokrane Bouzeghoub, Georges Gardarin, and Elisabeth Métais. Database design tools: An expert system approach. In Pirotte and Vassiliou [1892], pages 82–95. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        BouzeghoubGM85.html.

Broverman:1985:KBA

[493] Carol A. Broverman and W. Bruce Croft. A knowledge-based approach to data management for intelligent user interfaces. In Pirotte and Vassiliou [1892], pages 96–104. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/BrovermanC85.
html.

Buchmann:1985:ADM

[494] Alejandro P. Buchmann and Concepción Pérez de Célis. An architecture and data model for CAD databases. In Pirotte and Vassiliou [1892], pages 105–114. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Buchmann85.html.

Cellary:1985:LPC

[495] Wojciech Cellary and Tadeusz Morzy. Locking with prevention of cyclic and infinite restarting in distributed database systems. In Pirotte and Vassiliou [1892], pages 115–126. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CellaryM85.html.

Chou:1985:EBM

[496] Hong-Tai Chou and David J. De-Witt. An evaluation of buffer management strategies for relational database systems. In Pirotte and Vassiliou [1892], pages 127–141. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/ChouD85.html.

Dadam:1985:SDI

[497] Peter Dadam, Vincent Y. Lum, U. Prädel, and Gunter Schlageter. Selective deferred index maintenance & concurrency control in integrated information systems. In Pirotte and Vassiliou [1892], pages 142–150. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DadamLPS85.html.

DeWitt:1985:MHB

[498] David J. DeWitt and Robert H. Gerber. Multiprocessor hash-based join algorithms. In Pirotte and Vassiliou [1892], pages 151–164. ISBN 0-934613-17-6. LCCN QA 76.9 D3

I61 1985. URL http://www.vldb.
org/dblp/db/conf/vldb/DeWittG85.
html.

Faloutsos:1985:DSF

[499] Christos Faloutsos and Stavros Christodoulakis. Design of a signature file method that accounts for non-uniform occurrence and query frequencies. In Pirotte and Vassiliou [1892], pages 165–170. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/FaloutsosC85.
html.

Fischer:1985:DWS

[500] Patrick C. Fischer and Dirk Van Gucht. Determining when a structure is a nested relation. In Pirotte and Vassiliou [1892], pages 171–180. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FischerG85.html.

Furman:1985:EIR

[501] Marian S. Furman. An efficient implementation of a relational data base. In Pirotte and Vassiliou [1892], pages 181–191. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/Furman85.html.

Gold:1985:PWM

[502] Israel Gold, Oded Shmueli, and Micha Hofri. The private workspace model feasibility and applications to 2PL performance improvements. In Pirotte and Vassiliou [1892], pages 192–208. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/GoldSH85.html.

Heiler:1985:GWV

[503] Sandra Heiler and Arnon Rosenthal. G-WHIZ, a visual interface for the functional model with recursion. In Pirotte and Vassiliou [1892], pages 209–218. ISBN 0-934613-17-6. LCCN QA 76.9

D3 I61 1985. URL http://www.vldb.
org/dblp/db/conf/vldb/HeilerR85.
html.

Ioannidis:1985:TBM

[504] Yannis E. Ioannidis. A time bound on the materialization of some recursively defined views. In Pirotte and Vassiliou [1892], pages 219–226. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ioannidis85.html.

Jarke:1985:DCI

[505] Matthias Jarke, Volker Linnemann, and Joachim W. Schmidt. Data constructors: On the integration of rules and relations. In Pirotte and Vassiliou [1892], pages 227–240. ISBN 0-934613-17-6. LCCN QA 76.9 D3

I61 1985. URL http://www.vldb.
org/dblp/db/conf/vldb/JarkeLS85.
html.

Kiessling:1985:SRE

[506] Werner Kießling. On semantic reefs and efficient processing of correlation queries with aggregates. In Pirotte and Vassiliou [1892], pages 241– 250. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Kiessling85.html.

Klausner:1985:MSL

[507] Aviel Klausner and Nathan Goodman. Multirelations — semantics and languages. In Pirotte and Vassiliou [1892], pages 251–258. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985.

URL http://www.vldb.org/dblp/
db/conf/vldb/KlausnerG85.html.

Larson:1985:CQD

[508] Per-Åke Larson and H. Z. Yang. Computing queries from derived relations. In Pirotte and Vassiliou [1892], pages 259–269. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LarsonY85.html.

Lausen:1985:PAL

[509] Georg Lausen, Eljas Soisalon-Soininen, and Peter Widmayer. Pre-analysis locking: a safe and deadlock free locking policy. In Pirotte and Vassiliou [1892], pages 270–281. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985.

URL http://www.vldb.org/dblp/
db/conf/vldb/LausenSW85.html.

Leung:1985:FPE

[510] Clement H. C. Leung and K. S. Wong. File processing efficiency on the content addressable file store. In Pirotte and Vassiliou [1892], pages 282–291. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/LeungW85.html.

Lu:1985:SER

[511] Hongjun Lu and Michael J. Carey. Some experimental results on distributed join algorithms in a local network. In Pirotte and Vassiliou [1892], pages 292–304. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/LuC85.html.

MacGregor:1985:ASF

[512] Robert M. MacGregor. ARIEL — A semantic front-end to relational DBMSs. In Pirotte and Vassiliou [1892], pages 305–315. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/Gregor85.html.

Medeiros:1985:UIV

[513] Claudia Bauzer Medeiros and Frank Wm. Tompa. Understanding the implications of view update policies. In Pirotte and Vassiliou [1892], pages 316– 323. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MedeirosT85.html.

Mendelzon:1985:FDL

[514] Alberto O. Mendelzon. Functional dependencies in logic programs. In Pirotte and Vassiliou [1892], pages 324– 330. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Mendelzon85.html.

Mond:1985:CCB

[515] Y. Mond and Yoav Raz. Concurrency control in B+-trees databases using preparatory operations. In Pirotte and Vassiliou [1892], pages 331–334. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

                                        1985. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MondR85.html.

Nierstrasz:1985:OOE

[516] Oscar Nierstrasz and Dennis Tsichritzis. An object-oriented environment for OIS applications. In Pirotte and Vassiliou [1892], pages 335–345. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

                                        1985. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/NierstraszT85.
                                        html.

Omiecinski:1985:IFR

[517] Edward Omiecinski. Incremental file reorganization schemes. In Pirotte and Vassiliou [1892], pages 346–357. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

                                        1985. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Omiecinski85.
                                        html.

Ozkarahan:1985:DOP

[518] Esen A. Ozkarahan and Aris M. Ouksel. Dynamic and order preserving data partitioning for database machines. In Pirotte and Vassiliou [1892], pages 358– 368. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        OzkarahanO85.html.

Pu:1985:FIC

[519] Calton Pu. On-the-fly, incremental, consistent reading of entire databases. In Pirotte and Vassiliou [1892], pages 369–375. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Pu85.html.

Ramakrishnan:1985:MDM

[520] Raghu Ramakrishnan and Abraham Silberschatz. The MR diagram — A model for conceptual database design. In Pirotte and Vassiliou [1892], pages 376–393. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RamakrishnanS85.html.

Rowe:1985:SFP

[521] Lawrence A. Rowe. “fill-in-the-form” programming. In Pirotte and Vassiliou [1892], pages 394–404. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/Rowe85.html.

Satoh:1985:LGQ

[522] Kazuhiro Satoh, Masashi Tsuchida, Fumio Nakamura, and Kazuhiko Oomachi. Local and global query optimization mechanisms for relational databases. In Pirotte and Vassiliou [1892], pages 405–417. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985.

URL http://www.vldb.org/dblp/
db/conf/vldb/SatohTNO85.html.

Shasha:1985:NDM

[523] Dennis Shasha. NetBook — a data model to support knowledge exploration. In Pirotte and Vassiliou [1892], pages 418–425. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61

1985. URL http://www.vldb.org/
dblp/db/conf/vldb/Shasha85.html.

Sheard:1985:CCA

[524] Tim Sheard and David W. Stemple. Coping with complexity in automated reasoning about database systems. In Pirotte and Vassiliou [1892], pages 426– 435. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SheardS85.html.

Tucherman:1985:TMD

[525] Luiz Tucherman, Antonio L. Furtado, and Marco A. Casanova. A tool for modular database design. In Pirotte and Vassiliou [1892], pages 436– 447. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TuchermanFC85.html.

Wong:1985:BTF

[526] Harry K. T. Wong, Hsiu-Fen Liu, Frank Olken, Doron Rotem, and Linda Wong. Bit transposed files. In Pirotte and Vassiliou [1892], pages 448– 457. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WongLORW85.html.

Zaniolo:1985:RDR

[527] Carlo Zaniolo. The representation and deductive retrieval of complex objects. In Pirotte and Vassiliou [1892], pages 458–469. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Zaniolo85.html.

Zlatuska:1985:HDM

[528] Jirí Zlatuska. Hit data model data bases from the functional point of view. In Pirotte and Vassiliou [1892], pages 470–477. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Zlatuska85.html.

Qian:1986:KBI

[529] Xiaolei Qian and Gio Wiederhold. Knowledge-based integrity constraint validation. In Kambayashi et al. [1893], pages 3–12. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/QianW86.html.

Bry:1986:CCD

[530] François Bry and Rainer Manthey. Checking consistency of database constraints: a logical basis. In Kambayashi et al. [1893], pages 13–20. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/BryM86.html.

Bottcher:1986:APM

[531] Stefan Böttcher, Matthias Jarke, and Joachim W. Schmidt. Adaptive predicate managers in database systems. In Kambayashi et al. [1893], pages 21–29. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BottcherJS86.html.

Tsur:1986:LLB

[532] Shalom Tsur and Carlo Zaniolo. LDL: a logic-based data language. In Kambayashi et al. [1893], pages 33–41. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/TsurZ86.html.

Kellogg:1986:ORD

[533] Charles Kellogg, Anthony B. O’Hare, and Larry Travis. Optimizing the rule-data interface in a KMS. In Kambayashi et al. [1893], pages 42– 51. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KelloggOT86.html.

Morita:1986:RUO

[534] Yukihiro Morita, Haruo Yokota, Kenji Nishida, and Hidenori Itoh. Retrievalby-unification operation on a relational knowledge base. In Kambayashi et al. [1893], pages 52–59. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MoritaYNI86.html.

Adiba:1986:HMM

[535] Michel E. Adiba and N. Bui Quang. Historical multi-media databases. In Kambayashi et al. [1893], pages 63– 70. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AdibaQ86.html.

Sarin:1986:UHI

[536] Sunil K. Sarin, Charles W. Kaufman, and Janet E. Somers. Using history information to process delayed database updates. In Kambayashi et al. [1893], pages 71–78. ISBN 0-934613-18-4. LCCN QA 76.9 D3

I61 1986. URL http://www.vldb.
org/dblp/db/conf/vldb/SarinKS86.
html.

Shoshani:1986:TDM

[537] Arie Shoshani and Kyoji Kawagoe. Temporal data management. In Kambayashi et al. [1893], pages 79–88. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShoshaniK86.html.

Carey:1986:OFM

[538] Michael J. Carey, David J. DeWitt, Joel E. Richardson, and Eugene J. Shekita. Object and file management in the EXODUS extensible database system. In Kambayashi et al. [1893], pages 91–100. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CareyDRS86.html.

Valduriez:1986:ITC

[539] Patrick Valduriez, Setrag Khoshafian, and George P. Copeland. Implementation techniques of complex objects. In Kambayashi et al. [1893], pages 101–110. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ValduriezKC86.html.

Pavlovic-Lazetic:1986:MTD

[540] Gordana Pavlovic-Lazetic and Eugene Wong. Managing text as data. In Kambayashi et al. [1893], pages 111– 116. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Pavlovic-LazeticW86.html.

VanderZanden:1986:EBA

[541] Brad T. Vander Zanden, Howard M. Taylor, and Dina Bitton. Estimating block accesses when attributes are correlated. In Kambayashi et al. [1893], pages 119–127. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

URL http://www.vldb.org/dblp/
db/conf/vldb/ZandenTB86.html.

Krishnamurthy:1986:ONQ

[542] Ravi Krishnamurthy, Haran Boral, and Carlo Zaniolo. Optimization of nonrecursive queries. In Kambayashi et al. [1893], pages 128– 137. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KrishnamurthyBZ86.html.

Freytag:1986:TAQ

[543] Johann Christoph Freytag and Nathan Goodman. Translating aggregate queries into iterative programs. In Kambayashi et al. [1893], pages 138– 146. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Freytag86.html.

Mackert:1986:ROV

[544] Lothar F. Mackert and Guy M. Lohman. R* optimizer validation and performance evaluation for distributed queries. In Kambayashi et al. [1893], pages 149–159. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

URL http://www.vldb.org/dblp/
db/conf/vldb/MackertL86.html.

Olken:1986:SRS

[545] Frank Olken and Doron Rotem. Simple random sampling from relational databases. In Kambayashi et al. [1893], pages 160–169. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/OlkenR86.html.

Motro:1986:CIA

[546] Amihai Motro. Completeness information and its application to query processing. In Kambayashi et al. [1893], pages 170–178. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/Motro86.html.

Fang:1986:ICA

[547] M. T. Fang, R. C. T. Lee, and C. C. Chang. The idea of de-clustering and its applications. In Kambayashi et al. [1893], pages 181–188. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/FangLC86.html.

Thom:1986:SAD

[548] James A. Thom, Kotagiri Ramamohanarao, and Lee Naish. A superjoin algorithm for deductive databases. In Kambayashi et al. [1893], pages 189– 196. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ThomRN86.html.

Menon:1986:SSA

[549] Jai Menon. A study of sort algorithms for multiprocessor database machines. In Kambayashi et al. [1893], pages 197–206. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/Menon86.html.

Fushimi:1986:OSS

[550] Shinya Fushimi, Masaru Kitsuregawa, and Hidehiko Tanaka. An overview of the system software of A parallel relational database machine GRACE. In Kambayashi et al. [1893], pages 209– 219. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FushimiKT86.html.

Cheiney:1986:RBU

[551] Jean-Pierre Cheiney, Pascal Faudemay, Rodolphe Michel, and Jean-Marc Thévenin. A reliable backend using multiattribute clustering and select-join operator. In Kambayashi et al. [1893], pages 220–227. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/CheineyFMT86.
html.

DeWitt:1986:GHP

[552] David J. DeWitt, Robert H. Gerber, Goetz Graefe, Michael L. Heytens, Krishna B. Kumar, and M. Muralikrishna. GAMMA — A high performance dataflow database machine. In Kambayashi et al. [1893], pages 228– 237. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DeWittGGHKM86.html.

Hsu:1986:COE

[553] Meichun Hsu and Wei-Pang Yang. Concurrent operations in extendible hashing. In Kambayashi et al. [1893], pages 241–247. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/HsuY86.html.

Chen:1986:MDD

[554] Qiming Chen. The management of dynamically distributed database windows (extended abstract). In Kambayashi et al. [1893], page 248. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/Chen86.html.

Yu:1986:ABR

[555] Philip S. Yu, Douglas W. Cornell, Daniel M. Dias, and Balakrishna R. Iyer. On affinity based routing in multisystem data sharing. In Kambayashi et al. [1893], pages 249–256. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/YuCDI86.html.

Junet:1986:EEE

[556] Marc Junet, Gilles Falquet, and Michel Léonard. ECRINS/86: An extended entity-relationship data base management system and its semantic query language. In Kambayashi et al. [1893], pages 259–266. ISBN 0-934613-18-4. LCCN QA 76.9 D3

I61 1986. URL http://www.vldb.
org/dblp/db/conf/vldb/JunetFL86.
html.

Udagawa:1986:ERD

[557] Yosihisa Udagawa and Tetsuo Mizoguchi. An extended relational database system and it’s application to management of logic diagrama. In Kambayashi et al. [1893], pages 267– 277. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
UdagawaM86.html.

Pistor:1986:DGN

[558] Peter Pistor and F. Andersen. Designing A generalized NF2 model with an SQL-type language interface. In Kambayashi et al. [1893], pages 278–285.

ISBN 0-934613-18-4. LCCN QA 76.9

                                        D3 I61 1986. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/PistorA86.
                                        html.

Hagmann:1986:ODB

[559] Robert B. Hagmann. An observation on database buffering performance metrics. In Kambayashi et al. [1893], pages 289–293. ISBN 0-934613-18-4. LCCN QA 76.9 D3

                                        I61 1986. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Hagmann86.
                                        html.

Lehman:1986:SIS

[560] Tobin J. Lehman and Michael J. Carey. A study of index structures for main memory database management systems. In Kambayashi et al. [1893], pages 294–303. ISBN 0-934613-18-4. LCCN QA 76.9 D3

                                        I61 1986. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/LehmanC86.
                                        html.

Wong:1986:TAV

[561] Harry K. T. Wong and J. Z. Li. Transposition algorithms on very large compressed databases. In Kambayashi et al. [1893], pages 304–311. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

                                        1986. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/WongL86.html.

Parker:1986:KBD

[562] D. Stott Parker. Knowledge-bases and database engineering. In Kambayashi et al. [1893], page 315. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

                                        1986. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Parker86.html.

Klahold:1986:GMV

[563] Peter Klahold, Gunter Schlageter, and Wolfgang Wilkes. A general model for version management in databases. In Kambayashi et al. [1893], pages 319– 327. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KlaholdSW86.html.

Manola:1986:TGS

[564] Frank Manola and Jack A. Orenstein. Toward a general spatial data model for an object-oriented DBMS. In Kambayashi et al. [1893], pages 328–335. ISBN 0-934613-18-4. LCCN QA 76.9

D3 I61 1986. URL http://www.vldb.
org/dblp/db/conf/vldb/ManolaO86.
html.

Chou:1986:UFV

[565] Hong-Tai Chou and Won Kim. A unifying framework for version control in a CAD environment. In Kambayashi et al. [1893], pages 336–344. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/ChouK86.html.

Spooner:1986:FSC

[566] David L. Spooner, Arthur M. Keller, Gio Wiederhold, John Salasin, and Deborah Heystek. Framework for the security component of an Ada DBMS. In Kambayashi et al. [1893], pages 347–354. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SpoonerKWSH86.html.

Roussopoulos:1986:PDA

[567] Nick Roussopoulos and Hyunchul Kang. Preliminary design of ADMS±: a workstation-mainframe integrated architecture for database management systems. In Kambayashi et al. [1893], pages 355–364. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http://

www.vldb.org/dblp/db/conf/vldb/
RoussopoulosK86.html.

Lowden:1986:RSP

[568] Barry G. T. Lowden and A. N. De Roeck. The REMIT system for paraphrasing relational query expressions into natural language. In Kambayashi et al. [1893], pages 365–371. ISBN 0-934613-18-4. LCCN QA 76.9 D3

I61 1986. URL http://www.vldb.
org/dblp/db/conf/vldb/LowdenR86.
html.

Yao:1986:RDM

[569] Qingda Yao and Rong Yi. Relational data model with cause-effect association (abstract). In Kambayashi et al. [1893], page 375. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/YaoY86.html.

Yokomori:1986:AQP

[570] Takashi Yokomori. On analogical query processing in logic database. In Kambayashi et al. [1893], pages 376– 383. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Yokomori86.html.

Chakravarthy:1986:MQP

[571] Upen S. Chakravarthy and Jack Minker. Multiple query processing in deductive databases using query graphs. In Kambayashi et al. [1893], pages 384–391. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http://

www.vldb.org/dblp/db/conf/vldb/
ChakravarthyM86.html.

Ceri:1986:TOL

[572] Stefano Ceri, Georg Gottlob, and Luigi Lavazza. Translation and optimization of logic queries: The algebraic approach. In Kambayashi et al. [1893], pages 395–402. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/CeriGL86.html.

Ioannidis:1986:CTC

[573] Yannis E. Ioannidis. On the computation of the transitive closure of relational operators. In Kambayashi et al. [1893], pages 403–411. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

URL http://www.vldb.org/dblp/
db/conf/vldb/Ioannidis86.html.

Raschid:1986:PPS

[574] Louiqa Raschid and Stanley Y. W. Su. A parallel processing strategy for evaluating recursive queries. In Kambayashi et al. [1893], pages 412– 419. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RaschidS86.html.

Abiteboul:1986:TDS

[575] Serge Abiteboul, Michel Scholl, Georges Gardarin, and Eric Simon. Towards DBMSs for supporting new applications. In Kambayashi et al. [1893], pages 423– 435. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AbiteboulSGS86.html.

Lowenthal:1986:DRM

[576] Eugene I. Lowenthal. Database research at MCC (abstract). In Kambayashi et al. [1893], page 436. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Lowenthal86.html.

Itoh:1986:RDK

[577] Hidenori Itoh. Research and development on knowledge bases systems at ICOT. In Kambayashi et al. [1893], pages 437–445. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

                                        1986. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Itoh86.html.

Arisawa:1986:PEI

[578] Hiroshi Arisawa and Takao Miura. On the properties of extended inclusion dependencies. In Kambayashi et al. [1893], pages 449–456. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/ArisawaM86.html.

Blakeley:1986:UDR

[579] José A. Blakeley, Neil Coburn, and Per-Åke Larson. Updating derived relations: Detecting irrelevant and autonomously computable updates. In Kambayashi et al. [1893], pages 457– 466. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BlakeleyCL86.html.

Keller:1986:CVU

[580] Arthur M. Keller. Choosing a view update translator by dialog at view definition time. In Kambayashi et al. [1893], pages 467–474. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/Keller86.html.

Deen:1986:AVY

[581] S. Misbah Deen. Anyone for a VLDB in the year 2000? In Kambayashi et al. [1893], page 477. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61

1986. URL http://www.vldb.org/
dblp/db/conf/vldb/Deen86.html.

Kurokawa:1986:CCS

[582] Tsuneo Kurokawa. Computer and communication system development among financial institutions in japan. In Kambayashi et al. [1893], pages 481– 495. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Kurokawa86.html.

Mori:1986:VLD

[583] Michinao Mori, Kenji Suzuki, Hideo Abe, and Kenji Itoh. A very large database system to serve national welfare. In Kambayashi et al. [1893], pages 496–501. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MoriSAI86.html.

Seki:1986:NSR

[584] Eishiro Seki. New seat reservation system for Japanese National Railways — distributed processing network and high efficiency databases. In Kambayashi et al. [1893], pages 502– 510. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Seki86.html.

Walter:1987:DSK

[585] Ingrid M. Walter, Peter C. Lockemann, and Hans-Hellmut Nagel. Database support for knowledge-based image evaluation. In Stocker et al. [1894], pages 3–11. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/WalterLN87.html.

Bein:1987:MAD

[586] Jonathan Bein and Roger King. MOBY: An architecture for distributed expert database systems. In Stocker et al. [1894], pages 13–20. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BeinK87.html.

Gardarin:1987:MFT

[587] Georges Gardarin. Magic functions: a technique to optimize extended datalog recursive programs. In Stocker et al. [1894], pages 21–30. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Gardarin87.html.

Ceri:1987:OSA

[588] Stefano Ceri and Letizia Tanca. Optimization of systems of algebraic equations for evaluating datalog queries. In Stocker et al. [1894], pages 31–41. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/CeriT87.html.

Nejdl:1987:RSA

[589] Wolfgang Nejdl. Recursive strategies for answering recursive queries — the RQA/FQI strategy. In Stocker et al. [1894], pages 43–50. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/Nejdl87.html.

Siebes:1987:UDA

[590] Arno Siebes and Martin L. Kersten. Using design axioms and topology to model database semantics. In Stocker et al. [1894], pages 51–59. ISBN 0-934613-46-X. LCCN QA 76.9 D3

I61 1987. URL http://www.vldb.
org/dblp/db/conf/vldb/SiebesK87.
html.

Kowalski:1987:ICD

[591] Robert A. Kowalski, Fariba Sadri, and Paul Soper. Integrity checking in deductive databases. In Stocker et al. [1894], pages 61–69. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/KowalskiSS87.
html.

Cavallo:1987:TPD

[592] Roger Cavallo and Michael Pittarelli. The theory of probabilistic databases. In Stocker et al. [1894], pages 71– 81. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CavalloP87.html.

Rowe:1987:PDM

[593] Lawrence A. Rowe and Michael Stonebraker. The POSTGRES data model. In Stocker et al. [1894], pages 83– 96. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RoweS87.html.

Bancilhon:1987:FPS

[594] François Bancilhon, Ted Briggs, Setrag Khoshafian, and Patrick Valduriez.

FAD, a powerful and simple database language. In Stocker et al. [1894], pages 97–105. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BancilhonBKV87.html.

Sernadas:1987:OOS

[595] Amílcar Sernadas, Cristina Sernadas, and Hans-Dieter Ehrich. Objectoriented specification of databases: An algebraic approach. In Stocker et al. [1894], pages 107–116. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/SernadasSE87.
html.

Cooper:1987:CDS

[596] Richard Cooper, Malcolm P. Atkinson, Alan Dearle, and D. Abderrahmane. Constructing database systems in a persistent environment. In Stocker et al. [1894], pages 117– 125. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CooperADA87.html.

Markowitz:1987:IRR

[597] Victor M. Markowitz and Johann A. Makowsky. Incremental reorganization of relational databases. In Stocker et al. [1894], pages 127–135. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/MarkowitzM87.
html.

Scholl:1987:SFR

[598] Marc H. Scholl, H.-Bernhard Paul, and Hans-Jörg Schek. Supporting flat relations by a nested relational kernel.

In Stocker et al. [1894], pages 137– 146. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SchollPS87.html.

Lenzerini:1987:SDC

[599] Maurizio Lenzerini and Paolo Nobili. On the satisfiability of dependency constraints in entity-relationship schemata. In Stocker et al. [1894], pages 147–154. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

URL http://www.vldb.org/dblp/
db/conf/vldb/LenzeriniN87.html.

Mannila:1987:DI

[600] Heikki Mannila and Kari-Jouko Räihä. Dependency inference. In Stocker et al. [1894], pages 155–158. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

URL http://www.vldb.org/dblp/
db/conf/vldb/MannilaR87.html.

Atzeni:1987:IDS

[601] Paolo Atzeni and Edward P. F. Chan. Independent database schemes under functional and inclusion dependencies. In Stocker et al. [1894], pages 159–166. ISBN 0-934613-46-X. LCCN QA 76.9

D3 I61 1987. URL http://www.vldb.
org/dblp/db/conf/vldb/AtzeniC87.
html.

Abbod:1987:PTR

[602] T. Abbod, K. Brown, and H. Noble. Providing time-related constraints for conventional database systems. In Stocker et al. [1894], pages 167–175. ISBN 0-934613-46-X. LCCN QA 76.9

D3 I61 1987. URL http://www.vldb.
org/dblp/db/conf/vldb/AbbodBN87.
html.

Martin:1987:DTS

[603] N. G. Martin, Shamkant B. Navathe, and Rafi Ahmed. Dealing with temporal schema anomalies in history databases. In Stocker et al. [1894], pages 177–184. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/MartinNA87.html.

Lingat:1987:BMD

[604] Jean-Yves Lingat, P. Nobecourt, and Colette Rolland. Behaviour management in database applications. In Stocker et al. [1894], pages 185– 196. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LingatNR87.html.

Dayal:1987:NTU

[605] Umeshwar Dayal. Of nests and trees: a unified approach to processing queries that contain nested subqueries, aggregates, and quantifiers. In Stocker et al. [1894], pages 197–208. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Dayal87.html.

Christensen:1987:CSC

[606] A. Christensen and Torben U. Zahle. A comparison of self-contained and embedded database languages. In Stocker et al. [1894], pages 209– 216. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChristensenZ87.html.

Lacroix:1987:PPM

[607] M. Lacroix and Pierre Lavency. Preferences; putting more knowledge into queries. In Stocker et al. [1894], pages 217–225. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LacroixL87.html.

Howells:1987:SSM

[608] D. I. Howells, N. J. Fiddian, and W. A. Gray. A source-to-source metatranslation system for relational query languages. In Stocker et al. [1894], pages 227–234. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

URL http://www.vldb.org/dblp/
db/conf/vldb/HowellsFG87.html.

vonBultzingsloewen:1987:TOS

[609] Günter von Bültzingsloewen. Translating and optimizing SQL queries having aggregates. In Stocker et al. [1894], pages 235–243. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

URL http://www.vldb.org/dblp/
db/conf/vldb/Bultzingsloewen87.
html.

Yang:1987:QTP

[610] H. Z. Yang and Per-Åke Larson. Query transformation for PSJ-queries. In Stocker et al. [1894], pages 245– 254. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
YangL87.html.

Agrawal:1987:DAC

[611] Rakesh Agrawal and H. V. Jagadish. Direct algorithms for computing the transitive closure of database relations. In Stocker et al. [1894], pages 255– 266. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AgrawalJ87.html.

Lu:1987:NSC

[612] Hongjun Lu. New strategies for computing the transitive closure of a database relation. In Stocker et al. [1894], pages 267–274. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Lu87.html.

Whang:1987:EDN

[613] Kyu-Young Whang and Shamkant B. Navathe. An extended disjunctive normal form approach for optimizing recursive logic queries in loosely coupled environments. In Stocker et al. [1894], pages 275–287. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/WhangN87.html.

Stonebraker:1987:DPS

[614] Michael Stonebraker. The design of the POSTGRES storage system. In Stocker et al. [1894], pages 289–300. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Stonebraker87.
                                        html.

Sacco:1987:IAF

[615] Giovanni Maria Sacco. Index access with a finite buffer. In Stocker et al. [1894], pages 301–309. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Sacco87.html.

Li:1987:NCM

[616] Jianzhong Li, Doron Rotem, and Harry K. T. Wong. A new compression method with fast searching on large databases. In Stocker et al.

[1894], pages 311–318. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/LiRW87.html.

Woelk:1987:MIM

[617] Darrell Woelk and Won Kim. Multimedia information management in an object-oriented database system. In Stocker et al. [1894], pages 319– 329. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WoelkK87.html.

King:1987:FUA

[618] Roger King and Michael Novak. Freeform: a user-adaptable form management system. In Stocker et al. [1894], pages 331–338. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/KingN87.html.

Gonnet:1987:MYG

[619] Gaston H. Gonnet and Frank Wm. Tompa. Mind your grammar: a new approach to modelling text. In Stocker et al. [1894], pages 339–346. ISBN 0-934613-46-X. LCCN QA 76.9 D3

I61 1987. URL http://www.vldb.
org/dblp/db/conf/vldb/GonnetT87.
html.

Hwang:1987:AME

[620] Hai-Yann Hwang and Yao-Tin Yu. An analytical method for estimating and interpreting query time. In Stocker et al. [1894], pages 347–358. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/HwangY87.html.

Noe:1987:MPT

[621] Jerre D. Noe and David B. Wagner. Measured performance of time interval concurrency control techniques. In Stocker et al. [1894], pages 359– 367. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NoeW87.html.

Subieta:1987:QOS

[622] Kazimierz Subieta and Wiktor Rzeczkowski. Query optimization by stored queries. In Stocker et al. [1894], pages 369– 380. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SubietaR87.html.

Blakey:1987:BPI

[623] Mark Blakey. Basis of a partially informed distributed database. In Stocker et al. [1894], pages 381–388. ISBN 0-934613-46-X. LCCN QA 76.9

                                        D3 I61 1987. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Blakey87.
                                        html.

Kahler:1987:ELD

[624] Bo Kähler and Oddvar Risnes. Extending logging for database snapshot refresh. In Stocker et al. [1894], pages 389–398. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KahlerR87.html.

Jajodia:1987:EVA

[625] Sushil Jajodia and David Mutchler. Enhancements to the voting algorithm. In Stocker et al. [1894], pages 399– 406. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JajodiaM87.html.

Freytag:1987:MSC

[626] Johann Christoph Freytag, Flaviu Cristian, and Bo Kähler. Masking system crashes in database application programs. In Stocker et al. [1894], pages 407–416. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FreytagCK87.html.

Qian:1987:ICR

[627] Xiaolei Qian and Douglas R. Smith. Integrity constraint reformulation for efficient validation. In Stocker et al. [1894], pages 417–425. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/QianS87.html.

Moss:1987:LBR

[628] J. Eliot B. Moss. Log-based recovery for nested transactions. In Stocker et al. [1894], pages 427–432. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

1987. URL http://www.vldb.org/
dblp/db/conf/vldb/Moss87.html.

Harder:1987:PDP

[629] Theo Härder, Klaus Meyer-Wegener, Bernhard Mitschang, and Andrea Sikeler. PRIMA — a DBMS prototype supporting engineering applications. In Stocker et al. [1894], pages 433–442. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HarderMMS87.html.

Ecklund:1987:DDV

[630] Denise J. Ecklund, Earl F. Ecklund, Robert O. Eifrig, and Fred M. Tonge. DVSS: a distributed version storage server for CAD applications. In Stocker et al. [1894], pages 443–454. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/EcklundEET87.
                                        html.

Katz:1987:MCC

[631] Randy H. Katz and Ellis E. Chang. Managing change in a computer-aided design database. In Stocker et al. [1894], pages 455–462. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KatzC87.html.

Dias:1987:DAI

[632] Daniel M. Dias, Balakrishna R. Iyer, John T. Robinson, and Philip S. Yu. Design and analysis of integrated concurrency-coherence controls. In Stocker et al. [1894], pages 463–471. ISBN 0-934613-46-X. LCCN QA 76.9

                                        D3 I61 1987. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/DiasIRY87.
                                        html.

Kumar:1987:PEO

[633] Akhil Kumar and Michael Stonebraker. Performance evaluation of an operating system transaction manager. In Stocker et al. [1894], pages 473–481. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61

                                        1987. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KumarS87.html.

Kuspert:1987:COB

[634] Klaus Küspert, Peter Dadam, and Jürgen Günauer. Cooperative object buffer management in the advanced information management prototype. In Stocker et al. [1894], pages 483– 492. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KuspertDG87.html.

Nguyen:1987:EDS

[635] Gia Toan Nguyen and Dominique Rieu. Expert database support for consistent dynamic objects. In Stocker et al. [1894], pages 493–500. ISBN 0-934613-46-X. LCCN QA 76.9 D3

I61 1987. URL http://www.vldb.
org/dblp/db/conf/vldb/NguyenR87.
html.

Gunther:1987:DSR

[636] Oliver Günther and Eugene Wong. A dual space representation for geometric data. In Stocker et al. [1894], pages 501–506. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GuntherW87.html.

Sellis:1987:RTD

[637] Timos K. Sellis, Nick Roussopoulos, and Christos Faloutsos. The R+-tree: a dynamic index for multi-dimensional objects. In Stocker et al. [1894], pages 507–518. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SellisRF87.html.

Abbott:1988:SRT

[638] Robert K. Abbott and Hector Garcia-Molina. Scheduling real-time transactions: a performance evaluation. In Bancilhon and DeWitt [1895], pages 1–12. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3

I61 1988. URL http://www.vldb.
org/dblp/db/conf/vldb/AbbottG88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Carey:1988:DCC

[639] Michael J. Carey and Miron Livny. Distributed concurrency control performance: a study of algorithms, distribution, and replication. In Bancilhon and DeWitt [1895], pages 13–25. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

CareyL88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Pu:1988:STO

[640] Calton Pu, Gail E. Kaiser, and Norman C. Hutchinson. Split-transactions for open-ended activities. In Bancilhon and DeWitt [1895], pages 26–37. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

PuKH88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Casanova:1988:EID

[641] Marco A. Casanova, Luiz Tucherman, and Antonio L. Furtado. Enforcing inclusion dependencies and referencial integrity. In Bancilhon and DeWitt [1895], pages 38–49. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61

1988. URL http://www.vldb.org/

dblp/db/conf/vldb/CasanovaTF88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Gottlob:1988:CWD

[642] Georg Gottlob and Roberto Zicari. Closed world databases opened through null values. In Bancilhon and DeWitt [1895], pages 50–61. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GottlobZ88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lecluse:1988:IQU

[643] Christophe Lécluse and Nicolas Spyratos. Implementing queries and updates on universal scheme interfaces. In Bancilhon and DeWitt [1895], pages 62–75. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988.

URL http://www.vldb.org/dblp/
db/conf/vldb/LecluseS88.html. Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Deshpande:1988:INR

[644] Anand Deshpande and Dirk Van Gucht. An implementation for nested relational databases. In Bancilhon and DeWitt [1895], pages 76–87. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DeshpandeG88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Jhingran:1988:PSQ

[645] Anant Jhingran. A performance study of query optimization algorithms on a database system supporting procedures. In Bancilhon and DeWitt [1895], pages 88–99. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Jhingran88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Hafez:1988:PNS

[646] Aladdin Hafez and Gultekin Özsoyoglu. The partial normalized storage model of nested relations. In Bancilhon and DeWitt [1895], pages 100–111. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

HafezO88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Kent:1988:BEM

[647] Jack Kent, Douglas B. Terry, and Willie-Sue Orr. Browsing electronic mail: Experiences interfacing a mail system to a DBMS. In Bancilhon and DeWitt [1895], pages 112–123. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

KentTO88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Egenhofer:1988:TSQ

[648] Max J. Egenhofer and Andrew U. Frank. Towards a spatial query language: User interface considerations.

In Bancilhon and DeWitt [1895], pages 124–133. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61

1988. URL http://www.vldb.org/
dblp/db/conf/vldb/EgenhoferF88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Adiba:1988:MCO

[649] Michel E. Adiba and Christine Collet. Management of complex objects as dynamic forms. In Bancilhon and DeWitt [1895], pages 134–147. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

AdibaC88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Kim:1988:VSO

[650] Won Kim and Hong-Tai Chou. Versions of schema for object-oriented databases. In Bancilhon and De-Witt [1895], pages 148–159. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

KimC88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Chaudhuri:1988:TRD

[651] Surajit Chaudhuri. Temporal relationships in databases. In Bancilhon and DeWitt [1895], pages 160–170. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Chaudhuri88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Sripada:1988:LFT

[652] Suryanarayana M. Sripada. A logical framework for temporal deductive databases. In Bancilhon and DeWitt [1895], pages 171–182. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3

                                        I61 1988. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Sripada88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Neuhold:1988:DDP

[653] Erich J. Neuhold and Michael Schrefl. Dynamic derivation of personalized views. In Bancilhon and DeWitt [1895], pages 183–194. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NeuholdS88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Abbott:1988:AAR

[654] Kenneth R. Abbott and Dennis R. Mc-Carthy. Administration and autonomy in a replication-transparent distributed DBMS. In Bancilhon and DeWitt [1895], pages 195–205. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3

                                        I61 1988. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/AbbottM88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lingat:1988:RAP

[655] Jean-Yves Lingat, Pierre Colignon, and Colette Rolland. Rapid application prototyping the PROQUEL language. In Bancilhon and De-Witt [1895], pages 206–217. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LingatCR88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lee:1988:IIF

[656] Mavis K. Lee, Johann Christoph Freytag, and Guy M. Lohman. Implementing an interpreter for functional rules in a query optimizer. In Bancilhon and DeWitt [1895], pages 218–229. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LeeFL88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Rosenthal:1988:AMM

[657] Arnon Rosenthal and Upen S. Chakravarthy. Anatomy of a mudular multiple query optimizer. In Bancilhon and DeWitt [1895], pages 230–239. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61

1988. URL http://www.vldb.org/
dblp/db/conf/vldb/RosenthalC88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lynch:1988:SEQ

[658] Clifford A. Lynch. Selectivity estimation and query optimization in large databases with highly skewed distribution of column values. In Bancilhon and DeWitt [1895], pages 240–251. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Lynch88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Krishnamurthy:1988:TRH

[659] Ravi Krishnamurthy and Shamim A. Naqvi. Towards a real horn clause language. In Bancilhon and De-Witt [1895], pages 252–263. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KrishnamurthyN88.html.  Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

deRougemont:1988:FPS

[660] Michel de Rougemont. Fixed-point semantics and the representation of algorithms on large data. In Bancilhon and DeWitt [1895], pages 264–272. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Rougemont88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lozinskii:1988:CFN

[661] Eliezer L. Lozinskii. Computing facts in non-Horn deductive systems. In Bancilhon and DeWitt [1895], pages 273–279. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Lozinskii88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Faloutsos:1988:FTA

[662] Christos Faloutsos and Raphael Chan. Fast text access methods for optical and large magnetic disks: Designs and performance comparison. In Bancilhon and DeWitt [1895], pages 280–293. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61

1988. URL http://www.vldb.org/
dblp/db/conf/vldb/FaloutsosC88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Linnemann:1988:DIE

[663] Volker Linnemann, Klaus Küspert, Peter Dadam, Peter Pistor, R. Erbe, Alfons Kemper, Norbert Südkamp, Georg Walch, and Mechtild Wallrath. Design and implementation of an extensible database management system supporting user defined data types and functions. In Bancilhon and De-Witt [1895], pages 294–305. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LinnemannKDPEKSWW88.html. Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Lynch:1988:EUD

[664] Clifford A. Lynch and Michael Stonebraker. Extended user-defined indexing with application to textual databases. In Bancilhon and De-Witt [1895], pages 306–317. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LynchS88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Stonebraker:1988:DX

[665] Michael Stonebraker, Randy H. Katz, David A. Patterson, and John K. Ousterhout. The design of XPRS. In Bancilhon and DeWitt [1895], pages 318–330. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        StonebrakerKPO88.html.  Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Bitton:1988:DS

[666] Dina Bitton and Jim Gray. Disk shadowing. In Bancilhon and DeWitt [1895], pages 331–338. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3

                                        I61 1988. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/BittonG88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Bhide:1988:ATT

[667] Anupam Bhide. An analysis of three transaction processing architectures. In Bancilhon and DeWitt [1895], pages 339–350. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Bhide88.html. Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Kent:1988:SCS

[668] Alan J. Kent, Ron Sacks-Davis, and Kotagiri Ramamohanarao. A superimposed coding scheme based on multiple block descriptor files for indexing very large data bases. In Bancilhon and DeWitt [1895], pages 351–359. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

KentSR88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Seeger:1988:TDI

[669] Bernhard Seeger and Hans-Peter Kriegel. Techniques for design and implementation of efficient spatial access methods. In Bancilhon and DeWitt [1895], pages 360–371. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3

I61 1988. URL http://www.vldb.
org/dblp/db/conf/vldb/SeegerK88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Orlandic:1988:CCT

[670] Ratko Orlandic and John L. Pfaltz. Compact 0-complete trees. In Bancilhon and DeWitt [1895], pages 372– 381. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988.

URL http://www.vldb.org/dblp/
db/conf/vldb/OrlandicP88.html. Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Ioannidis:1988:ETC

[671] Yannis E. Ioannidis and Raghu Ramakrishnan. Efficient transitive closure algorithms. In Bancilhon and DeWitt [1895], pages 382–394. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
IoannidisR88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

deMaindreville:1988:MND

[672] Christophe de Maindreville and Eric Simon. Modelling non deterministic queries and updates in deductive databases. In Bancilhon and De-Witt [1895], pages 395–406. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        MaindrevilleS88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Agrawal:1988:ESV

[673] Rakesh Agrawal and H. V. Jagadish. Efficient search in very large databases. In Bancilhon and De-Witt [1895], pages 407–418. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AgrawalJ88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Agrawal:1988:RSQ

[674] Divyakant Agrawal and Amr El Abbadi. Reducing storage for quorum consensus algorithms. In Bancilhon and DeWitt [1895], pages 419–430. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AgrawalA88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Aspnes:1988:TTB

[675] James Aspnes, Alan Fekete, Nancy A. Lynch, Michael Merritt, and William E. Weihl. A theory of timestamp-based concurrency control for nested transactions. In Bancilhon and DeWitt [1895], pages 431–444. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61

1988. URL http://www.vldb.org/
dblp/db/conf/vldb/AspnesFLMW88.

html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Pons:1988:MCC

[676] Jean-François Pons and J. F. Vilarem. Mixed concurrency control: Dealing with heterogeneity in distributed database systems. In Bancilhon and DeWitt [1895], pages 445–456. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

PonsV88.html. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Shekhar:1988:FMT

[677] Shashi Shekhar, Jaideep Srivastava, and Soumitra Dutta. A formal model of trade-off between optimization and execution costs in semantic query optimization. In Bancilhon and De-Witt [1895], pages 457–467. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShekarSD88.html.  Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Nakayama:1988:HPJ

[678] Masaya Nakayama, Masaru Kitsuregawa, and Mikio Takagi. Hashpartitioned join method using dynamic destaging strategy. In Bancilhon and DeWitt [1895], pages 468–478. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        NakayamaK88.html. Sponsored by

VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Shum:1988:ITS

[679] Chung-Dak Shum and Richard R. Muntz. An information-theoretic study on aggregate responses. In Bancilhon and DeWitt [1895], pages 479– 490. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/ShumM88.html. Spon-

sored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Gallaire:1989:LOW

[680] Hervé Gallaire. From a laguna to open waters: Another view on the next generations of databases. In Apers and Wiederhold [1896], page 1. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Gallaire89.html.

Wiederhold:1989:KMU

[681] Gio Wiederhold. Knowledge to mediate from user’s workstations to databases. In Apers and Wiederhold [1896], page 2. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Wiederhold89.html.

Olive:1989:DII

[682] Antoni Olivé. On the design and implementation of information systems from deductive conceptual models. In Apers and Wiederhold [1896], pages 3–11. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Olive89.html.

DiBattista:1989:DME

[683] Giuseppe Di Battista and Maurizio Lenzerini. A deductive method for entity-relationship modeling. In Apers and Wiederhold [1896], pages 13– 21. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BattistaL89.html.

Ola:1989:FIR

[684] Adegbemiga Ola and Gultekin Özsoyoglu. A family of incomplete relational database models. In Apers and Wiederhold [1896], pages 23–31. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/OlaO89.html.

Guting:1989:GER

[685] Ralf Hartmut Güting. Gral: An extensible relational database system for geometric applications. In Apers and Wiederhold [1896], pages 33–44. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/Guting89.html.

Henrich:1989:LTS

[686] Andreas Henrich, Hans-Werner Six, and Peter Widmayer. The LSD tree: Spatial access to multidimensional point and nonpoint objects. In Apers and Wiederhold [1896], pages 45–53. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HenrichSW89.html.

Gardarin:1989:MCO

[687] Georges Gardarin, Jean-Pierre Cheiney, Gerald Kiernan, Dominique Pastre, and Hervé Stora. Managing complex objects in an extensible relational DBMS. In Apers and Wiederhold [1896], pages 55– 65. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GardarinCKPS89.html.

Murphy:1989:ERU

[688] Marguerite C. Murphy and Doron Rotem. Effective resource utilization for multiprocessor join execution. In Apers and Wiederhold [1896], pages 67–75. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        MurphyR89.html.

Muralikrishna:1989:ODA

[689] M. Muralikrishna. Optimization and dataflow algorithms for nested tree queries. In Apers and Wiederhold [1896], pages 77–85. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        Muralikrishna89.html.

Hulin:1989:PPR

[690] Guy Hulin. Parallel processing of recursive queries in distributed architectures. In Apers and Wiederhold [1896], pages 87–96. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

                                        1989. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Hulin89.html.

Kuntz:1989:PGQ

[691] Michel Kuntz and Rainer Melchert. Pasta-3’s graphical query language: Direct manipulation, cooperative queries, full expressive power. In Apers and Wiederhold [1896], pages 97–105. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/KuntzM89.html.

Creasy:1989:EMC

[692] Peter N. Creasy. ENIAM: a more complete conceptual schema language. In Apers and Wiederhold [1896], pages 107–114. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Creasy89.html.

King:1989:FDI

[693] Roger King and Michael Novak. FaceKit: a database interface design toolkit. In Apers and Wiederhold [1896], pages 115–123. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/KingN89.html.

Lorie:1989:LCS

[694] Raymond A. Lorie and Honesty C. Young. A low communication sort algorithm for a parallel database machine. In Apers and Wiederhold [1896], pages 125–134. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/LorieY89.html.

Iyer:1989:PFA

[695] Balakrishna R. Iyer, Gray R. Ricard, and Peter J. Varman. Percentile finding algorithm for multiple sorted runs. In Apers and Wiederhold [1896], pages 135–144. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/IyerRV89.html.

Chang:1989:SAM

[696] Walter W. Chang and Hans-Jörg Schek. A signature access method for the Starburst database system. In Apers and Wiederhold [1896], pages 145–153. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChangS89.html.

Ioannidis:1989:CRP

[697] Yannis E. Ioannidis. Commutativity and its role in the processing of linear recursion. In Apers and Wiederhold [1896], pages 155– 163. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ioannidis89.html.

Lipton:1989:ESG

[698] Richard J. Lipton and Jeffrey F. Naughton. Estimating the size of generalized transitive closures. In Apers and Wiederhold [1896], pages 165–171. ISBN 1-55860-101-5. LCCN QA 76.9

D3 I61 1989. URL http://www.vldb.
org/dblp/db/conf/vldb/LiptonN89.
html.

Naughton:1989:ARF

[699] Jeffrey F. Naughton, Raghu Ramakrishnan, Yehoshua Sagiv, and Jeffrey D. Ullman. Argument reduction by factoring. In Apers and Wiederhold [1896], pages 173–182. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/NaughtonRSU89.
html.

Rossiter:1989:DSH

[700] Nick Rossiter. Database support for hypertext. In Apers and Wiederhold [1896], page 183. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

URL http://www.vldb.org/dblp/
db/conf/vldb/Rossiter89.html.

Mendelzon:1989:FRS

[701] Alberto O. Mendelzon and Peter T. Wood. Finding regular simple paths in graph databases. In Apers and Wiederhold [1896], pages 185–193. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/MendelzonW89.
html.

Chimenti:1989:TOA

[702] Danette Chimenti, Ruben Gamboa, and Ravi Krishnamurthy. Towards on open architecture for LDL. In Apers and Wiederhold [1896], pages 195–203. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChimentiGK89.html.

Segev:1989:EJO

[703] Arie Segev and Himawan Gunadhi. Event-join optimization in temporal relational databases. In Apers and Wiederhold [1896], pages 205–215. ISBN 1-55860-101-5. LCCN QA 76.9

D3 I61 1989. URL http://www.vldb.
org/dblp/db/conf/vldb/SegevG89.
html.

Bhargava:1989:AZI

[704] Gautam Bhargava and Shashi K. Gadia. Achieving zero information-loss in a classical database environment. In Apers and Wiederhold [1896], pages 217–224. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BhargavaG89.html.

Chen:1989:DDU

[705] I-Min A. Chen and Dennis McLeod. Derived data update in semantic databases. In Apers and Wiederhold [1896], pages 225–235. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

                                        1989. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ChenL89.html.

Motro:1989:UIC

[706] Amihai Motro. Using integrity constraints to provide intensional answers to relational queries. In Apers and Wiederhold [1896], pages 237– 246. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Motro89.html.

Cornell:1989:IBM

[707] Douglas W. Cornell and Philip S. Yu. Integration of buffer management and query optimization in relational database environment. In Apers and Wiederhold [1896], pages 247–255. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        CornellY89.html.

Kitsuregawa:1989:EBS

[708] Masaru Kitsuregawa, Masaya Nakayama, and Mikio Takagi. The effect of bucket size tuning in the dynamic hybrid GRACE hash join method. In Apers and Wiederhold [1896], pages 257–266. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KitsuregawaNT89.html.

Agrawal:1989:BKB

[709] Rakesh Agrawal. Building knowledgebased applications with cooperating databases. In Apers and Wiederhold [1896], page 267. ISBN 1-55860-101-5. LCCN QA 76.9 D3

I61 1989. URL http://www.vldb.
org/dblp/db/conf/vldb/Agrawal89.
html.

Olken:1989:RSB

[710] Frank Olken and Doron Rotem. Random sampling from B+ trees. In Apers and Wiederhold [1896], pages 269–277. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
OlkenR89.html.

Malvestuto:1989:AES

[711] Francesco M. Malvestuto and Marina Moscarini. Aggregate evaluability in statistical databases. In Apers and Wiederhold [1896], pages 279– 286. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MalvestutoM89.html.

Rundensteiner:1989:APD

[712] Elke A. Rundensteiner and Lubomir Bic. Aggregates in possibilistic databases. In Apers and Wiederhold [1896], pages 287–295. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        RundensteinerB89.html.

Mitschang:1989:ERA

[713] Bernhard Mitschang. Extending the relational algebra to capture complex objects. In Apers and Wiederhold [1896], pages 297–305. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Mitschang89.html.

Saake:1989:SGD

[714] Gunter Saake, Volker Linnemann, Peter Pistor, and Lutz Michael Wegner. Sorting, grouping and duplicate elimination in the advanced information management prototype. In Apers and Wiederhold [1896], pages 307–316. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SaakeLPW89.html.

Casanova:1989:ORS

[715] Marco A. Casanova, Luiz Tucherman, Antonio L. Furtado, and Anelise P. Braga. Optimization of relational schemas containing inclusion dependencies. In Apers and Wiederhold [1896], pages 317–325. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

                                        1989. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/CasanovaTFB89.
                                        html.

Copeland:1989:CSR

[716] George P. Copeland, Tom Keller, Ravi Krishnamurthy, and Marc Smith. The case for safe RAM. In Apers and Wiederhold [1896], pages 327– 335. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CopelandKKS89.html.

Rothermel:1989:ANR

[717] Kurt Rothermel and C. Mohan. ARIES/NT: a recovery method based on write-ahead logging for nested transactions. In Apers and Wiederhold [1896], pages 337–346. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/RothermelM89.
html.

Du:1989:QSC

[718] Weimin Du and Ahmed K. Elmagarmid. Quasi serializability: a correctness criterion for global concurrency control in InterBase. In Apers and Wiederhold [1896], pages 347– 355. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DuE89.html.

Velez:1989:OOM

[719] Fernando Vélez, Guy Bernard, and Vineeta Darnis. The O2 object manager: an overview. In Apers and Wiederhold [1896], pages 357–366. ISBN 1-55860-101-5. LCCN QA 76.9 D3

I61 1989. URL http://www.vldb.
org/dblp/db/conf/vldb/VelezBD89.
html.

Agrawal:1989:CCV

[720] Rakesh Agrawal and H. V. Jagadish. On correctly configuring versioned objects. In Apers and Wiederhold [1896], pages 367–374. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

URL http://www.vldb.org/dblp/
db/conf/vldb/AgrawalJ89.html.

Lehman:1989:SLF

[721] Tobin J. Lehman and Bruce G. Lindsay. The Starburst long field manager. In Apers and Wiederhold [1896], pages 375–383. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LehmanL89.html.

Abbott:1989:SRT

[722] Robert K. Abbott and Hector Garcia-Molina. Scheduling real-time transactions with disk resident data. In Apers and Wiederhold [1896], pages 385–396. ISBN 1-55860-101-5. LCCN QA 76.9

                                        D3 I61 1989. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/AbbottG89.
                                        html.

Carey:1989:PDR

[723] Michael J. Carey, Rajiv Jauhari, and Miron Livny. Priority in DBMS resource scheduling. In Apers and Wiederhold [1896], pages 397–410. ISBN 1-55860-101-5. LCCN QA 76.9

                                        D3 I61 1989. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/CareyJL89.
                                        html.

Lecluse:1989:ODP

[724] Christophe Lécluse and Philippe Richard. The O2 database programming language. In Apers and Wiederhold [1896], pages 411–422. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LecluseR89.html.

Kim:1989:MQO

[725] Won Kim. A model of queries for object-oriented databases. In Apers and Wiederhold [1896], pages 423– 432. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Kim89.html.

Alashqur:1989:OQL

[726] A. M. Alashqur, Stanley Y. W. Su, and Herman Lam. OQL: a query language for manipulating object-oriented databases. In Apers and Wiederhold [1896], pages 433–442. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/AlashqurSL89.
html.

Bitton:1989:DTI

[727] Dina Bitton. Database tools and interfaces. In Apers and Wiederhold [1896], page 443. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/Bitton89.html.

Risch:1989:MDO

[728] Tore Risch. Monitoring database objects. In Apers and Wiederhold [1896], pages 445–453. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

1989. URL http://www.vldb.org/
dblp/db/conf/vldb/Risch89.html.

Rosenthal:1989:SMA

[729] Arnon Rosenthal, Sharma Chakravarthy, Barbara T. Blaustein, and José A. Blakeley. Situation monitoring for active databases. In Apers and Wiederhold [1896], pages 455–464. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RosenthalCBB89.html.

Apers:1989:FRD

[730] Peter M. G. Apers. Future research directions: Evidence from this conference. In Apers and Wiederhold [1896], page 465. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61

                                        1989. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Apers89.html.

Elmasri:1990:TIA

[731] Ramez Elmasri, Gene T. J. Wuu, and Yeong-Joon Kim. The time index: An access structure for temporal data. In McLeod et al. [1897], pages 1–12. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ElmasriWK90.html.

Tuzhilin:1990:TRA

[732] Alexander Tuzhilin and James Clifford. A temporal relational algebra as basis for temporal relational completeness. In McLeod et al. [1897], pages 13–23. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        TuzhilinC90.html.

Kafer:1990:STD

[733] Wolfgang Käfer, Norbert Ritter, and Harald Schöning. Support for temporal data by complex objects. In McLeod et al. [1897], pages 24– 35. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KaferRS90.html.

Clifton:1990:IHD

[734] Chris Clifton and Hector Garcia-Molina. Indexing in a hypertext database. In McLeod et al. [1897], pages 36–49. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/CliftonG90.html.

Hardjono:1990:SKS

[735] Thomas Hardjono and Jennifer Seberry. Search key substitution in the encipherment of B-trees. In McLeod et al. [1897], pages 50– 58. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HardjonoS90.html.

Harada:1990:QPM

[736] Lilian Harada, Miyuki Nakano, Masaru Kitsuregawa, and Mikio Takagi. Query processing for multi-attribute clustered records. In McLeod et al. [1897], pages 59–70. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HaradaNKT90.html.

Korth:1990:TRT

[737] Henry F. Korth, Nandit Soparkar, and Abraham Silberschatz. Triggered realtime databases with consistency constraints. In McLeod et al. [1897], pages 71–82. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KorthSS90.html.

Nodine:1990:CTH

[738] Marian H. Nodine and Stanley B. Zdonik. Cooperative transaction hierarchies: a transaction model to support design applications. In McLeod et al. [1897], pages 83–94. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/NodineZ90.html.

Korth:1990:FAR

[739] Henry F. Korth, Eliezer Levy, and Abraham Silberschatz. A formal approach to recovery by compensating transactions. In McLeod et al. [1897], pages 95–106. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/KorthLS90.html.

DeWitt:1990:STA

[740] David J. DeWitt, Philippe Futtersack, David Maier, and Fernando Vélez. A study of three alternative workstation-server architectures for object oriented database systems. In McLeod et al. [1897], pages 107– 121. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DeWittFMV90.html.

Wilkinson:1990:MCC

[741] W. Kevin Wilkinson and Marie-Anne Neimat. Maintaining consistency of client-cached data. In McLeod et al. [1897], pages 122–133. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/WilkinsonN90.
                                        html.

Drew:1990:PUC

[742] Pamela Drew, Roger King, and Scott E. Hudson. The performance and utility of the cactis implementation algorithms. In McLeod et al. [1897], pages 135–147. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/DrewKH90.html.

Gray:1990:PSD

[743] Jim Gray, Bob Horst, and Mark Walker. Parity striping of disk arrays: Low-cost reliable storage with acceptable throughput. In McLeod et al. [1897], pages 148–161. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/GrayHW90.html.

Muntz:1990:PAD

[744] Richard R. Muntz and John C. S. Lui. Performance analysis of disk arrays under failure. In McLeod et al. [1897], pages 162–173. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/MuntzL90.html.

Seltzer:1990:TSR

[745] Margo I. Seltzer and Michael Stonebraker. Transaction support in read optimized and write optimized file systems. In McLeod et al. [1897], pages 174–185. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SeltzerS90.html.

Zeller:1990:AHJ

[746] Hansjörg Zeller and Jim Gray. An adaptive hash join algorithm for multiuser environments. In McLeod et al. [1897], pages 186–197. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/ZellerG90.html.

Lu:1990:HBJ

[747] Hongjun Lu, Kian-Lee Tan, and Ming-Chien Shan. Hash-based join algorithms for multiprocessor computers.

In McLeod et al. [1897], pages 198– 209. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LuTS90.html.

Kitsuregawa:1990:BSP

[748] Masaru Kitsuregawa and Yasushi Ogawa. Bucket spreading parallel hash: a new, robust, parallel hash join method for data skew in the super database computer (SDC). In McLeod et al. [1897], pages 210– 221. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KitsuregawaO90.html.

Garcia-Molina:1990:TEA

[749] Hector Garcia-Molina, Christos A. Polyzois, and Robert B. Hagmann. Two epoch algorithms for disaster recovery. In McLeod et al. [1897], pages 222–230. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Garcia-MolinaPH90.html.

Bhargava:1990:ARU

[750] Bharat K. Bhargava and Shirley Browne. Adaptable recovery using dynamic quorum assignments. In McLeod et al. [1897], pages 231– 242. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BhargavaB90.html.

Agrawal:1990:TQP

[751] Divyakant Agrawal and Amr El Abbadi. The tree quorum protocol: An efficient approach for managing replicated data. In McLeod et al.

[1897], pages 243–254. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/AgrawalA90.html.

Wood:1990:FAR

[752] Peter T. Wood. Factoring augmented regular chain programs. In McLeod et al. [1897], pages 255–263. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/Wood90.html.

Mumick:1990:MDA

[753] Inderpal Singh Mumick, Hamid Pirahesh, and Raghu Ramakrishnan. The magic of duplicates and aggregates. In McLeod et al. [1897], pages 264– 277. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MumickPR90.html.

Naughton:1990:HFP

[754] Jeffrey F. Naughton and Raghu Ramakrishnan. How to forget the past without repeating it. In McLeod et al. [1897], pages 278–289. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/NaughtonR90.html.

Kemper:1990:AQP

[755] Alfons Kemper and Guido Moerkotte. Advanced query processing in object bases using access support relations. In McLeod et al. [1897], pages 290– 301. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KemperM90.html.

Sudkamp:1990:EVR

[756] Norbert Südkamp and Volker Linnemann. Elimination of view and redundant variables in a SQL-like database language for extended NF2 structures. In McLeod et al. [1897], pages 302– 313. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SudkampL90.html.

Ono:1990:MCJ

[757] Kiyoshi Ono and Guy M. Lohman. Measuring the complexity of join enumeration in query optimization. In McLeod et al. [1897], pages 314– 325. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        OnoL90.html.

Agrawal:1990:HTC

[758] Rakesh Agrawal and H. V. Jagadish. Hybrid transitive closure algorithms. In McLeod et al. [1897], pages 326– 334. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AgrawalJ90.html.

Houtsma:1990:DTC

[759] Maurice A. W. Houtsma, Peter M. G. Apers, and Stefano Ceri. Distributed transitive closure computations: The disconnection set approach. In McLeod et al. [1897], pages 335– 346. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HoutsmaAC90.html.

Cheiney:1990:PST

[760] Jean-Pierre Cheiney and Christophe de Maindreville. A parallel strategy for transitive closure usind double hashbased clustering. In McLeod et al. [1897], pages 347–358. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/CheineyM90.html.

Ramakrishnan:1990:ROB

[761] Raghu Ramakrishnan, Divesh Srivastava, and S. Sudarshan. Rule ordering in bottom-up fixpoint evaluation of logic programs. In McLeod et al. [1897], pages 359– 371. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RamakrishnanSS90.html.

Kuittinen:1990:EIL

[762] Juhani Kuittinen, Otto Nurmi, Seppo Sippu, and Eljas Soisalon-Soininen. Efficient implementation of loops in bottom-up evaluation of logic queries. In McLeod et al. [1897], pages 372– 379. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KuittinenNSS90.html.

Kemp:1990:RLM

[763] David B. Kemp, Kotagiri Ramamohanarao, and Zoltan Somogyi. Right-, left- and multi-linear rule transformations that maintain context information. In McLeod et al. [1897], pages 380–391. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KempRS90.html.

Mohan:1990:AKK

[764] C. Mohan. ARIES/KVL: a key-value locking method for concurrency control of multiaction transactions operating on B-tree indexes. In McLeod et al. [1897], pages 392–405. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Mohan90.html.

Mohan:1990:CNS

[765] C. Mohan. Commit LSN: a novel and simple method for reducing locking and latching in transaction processing systems. In McLeod et al. [1897], pages 406–418. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Mohan90a.html.

Dan:1990:ESD

[766] Asit Dan, Daniel M. Dias, and Philip S. Yu. The effect of skewed data access on buffer hits and data contention an a data sharing environment. In McLeod et al. [1897], pages 419–431. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/DanDY90.html.

Cellary:1990:CVO

[767] Wojciech Cellary and Geneviève Jomier. Consistency of versions in object-oriented databases. In McLeod et al. [1897], pages 432– 441. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        CellaryJ90.html.

Jeusfeld:1990:DAD

[768] Manfred A. Jeusfeld, Michael Mertikas, Ingrid Wetzel, Matthias Jarke, and Joachim W. Schmidt. Database application development as an object modeling activity. In McLeod et al. [1897], pages 442–454. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/JeusfeldMWJS90.html.

Hull:1990:IDC

[769] Richard Hull and Masatoshi Yoshikawa. ILOG: Declarative creation and manipulation of object identifiers. In McLeod et al. [1897], pages 455–468. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/HullY90.html.

Schneider:1990:TPC

[770] Donovan A. Schneider and David J. De-Witt. Tradeoffs in processing complex join queries via hashing in multiprocessor database machines. In McLeod et al. [1897], pages 469–480. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/SchneiderD90.
html.

Ghandeharizadeh:1990:HRP

[771] Shahram Ghandeharizadeh and David J. DeWitt. Hybrid-range partitioning strategy: a new declustering strategy for multiprocessor database machines. In McLeod et al. [1897], pages 481– 492. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GhandeharizadehD90.html.

Hua:1990:ADP

[772] Kien A. Hua and Chiang Lee. An adaptive data placement scheme for parallel database computer systems.

In McLeod et al. [1897], pages 493– 506. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HuaL90.html.

Elmagarmid:1990:MTM

[773] Ahmed K. Elmagarmid, Yungho Leu, Witold Litwin, and Marek Rusinkiewicz. A multidatabase transaction model for InterBase. In McLeod et al. [1897], pages 507– 518. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ElmagarmidLLR90.html.

Wang:1990:PMH

[774] Y. Richard Wang and Stuart E. Madnick. A polygen model for heterogeneous database systems: The source tagging perspective. In McLeod et al. [1897], pages 519–538. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/WangM90.html.

Aoshima:1990:CBD

[775] M. Aoshima, Y. Izumida, Akifumi Makinouchi, F. Suzuki, and Yasuo Yamane. The C-based database programming language Jasmine/C. In McLeod et al. [1897], pages 539–551. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

                                        1990. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/AoshimaIMSY90.
                                        html.

Qian:1990:SDT

[776] Xiaolei Qian. Synthesizing database transactions. In McLeod et al. [1897], pages 552–565. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/Qian90.html.

Ceri:1990:DPR

[777] Stefano Ceri and Jennifer Widom. Deriving production rules for constraint maintenance. In McLeod et al. [1897], pages 566–577. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/CeriW90.html.

Markowitz:1990:RIR

[778] Victor M. Markowitz. Referential integrity revisited: An object-oriented perspective. In McLeod et al. [1897], pages 578–589. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/Markowitz90.html.

Seeger:1990:BTE

[779] Bernhard Seeger and Hans-Peter Kriegel. The buddy-tree: An efficient and robust access method for spatial data base systems. In McLeod et al. [1897], pages 590–601. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

URL http://www.vldb.org/dblp/
db/conf/vldb/SeegerK90.html.

Wang:1990:QPD

[780] Jason Tsong-Li Wang and Dennis Shasha. Query processing for distance metrics. In McLeod et al. [1897], pages 602–613. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/WangS90.html.

Jagadish:1990:ILS

[781] H. V. Jagadish. On indexing line segments. In McLeod et al. [1897], pages 614–625. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Jagadish90.html.

Hulin:1990:RNR

[782] Guy Hulin. On restructuring nested relations in partitioned normal form. In McLeod et al. [1897], pages 626– 637. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hulin90.html.

Esculier:1990:NMK

[783] Christian Esculier. Non-monotonic knowledge evolution in VLKDBs. In McLeod et al. [1897], pages 638– 649. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Esculier90.html.

Kakas:1990:DUT

[784] Antonis C. Kakas and Paolo Mancarella. Database updates through abduction. In McLeod et al. [1897], pages 650–661. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KakasM90.html.

Carey:1990:IJA

[785] Michael J. Carey, Eugene J. Shekita, George Lapis, Bruce G. Lindsay, and John McPherson. An incremental join attachment for Starburst. In McLeod et al. [1897], pages 662– 673. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CareySLLM90.html.

Severance:1990:DLH

[786] Charles Severance, Sakti Pramanik, and P. Wolberg. Distributed linear hashing and parallel projection in main memory databases. In McLeod et al. [1897], pages 674–682. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/SeverancePW90.
html.

Pucheral:1990:EMM

[787] Philippe Pucheral, Jean-Marc Thévenin, and Patrick Valduriez. Efficient main memory data management using the DBGraph storage model. In McLeod et al. [1897], pages 683–695. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/PucheralTV90.
html.

Fuhr:1990:PFV

[788] Norbert Fuhr. A probabilistic framework for vague queries and imprecise information in databases. In McLeod et al. [1897], pages 696–707. ISBN 1-55860-149-X. LCCN QA76.9.D3I559

1990. URL http://www.vldb.org/
dblp/db/conf/vldb/Fuhr90.html.

Jauhari:1990:PHA

[789] Rajiv Jauhari, Michael J. Carey, and Miron Livny. Priority-Hints: An algorithm for priority-based buffer management. In McLeod et al. [1897], pages 708–721. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JauhariCL90.html.

Catarci:1990:CDL

[790] Tiziana Catarci, Giovanna D’Angiolini, and Maurizio Lenzerini. Concept description language for statistical data modeling. In McLeod et al. [1897], pages 722–729. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/CatarciDL90.html.

Bouzeghoub:1991:SMO

[791] Mokrane Bouzeghoub and Elisabeth Métais. Semantic modeling of object oriented databases. In Camps et al. [1898], pages 3–14. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BouzeghoubM91.
                                        html.

Siegelmann:1991:IIA

[792] Hava T. Siegelmann and B. R. Badrinath. Integrating implicit answers with object-oriented queries. In Camps et al. [1898], pages 15–24. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SiegelmannB91.
                                        html.

Eick:1991:MDT

[793] Christoph F. Eick. A methodology for the design and transformation of conceptual schemas. In Camps et al. [1898], pages 25–34. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Eick91.html.

Huang:1991:EER

[794] Jiandong Huang, John A. Stankovic, Krithi Ramamritham, and Donald F. Towsley. Experimental evaluation of real-time optimistic concurrency control schemes. In Camps et al. [1898], pages 35–46. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991.

URL http://www.vldb.org/dblp/
db/conf/vldb/HuangSRT91.html.

Heiss:1991:ALC

[795] Hans-Ulrich Heiss and Roger Wagner. Adaptive load control in transaction processing systems. In Camps et al. [1898], pages 47–54. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/HeissW91.html.

Hasse:1991:PEM

[796] Christof Hasse and Gerhard Weikum. A performance evaluation of multi-level transaction management. In Camps et al. [1898], pages 55–66. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/HasseW91.html.

Elmasri:1991:CMU

[797] Ramez Elmasri. Conceptual modeling using and extended E-R model (abstract). In Camps et al. [1898], page 67. ISBN 1-55860-150-3. LCCN QA 76.9

D3 I61 1991. URL http://www.vldb.
org/dblp/db/conf/vldb/Elmasri91.
html.

Gupta:1991:SQP

[798] Amarnath Gupta, Terry E. Weymouth, and Ramesh Jain. Semantic queries with pictures: The VIMSYS model. In Camps et al. [1898], pages 69–79. ISBN 1-55860-150-3. LCCN QA 76.9

D3 I61 1991. URL http://www.vldb.
org/dblp/db/conf/vldb/GuptaWJ91.
html.

Aref:1991:OSQ

[799] Walid G. Aref and Hanan Samet. Optimization for spatial query processing. In Camps et al. [1898], pages 81–90. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ArefS91.html.

Zobel:1991:ENR

[800] Justin Zobel, James A. Thom, and Ron Sacks-Davis. Efficiency of nested relational document database systems. In Camps et al. [1898], pages 91–102. ISBN 1-55860-150-3. LCCN QA 76.9

                                        D3 I61 1991. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/ZobelTS91.
                                        html.

Chrysanthis:1991:FET

[801] Panos K. Chrysanthis and Krithi Ramamritham. A formalism for extended transaction model. In Camps et al. [1898], pages 103–112. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChrysanthisR91.html.

Dayal:1991:TML

[802] Umeshwar Dayal, Meichun Hsu, and Rivka Ladin. A transactional model for long-running activities. In Camps et al. [1898], pages 113–122. ISBN 1-55860-150-3. LCCN QA 76.9 D3

                                        I61 1991. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/DayalHL91.
                                        html.

Markowitz:1991:SRS

[803] Victor M. Markowitz. Safe referential structures in relational databases.

In Camps et al. [1898], pages 123– 132. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Markowitz91.html.

Siegel:1991:MAR

[804] Michael Siegel and Stuart E. Madnick. A metadata approach to resolving semantic conflicts. In Camps et al. [1898], pages 133–145. ISBN 1-55860-150-3. LCCN QA 76.9 D3

I61 1991. URL http://www.vldb.
org/dblp/db/conf/vldb/SiegelM91.
html.

Kent:1991:SDM

[805] William Kent. Solving domain mismatch and schema mismatch problems with an object-oriented database programming language. In Camps et al. [1898], pages 147–160. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/Kent91.html.

Andany:1991:MSE

[806] José Andany, Michel Léonard, and Carole Palisser. Management of schema evolution in databases. In Camps et al. [1898], pages 161–170. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991.

URL http://www.vldb.org/dblp/
db/conf/vldb/AndanyLP91.html.

Sullivan:1991:UWP

[807] Mark Sullivan and Michael Stonebraker. Using write protected data structures to improve software fault tolerance in highly available database management systems. In Camps et al. [1898], pages 171–180. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/SullivanS91.html.

Joshi:1991:ALS

[808] Ashok M. Joshi. Adaptive locking strategies in a multi-node data sharing environment. In Camps et al. [1898], pages 181–191. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Joshi91.html.

Mohan:1991:RCC

[809] C. Mohan and Inderpal Narang. Recovery and coherency-control protocols for fast intersystem page transfer and finegranularity locking in a shared disks transaction environment. In Camps et al. [1898], pages 193–207. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MohanN91.html.

Dayal:1991:ADS

[810] Umeshwar Dayal and Klaus R. Dittrich. Active database systems (abstract). In Camps et al. [1898], page 209. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DayalD91.html.

Albert:1991:APB

[811] Joseph Albert. Algebraic properties of bag data types. In Camps et al. [1898], pages 211–219. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Albert91.html.

Denninghoff:1991:PMP

[812] Karl Denninghoff and Victor Vianu. The power of methods with parallel semantics. In Camps et al. [1898], pages 221–232. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DenninghoffV91.html.

Su:1991:DCO

[813] Jianwen Su. Dynamic constraints and object migration. In Camps et al. [1898], pages 233–242. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/Su91.html.

Ghandeharizadeh:1991:OPP

[814] Shahram Ghandeharizadeh, Luis Ramos, Zubair Asad, and Waheed Qureshi. Object placement in parallel hypermedia systems. In Camps et al. [1898], pages 243–254. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GhandeharizadehRAQ91.html.

Palmer:1991:FCL

[815] Mark Palmer and Stanley B. Zdonik. Fido: a cache that learns to fetch. In Camps et al. [1898], pages 255–264. ISBN 1-55860-150-3. LCCN QA 76.9

D3 I61 1991. URL http://www.vldb.
org/dblp/db/conf/vldb/PalmerZ91.
html.

Faloutsos:1991:PLC

[816] Christos Faloutsos, Raymond T. Ng, and Timos K. Sellis. Predictive load control for flexible buffer allocation. In Camps et al. [1898], pages 265– 274. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FaloutsosNS91.html.

Widom:1991:ISO

[817] Jennifer Widom, Roberta Cochrane, and Bruce G. Lindsay. Implementing set-oriented production rules as an extension to Starburst. In Camps et al. [1898], pages 275–285. ISBN 1-55860-150-3. LCCN QA 76.9 D3

                                        I61 1991. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/WidomCL91.
                                        html.

Brant:1991:EDS

[818] David A. Brant, Timothy Grose, Bernie Lofaso, and Daniel P. Miranker. Effects of database size on rule system performance: Five case studies. In Camps et al. [1898], pages 287– 296. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BrantGLM91.html.

Segev:1991:DML

[819] Arie Segev and J. Leon Zhao. Data management for large rule systems. In Camps et al. [1898], pages 297– 307. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SegevZ91.html.

Kamel:1991:DKB

[820] Nabil Kamel, M. Delobel, Thomas G. Marr, Robert Robbins, Jean Thierry-Mieg, and Akira Tsugita. Data and knowledge bases for genome mapping: What lies ahead? (panel). In Camps et al. [1898], page 309. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KamelDMRTT91.
                                        html.

Ozsu:1991:DDM

[821] M. Tamer Özsu. Distributed database management: Current state-of-the-art, unsolved problems, new issues (abstract). In Camps et al. [1898], page 311. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ozsu91.html.

Ramamritham:1991:RTD

[822] Krithi Ramamritham, Sang Hyuk Son, Alejandro P. Buchmann, Klaus R. Dittrich, and C. Mohan. Real-time databases (panel). In Camps et al. [1898], page 313. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991.

URL http://www.vldb.org/dblp/
db/conf/vldb/RamamrithamSBDM91.
html.

Spaccapietra:1991:CDD

[823] Stefano Spaccapietra, Shamkant B. Navathe, Erich J. Neuhold, and Amit P. Sheth. Cooperative database design (panel). In Camps et al. [1898], page 315. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SpaccapietraNNS91.html.

Diaz:1991:RMO

[824] Oscar Díaz, Norman W. Paton, and Peter M. D. Gray. Rule management in object oriented databases: a uniform approach. In Camps et al. [1898], pages 317–326. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/DiazPG91.html.

Gehani:1991:OAD

[825] Narain H. Gehani and H. V. Jagadish. Ode as an active database: Constraints and triggers. In Camps et al. [1898], pages 327–336. ISBN 1-55860-150-3. LCCN QA 76.9 D3

                                        I61 1991. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/GehaniJ91.
                                        html.

Beeri:1991:MAO

[826] Catriel Beeri and Tova Milo. A model for active object oriented databases. In Camps et al. [1898], pages 337– 349. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BeeriM91.html.

Cha:1991:KDM

[827] Sang Kyun Cha and Gio Wiederhold. Kaleidoscope data model for an english-like query language. In Camps et al. [1898], pages 351–361. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ChaW91.html.

Lanzelotte:1991:ESS

[828] Rosana S. G. Lanzelotte and Patrick Valduriez. Extending the search strategy in a query optimizer. In Camps et al. [1898], pages 363–373. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/LanzelotteV91.
                                        html.

Omiecinski:1991:PAL

[829] Edward Omiecinski. Performance analysis of a load balancing hash-join algorithm for a shared memory multiprocessor. In Camps et al. [1898], pages 375–385. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Omiecinski91.html.

Demolombe:1991:CAD

[830] Robert Demolombe. Cooperative access to data and knowledge baes (abstract). In Camps et al. [1898], page 387. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Demolombe91.html.

Blankinship:1991:IMD

[831] Rex Blankinship, Alan R. Hevner, and S. Bing Yao. An iterative method for distributed database design. In Camps et al. [1898], pages 389–400. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BlankinshipHY91.html.

Rozen:1991:FAP

[832] Steve Rozen and Dennis Shasha. A framework for automating physical database design. In Camps et al. [1898], pages 401–411. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/RozenS91.html.

Ford:1991:ORR

[833] Daniel Alexander Ford and Stavros Christodoulakis. Optimizing random retrievals from CLV format optical disks. In Camps et al. [1898], pages 413–422. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FordC91.html.

Gabbay:1991:TLH

[834] Dov M. Gabbay and Peter McBrien. Temporal logic & historical databases. In Camps et al. [1898], pages 423–430. ISBN 1-55860-150-3. LCCN QA 76.9

                                        D3 I61 1991. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/GabbayM91.
                                        html.

Su:1991:TKR

[835] Stanley Y. W. Su and Hsin-Hsing M. Chen. A temporal knowledge representation model OSAM*/T and its query language OQL/T. In Camps et al. [1898], pages 431–442. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SuC91.html.

DeWitt:1991:ENE

[836] David J. DeWitt, Jeffrey F. Naughton, and Donovan A. Schneider. An evaluation of non-equijoin algorithms. In Camps et al. [1898], pages 443– 452. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DeWittNS91.html.

Kamel:1991:DTB

[837] Magdi N. Kamel, Umeshwar Dayal, Rakesh Agrawal, Douglas Tolbert, and Gilbert Vidal. Database technologies for the 90’s and beyond (panel). In Camps et al. [1898], page 453. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KamelDATV91.html.

Hull:1991:LCP

[838] Richard Hull and Dean Jacobs. Language constructs for programming active databases. In Camps et al. [1898], pages 455–467. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/HullJ91.html.

Schreier:1991:AAT

[839] Ulf Schreier, Hamid Pirahesh, Rakesh Agrawal, and C. Mohan. Alert: An architecture for transforming a passive DBMS into an active DBMS. In Camps et al. [1898], pages 469–478. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/SchreierPAM91.
html.

Agrawal:1991:MPP

[840] Rakesh Agrawal, Roberta Cochrane, and Bruce G. Lindsay. On maintaining priorities in a production rule system. In Camps et al. [1898], pages 479–487. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AgrawalCL91.html.

Sheth:1991:FDS

[841] Amit P. Sheth. Federated database systems for managing distributed, heterogeneous, and autonomous databases. In Camps et al. [1898], page 489. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/Sheth91.html.

Poulovassilis:1991:FPA

[842] Alexandra Poulovassilis and Carol Small. A functional programming approach to deductive databases. In Camps et al. [1898], pages 491–500. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        PoulovassilisS91.html.

Sudarshan:1991:ARD

[843] S. Sudarshan and Raghu Ramakrishnan. Aggregation and relevance in deductive databases. In Camps et al. [1898], pages 501–511. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SudarshanR91.
                                        html.

Olive:1991:ICC

[844] Antoni Olivé. Integrity constraints checking in deductive databases. In Camps et al. [1898], pages 513–523. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Olive91.html.

Hua:1991:HDS

[845] Kien A. Hua and Chiang Lee. Handling data skew in multiprocessor database computers using partition tuning. In Camps et al. [1898], pages 525–535. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HuaL91.html.

Walton:1991:TPM

[846] Christopher B. Walton, Alfred G. Dale, and Roy M. Jenevein. A taxonomy and performance model of data skew effects in parallel joins. In Camps et al. [1898], pages 537– 548. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        WaltonDJ91.html.

Lu:1991:OMW

[847] Hongjun Lu, Ming-Chien Shan, and Kian-Lee Tan. Optimization of multiway join queries for parallel execution. In Camps et al. [1898], pages 549–560. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LuST91.html.

Breitbart:1991:IMS

[848] Yuri Breitbart, Hector Garcia-Molina, Witold Litwin, Nick Roussopoulos, Hans-Jörg Schek, and Gio Wiederhold. Interoperability in multidatabases: Semantic and system issues (panel). In Camps et al. [1898], page 561. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BreitbartGLRSW91.html.

Bocca:1991:LPE

[849] Jorge B. Bocca. Logic programming environments for large knowledge bases: a practical perspective (abstract). In Camps et al. [1898], page 563. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

1991. URL http://www.vldb.org/
dblp/db/conf/vldb/Bocca91.html.

Albano:1991:RMS

[850] Antonio Albano, Giorgio Ghelli, and Renzo Orsini. A relationship mechanism for a strongly typed objectoriented database programming language. In Camps et al. [1898], pages 565–575. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AlbanoGO91.html.

Ceri:1991:DPR

[851] Stefano Ceri and Jennifer Widom. Deriving production rules for incremental view maintenance. In Camps et al. [1898], pages 577–589. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61

                                        1991. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/CeriW91.html.

Li:1992:CMDa

[852] Jianzhong Li, Jaideep Srivastava, and Doron Rotem. CMD: a multidimensional declustering method for parallel data systems. In Yuan [1899], pages 3–14. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LiSR92.html.

Chen:1992:USR

[853] Ming-Syan Chen, Ming-Ling Lo, Philip S. Yu, and Honesty C. Young. Using segmented right-deep trees for the execution of pipelined hash joins. In Yuan [1899], pages 15–26. ISBN 1-55860-151-1. LCCN QA76.9.D3

                                        I61 1992. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/ChenLYY92.
                                        html.

DeWitt:1992:PSH

[854] David J. DeWitt, Jeffrey F. Naughton, Donovan A. Schneider, and S. Seshadri. Practical skew handling in parallel joins. In Yuan [1899], pages 27–40. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DeWittNSS92.html.

Imielinski:1992:QHM

[855] Tomasz Imielinski and B. R. Badrinath. Querying in highly mobile distributed environments. In Yuan [1899], pages 41–52. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/ImielinskiB92.html.

Chu:1992:TEO

[856] Wesley W. Chu, Ion Tim Ieong, Ricky K. Taira, and Claudine M. Breant. A temporal evolutionary objectoriented data model and its query language for medical image management. In Yuan [1899], pages 53–64. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/ChuITB92.html.

Ansari:1992:UFT

[857] Mansoor Ansari, Linda Ness, Marek Rusinkiewicz, and Amit P. Sheth. Using flexible transactions to support multi-system telecommunication applications. In Yuan [1899], pages 65– 76. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AnsariNRS92.html.

Valduriez:1992:OOD

[858] Patrick Valduriez. Object-oriented database systems. In Yuan [1899], page 77. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Valduriez92.html.

Kemper:1992:OBE

[859] Alfons Kemper, Guido Moerkotte, and Michael Steinbrunn. Optimizing Boolean expressions in objectbases. In Yuan [1899], pages 79– 90. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KemperMS92.html.

Muralikrishna:1992:IUA

[860] M. Muralikrishna. Improved unnesting algorithms for join aggregate SQL queries. In Yuan [1899], pages 91– 102. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Muralikrishna92.html.

Ioannidis:1992:PQO

[861] Yannis E. Ioannidis, Raymond T. Ng, Kyuseok Shim, and Timos K. Sellis. Parametric query optimization. In Yuan [1899], pages 103– 114. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        IoannidisNSS92.html.

Parker:1992:SMC

[862] D. Stott Parker, Eric Simon, and Patrick Valduriez. SVP: a model capturing sets, lists, streams, and parallelism. In Yuan [1899], pages 115– 126. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ParkerSV92.html.

Richardson:1992:SLD

[863] Joel Richardson. Supporting lists in a data model (A timely approach). In Yuan [1899], pages 127–138. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

                                        1992. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Richardson92.
                                        html.

Read:1992:MRR

[864] Robert L. Read, Donald S. Fussell, and Abraham Silberschatz. A multiresolution relational data model. In Yuan [1899], pages 139–150. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/ReadFS92.html.

Agrawal:1992:RLS

[865] Divyakant Agrawal and Amr El Abbadi. Resilient logical structures for efficient management of replicated data. In Yuan [1899], pages 151– 162. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AgrawalA92.html.

Chamberlin:1992:DDD

[866] Donald D. Chamberlin and Frank B. Schmuck. Dynamic data distribution (D 3 ) in a shared-nothing multiprocessor data store. In Yuan [1899], pages 163–174. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChamberlinS92.html.

Krishnakumar:1992:HTE

[867] Narayanan Krishnakumar and Arthur J. Bernstein. High throughput escrow algorithms for replicated databases. In Yuan [1899], pages 175–186. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/KrishnakumarB92.
html.

Rundensteiner:1992:MMS

[868] Elke A. Rundensteiner. Multiview: a methodology for supporting multiple views in object-oriented databases. In Yuan [1899], pages 187– 198. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Rundensteiner92.html.

Smith:1992:EMM

[869] Kenneth Smith and Marianne Winslett. Entity modeling in the MLS relational model. In Yuan [1899], pages 199– 210. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SmithW92.html.

Lee:1992:ERD

[870] Suk Kyoon Lee. An extended relational database model for uncertain and imprecise information. In Yuan [1899], pages 211–220. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

                                        1992. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Lee92.html.

Samet:1992:SD

[871] Hanan Samet. Spatial databases. In Yuan [1899], page 221. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

                                        1992. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Samet92.html.

Sorenson:1992:DMY

[872] Paul G. Sorenson, Felipe Cariño, Jnan R. Dash, and Patricia G. Selinger. Database management in the year 2000: Projections and star gazing (panel). In Yuan [1899], page 223. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SorensonCDS92.html.

Urpi:1992:MCC

[873] Toni Urpí and Antoni Olivé. A method for change computation in deductive databases. In Yuan [1899], pages 225– 237. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
UrpiO92.html.

Ramakrishnan:1992:CCR

[874] Raghu Ramakrishnan, Divesh Srivastava, and S. Sudarshan. CORAL — control, relations and logic. In Yuan [1899], pages 238–250. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/RamakrishnanSS92.html.

Kramer:1992:URB

[875] Michael Kramer, Georg Lausen, and Gunter Saake. Updates in a rulebased language for objects. In Yuan [1899], pages 251–262. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/KramerLS92.html.

Otoo:1992:GIS

[876] Ekow J. Otoo, Ron Lake, Wo-Shun Luk, T. H. Merrett, and Hanan Samet. Georgraphic information systems, A challenge for the 90’s (panel). In Yuan [1899], page 263. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/OtooLLMS92.html.

Jagadish:1992:PBM

[877] H. V. Jagadish and Oded Shmueli. Proclamation-based model for cooperating transactions. In Yuan [1899], pages 265–276. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/JagadishS92.html.

Du:1992:QOH

[878] Weimin Du, Ravi Krishnamurthy, and Ming-Chien Shan. Query optimization in a heterogeneous DBMS. In Yuan [1899], pages 277–291. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/DuKS92.html.

Raz:1992:PCO

[879] Yoav Raz. The principle of commitment ordering, or guaranteeing serializability in a heterogeneous environment of multiple autonomous resource mangers using atomic commitment. In Yuan [1899], pages 292–312. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/Raz92.html.

Shasha:1992:DT

[880] Dennis Shasha and Steve Rozen. Database tuning. In Yuan [1899], page 313. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShashaR92.html.

Simon:1992:IHL

[881] Eric Simon, Jerry Kiernan, and Christophe de Maindreville. Implementing high level active rules on top of a relational DBMS. In Yuan [1899], pages 315–326. ISBN 1-55860-151-1. LCCN QA76.9.D3

I61 1992. URL http://www.vldb.
org/dblp/db/conf/vldb/SimonKM92.
html.

Gehani:1992:CES

[882] Narain H. Gehani, H. V. Jagadish, and Oded Shmueli. Composite event specification in active databases: Model & implementation. In Yuan [1899], pages 327–338. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GehaniJS92.html.

Ceri:1992:PRP

[883] Stefano Ceri and Jennifer Widom. Production rules in parallel and distributed database environments. In Yuan [1899], pages 339–351. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/CeriW92.html.

Zobel:1992:EIT

[884] Justin Zobel, Alistair Moffat, and Ron Sacks-Davis. An efficient indexing technique for full text databases. In Yuan [1899], pages 352–362. ISBN 1-55860-151-1. LCCN QA76.9.D3

I61 1992. URL http://www.vldb.
org/dblp/db/conf/vldb/ZobelMS92.
html.

Faloutsos:1992:BTI

[885] Christos Faloutsos and H. V. Jagadish. On B-tree indices for skewed distributions. In Yuan [1899], pages 363– 374. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FaloutsosJ92.html.

Antoshenkov:1992:RSP

[886] Gennady Antoshenkov. Random sampling from pseudo-ranked B+ trees. In Yuan [1899], pages 375–382. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/Antoshenkov92.
html.

Leung:1992:TQP

[887] T. Y. Cliff Leung and Richard R. Muntz. Temporal query processing and optimization in multiprocessor database machines. In Yuan [1899], pages 383–394. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/LeungM92.html.

Gadia:1992:IIR

[888] Shashi K. Gadia, Sunil S. Nair, and Yiu-Cheong Poon. Incomplete information in relational temporal databases. In Yuan [1899], pages 395–406. ISBN 1-55860-151-1. LCCN QA76.9.D3

I61 1992. URL http://www.vldb.
org/dblp/db/conf/vldb/GadiaNP92.
html.

Dayal:1992:UAP

[889] Umeshwar Dayal and Gene T. J. Wuu. A uniform approach to processing temporal queries. In Yuan [1899], pages 407–418. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/DayalW92.html.

White:1992:PSA

[890] Seth J. White and David J. DeWitt. A performance study of alternative object faulting and pointer swizzling strategies. In Yuan [1899], pages 419– 431. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WhiteD92.html.

Moenkeberg:1992:PEA

[891] Axel Moenkeberg and Gerhard Weikum. Performance evaluation of an adaptive and robust load control method for the avoidance of data-contention thrashing. In Yuan [1899], pages 432– 443. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MoenkebergW92.html.

Chan:1992:EBM

[892] Chee Yong Chan, Beng Chin Ooi, and Hongjun Lu. Extensible buffer management of indexes. In Yuan [1899], pages 444–454. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/ChanOL92.html.

Mylopoulos:1992:SR

[893] John Mylopoulos and Thomas Rose. Software repositories. In Yuan [1899], page 455. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MylopoulosR92.html.

Li:1992:CMDb

[894] Qing Li and John L. Smith. A conceptual model for dynamic clustering in object databases. In Yuan [1899], pages 457–468. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/LiS92.html.

Jagadish:1992:IMO

[895] H. V. Jagadish and Xiaolei Qian. Integrity maintenance in object-oriented databases. In Yuan [1899], pages 469– 480. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JagadishQ92.html.

Liu:1992:AMD

[896] Ling Liu and Robert Meersman. Activity model: a declarative approach for capturing communication behavior in object-oriented databases. In Yuan [1899], pages 481–493. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/LiuM92.html.

Morgenstern:1992:DTR

[897] Matthew Morgenstern. Database technology for reliable systems: Issues, impact, and approaches (panel). In Yuan [1899], page 495. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/Morgenstern92.
html.

Bober:1992:MQL

[898] Paul M. Bober and Michael J. Carey. Multiversion query locking. In Yuan [1899], pages 497–510. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/BoberC92.html.

Salzberg:1992:PTB

[899] Betty Salzberg and Allyn Dimock. Principles of transaction-based on-line reorganization. In Yuan [1899], pages 511–520. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SalzbergD92.html.

Wilschut:1992:PMM

[900] Annita N. Wilschut, Jan Flokstra, and Peter M. G. Apers. Parallelism in a main-memory DBMS: The performance of PRISMA/DB. In Yuan [1899], pages 521–532. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/WilschutFA92.
html.

Gottemukkala:1992:LLM

[901] Vibby Gottemukkala and Tobin J. Lehman. Locking and latching in a memory-resident database system. In Yuan [1899], pages 533– 544. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GottemukkalaL92.html.

Rusinkiewicz:1992:MAS

[902] Marek Rusinkiewicz and Amit P. Sheth. Multidatabase applications: Semantic and system issues. In Yuan [1899], page 545. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/RusinkiewiczS92.html.

Han:1992:KDD

[903] Jiawei Han, Yandong Cai, and Nick Cercone. Knowledge discovery in databases: An attribute-oriented approach. In Yuan [1899], pages 547– 559. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HanCC92.html.

Agrawal:1992:ICD

[904] Rakesh Agrawal, Sakti P. Ghosh, Tomasz Imielinski, Balakrishna R. Iyer, and Arun N. Swami. An interval classifier for database mining applications. In Yuan [1899], pages 560–573. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/AgrawalGIIS92.
html.

Rabitti:1992:IRA

[905] Fausto Rabitti and Pasquale Savino. An information-retrieval approach for image databases. In Yuan [1899], pages 574–584. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RabittiS92.html.

Heimbigner:1992:EOM

[906] Dennis Heimbigner. Experiences with an object manager for a processcentered environment. In Yuan [1899], pages 585–595. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/Heimbigner92.
html.

Franklin:1992:GMM

[907] Michael J. Franklin, Michael J. Carey, and Miron Livny. Global memory management in client-server database architectures. In Yuan [1899], pages 596–609. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FranklinCL92.html.

Delis:1992:PSC

[908] Alex Delis and Nick Roussopoulos. Performance and scalability of clientserver database architectures. In Yuan [1899], pages 610–623. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61

1992. URL http://www.vldb.org/
dblp/db/conf/vldb/DelisR92.html.

Buchmann:1992:DMR

[909] Alejandro P. Buchmann. Data management for real-time systems. In Yuan [1899], page 625. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

URL http://www.vldb.org/dblp/
db/conf/vldb/Buchmann92.html.

Hachem:1993:MDD

[910] Nabil I. Hachem, Ke Qiu, Michael A. Gennert, and Matthew O. Ward. Managing derived data in the Gaea scientific DBMS. In Agrawal et al. [1900], pages 1–12. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/HachemQGW93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Wolniewicz:1993:AOC

[911] Richard H. Wolniewicz and Goetz Graefe. Algebraic optimization of computations over scientific databases. In Agrawal et al. [1900], pages 13– 24. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WolniewiczG93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Stonebraker:1993:TPD

[912] Michael Stonebraker, Jolly Chen, Nobuko Nathan, Caroline Paxson, and Jiang Wu. Tioga: Providing data management support for scientific visualization applications. In Agrawal et al. [1900], pages 25–38. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/StonebrakerCNPW93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Albano:1993:ODM

[913] Antonio Albano, Roberto Bergamini, Giorgio Ghelli, and Renzo Orsini. An object data model with roles. In Agrawal et al. [1900], pages 39– 51. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AlbanoBGO93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Gal-Oz:1993:MMA

[914] Nurith Gal-Oz, Ehud Gudes, and Eduardo B. Fernández. A model of methods access authorization in object- oriented databases. In Agrawal et al. [1900], pages 52–61. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Gal-OzGF93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Talens:1993:VSC

[915] G. Talens, Chabane Oussalah, and M. F. Colinas. Versions of simple and composite objects. In Agrawal et al. [1900], pages 62– 72. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        TalensOC93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Abiteboul:1993:QUF

[916] Serge Abiteboul, Sophie Cluet, and Tova Milo. Querying and updating the file. In Agrawal et al. [1900], pages 73–84. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AbiteboulCM93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Chen:1993:DOV

[917] Ling Tony Chen and Doron Rotem. Declustering objects for visualization. In Agrawal et al. [1900], pages 85– 96. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

ChenR93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Shoens:1993:RSI

[918] Kurt A. Shoens, Allen Luniewski, Peter M. Schwarz, James W. Stamos, and Joachim Thomas II. The Rufus system: Information organization for semi-structured data. In Agrawal et al. [1900], pages 97– 107. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SoensLSST93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Ceri:1993:MSH

[919] Stefano Ceri and Jennifer Widom. Managing semantic heterogeneity with production rules and persistent queues. In Agrawal et al. [1900], pages 108– 119. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

CeriW93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Miller:1993:UIC

[920] Renée J. Miller, Yannis E. Ioannidis, and Raghu Ramakrishnan. The use of information capacity in schema integration and translation. In Agrawal et al. [1900], pages 120– 133. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MillerIR93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Attie:1993:SEI

[921] Paul C. Attie, Munindar P. Singh, Amit P. Sheth, and Marek Rusinkiewicz. Specifying and enforcing intertask dependencies. In Agrawal et al. [1900], pages 134–145. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/AttieSSR93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Plexousakis:1993:ICR

[922] Dimitris Plexousakis. Integrity constraint and rule maintenance in temporal deductive knowledge bases. In Agrawal et al. [1900], pages 146– 157. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Plexousakis93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Srivastava:1993:CAO

[923] Divesh Srivastava, Raghu Ramakrishnan, Praveen Seshadri, and S. Sudarshan. Coral ++: Adding objectorientation to a logic database language. In Agrawal et al. [1900], pages 158–170. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SrivastavaRSS93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Levy:1993:QIU

[924] Alon Y. Levy and Yehoshua Sagiv. Queries independent of updates. In Agrawal et al. [1900], pages 171– 181. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LevyS93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Rahm:1993:ADL

[925] Erhard Rahm and Robert Marek. Analysis of dynamic load balancing strategies for parallel shared nothing database systems. In Agrawal et al. [1900], pages 182–193. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/RahmM93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Hong:1993:PCM

[926] Eui Kyeong Hong. Performance of catalog management schemes for running access modules in a locally distributed database system. In Agrawal et al. [1900], pages 194– 205. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Hong93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Houtsma:1993:IPE

[927] Maurice A. W. Houtsma, Annita N. Wilschut, and Jan Flokstra. Implementation and performance evaluation of a parallel transitive closure algorithm on PRISMA/DB. In Agrawal et al. [1900], pages 206– 217. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HoutsmaWF93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Bernstein:1993:SPL

[928] Philip A. Bernstein, Per O. Gyllstrom, and Tom Wimberg. STDL — A portable language for transaction processing. In Agrawal et al. [1900], pages 218–229. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BernsteinGW93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Witkowski:1993:NNG

[929] Andrew Witkowski, Felipe Cariño, and Pekka Kostamaa. NCR 3700 — the next-generation industrial database computer. In Agrawal et al. [1900], pages 230–243. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/WitkowskiCK93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Beech:1993:COS

[930] David Beech. Collections of objects in SQL3. In Agrawal et al. [1900], pages 244–255. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Beech93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Ioannidis:1993:USH

[931] Yannis E. Ioannidis. Universality of serial histograms. In Agrawal et al. [1900], pages 256– 267. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Ioannidis93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Subieta:1993:VDW

[932] Kazimierz Subieta, Florian Matthes, Joachim W. Schmidt, and Andreas Rudloff. Viewers: a data-world analogue of procedure calls. In Agrawal et al. [1900], pages 268–277. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

                                        1993. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SubietaMSR93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Poncelet:1993:TFA

[933] Pascal Poncelet, Maguelonne Teisseire, Rosine Cicchetti, and Lotfi Lakhal. Towards a formal approach for object database design. In Agrawal et al. [1900], pages 278–289. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

                                        1993. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/PonceletTCL93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Zobel:1993:SLL

[934] Justin Zobel, Alistair Moffat, and Ron Sacks-Davis. Searching large lexicons for partially specified terms using compressed inverted files. In Agrawal et al. [1900], pages 290–301. ISBN 1-55860-152-X. LCCN QA76.9.D3

I61 1993. URL http://www.vldb.
org/dblp/db/conf/vldb/ZobelMS93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Chandra:1993:MTF

[935] Rakesh Chandra and Arie Segev. Managing temporal financial data in an extensible database. In Agrawal et al. [1900], pages 302–313. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/ChandraS93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Zezula:1993:HFD

[936] Pavel Zezula, Paolo Ciaccia, and Paolo Tiberio. Hamming filter: a dynamic signature file organization for parallel stores. In Agrawal et al. [1900], pages 314–327. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ZezulaCT93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Brown:1993:MMM

[937] Kurt P. Brown, Michael J. Carey, and Miron Livny. Managing memory to meet multiclass workload response time goals. In Agrawal et al. [1900], pages 328–341. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

BrownCL93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Chen:1993:ADB

[938] Chung-Min M. Chen and Nick Roussopoulos. Adaptive database buffer allocation using query feedback. In Agrawal et al. [1900], pages 342– 353. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

ChenR93a.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Mehta:1993:DMA

[939] Manish Mehta and David J. De-Witt. Dynamic memory allocation for multiple-query workloads. In Agrawal et al. [1900], pages 354–367. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

                                        1993. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MehtaD93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Mohan:1993:CEM

[940] C. Mohan. A cost-effective method for providing improved data availability during DBMS restart recovery after a failure. In Agrawal et al. [1900], pages 368–379. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Mohan93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Lomet:1993:EHD

[941] David B. Lomet and Betty Salzberg. Exploiting A history database for backup. In Agrawal et al. [1900], pages 380–390. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LometS93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Jagadish:1993:RMM

[942] H. V. Jagadish, Abraham Silberschatz, and S. Sudarshan. Recovering from main-memory lapses. In Agrawal et al. [1900], pages 391–404. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/JagadishSS93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Amato:1993:DSA

[943] Giuseppe Amato, Fosca Giannotti, and Gianni Mainetto. Data sharing analysis for a database programming language via abstract interpretation. In Agrawal et al. [1900], pages 405–415. ISBN 1-55860-152-X. LCCN QA76.9.D3

I61 1993. URL http://www.vldb.
org/dblp/db/conf/vldb/AmatoGM93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Poulovassilis:1993:DTA

[944] Alexandra Poulovassilis and Carol Small. A domain-theoretic approach to integrating functional and logic database languages. In Agrawal et al. [1900], pages 416– 428. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PoulovassilisS93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Hosking:1993:ULP

[945] Antony L. Hosking, Eric W. Brown, and J. Eliot B. Moss. Update logging for persistent programming languages: A comparative performance evaluation. In Agrawal et al. [1900], pages 429– 440. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HoskingBM93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Ghandeharizadeh:1993:ILS

[946] Shahram Ghandeharizadeh, Richard Hull, Dean Jacobs, Jaime Castillo, Martha Escobar-Molano, Shih-Hui Lu, Junhui Luo, Chiu Tsang, and Gang Zhou. On implementing a language for specifying active database execution models. In Agrawal et al. [1900], pages 441–454. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GhandeharizadehHJCELLTZ93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Fabret:1993:AAI

[947] Françoise Fabret, Mireille Régnier, and Eric Simon. An adaptive algorithm for incremental evaluation of production rules in databases. In Agrawal et al. [1900], pages 455– 466. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FabretRS93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Ishikawa:1993:AOO

[948] Hiroshi Ishikawa and Kazumi Kubota. An active object-oriented database: a multi-paradigm approach to constraint management. In Agrawal et al. [1900], pages 467– 478. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
IshikawaK93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Shekita:1993:MJO

[949] Eugene J. Shekita, Honesty C. Young, and Kian-Lee Tan. Multi-join optimization for symmetric multiprocessors. In Agrawal et al. [1900], pages 479–492. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ShekitaYT93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Lanzelotte:1993:EOS

[950] Rosana S. G. Lanzelotte, Patrick Valduriez, and Mohamed Zaït. On the effectiveness of optimization search strategies for parallel execution spaces. In Agrawal et al. [1900], pages 493– 504. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LanzelotteVZ93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Chen:1993:AHF

[951] Ming-Syan Chen, Hui-I Hsiao, and Philip S. Yu. Applying hash filters to improving the execution of bushy trees. In Agrawal et al. [1900], pages 505–516. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

VhanHY93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Mitchell:1993:CEQ

[952] Gail Mitchell, Umeshwar Dayal, and Stanley B. Zdonik. Control of an extensible query optimizer: a planningbased approach. In Agrawal et al. [1900], pages 517–528. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/MitchellDZ93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Chaudhuri:1993:QOP

[953] Surajit Chaudhuri and Kyuseok Shim. Query optimization in the presence of foreign functions. In Agrawal et al. [1900], pages 529–542. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/ChaudhuriS93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Kemper:1993:BAQ

[954] Alfons Kemper, Guido Moerkotte, and Klaus Peithner. A blackboard architecture for query optimization in object bases. In Agrawal et al. [1900], pages 543–554. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KemperMP93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Thomas:1993:POC

[955] Joachim Thomas and Stefan Deßloch. A plan-operator concept for clientbased knowledge processing. In Agrawal et al. [1900], pages 555– 566. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

ThomasD93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Brodsky:1993:TPC

[956] Alexander Brodsky, Joxan Jaffar, and Michael J. Maher. Toward practical constraint databases. In Agrawal et al. [1900], pages 567– 580. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BrodskyJM93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Grefen:1993:CTP

[957] Paul W. P. J. Grefen. Combining theory and practice in integrity control: A declarative approach to the specification of a transaction modification subsystem. In Agrawal et al. [1900], pages 581–591. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Grefen93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Seeger:1993:RSD

[958] Bernhard Seeger, Per-Åke Larson, and Ron McFayden. Reading a set of disk pages. In Agrawal et al. [1900], pages 592–603. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SeegerLM93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Polyzois:1993:DMA

[959] Christos A. Polyzois, Anupam Bhide, and Daniel M. Dias. Disk mirroring with alternating deferred updates… (prize paper). In Agrawal et al. [1900], pages 604–617. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/PolyzoisBD93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Pang:1993:MAE

[960] HweeHwa Pang, Michael J. Carey, and Miron Livny. Memory-adaptive external sorting. In Agrawal et al. [1900], pages 618–629. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/PangCL93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Lampson:1993:NPC

[961] Butler W. Lampson and David B. Lomet. A new presumed commit optimization for two phase commit. In Agrawal et al. [1900], pages 630– 640. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LampsonL93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Franklin:1993:LDC

[962] Michael J. Franklin, Michael J. Carey, and Miron Livny. Local disk caching for client-server database systems. In Agrawal et al. [1900], pages 641– 654. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FranklinCL93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Lomet:1993:KRL

[963] David B. Lomet. Key range locking strategies for improved concurrency. In Agrawal et al. [1900], pages 655– 664. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Lomet93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Selinger:1993:PCD

[964] Patricia G. Selinger. Predictions and challenges for database systems in the year 2000. In Agrawal et al. [1900], pages 667–675. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/Selinger93.html. Co-

sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Kim:1993:OOD

[965] Won Kim. Object-oriented database systems: Promises, reality, and future. In Agrawal et al. [1900], pages 676– 687. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Kim93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Stonebraker:1993:DRC

[966] Michael Stonebraker, Rakesh Agrawal, Umeshwar Dayal, Erich J. Neuhold, and Andreas Reuter. DBMS research at a crossroads: The Vienna update. In Agrawal et al. [1900], pages 688–692. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993.

URL http://www.vldb.org/dblp/
db/conf/vldb/StonebrakerADNR93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Bancilhon:1993:ODM

[967] François Bancilhon. Object database morphology. In Agrawal et al. [1900], page 693. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bancilhon93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Jeffery:1993:DRS

[968] Keith G. Jeffery. Database research strategies of funding agencies (panel). In Agrawal et al. [1900], pages 697– 698. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

Jeffery93.html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Korth:1993:MCF

[969] Henry F. Korth and Tomasz Imielinski. Mobile computing: Fertile research area or black hole? (panel). In Agrawal et al. [1900], pages 699–700. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

                                        1993. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KorthI93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Madnick:1993:VCI

[970] Stuart E. Madnick. The voice of the customer: Innovative and useful research directions (panel). In Agrawal et al. [1900], pages 701–704. ISBN 1-55860-152-X. LCCN QA76.9.D3

                                        I61 1993. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Madnick93.

html. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Cohen:1993:PIC

[971] Gerald Cohen. A practical issue concerning very large data bases: The need for query governors. In Agrawal et al. [1900], pages 705–706. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61

1993. URL http://www.vldb.org/
dblp/db/conf/vldb/Cohen93.html.

Co-sponsored by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.

Fitzgerald:1993:PCF

[972] Kevin Fitzgerald. Problems/challenges facing industry data base users. In Agrawal et al. [1900], pages 707– 708. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Fitzgerald93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Patterson:1993:NDQ

[973] Blake Patterson. The need for data quality. In Agrawal et al. [1900], page 709. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Patterson93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Schreier:1993:DRK

[974] Ulf Schreier. Database requirements of knowledge-based production scheduling and control: a CIM perspective.

In Agrawal et al. [1900], pages 710– 711. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Schreier93.html. Co-sponsored by

VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Li:1994:SIH

[975] Wen-Syan Li and Chris Clifton. Semantic integration in heterogeneous databases using neural networks. In Bocca et al. [1901], pages 1–12. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559

1994. URL http://www.vldb.org/
dblp/db/conf/vldb/vldb94-1.html.

Idris:1994:PDS

[976] N. B. Idris, W. A. Gray, and R. F. Churchhouse. Providing dynamic security control in a federated database. In Bocca et al. [1901], pages 13– 23. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-13.html.

Jonscher:1994:ABS

[977] Dirk Jonscher and Klaus R. Dittrich. An approach for building secure database federations. In Bocca et al. [1901], pages 24–35. ISBN 1-55860-153-8. LCCN QA76.9.D3

I559 1994. URL http://www.vldb.
org/dblp/db/conf/vldb/vldb94-24.
html.

Hasan:1994:OAE

[978] Waqar Hasan and Rajeev Motwani. Optimization algorithms for exploiting the parallelism-communication tradeoff in pipelined parallelism. In Bocca et al. [1901], pages 36–47. ISBN 1-55860-153-8. LCCN QA76.9.D3

I559 1994. URL http://www.vldb.
org/dblp/db/conf/vldb/vldb94-36.
html.

Jagadish:1994:DHP

[979] H. V. Jagadish, Daniel F. Lieuwen, Rajeev Rastogi, Abraham Silberschatz, and S. Sudarshan. Dalí: a high performance main memory storage manager. In Bocca et al. [1901], pages 48–59. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-48.html.

Mohania:1994:SID

[980] Mukesh K. Mohania and Nandlal L. Sarda. Some issues in design of distributed deductive databases. In Bocca et al. [1901], pages 60–71. ISBN 1-55860-153-8. LCCN QA76.9.D3

I559 1994. URL http://www.vldb.
org/dblp/db/conf/vldb/vldb94-60.
html.

Brown:1994:TAP

[981] Kurt P. Brown, Manish Mehta, Michael J. Carey, and Miron Livny. Towards automated performance tuning for complex workloads. In Bocca et al. [1901], pages 72–84. ISBN 1-55860-153-8. LCCN QA76.9.D3

I559 1994. URL http://www.vldb.
org/dblp/db/conf/vldb/vldb94-72.
html.

Galindo-Legaria:1994:FRJ

[982] César A. Galindo-Legaria, Arjan Pellenkoft, and Martin L. Kersten. Fast, randomized join-order selection — why use transformations? In Bocca et al. [1901], pages 85–95. ISBN 1-55860-153-8. LCCN QA76.9.D3

                                        I559 1994. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/vldb94-85.
                                        html.

Levy:1994:QOP

[983] Alon Y. Levy, Inderpal Singh Mumick, and Yehoshua Sagiv. Query optimization by predicate move-around. In Bocca et al. [1901], pages 96– 107. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-96.html.

Amiel:1994:SEB

[984] Eric Amiel, Marie-Jo Bellosta, Eric Dujardin, and Eric Simon. Supporting exceptions to behavioral schema consistency to ease schema evolution in OODBMS. In Bocca et al. [1901], pages 108–119. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-108.html.

Wiener:1994:BLO

[985] Janet L. Wiener and Jeffrey F. Naughton. Bulk loading into an OODB: a performance study. In Bocca et al. [1901], pages 120– 131. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-120.html.

Collet:1994:NEM

[986] Christine Collet, Thierry Coupaye, and T. Svensen. NAOS — efficient and modular reactive capabilities in an object-oriented database system. In Bocca et al. [1901], pages 132– 143. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-132.html.

Ng:1994:EEC

[987] Raymond T. Ng and Jiawei Han. Efficient and effective clustering methods for spatial data mining. In Bocca et al. [1901], pages 144–155. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-144.html.

Hoel:1994:PDP

[988] Erik G. Hoel and Hanan Samet. Performance of data-parallel spatial operations. In Bocca et al. [1901], pages 156–167. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-156.html.

Brinkhoff:1994:IGC

[989] Thomas Brinkhoff and Hans-Peter Kriegel. The impact of global clustering on spatial database systems. In Bocca et al. [1901], pages 168– 179. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-168.html.

Kilger:1994:IMS

[990] Christoph Kilger and Guido Moerkotte. Indexing multiple sets. In Bocca et al. [1901], pages 180– 191. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-180.html.

Brown:1994:FII

[991] Eric W. Brown, James P. Callan, and W. Bruce Croft. Fast incremental indexing for full-text information retrieval. In Bocca et al. [1901], pages 192–202. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-192.html.

Sreenath:1994:HTE

[992] B. Sreenath and S. Seshadri. The hcC-tree: An efficient index structure for object oriented databases. In Bocca et al. [1901], pages 203– 213. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-203.html.

Fu:1994:TRS

[993] Ada Wai-Chee Fu and David Wai-Lok Cheung. A transaction replication scheme for a replicated database with node autonomy. In Bocca et al. [1901], pages 214–225. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-214.html.

Ouzzani:1994:TAT

[994] M. Ouzzani, M. A. Atroun, and N. L. Belkhodja. A top-down approach for two level serializability. In Bocca et al. [1901], pages 226– 237. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-226.html.

Setzer:1994:NCC

[995] V. W. Setzer and Andrea Zisman. New concurrency control algorithms for accessing and compacting B-trees. In Bocca et al. [1901], pages 238– 248. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-238.html.

Gehani:1994:OFS

[996] Narain H. Gehani, H. V. Jagadish, and William D. Roome. OdeFS: a file system interface to an object-oriented database. In Bocca et al. [1901], pages 249–260. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-249.html.

Ferrandina:1994:ILD

[997] Fabrizio Ferrandina, Thorsten Meyer, and Roberto Zicari. Implementing lazy database updates for an object database system. In Bocca et al. [1901], pages 261–272. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-261.html.

Frohn:1994:AOP

[998] Jürgen Frohn, Georg Lausen, and Heinz Uphoff. Access to objects by path expressions and rules. In Bocca et al. [1901], pages 273– 284. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-273.html.

Teisseire:1994:TED

[999] Maguelonne Teisseire, Pascal Poncelet, and Rosine Cicchetti. Towards eventdriven modelling for database design. In Bocca et al. [1901], pages 285– 296. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-285.html.

Guting:1994:GMQ

[1000] Ralf Hartmut Güting. GraphDB: Modeling and querying graphs in databases. In Bocca et al. [1901], pages 297– 308. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-297.html.

Gaasterland:1994:QAR

[1001] Terry Gaasterland and Jorge Lobo. Qualified answers that reflect user needs and preferences. In Bocca et al. [1901], pages 309–320. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/vldb94-309.html.

Mediano:1994:VTS

[1002] Maurìcio R. Mediano, Marco A. Casanova, and Marcelo Dreux. V-trees — A storage method for long vector data. In Bocca et al. [1901], pages 321–330. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-321.html.

Chawathe:1994:ISS

[1003] Sudarshan S. Chawathe, Ming-Syan Chen, and Philip S. Yu. On index selection schemes for nested object hierarchies. In Bocca et al. [1901], pages 331–341. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-331.html.

Litwin:1994:RFO

[1004] Witold Litwin, Marie-Anne Neimat, and Donovan A. Schneider. RP*: a family of order preserving scalable distributed data structures. In Bocca et al.

[1901], pages 342–353. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-342.html.

Chaudhuri:1994:IGQ

[1005] Surajit Chaudhuri and Kyuseok Shim.

Including group-by in query optimization. In Bocca et al. [1901], pages 354–366. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-354.html.

Tsatalos:1994:GVT

[1006] Odysseas G. Tsatalos, Marvin H.

Solomon, and Yannis E. Ioannidis. The GMAP: a versatile tool for physical data independence. In Bocca et al. [1901], pages 367–378. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-367.html.

Davison:1994:MCR

[1007] Diane L. Davison and Goetz Graefe.

Memory-contention responsive hash joins. In Bocca et al. [1901], pages 379–390. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-379.html.

Gutierrez:1994:DGV

[1008] Alejandro Gutiérrez, Philippe Pucheral, Hermann Steffen, and Jean-Marc Thévenin. Database graph views: a practical model to manage persistent graphs. In Bocca et al. [1901], pages 391–402. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-391.html.

Matthes:1994:PT

[1009] Florian Matthes and Joachim W. Schmidt. Persistent threads. In Bocca et al. [1901], pages 403– 414. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-403.html.

Poulovassilis:1994:IAQ

[1010] Alexandra Poulovassilis and Carol Small. Investigation of algebraic query optimisation techniques for database programming languages. In Bocca et al. [1901], pages 415–426. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/vldb94-415.html.

Kemper:1994:DBS

[1011] Alfons Kemper and Donald Kossmann. Dual-buffering strategies in object bases. In Bocca et al. [1901], pages 427–438. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-427.html.

Johnson:1994:LOH

[1012] Theodore Johnson and Dennis Shasha. 2Q: a low overhead high performance buffer management replacement algorithm. In Bocca et al. [1901], pages 439–450. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-439.html.

Ng:1994:MBD

[1013] Raymond T. Ng and Jinhai Yang. Maximizing buffer and disk utilizations for news on- demand. In Bocca et al.

[1901], pages 451–462. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-451.html.

Barja:1994:EDO

[1014] Maria L. Barja, Norman W. Paton, Alvaro A. A. Fernandes, M. Howard Williams, and Andrew Dinn. An effective deductive object-oriented database through language integration. In Bocca et al. [1901], pages 463– 474. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-463.html.

Baralis:1994:AAR

[1015] Elena Baralis and Jennifer Widom.

An algebraic approach to rule analysis in expert database systems. In Bocca et al. [1901], pages 475– 486. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-475.html.

Agrawal:1994:FAM

[1016] Rakesh Agrawal and Ramakrishnan Srikant. Fast algorithms for mining association rules in large databases. In Bocca et al. [1901], pages 487– 499. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-487.html.

Kamel:1994:HRT

[1017] Ibrahim Kamel and Christos Faloutsos. Hilbert R-tree: An improved Rtree using fractals. In Bocca et al. [1901], pages 500–509. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/vldb94-500.html.

Shatdal:1994:CCA

[1018] Ambuj Shatdal, Chander Kant, and Jeffrey F. Naughton. Cache conscious algorithms for relational query processing. In Bocca et al. [1901], pages 510–521. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-510.html.

Xie:1994:JIH

[1019] Zhaohui Xie and Jiawei Han. Join index hierarchies for supporting efficient navigations in object-oriented databases. In Bocca et al. [1901], pages 522–533. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-522.html.

Sripada:1994:CU

[1020] Suryanarayana M. Sripada and Beat Wüthrich. Cumulative updates. In Bocca et al. [1901], pages 534– 545. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-534.html.

Lu:1994:SPT

[1021] Hongjun Lu, Beng Chin Ooi, and Kian-Lee Tan. On spatially partitioned temporal join. In Bocca et al. [1901], pages 546–557. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        vldb94-546.html.

DeWitt:1994:CSP

[1022] David J. DeWitt, Navin Kabra, Jun Luo, Jignesh M. Patel, and Jie-Bing Yu. Client-server paradise. In Bocca et al. [1901], pages 558– 569. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-558.html.

Sistla:1994:RAS

[1023] A. Prasad Sistla, Clement T. Yu, and R. Haddad. Reasoning about spatial relationships in picture retrieval systems. In Bocca et al. [1901], pages 570–581. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-570.html.

Chiueh:1994:CBI

[1024] Tzi cker Chiueh. Content-based image indexing. In Bocca et al. [1901], pages 582–593. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-582.html.

Ozden:1994:LCS

[1025] Banu Özden, Alexandros Biliris, Rajeev Rastogi, and Abraham Silberschatz. A low-cost storage server for movie on demand databases. In Bocca et al. [1901], pages 594– 605. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-594.html.

Chakravarthy:1994:CEA

[1026] Sharma Chakravarthy, V. Krishnaprasad, Eman Anwar, and S.-K. Kim. Composite events for active databases: Semantics, contexts and detection. In Bocca et al. [1901], pages 606–617. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-606.html.

Steenhagen:1994:NLJ

[1027] Hennie J. Steenhagen, Peter M. G. Apers, Henk M. Blanken, and Rolf A. de By. From nested-loop to join queries in OODB. In Bocca et al. [1901], pages 618–629. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-618.html.

Pirotte:1994:MPU

[1028] Alain Pirotte, Esteban Zimányi, David Massart, and Tatiana Yakusheva. Materialization: a powerful and ubiquitous abstraction pattern. In Bocca et al. [1901], pages 630– 641. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-630.html.

Muntz:1994:RBA

[1029] Alice H. Muntz and Christian T. Ramiller. A requirement-based approach to data modeling and reengineering. In Bocca et al. [1901], pages 643–654. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-643.html.

Lesk:1994:EAD

[1030] Michael Lesk. Experiments on access to digital libraries: How can images and text be used together. In Bocca et al. [1901], pages 655– 667. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-655.html.

Rosenthal:1994:DIL

[1031] Arnon Rosenthal and Leonard J. Seligman. Data integration in the large: The challenge of reuse. In Bocca et al. [1901], pages 669–675. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-669.html.

Anand:1994:EPS

[1032] Sarabjot S. Anand, David A. Bell, and John G. Hughes. An empirical performance study of the Ingres search accelerator for a large property management database system. In Bocca et al. [1901], pages 676– 685. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-676.html.

Hjelsvold:1994:MQV

[1033] Rune Hjelsvold and Roger Midtstraum.

Modelling and querying video data. In Bocca et al. [1901], pages 686– 694. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-686.html.

Iyer:1994:DCS

[1034] Balakrishna R. Iyer and David Wilhite. Data compression support in databases. In Bocca et al. [1901], pages 695–704. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-695.html.

Bernstein:1994:ORT

[1035] Philip A. Bernstein and Umeshwar Dayal. An overview of repository technology. In Bocca et al. [1901], pages 705–713. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-705.html.

Cochinwala:1994:MST

[1036] Munir Cochinwala and John Bradley. A multidatabase system for tracking and retrieval of financial data. In Bocca et al. [1901], pages 714– 721. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-714.html.

Goodman:1994:BLI

[1037] Nathan Goodman, Steve Rozen, and Lincoln Stein. Building a laboratory information system around a C++-based object-oriented DBMS. In Bocca et al. [1901], pages 722– 729. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-722.html.

Martin:1994:RDO

[1038] Bruce E. Martin and R. G. G. Cattell. Relating distributed objects. In Bocca et al. [1901], pages 730– 739. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-730.html.

Yan:1994:IST

[1039] Tak W. Yan and Jurgen Annevelink. Integrating a structured-text retrieval system with an object-oriented database system. In Bocca et al. [1901], pages 740–749. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-740.html.

Spaccapietra:1994:UIW

[1040] Stefano Spaccapietra. User interfaces; who cares? In Bocca et al. [1901], page 751. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-751.html.

Zemankova:1994:SDS

[1041] Maria Zemankova and Yannis E. Ioannidis. Scientific databases — state of the art and future directions. In Bocca et al. [1901], pages 752– 753. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-752.html.

Bernstein:1994:PDS

[1042] Philip A. Bernstein. PC database systems — present and future. In Bocca et al. [1901], page 754. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

URL http://www.vldb.org/dblp/
db/conf/vldb/vldb94-754.html.

Levy:1994:CGI

[1043] Alon Y. Levy, Abraham Silberschatz, Divesh Srivastava, and Maria Zemankova. Challenges for global information systems. In Bocca et al. [1901], page 755. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
vldb94-755.html.

Korth:1995:DLT

[1044] Henry F. Korth. The double life of the transaction abstraction: Fundamental principle and evolving system concept. In Dayal et al. [1902], pages 2–6. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/Korth95.html.

Bayer:1995:DMD

[1045] Rudolf Bayer. Document management as a database problem. In Dayal et al. [1902], pages 7–10. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Bayer95.html.

Madnick:1995:VVV

[1046] Stuart E. Madnick. From VLDB to VMLDB (Very MANY Large Data Bases): Dealing with large-scale semantic heterogenity. In Dayal et al. [1902], pages 11–16. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Madnick95.html.

Eickler:1995:PEO

[1047] André Eickler, Carsten Andreas Gerlhof, and Donald Kossmann. A performance evaluation of OID mapping techniques. In Dayal et al. [1902], pages 18–29. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        EicklerGK95.html.

Wiener:1995:OBL

[1048] Janet L. Wiener and Jeffrey F. Naughton. OODB bulk loading revisited: The partitioned-list approach. In Dayal et al. [1902], pages 30–41. ISBN 1-55860-379-4. LCCN QA76.9.D3

                                        I61 1995. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/WienerN95.
                                        html.

Amsaleg:1995:EIG

[1049] Laurent Amsaleg, Michael J. Franklin, and Olivier Gruber. Efficient incremental garbage collection for client-server object database systems. In Dayal et al.

[1902], pages 42–53. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995.

URL http://www.vldb.org/dblp/
db/conf/vldb/AmsalegFG95.html.

Konopnicki:1995:WQS

[1050] David Konopnicki and Oded Shmueli.

W3QS: a query system for the World-Wide Web. In Dayal et al. [1902], pages 54–65. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KonopnickiS95.html.

Yan:1995:DRI

[1051] Tak W. Yan and Hector Garcia-Molina. Duplicate removal in information dissemination. In Dayal et al. [1902], pages 66–77. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/YanG95.html.

Gravano:1995:GGV

[1052] Luis Gravano and Hector Garcia-Molina. Generalizing GlOSS to vectorspace databases and broker hierarchies. In Dayal et al. [1902], pages 78–89. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GravanoG95.html.

Mogi:1995:HBC

[1053] Kazuhiko Mogi and Masaru Kitsuregawa. Hot block clustering for disk arrays with dynamic striping. In Dayal et al. [1902], pages 90–99. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/MogiK95.html.

Wu:1995:PSW

[1054] Kun-Lung Wu, Philip S. Yu, Jen-Yao Chung, and James Z. Teng. A performance study of workfile disk management for concurrent mergesorts in a multiprocessor database system. In Dayal et al. [1902], pages 100– 109. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WuYCT95.html.

Chen:1995:DDH

[1055] Ling Tony Chen, Doron Rotem, and Sidhar Seshadri. Declustering databases on heterogeneous disk systems. In Dayal et al. [1902], pages 110–121. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChenRS95.html.

Bestavros:1995:VCS

[1056] Azer Bestavros and Spyridon Braoudakis. Value-cognizant speculative concurrency control. In Dayal et al. [1902], pages 122–133. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/BestavrosB95.
html.

Kornacker:1995:HCL

[1057] Marcel Kornacker and Douglas Banks. High-concurrency locking in R-trees. In Dayal et al. [1902], pages 134– 145. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KornackerB95.html.

Goyal:1995:ICC

[1058] Brajesh Goyal, Jayant R. Haritsa, S. Seshadri, and V. Srinivasan. Index concurrency control in firm real-time DBMS. In Dayal et al. [1902], pages 146–157. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GoyalHSS95.html.

Buneman:1995:DTS

[1059] Peter Buneman, Susan B. Davidson, Kyle Hart, G. Christian Overton, and Limsoon Wong. A data transformation system for biological data sources. In Dayal et al. [1902], pages 158– 169. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BunemanDHOW95.html.

Ferrandina:1995:SDE

[1060] Fabrizio Ferrandina, Thorsten Meyer, Roberto Zicari, Guy Ferran, and Joëlle Madec. Schema and database evolution in the O2 object database system. In Dayal et al. [1902], pages 170–181. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FerrandinaMZFM95.html.

Subieta:1995:POO

[1061] Kazimierz Subieta, Yahiko Kambayashi, and Jacek Leszczylowski. Procedures in object-oriented query languages. In Dayal et al. [1902], pages 182–193. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SubietaKL95.html.

Rusinkiewicz:1995:TCT

[1062] Marek Rusinkiewicz, Wolfgang Klas, Thomas Tesch, Jürgen Wäsch, and Peter Muth. Towards a cooperative transaction model — the cooperative activity model. In Dayal et al. [1902], pages 194–205. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        RusinkiewiczKTWM95.html.

Barga:1995:PMI

[1063] Roger S. Barga and Calton Pu. A practical and modular method to implement extended transaction models. In Dayal et al. [1902], pages 206– 217. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BargaP95.html.

Ammann:1995:UFM

[1064] Paul Ammann, Sushil Jajodia, and Indrakshi Ray. Using formal methods to reason about semantics-based decompositions of transactions. In Dayal et al. [1902], pages 218–227. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/AmmannJR95.html.

Steinbrunn:1995:BJD

[1065] Michael Steinbrunn, Klaus Peithner, Guido Moerkotte, and Alfons Kemper. Bypassing joins in disjunctive queries. In Dayal et al. [1902], pages 228–238. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SteinbrunnPMK95.html.

Hasan:1995:CCP

[1066] Waqar Hasan and Rajeev Motwani. Coloring away communication in parallel query optimization. In Dayal et al. [1902], pages 239–250. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/HasanM95.html.

Lu:1995:FSA

[1067] Hongjun Lu, Kian-Lee Tan, and Son Dao. The fittest survives: An adaptive approach to query optimization. In Dayal et al. [1902], pages 251– 262. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LuTD95.html.

Tresch:1995:TCS

[1068] Markus Tresch, Neal Palmer, and Allen Luniewski. Type classification of semi-structured documents. In Dayal et al. [1902], pages 263– 274. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
TreschPL95.html.

Moser:1995:MBM

[1069] Frank Moser, Achim Kraiss, and Wolfgang Klas. L/MRP: a buffer management strategy for interactive continuous data flows in a multimedia DBMS. In Dayal et al. [1902], pages 275–286. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MoserKK95.html.

Chaudhuri:1995:ARC

[1070] Surajit Chaudhuri, Shahram Ghandeharizadeh, and Cyrus Shahabi. Avoiding retrieval contention for composite multimedia objects. In Dayal et al. [1902], pages 287–298. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/ChaudhuriGS95.
html.

Belussi:1995:ESS

[1071] Alberto Belussi and Christos Faloutsos. Estimating the selectivity of spatial queries using the ‘correlation’ fractal dimension. In Dayal et al. [1902], pages 299–310. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BelussiF95.html.

Haas:1995:SBE

[1072] Peter J. Haas, Jeffrey F. Naughton, S. Seshadri, and Lynne Stokes. Sampling-based estimation of the number of distinct values of an attribute. In Dayal et al. [1902], pages 311– 322. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HaasNSS95.html.

Gardarin:1995:CMC

[1073] Georges Gardarin, Jean-Robert Gruser, and Zhao-Hui Tang. A cost model for clustered object-oriented databases. In Dayal et al. [1902], pages 323– 334. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GardarinGT95.html.

Flodin:1995:POO

[1074] Staffan Flodin and Tore Risch. Processing object-oriented queries with invertible late bound functions. In Dayal et al. [1902], pages 335–344. ISBN 1-55860-379-4. LCCN QA76.9.D3

                                        I61 1995. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/FlodinR95.
                                        html.

Yan:1995:EAL

[1075] Weipeng P. Yan and Per-Åke Larson. Eager aggregation and lazy aggregation. In Dayal et al. [1902], pages 345–357. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
YanL95.html.

Gupta:1995:AQP

[1076] Ashish Gupta, Venky Harinarayan, and Dallan Quass. Aggregate-query processing in data warehousing environments. In Dayal et al. [1902], pages 358–369. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GuptaHQ95.html.

Bassiliades:1995:NUD

[1077] N. Bassiliades and Ioannis P. Vlahavas. A non-uniform data fragmentation strategy for parallel mainmemory database systems. In Dayal et al. [1902], pages 370– 381. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BassiliadesV95.html.

Mehta:1995:MIO

[1078] Manish Mehta and David J. DeWitt.

Managing intra-operator parallelism in parallel database systems. In Dayal et al. [1902], pages 382–394. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/MehtaD95.html.

Rahm:1995:DMR

[1079] Erhard Rahm and Robert Marek. Dynamic multi-resource load balancing in parallel database systems. In Dayal et al. [1902], pages 395–406. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/RahmM95.html.

Srikant:1995:MGA

[1080] Ramakrishnan Srikant and Rakesh Agrawal. Mining generalized association rules. In Dayal et al. [1902], pages 407–419. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SrikantA95.html.

Han:1995:DML

[1081] Jiawei Han and Yongjian Fu. Discovery of multiple-level association rules from large databases. In Dayal et al. [1902], pages 420–431. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/HanF95.html.

Savasere:1995:EAM

[1082] Ashoka Savasere, Edward Omiecinski, and Shamkant B. Navathe. An efficient algorithm for mining association rules in large databases. In Dayal et al. [1902], pages 432–444. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SavasereON95.
                                        html.

Gallersdorfer:1995:IPR

[1083] Rainer Gallersdörfer and Matthias Nicola. Improving performance in replicated databases through relaxed coherency. In Dayal et al. [1902], pages 445–456. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GallersdorferN95.html.

Lomet:1995:RRA

[1084] David B. Lomet and Mark R. Tuttle. Redo recovery after system crashes.

In Dayal et al. [1902], pages 457– 468. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LometT95.html.

Hvasshovd:1995:CTD

[1085] Svein-Olaf Hvasshovd, Øystein Torbjørnsen, Svein Erik Bratsberg, and Per Holager. The ClustRa telecom database: High availability, high throughput, and real-time response. In Dayal et al. [1902], pages 469– 477. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HvasshovdTBH95.html.

Lu:1995:NCA

[1086] Hongjun Lu, Rudy Setiono, and Huan Liu. NeuroRule: a connectionist approach to data mining. In Dayal et al. [1902], pages 478–489. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/LuSL95.html.

Agrawal:1995:FSS

[1087] Rakesh Agrawal, King-Ip Lin, Harpreet S.

Sawhney, and Kyuseok Shim. Fast similarity search in the presence of noise, scaling, and translation in time-series databases. In Dayal et al. [1902], pages 490–501. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AgrawalLSS95.html.

Agrawal:1995:QSH

[1088] Rakesh Agrawal, Giuseppe Psaila, Edward L. Wimmers, and Mohamed Zaït. Querying shapes of histories. In Dayal et al. [1902], pages 502–514. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/AgrawalPWZ95.
                                        html.

Milliner:1995:SAA

[1089] Steven Milliner, Athman Bouguettaya, and Mike P. Papazoglou. A scalable architecture for autonomous heterogeneous database interactions. In Dayal et al. [1902], pages 515–526. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MillinerBP95.
                                        html.

Haber:1995:ODT

[1090] Eben M. Haber, Yannis E. Ioannidis, and Miron Livny. OPOSSUM: Desk-top schema management through customizable visualization. In Dayal et al. [1902], pages 527–538. ISBN 1-55860-379-4. LCCN QA76.9.D3

                                        I61 1995. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/HaberIL95.
                                        html.

Kalus:1995:FRO

[1091] Christian Kalus and Peter Dadam. Flexible relations — operational support of variant relational structures. In Dayal et al. [1902], pages 539– 550. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KalusD95.html.

Evangelidis:1995:HTM

[1092] Georgios Evangelidis, David B. Lomet, and Betty Salzberg. The hB^{Π}-tree: a modified hB-tree supporting concurrency, recovery and node consolidation. In Dayal et al. [1902], pages 551–561. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
EvangelidisLS95.html.

Hellerstein:1995:GST

[1093] Joseph M. Hellerstein, Jeffrey F.

Naughton, and Avi Pfeffer. Generalized search trees for database systems. In Dayal et al. [1902], pages 562–573. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HellersteinNP95.html.

Brin:1995:NNS

[1094] Sergey Brin. Near neighbor search in large metric spaces. In Dayal et al. [1902], pages 574–584. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/Brin95.html.

Sarawagi:1995:QPT

[1095] Sunita Sarawagi. Query processing in tertiary memory databases. In Dayal et al. [1902], pages 585– 596. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sarawagi95.html.

Cabrera:1995:ADT

[1096] Luis-Felipe Cabrera, Robert Rees, and Wayne Hineman. Applying database technology in the ADSM mass storage system. In Dayal et al. [1902], pages 597–605. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CabreraRH95.html.

Hoel:1995:BSJ

[1097] Erik G. Hoel and Hanan Samet. Benchmarking spatial join operations with spatial output. In Dayal et al. [1902], pages 606–618. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/HoelS95.html.

Sistla:1995:SBR

[1098] A. Prasad Sistla, Clement T. Yu, Chengwen Liu, and King Liu. Similarity based retrieval of pictures using indices on spatial relationships. In Dayal et al. [1902], pages 619– 639. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SistlaYLL95.html.

Ramakrishnan:1995:SJE

[1099] Raghu Ramakrishnan, Hector Garcia-Molina, Gerhard Rossbach, Abraham Silberschatz, Gio Wiederhold, and Jaco Zijlstra. Scientific journals: Extinction or explosion? (panel). In Dayal et al. [1902], page 631. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995.

URL http://www.vldb.org/dblp/
db/conf/vldb/RamakrishnanGRSWZ95.
html.

Reuter:1995:DWM

[1100] Andreas Reuter, Stefano Ceri, Jim Gray, Betty Salzberg, and Gerhard Weikum. Databases and workflow management: What is it all about? (panel). In Dayal et al. [1902], page 632. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ReuterCGSW95.html.

Seligman:1995:MAH

[1101] Leonard J. Seligman, Nicholas J. Belkin, Erich J. Neuhold, Michael Stonebraker, and Gio Wiederhold.

Metrics for accessing heterogeneous data: Is there any hope? (panel). In Dayal et al. [1902], page 633. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/SeligmanBNSW95.
html.

Meier:1995:PDM

[1102] Andreas Meier. Providing database migration tools — A practitioner’s approach. In Dayal et al. [1902], pages 635–641. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Meier95.html.

Simon:1995:PRA

[1103] Eric Simon and Angelika Kotz Dittrich. Promises and realities of active database systems. In Dayal et al. [1902], pages 642–653. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

1995. URL http://www.vldb.org/
dblp/db/conf/vldb/SimonD95.html.

Didriksen:1995:DCP

[1104] Tor Didriksen, César A. Galindo-Legaria, and Eirik Dahle. Database de-centralization — A practical approach. In Dayal et al. [1902], pages 654–665. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DidriksenGD95.html.

Sumiya:1995:PSD

[1105] Kazutoshi Sumiya, Kouichi Yasutake, Hirohiko Tanaka, Norio Sanada, and Yoshihiko Imai. A product specification database for visual prototyping. In Dayal et al. [1902], pages 666–676. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SumiyaYTSI95.html.

Vaskevitch:1995:VLD

[1106] David Vaskevitch. Very large databases: How large, how different? In Dayal et al. [1902], pages 677– 685. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Vaskevitch95.html.

Wang:1995:DQP

[1107] Yun Wang. DB2 query parallelism: Staging and implementation. In Dayal et al. [1902], pages 686–691. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Wang95.html.

Fecteau:1995:MDP

[1108] Gilles Fecteau. Managing a DB2 parallel edition database. In Dayal et al. [1902], pages 692–693. ISBN 1-55860-379-4. LCCN QA76.9.D3

                                        I61 1995. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Fecteau95.
                                        html.

Lindsay:1995:DCS

[1109] Bruce G. Lindsay. DB2 Common Server: Technology, progress, & directions. In Dayal et al. [1902], pages 694–695. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Lindsay95.html.

Catozzi:1995:SVU

[1110] John Catozzi and Sorana Rabinovici. OS support for VLDBs: Unix enhancements for the Teradata database.

In Dayal et al. [1902], pages 696– 701. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CatozziR95.html.

Orenstein:1995:ARD

[1111] Jack A. Orenstein and D. N. Kamber. Accessing a relational database through an object-oriented database interface. In Dayal et al. [1902], pages 702–705. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
OrensteinK95.html.

Spintzik:1995:IOX

[1112] Hannes Spintzik. Informix-Online XPS: a dynamically scalable RDBMS for open parallel platforms. In Dayal et al. [1902], page 706. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995.

URL http://www.vldb.org/dblp/
db/conf/vldb/Spintzik95.html.

Hallmark:1995:OW

[1113] Gary Hallmark. The Oracle warehouse. In Dayal et al. [1902], pages 707–709. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hallmark95.html.

Leslie:1995:ESM

[1114] Harry Leslie, Rohit Jain, Dave Birdsall, and Hedieh Yaghmai. Efficient search of multi-dimensional B-trees. In Dayal et al. [1902], pages 710– 719. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61 1995. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LeslieJBY95.html.

Brown:1995:BSM

[1115] Paul Brown and Michael Stonebraker. BigSur: a system for the management of earth science data. In Dayal et al. [1902], pages 720–728. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

                                        1995. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BrownS95.html.

Hess:1996:VLD

[1116] Karl-Heinz Hess. Very large databases in a commercial application environment. In Vijayaraman et al. [1904], page 2. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Hess96.html.

Mannila:1996:DMM

[1117] H. Mannila. Data mining: Machine learning, statistics, and databases. In Svensson and French [1903], pages 2– 11. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Gupta:1996:KAC

[1118] Umang Gupta. The changing landscape of the software industry and its implications for India. In Vijayaraman et al. [1904], page 2. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Gupta96.html.

Carey:1996:ODD

[1119] Michael J. Carey and David J. De-Witt. Of objects and databases: a decade of turmoil. In Vijayaraman et al. [1904], pages 3–14. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/CareyD96.html.

Malvestuto:1996:CST

[1120] F. M. Malvestuto and M. Moscarini.

Censoring statistical tables to protect sensitive information: Easy and hard problems. In Svensson and French [1903], pages 12–21. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Sevcik:1996:FTM

[1121] Kenneth C. Sevcik and Nick Koudas.

Filter trees for managing spatial data over a range of size granularities. In Vijayaraman et al. [1904], pages 16–27. ISBN 1-55860-382-4. LCCN QA76.9D3

I559 1996. URL http://www.vldb.
org/dblp/db/conf/vldb/SevcikK96.
html.

Bezenchek:1996:DSR

[1122] A. Bezenchek, M. Rafanelli, and L. Tininini. A data structure for representing aggregate data. In Svensson and French [1903], pages 22–31. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Berchtold:1996:XTI

[1123] Stefan Berchtold, Daniel A. Keim, and Hans-Peter Kriegel. The Xtree: An index structure for highdimensional data. In Vijayaraman et al. [1904], pages 28–39. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/BerchtoldKK96.
html.

Cotofrei:1996:SD

[1124] P. Cotofrei and H. Luchian. Statistical dependencies. In Svensson and French [1903], pages 32–43. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Faloutsos:1996:ADQ

[1125] Christos Faloutsos and Volker Gaede. Analysis of n-dimensional quadtrees using the Hausdorff fractal dimension. In Vijayaraman et al. [1904], pages 40–50. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        FaloutsosG96.html.

Page:1996:ATI

[1126] C. G. Page. Astronomical tables, 2-D indexing, and fuzzy-joins. In Svensson and French [1903], pages 44–52. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Scheuermann:1996:WDW

[1127] Peter Scheuermann, Junho Shim, and Radek Vingralek. WATCH-MAN: a data warehouse intelligent cache manager. In Vijayaraman et al. [1904], pages 51– 62. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ScheuermannSV96.html.

Yazdani:1996:SMI

[1128] N. Yazdani and Z. M. Oezsoyoglu. Sequence matching of images. In Svensson and French [1903], pages 53– 62. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Labio:1996:ESD

[1129] Wilburt J. Labio and Hector Garcia-Molina. Efficient snapshot differential algorithms for data warehousing. In Vijayaraman et al. [1904], pages 63–74. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LabioG96.html.

Hachem:1996:AQA

[1130] N. Hachem, C. Bao, and S. Taylor.

Approximate query answering in numerical databases. In Svensson and French [1903], pages 63–75. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Staudt:1996:IME

[1131] Martin Staudt and Matthias Jarke.

Incremental maintenance of externally materialized views. In Vijayaraman et al. [1904], pages 75–86. ISBN 1-55860-382-4. LCCN QA76.9D3

I559 1996. URL http://www.vldb.
org/dblp/db/conf/vldb/StaudtJ96.
html.

Anjur:1996:FTV

[1132] V. Anjur, Y. E. Ioannidis, and M. Livny. FROG and TURTLE: Visual bridges between files and objectoriented data. In Svensson and French [1903], pages 76–85. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Lee:1996:SMD

[1133] T. Lee, T. Bozkaya, H.-C. Kuo, and G. Oezsoyoglu. A scientific multimedia database system for polymer science experiments. In Svensson and French [1903], pages 86–97. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Chaudhuri:1996:OQU

[1134] Surajit Chaudhuri and Kyuseok Shim. Optimization of queries with userdefined predicates. In Vijayaraman et al. [1904], pages 87–98. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/ChaudhuriS96.
html.

Seamons:1996:PAA

[1135] K. E. Seamons, Y. Chen, M. Winslett, and Y. Cho. Persistent array access using server-directed I/O. In Svensson and French [1903], pages 98–107. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Seshadri:1996:DIS

[1136] Praveen Seshadri, Miron Livny, and Raghu Ramakrishnan. The design and implementation of a sequence database system. In Vijayaraman et al. [1904], pages 99–110. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/SeshadriLR96.
html.

Rotem:1996:EAS

[1137] D. Rotem and J. L. Zhao. Extendible arrays for statistical databases and OLAP applications. In Svensson and French [1903], pages 108–117. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

McKenna:1996:ETB

[1138] William J. McKenna, Louis Burger, Chi Hoang, and Melissa Truong. EROC: a toolkit for building NEATO query optimizers. In Vijayaraman et al. [1904], pages 111–121. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/McKennaBHT96.
html.

Li:1996:OSQ

[1139] C. Li and X. S. Wang. Optimizing statistical queries by exploiting orthogonality and interval properties of grouping relations. In Svensson and French [1903], pages 118–129. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Meo:1996:NSL

[1140] Rosa Meo, Giuseppe Psaila, and Stefano Ceri. A new SQL-like operator for mining association rules. In Vijayaraman et al. [1904], pages 122–133. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/MeoPC96.html.

Olenski:1996:PPI

[1141] J. Olenski. Practical problems of implementing metadata standards in official statistics. In Svensson and French [1903], pages 130–149. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Toivonen:1996:SLD

[1142] Hannu Toivonen. Sampling large databases for association rules. In Vijayaraman et al. [1904], pages 134– 145. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Toivonen96.html.

Fukuda:1996:CED

[1143] Takeshi Fukuda, Yasuhiko Morimoto, Shinichi Morishita, and Takeshi Tokuyama. Constructing efficient decision trees by using optimized numeric association rules. In Vijayaraman et al. [1904], pages 146– 155. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FukudaMMT96.html.

Gillman:1996:DPU

[1144] D. W. Gillman, M. V. Appel, and W. P. LaPlant. Design principles for a unified statistical data/metadata system. In Svensson and French [1903], pages 150–155. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Dean:1996:QAM

[1145] P. Dean and B. Sundgren. Quality aspects of a modern database service. In Svensson and French [1903], pages 156– 161. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Sarawagi:1996:RQE

[1146] Sunita Sarawagi and Michael Stonebraker. Reordering query execution in tertiary memory databases. In Vijayaraman et al. [1904], pages 156– 167. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SarawagiS96.html.

Lewis:1996:WCD

[1147] P. Lewis. What clinical database management system does the NHS need? In Svensson and French [1903], pages 162–167. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

DeGiacomo:1996:CDM

[1148] G. De Giacomo and P. Naggar. Conceptual data model with structured objects for statistical databases. In Svensson and French [1903], pages 168– 175. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

VandenBercken:1996:QPT

[1149] Jochen Van den Bercken and Bernhard Seeger. Query processing techniques for multiversion access methods. In Vijayaraman et al. [1904], pages 168–179. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BerckenS96.html.

Kemp:1996:UFD

[1150] G. J. L. Kemp, J. Dupont, and P. M. D.

Gray. Using the functional data model to integrate distributed biological data sources. In Svensson and French [1903], pages 176–185. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Boehlen:1996:CTD

[1151] M. H. Boehlen, R. T. Snodgrass, and M. D. Soo. Coalescing in temporal databases. In Vijayaraman et al. [1904], pages 180–191. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

Bohlen:1996:CTD

[1152] Michael H. Böhlen, Richard T. Snodgrass, and Michael D. Soo. Coalescing in temporal databases. In Vijayaraman et al. [1904], pages 180– 191. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BohlenSS96.html.

Washburn:1996:RHI

[1153] G. A. Washburn, L. M. L. Delcambre, and M. A. Whiting. Reuse of highlevel information requests: Leveraging the investment. In Svensson and French [1903], pages 186–197. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Ogawa:1996:EED

[1154] Yasushi Ogawa. Effective and efficient document ranking without using a large lexicon. In Vijayaraman et al. [1904], pages 192–202. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Ogawa96.html.

Kapetanios:1996:EPK

[1155] E. Kapetanios. Extracting and providing knowledge within an objectcentered scientific information system for atmospheric research. In Svensson and French [1903], pages 198– 207. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Carey:1996:PIQ

[1156] Michael J. Carey, Laura M. Haas, Vivekananda Maganty, and John H. Williams. PESTO: An integrated query/browser for object databases. In Vijayaraman et al. [1904], pages 203–214. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CareyHMW96.html.

Dreyer:1996:RDR

[1157] W. Dreyer, D. Schmidt, A. K. Dittrich, and M. Bleichenbacher. Requirements and design of replication services for a time series management system. In Svensson and French [1903], pages 208– 219. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Korn:1996:FNN

[1158] Flip Korn, Nikolaos Sidiropoulos, Christos Faloutsos, Eliot Siegel, and Zenon Protopapas. Fast nearest neighbor search in medical image databases. In Vijayaraman et al. [1904], pages 215–226. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KornSFSP96.html.

Sargent:1996:DID

[1159] R. Sargent, D. Fuhrman, T. Critchlow, and T. Di Sera. The design and implementation of a database for human genome research. In Svensson and French [1903], pages 220–225. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Kerschberg:1996:DIA

[1160] L. Kerschberg, H. Gomaa, D. Menasce, and J. P. Yoon. Data and information architectures for large-scale distributed data intensive information systems. In Svensson and French [1903], pages 226– 233. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Suciu:1996:QDV

[1161] Dan Suciu. Query decomposition and view maintenance for query languages for unstructured data. In Vijayaraman et al. [1904], pages 227–238. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Suciu96.html.

Shoshani:1996:CTS

[1162] A. Shoshani, P. Holland, J. Jacobsen, and D. Mitra. Characterization of temporal sequences in geophysical databases. In Svensson and French [1903], pages 234–241. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Lakshmanan:1996:SLI

[1163] Laks V. S. Lakshmanan, Fereidoon Sadri, and Iyer N. Subramanian. SchemaSQL — A language for interoperability in relational multi-database systems. In Vijayaraman et al. [1904], pages 239– 250. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LakshmananSS96.html.

Falkenroth:1996:CIT

[1164] E. Falkenroth. Computational indexes for time series. In Svensson and French [1903], pages 242–251. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Levy:1996:QHI

[1165] Alon Y. Levy, Anand Rajaraman, and Joann J. Ordille. Querying heterogeneous information sources using source descriptions. In Vijayaraman et al. [1904], pages 251–262. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/LevyRO96.html.

Son:1996:ETJ

[1166] D. Son and R. Elmasri. Efficient temporal join processing using time index. In Svensson and French [1903], pages 252–261. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Sheth:1996:SSW

[1167] Amit P. Sheth, Krys Kochut, John A.

Miller, Devashish Worah, Souvik Das, Chenye Lin, Devanand Palaniswami, John Lynch, and Ivan Shevchenko. Supporting state-wide immunization tracking using multi-paradigm workflow technology. In Vijayaraman et al. [1904], pages 263– 273. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShethKMWDLPLS96.html.

Ioannidis:1996:ZDE

[1168] Yannis E. Ioannidis, Miron Livny, S. Gupta, and Nagavamsi Ponnekanti. ZOO: a desktop experiment management environment. In Vijayaraman et al. [1904], pages 274– 285. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
IoannidisLGP96.html.

Viveros:1996:ADM

[1169] Marisa S. Viveros, John P. Nearhos, and Michael J. Rothman. Applying data mining techniques to a health insurance information system. In Vijayaraman et al. [1904], pages 286– 294. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ViverosNR96.html.

Chatziantoniou:1996:QMF

[1170] Damianos Chatziantoniou and Kenneth A. Ross. Querying multiple features of groups in relational databases. In Vijayaraman et al. [1904], pages 295–306. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChatziantoniouR96.html.

Faloutsos:1996:MSD

[1171] Christos Faloutsos, Yossi Matias, and Abraham Silberschatz. Modeling skewed distribution using multifractals and the ‘80–20’ law. In Vijayaraman et al. [1904], pages 307–317. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/FaloutsosMS96.
                                        html.

Srivastava:1996:AQA

[1172] Divesh Srivastava, Shaul Dar, H. V. Jagadish, and Alon Y. Levy. Answering queries with aggregation using views. In Vijayaraman et al. [1904], pages 318–329. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SrivastavaDJL96.html.

Dar:1996:SDC

[1173] Shaul Dar, Michael J. Franklin, Björn Tór Jónsson, Divesh Srivastava, and Michael Tan. Semantic data caching and replacement. In Vijayaraman et al. [1904], pages 330–341. ISBN 1-55860-382-4. LCCN QA76.9D3

I559 1996. URL http://www.vldb.
org/dblp/db/conf/vldb/DarFJST96.
html.

Diwan:1996:CTM

[1174] A. A. Diwan, Sanjeeva Rane, S. Seshadri, and S. Sudarshan. Clustering techniques for minimizing external path length. In Vijayaraman et al. [1904], pages 342–353. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

URL http://www.vldb.org/dblp/
db/conf/vldb/DiwanRSS96.html.

Acharya:1996:DUB

[1175] Swarup Acharya, Michael J. Franklin, and Stanley B. Zdonik. Disseminating updates on broadcast disks. In Vijayaraman et al. [1904], pages 354– 365. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AcharyaFZ96.html.

Suciu:1996:IAP

[1176] Dan Suciu. Implementation and analysis of a parallel collection query language. In Vijayaraman et al. [1904], pages 366–377. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Suciu96a.html.

Gardarin:1996:CQO

[1177] Georges Gardarin, Fei Sha, and Zhao-Hui Tang. Calibrating the query optimizer cost model of IRO-DB, an objectoriented federated database system. In Vijayaraman et al. [1904], pages 378– 389. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GardarinST96.html.

Gardarin:1996:CBS

[1178] Georges Gardarin, Jean-Robert Gruser, and Zhao-Hui Tang. Cost-based selection of path expression processing algorithms in object-oriented databases. In Vijayaraman et al. [1904], pages 390–401. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GardarinGT96.html.

Levy:1996:OCA

[1179] Alon Y. Levy. Obtaining complete answers from incomplete databases. In Vijayaraman et al. [1904], pages 402– 412. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Levy96.html.

Papakonstantinou:1996:OFM

[1180] Yannis Papakonstantinou, Serge Abiteboul, and Hector Garcia-Molina. Object fusion in mediator systems. In Vijayaraman et al. [1904], pages 413– 424. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PapakonstantinouAG96.html.

Vermeer:1996:RIC

[1181] Mark W. W. Vermeer and Peter M. G. Apers. The role of integrity constraints in database interoperation. In Vijayaraman et al. [1904], pages 425– 435. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
VermeerA96.html.

Bouganim:1996:DLB

[1182] Luc Bouganim, Daniela Florescu, and Patrick Valduriez. Dynamic load balancing in hierarchical parallel database systems. In Vijayaraman et al. [1904], pages 436–447. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/BouganimFV96.
html.

Poosala:1996:EQR

[1183] Viswanath Poosala and Yannis E.

Ioannidis. Estimation of query-result distribution and its application in parallel-join load balancing. In Vijayaraman et al. [1904], pages 448– 459. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PoosalaI96.html.

Rys:1996:ITP

[1184] Michael Rys, Moira C. Norrie, and Hans-Jörg Schek. Intra-transaction parallelism in the mapping of an object model to a relational multiprocessor system. In Vijayaraman et al. [1904], pages 460–471. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/RysNS96.html.

Bertino:1996:SPA

[1185] Elisa Bertino, Claudio Bettini, Elena Ferrari, and Pierangela Samarati. Supporting periodic authorizations and temporal reasoning in database access control. In Vijayaraman et al.

[1904], pages 472–483. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BertinoBFS96.
                                        html.

Cuppens:1996:QMD

[1186] Frédéric Cuppens. Querying a multilevel database: a logical analysis. In Vijayaraman et al. [1904], pages 484–494. ISBN 1-55860-382-4. LCCN QA76.9D3

                                        I559 1996. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Cuppens96.
                                        html.

Lee:1996:FII

[1187] Sin Yeung Lee and Tok Wang Ling. Further improvement on integrity constraint checking for stratifiable deductive databases. In Vijayaraman et al. [1904], pages 495–505. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/LeeL96.html.

Agarwal:1996:CMA

[1188] Sameet Agarwal, Rakesh Agrawal, Prasad Deshpande, Ashish Gupta, Jeffrey F. Naughton, Raghu Ramakrishnan, and Sunita Sarawagi. On the computation of multidimensional aggregates. In Vijayaraman et al. [1904], pages 506– 521. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AgarwalADGNRS96.html.

Shukla:1996:SEM

[1189] Amit Shukla, Prasad Deshpande, Jeffrey F. Naughton, and Karthikeyan Ramasamy. Storage estimation for multidimensional aggregates in the presence of hierarchies. In Vijayaraman et al. [1904], pages 522– 531. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShuklaDNR96.html.

Dyreson:1996:IRI

[1190] Curtis E. Dyreson. Information retrieval from an incomplete data cube. In Vijayaraman et al. [1904], pages 532–543. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Dyreson96.html.

Shafer:1996:SSP

[1191] John C. Shafer, Rakesh Agrawal, and Manish Mehta. SPRINT: a scalable parallel classifier for data mining. In Vijayaraman et al. [1904], pages 544– 555. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShaferAM96.html.

Ramakrishnan:1996:MDV

[1192] R. Ramakrishnan and D. Janaki Ram.

Modeling design versions. In Vijayaraman et al. [1904], pages 556– 566. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RamakrishnanR96.html.

Cochrane:1996:ITD

[1193] Roberta Cochrane, Hamid Pirahesh, and Nelson Mendonça Mattos. Integrating triggers and declarative constraints in SQL database systems. In Vijayaraman et al. [1904], pages 567– 578. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        CochranePM96.html.

Kleewein:1996:PIC

[1194] Jim Kleewein. Practical issues with commercial use of federated databases. In Vijayaraman et al. [1904], page 580. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Kleewein96.html.

Mattos:1996:SPC

[1195] Nelson Mendonça Mattos. Supporting procedural constructs in SQL compilers. In Vijayaraman et al. [1904], page 581. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Mattos96.html.

Padmanabhan:1996:ELD

[1196] Sriram Padmanabhan. Extracting large data sets using DB2 parallel edition. In Vijayaraman et al. [1904], page 582. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Padmanabhan96.html.

Klots:1996:CCO

[1197] Boris Klots. Cache coherency in Oracle parallel server. In Vijayaraman et al. [1904], page 583. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Klots96.html.

Gawlick:1996:SAO

[1198] Dieter Gawlick. Scalablity and availability in Oracle7 7.3. In Vijayaraman et al. [1904], page 584. ISBN 1-55860-382-4. LCCN QA76.9D3

I559 1996. URL http://www.vldb.
org/dblp/db/conf/vldb/Gawlick96.
html.

Sheth:1996:BAX

[1199] Amit P. Sheth. Bellcore’s ADAPT/X harness system for managing information on Internet and intranets. In Vijayaraman et al. [1904], page 585. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Sheth96.html.

Moni:1996:DPN

[1200] M. Moni. DISNIC-PLAN: a NICNET based distributed database for microlevel planning in India. In Vijayaraman et al. [1904], page 586. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Moni96.html.

Dasgupta:1996:LDR

[1201] A. R. Dasgupta. Large databases for remote sensing and GIS. In Vijayaraman et al. [1904], page 587. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

URL http://www.vldb.org/dblp/
db/conf/vldb/Dasgupta96.html.

Mohan:1996:DDW

[1202] Narendra Mohan. DWMS: Data warehouse management system. In Vijayaraman et al. [1904], page 588. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Mohan96.html.

Atal:1996:XAL

[1203] Sanket Atal. The XPS approach to loading and unloading terabyte databases. In Vijayaraman et al. [1904], page 589. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Atal96.html.

Rengarajan:1996:HSS

[1204] T. K. Rengarajan. How System 11 SQL server became fast. In Vijayaraman et al. [1904], page 590. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Rengarajan96.
html.

Hammond:1996:URI

[1205] Brad Hammond. Using referential integrity to easily define consistent subset replicas. In Vijayaraman et al. [1904], page 591. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Hammond96.html.

Celis:1996:QOT

[1206] Pedro Celis. The query optimizer in Tandem’s new ServerWare SQL product. In Vijayaraman et al. [1904], page 592. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Celis96.html.

Bhashyam:1996:TDC

[1207] Ramesh Bhashyam. TPC-D: The challenges, issues and results. In Vijayaraman et al. [1904], page 593. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bhashyam96.html.

Sullivan:1996:TSD

[1208] Mark Sullivan. Tribeca: a stream database manager for network traffic analysis. In Vijayaraman et al. [1904], page 594. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sullivan96.html.

DMVG-SGI:1996:MTS

[1209] Database Mining and Visualisation Group — SGI Inc. MineSet(tm): a system for high-end data mining and visualization. In Vijayaraman et al. [1904], page 595. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MiningGroup96.html.

Sacks-Davis:1996:SIM

[1210] Ron Sacks-Davis. The structured information manager: a database system for SGML documents. In Vijayaraman et al. [1904], page 596. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Sacks-Davis96.
html.

Raghavan:1996:LDW

[1211] Vijay V. Raghavan. Loading the data warehouse across various parallel architectures. In Vijayaraman et al. [1904], page 597. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Raghavan96.html.

Baru:1996:PFD

[1212] Chaitanya K. Baru. Performance of future database systems: Bottlenecks and bonananzas. In Vijayaraman et al. [1904], page 599. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

1996. URL http://www.vldb.org/
dblp/db/conf/vldb/Baru96.html.

Malaika:1996:DMS

[1213] Susan Malaika. DataBase management systems and the Internet. In Vijayaraman et al. [1904], pages 600–601. ISBN 1-55860-382-4. LCCN QA76.9D3

                                        I559 1996. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/Malaika96.
                                        html.

Gupta:1996:WDW

[1214] Ashish Gupta and Inderpal Singh Mumick. What is the data warehousing problem? (are materialized views the answer?). In Vijayaraman et al. [1904], page 602. ISBN 1-55860-382-4. LCCN QA76.9D3 I559

                                        1996. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/GuptaM96.html.

Jacobs:1997:IDM

[1215] Kenneth R. Jacobs. Innovation in database management: Computer science vs. engineering. In Jarke et al. [1905], page 2. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/Jacobs97.html.

Also known as VLDB’97.

Bernstein:1997:MR

[1216] Philip A. Bernstein, Brian Harry, Paul Sanders, David Shutt, and Jason Zander. The Microsoft repository. In Jarke et al. [1905], pages 3–12. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BernsteinHSSZ97.html. Also known

as VLDB’97.

Sellis:1997:MAM

[1217] Timos K. Sellis, Nick Roussopoulos, and Christos Faloutsos. Multidimensional access methods: Trees have grown everywhere. In Jarke et al. [1905], pages 13–14. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http://

www.vldb.org/dblp/db/conf/vldb/
SellisRF97.html. Also known as

VLDB’97.

Jagadish:1997:IOD

[1218] H. V. Jagadish, P. P. S. Narayan, S. Seshadri, S. Sudarshan, and Rama Kanneganti. Incremental organization for data recording and warehousing. In Jarke et al. [1905], pages 16–25. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JagadishNSSK97.html. Also known as

VLDB’97.

Huyn:1997:MVS

[1219] Nam Huyn. Multiple-view selfmaintenance in data warehousing environments. In Jarke et al. [1905], pages 26–35. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997.

URL http://www.vldb.org/dblp/
db/conf/vldb/Huyn97.html.  Also

known as VLDB’97.

Faloutsos:1997:RIS

[1220] Christos Faloutsos, H. V. Jagadish, and Nikolaos D. Sidiropoulos. Recovering information from summary data. In Jarke et al. [1905], pages 36–45. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

1997. URL http://www.vldb.org/
dblp/db/conf/vldb/FaloutsosJS97.

html. Also known as VLDB’97.

Marathe:1997:LMA

[1221] Arunprasad P. Marathe and Kenneth Salem. A language for manipulating arrays. In Jarke et al. [1905], pages 46–55. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        MaratheS97.html. Also known as

VLDB’97.

Jamil:1997:IAO

[1222] Hasan M. Jamil. Implementing abstract objects with inheritance in datalog neg. In Jarke et al. [1905], pages 56–65. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Jamil97.html.  Also known as

VLDB’97.

Seshadri:1997:CEA

[1223] Praveen Seshadri, Miron Livny, and Raghu Ramakrishnan. The case for enhanced abstract data types. In Jarke et al. [1905], pages 66–75. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/SeshadriLR97.

html. Also known as VLDB’97.

Ng:1997:IRM

[1224] Wee Teck Ng and Peter M. Chen. Integrating reliable memory in databases. In Jarke et al. [1905], pages 76– 85. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

NgC97.html. Also known as VLDB’97.

Rastogi:1997:LPV

[1225] Rajeev Rastogi, S. Seshadri, Philip Bohannon, Dennis Leinbaugh, Avi Silberschatz, and S. Sudarshan. Logical and physical versioning in main memory databases. In Jarke et al. [1905], pages 86–95. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BohannonLRSSS97.html. Also known

as VLDB’97.

Llirbat:1997:UVU

[1226] François Llirbat, Eric Simon, and Dimitri Tombroff. Using versions in update transactions: Application to integrity checking. In Jarke et al. [1905], pages 96–105. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LlirbatST97.html. Also known as

VLDB’97.

Gyssens:1997:FMD

[1227] Marc Gyssens and Laks V. S.

Lakshmanan. A foundation for multi-dimensional databases. In Jarke et al. [1905], pages 106– 115. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GyssensL97.html. Also known as

VLDB’97.

Ross:1997:FCS

[1228] Kenneth A. Ross and Divesh Srivastava. Fast computation of sparse datacubes. In Jarke et al. [1905], pages 116–125. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997.

URL http://www.vldb.org/dblp/
db/conf/vldb/RossS97.html. Also

known as VLDB’97.

Theodoratos:1997:DWC

[1229] Dimitri Theodoratos and Timos K.

Sellis. Data warehouse configuration. In Jarke et al. [1905], pages 126– 135. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

                                        TheodoratosS97.html. Also known as

VLDB’97.

Yang:1997:AMV

[1230] Jian Yang, Kamalakar Karlapalem, and Qing Li. Algorithms for materialized view design in data warehousing environment. In Jarke et al. [1905], pages 136–145. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/YangKL97.html.

Also known as VLDB’97.

Chaudhuri:1997:ECD

[1231] Surajit Chaudhuri and Vivek R. Narasayya. An efficient, cost-driven index selection tool for Microsoft SQL server. In Jarke et al. [1905], pages 146–155. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChaudhuriN97.html. Also known as

VLDB’97.

Baralis:1997:MVS

[1232] Elena Baralis, Stefano Paraboschi, and Ernest Teniente. Materialized view selection in a multidimensional database. In Jarke et al. [1905], pages 156– 165. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BaralisPT97.html. Also known as

VLDB’97.

Morimoto:1997:ECR

[1233] Yasuhiko Morimoto, Hiromu Ishii, and Shinichi Morishita. Efficient construction of regression trees with range and region splitting. In Jarke et al. [1905], pages 166–175. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

1997. URL http://www.vldb.org/
dblp/db/conf/vldb/MorimotoIM97.

html. Also known as VLDB’97.

Shafer:1997:PAH

[1234] John C. Shafer and Rakesh Agrawal.

Parallel algorithms for high-dimensional proximity joins. In Jarke et al. [1905], pages 176–185. ISBN 1-55860-470-7. LCCN QA76.9 D3

I61 1997. URL http://www.vldb.
org/dblp/db/conf/vldb/ShaferA97.

html. Also known as VLDB’97.

Wang:1997:SSI

[1235] Wei Wang, Jiong Yang, and Richard R.

Muntz. STING: a statistical information grid approach to spatial data mining. In Jarke et al. [1905], pages 186–195. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WangYM97.html.  Also known as

VLDB’97.

Gravano:1997:MRH

[1236] Luis Gravano and Hector Garcia-Molina. Merging ranks from heterogeneous Internet sources. In Jarke et al. [1905], pages 196– 205. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GravanoG97.html. Also known as

VLDB’97.

Atzeni:1997:WW

[1237] Paolo Atzeni, Giansalvatore Mecca, and Paolo Merialdo. To weave the Web. In Jarke et al. [1905], pages 206–215. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

                                        AtzeniMM97.html. Also known as

VLDB’97.

Florescu:1997:UPI

[1238] Daniela Florescu, Daphne Koller, and Alon Levy. Using probabilistic information in data integration. In Jarke et al. [1905], pages 216–225. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/FlorescuKLP97.

html. Also known as VLDB’97.

Triantafillou:1997:DDE

[1239] Peter Triantafillou and Thomas Papadakis. On-demand data elevation in a hierarchical multimedia storage server. In Jarke et al. [1905], pages 226–235. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        TriantafillouP97.html. Also known

as VLDB’97.

Christodoulakis:1997:POP

[1240] Stavros Christodoulakis, Peter Triantafillou, and Fenia Zioga. Principles of optimally placing data in tertiary storage libraries. In Jarke et al. [1905], pages 236– 245. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChristodoulakisTZ97.html.  Also

known as VLDB’97.

Kraiss:1997:VDM

[1241] Achim Kraiss and Gerhard Weikum. Vertical data migration in large nearline document archives based on Markov-chain predictions. In Jarke et al. [1905], pages 246–255. ISBN 1-55860-470-7. LCCN QA76.9 D3

I61 1997. URL http://www.vldb.
org/dblp/db/conf/vldb/KraissW97.

html. Also known as VLDB’97.

Vassalos:1997:DUQ

[1242] Vasilis Vassalos and Yannis Papakonstantinou. Describing and using the query capabilities of heterogeneous sources. In Jarke et al. [1905], pages 256–265. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
VassalosP97.html. Also known as

VLDB’97.

Roth:1997:DSI

[1243] Mary Tork Roth and Peter M. Schwarz.

Don’t scrap it, wrap it! A wrapper architecture for legacy data sources. In Jarke et al. [1905], pages 266– 275. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RothS97.html.  Also known as

VLDB’97.

Haas:1997:OQA

[1244] Laura M. Haas, Donald Kossmann, Edward L. Wimmers, and Jun Yang. Optimizing queries across diverse data sources. In Jarke et al. [1905], pages 276–285. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HaasKWY97.html. Also known as

VLDB’97.

Claussen:1997:OQU

[1245] Jens Claußen, Alfons Kemper, Guido Moerkotte, and Klaus Peithner. Optimizing queries with universal quantification in object-oriented and objectrelational databases. In Jarke et al.

[1905], pages 286–295. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

1997. URL http://www.vldb.org/
dblp/db/conf/vldb/ClaussenKMP97.

html. Also known as VLDB’97.

Garofalakis:1997:PQS

[1246] Minos N. Garofalakis and Yannis E. Ioannidis. Parallel query scheduling and optimization with timeand space-shared resources. In Jarke et al. [1905], pages 296– 305. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GarofalakisI97.html. Also known as

VLDB’97.

Pellenkoft:1997:CTB

[1247] Arjan Pellenkoft, César A. Galindo-Legaria, and Martin L. Kersten. The complexity of transformationbased join enumeration. In Jarke et al. [1905], pages 306–315. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http://

www.vldb.org/dblp/db/conf/vldb/
PellenkoftLK97.html. Also known as

VLDB’97.

Humborstad:1997:SAS

[1248] Rune Humborstad, Maitrayi Sabaratnam, Svein-Olaf Hvasshovd, and Øystein Torbjørnsen. 1-safe algorithms for symmetric site configurations. In Jarke et al. [1905], pages 316– 325. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HumborstadSHT97.html. Also known

as VLDB’97.

Stathatos:1997:ADB

[1249] Konstantinos Stathatos, Nick Roussopoulos, and John S. Baras. Adaptive data broadcast in hybrid networks. In Jarke et al. [1905], pages 326–335. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
StathatosRB97.html. Also known as

VLDB’97.

Eickler:1997:FDN

[1250] André Eickler, Alfons Kemper, and Donald Kossmann. Finding data in the neighborhood. In Jarke et al. [1905], pages 336–345. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http://

www.vldb.org/dblp/db/conf/vldb/
EicklerKK97.html. Also known as

VLDB’97.

Alsabti:1997:OPA

[1251] Khaled Alsabti, Sanjay Ranka, and Vineet Singh. A one-pass algorithm for accurately estimating quantiles for disk-resident data. In Jarke et al. [1905], pages 346– 355. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AlsabtiRS97.html. Also known as

VLDB’97.

Skubiszewski:1997:CGC

[1252] Marcin Skubiszewski and Patrick Valduriez. Concurrent garbage collection in O2. In Jarke et al. [1905], pages 356–365. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SkubiszewskiV97.html. Also known

as VLDB’97.

Ashwin:1997:GCO

[1253] S. Ashwin, Prasan Roy, S. Seshadri, Abraham Silberschatz, and S. Sudarshan. Garbage collection in object oriented databases using transactional cyclic reference counting. In Jarke et al. [1905], pages 366–375. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/AshwinRSSS97.

html. Also known as VLDB’97.

Zhang:1997:DMA

[1254] Weiye Zhang and Per-Åke Larson. Dynamic memory adjustment for external mergesort. In Jarke et al. [1905], pages 376–385. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ZhangL97.html.

Also known as VLDB’97.

Helmer:1997:EMM

[1255] Sven Helmer and Guido Moerkotte. Evaluation of main memory join algorithms for joins with subset join predicates. In Jarke et al. [1905], pages 386–395. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HelmerM97.html. Also known as

VLDB’97.

Huang:1997:SJU

[1256] Yun-Wu Huang, Ning Jing, and Elke A. Rundensteiner. Spatial joins using R-trees: Breadth-first traversal with global optimizations. In Jarke et al. [1905], pages 396–405. ISBN 1-55860-470-7. LCCN QA76.9 D3

                                        I61 1997. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/HuangJR97.

html. Also known as VLDB’97.

VandenBercken:1997:GAB

[1257] Jochen Van den Bercken, Bernhard Seeger, and Peter Widmayer. A generic approach to bulk loading multidimensional index structures. In Jarke et al. [1905], pages 406– 415. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BerckenSW97.html. Also known as

VLDB’97.

Lee:1997:RSS

[1258] Jong-Hak Lee, Young-Koo Lee, Kyu-Young Whang, and Il-Yeol Song. A region splitting strategy for physical database design of multidimensional file organizations. In Jarke et al. [1905], pages 416–425. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

1997. URL http://www.vldb.org/
dblp/db/conf/vldb/LeeLWS97.html.

Also known as VLDB’97.

Ciaccia:1997:MTE

[1259] Paolo Ciaccia, Marco Patella, and Pavel Zezula. M-tree: An efficient access method for similarity search in metric spaces. In Jarke et al. [1905], pages 426– 435. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CiacciaPZ97.html. Also known as

VLDB’97.

Goldman:1997:DEQ

[1260] Roy Goldman and Jennifer Widom.

DataGuides: Enabling query formulation and optimization in semistructured databases. In Jarke et al. [1905], pages 436–445. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http://

                                        www.vldb.org/dblp/db/conf/vldb/
                                        GoldmanW97.html. Also known as

VLDB’97.

Chakrabarti:1997:UTD

[1261] Soumen Chakrabarti, Byron Dom, Rakesh Agrawal, and Prabhakar Raghavan. Using taxonomy, discriminants, and signatures for navigating in text databases. In Jarke et al. [1905], pages 446– 455. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChakrabartiDAR97.html. Also known

as VLDB’97.

Chavda:1997:TOC

[1262] Manoj Chavda and Peter T. Wood. Towards an ODMG-compliant visual object query language. In Jarke et al. [1905], pages 456–465. ISBN 1-55860-470-7. LCCN QA76.9 D3

                                        I61 1997. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/ChavdaW97.

html. Also known as VLDB’97.

Gibbons:1997:FIM

[1263] Phillip B. Gibbons, Yossi Matias, and Viswanath Poosala. Fast incremental maintenance of approximate histograms. In Jarke et al. [1905], pages 466–475. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GibbonsMP97.html. Also known as

VLDB’97.

Chatziantoniou:1997:GPR

[1264] Damianos Chatziantoniou and Kenneth A. Ross. Groupwise processing of relational queries. In Jarke et al.

[1905], pages 476–485. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997.

URL http://www.vldb.org/dblp/
db/conf/vldb/ChatziantoniouR97.

html. Also known as VLDB’97.

Poosala:1997:SEA

[1265] Viswanath Poosala and Yannis E. Ioannidis. Selectivity estimation without the attribute value independence assumption. In Jarke et al. [1905], pages 486–495. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PoosalaI97.html. Also known as

VLDB’97.

Chang:1997:EMU

[1266] Edward Chang and Hector Garcia-Molina. Effective memory use in a media server. In Jarke et al. [1905], pages 496–505. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChangG97.html.  Also known as

VLDB’97.

Seidl:1997:EUA

[1267] Thomas Seidl and Hans-Peter Kriegel.

Efficient user-adaptable similarity search in large multimedia databases. In Jarke et al. [1905], pages 506–515. ISBN 1-55860-470-7. LCCN QA76.9

D3 I61 1997. URL http://www.vldb.
org/dblp/db/conf/vldb/SeidlK97.

html. Also known as VLDB’97.

Garofalakis:1997:RSE

[1268] Minos N. Garofalakis, Banu Özden, and Abraham Silberschatz. Resource scheduling in enhanced payper-view continuous media databases. In Jarke et al. [1905], pages 516– 525. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GarofalakisOS97.html. Also known

as VLDB’97.

Dessloch:1997:ISD

[1269] Stefan Deßloch and Nelson Mendonça Mattos. Integrating SQL databases with content-specific search engines. In Jarke et al. [1905], pages 528– 537. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DesslochM97.html. Also known as

VLDB’97.

Li:1997:FMD

[1270] Wen-Syan Li, K. Selçuk Candan, Kyoji Hirata, and Yoshinori Hara. Facilitating multimedia database exploration through visual interfaces and perpetual query reformulations. In Jarke et al. [1905], pages 538–547. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/LiCHH97.html. Also

known as VLDB’97.

Baumann:1997:GEM

[1271] Peter Baumann, Paula Furtado, Roland Ritsch, and Norbert Widmann. Geo/environmental and medical data management in the RasDaMan system. In Jarke et al. [1905], pages 548–552. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BaumannFRW97.html. Also known as

VLDB’97.

Whitehead:1997:GSU

[1272] Steven D. Whitehead, Himanshu Sinha, and Michael Murphy. GTE SuperPages: Using IR techniques for searching complex objects. In Jarke et al. [1905], pages 553–557. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

1997. URL http://www.vldb.org/
dblp/db/conf/vldb/WhiteheadSM97.

html. Also known as VLDB’97.

Hughes:1997:DME

[1273] Eric Hughes, Roman Ginis, Bhavani M. Thuraisingham, Peter C. Krupp, and John A. Maurer. Data manager for evolvable real-time command and control systems. In Jarke et al. [1905], pages 558– 561. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GinisHKMT97.html. Also known as

VLDB’97.

Kietz:1997:MID

[1274] Jörg-Uwe Kietz, Ulrich Reimer, and Martin Staudt. Mining insurance data at Swiss Life. In Jarke et al. [1905], pages 562–566. ISBN 1-55860-470-7. LCCN QA76.9 D3

I61 1997. URL http://www.vldb.
org/dblp/db/conf/vldb/KietzRS97.

html. Also known as VLDB’97.

Lockemann:1997:NGD

[1275] Peter C. Lockemann, Ulrike Kölsch, Arne Koschel, Ralf Kramer, Ralf Nikolai, Mechtild Wallrath, and Hans-Dirk Walter. The network as a global database: Challenges of interoperability, proactivity, interactiveness, legacy. In Jarke et al. [1905], pages 567– 574. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LockemannKKKN97.html. Also known

as VLDB’97.

Alonso:1997:DPS

[1276] Gustavo Alonso, Claus Hagen, Hans-Jörg Schek, and Markus Tresch. Distributed processing over standalone systems and applications. In Jarke et al. [1905], pages 575– 579. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AlonsoHST97.html. Also known as

VLDB’97.

Malon:1997:CDT

[1277] David M. Malon and Edward N. May. Critical database technologies for high energy physics. In Jarke et al. [1905], pages 580–584. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/MalonM97.html.

Also known as VLDB’97.

Kumar:1997:CAF

[1278] Suresh Kumar, Eng-Kee Kwang, and Divyakant Agrawal. Caprera: An activity framework for transaction processing on wide-area networks. In Jarke et al. [1905], pages 585–589. ISBN 1-55860-470-7. LCCN QA76.9

                                        D3 I61 1997. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/KumarKA97.

html. Also known as VLDB’97.

Bridge:1997:OUS

[1279] William Bridge, Ashok Joshi, M. Keihl, Tirthankar Lahiri, Juan Loaiza, and N. MacNaughton. The Oracle Universal Server Buffer Manager. In Jarke et al. [1905], pages 590–594. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61

                                        1997. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/BridgeJKLLM97.

html. Also known as VLDB’97.

Wildfogel:1997:ETH

[1280] D. Wildfogel and Ramana Yerneni.

Efficient testing of high performance transaction processing systems. In Jarke et al. [1905], pages 595– 598. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WildfogelY97.html. Also known as

VLDB’97.

Tygar:1998:AVA

[1281] J. D. Tygar. Atomicity versus anonymity: Distributed transactions for electronic commerce. In Gupta et al. [1906], pages 1–12. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

URL http://www.vldb.org/dblp/
db/conf/vldb/Tygar98.html. Also

known as VLDB.

Shaw:1998:TFC

[1282] David Elliot Shaw. Technology and the future of commerce and finance (abstract). In Gupta et al. [1906], page 13. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Shaw98.html. Also known as VLDB.

Meng:1998:DTD

[1283] Weiyi Meng, King-Lup Liu, Clement T.

Yu, Xiaodong Wang, Yuhsi Chang, and Naphtali Rishe. Determining text databases to search in the Internet. In Gupta et al. [1906], pages 14–25. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MengLYWCR98.html. Also known as

VLDB.

Goldman:1998:PSD

[1284] Roy Goldman, Narayanan Shivakumar, Suresh Venkatasubramanian, and Hector Garcia-Molina. Proximity search in databases. In Gupta et al. [1906], pages 26–37. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GoldmanSVG98.html. Also known as

VLDB.

Abiteboul:1998:IMM

[1285] Serge Abiteboul, Jason McHugh, Michael Rys, Vasilis Vassalos, and Janet L. Wiener. Incremental maintenance for materialized views over semistructured data. In Gupta et al. [1906], pages 38–49. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

URL http://www.vldb.org/dblp/
db/conf/vldb/AbiteboulMRVW98.html.

Also known as VLDB.

Johnson:1998:PMT

[1286] Theodore Johnson and Ethan L. Miller. Performance measurements of tertiary storage devices. In Gupta et al. [1906], pages 50– 61. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JohnsonM98.html. Also known as

VLDB.

Riedel:1998:ASL

[1287] Erik Riedel, Garth A. Gibson, and Christos Faloutsos. Active storage for large-scale data mining and multimedia. In Gupta et al. [1906], pages 62–73. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

RiedelGF98.html. Also known as

VLDB.

Garofalakis:1998:RSC

[1288] Minos N. Garofalakis, Yannis E. Ioannidis, and Banu Özden. Resource scheduling for composite multimedia objects. In Gupta et al. [1906], pages 74–85. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GarofalakisIO98.html. Also known

as VLDB.

Graefe:1998:HJH

[1289] Goetz Graefe, Ross Bunker, and Shaun Cooper. Hash joins and hash teams in Microsoft SQL server. In Gupta et al. [1906], pages 86– 97. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GraefeBC98.html. Also known as

VLDB.

Helmer:1998:DJO

[1290] Sven Helmer, Till Westmann, and Guido Moerkotte. Diag-Join: An opportunistic join algorithm for 1: N relationships. In Gupta et al. [1906], pages 98–109. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HelmerWM98.html. Also known as

VLDB.

Braumandl:1998:EFJ

[1291] Reinhard Braumandl, Jens Claußen, and Alfons Kemper. Evaluating functional joins along nested reference sets in object-relational and object-oriented databases. In Gupta et al. [1906], pages 110–121. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BraumandlCK98.html. Also known as

VLDB.

Milo:1998:USM

[1292] Tova Milo and Sagit Zohar. Using schema matching to simplify heterogeneous data translation. In Gupta et al. [1906], pages 122–133. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/MiloZ98.html. Also

known as VLDB.

Gingras:1998:NSM

[1293] Frédéric Gingras and Laks V. S. Lakshmanan. nD-SQL: a multi-dimensional language for interoperability and OLAP. In Gupta et al. [1906], pages 134–145. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GingrasL98.html. Also known as

VLDB.

deFerreiraRezende:1998:HPM

[1294] Fernando de Ferreira Rezende and Klaudia Hergula. The heterogeneity problem and middleware technology: Experiences with and performance of database gateways. In Gupta et al. [1906], pages 146– 157. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        RezendeH98.html. Also known as

VLDB.

Carey:1998:RBD

[1295] Michael J. Carey and Donald Kossmann. Reducing the braking distance of an SQL query engine. In Gupta et al. [1906], pages 158–169. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/CareyK98.html.

Also known as VLDB.

Lin:1998:QCT

[1296] Ling Lin and Tore Risch. Querying continuous time sequences. In Gupta et al. [1906], pages 170– 181. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

LinR98.html. Also known as VLDB.

Grovlen:1998:LCC

[1297] Øystein Grøvlen, Svein-Olaf Hvasshovd, and Øystein Torbjørnsen. Low-cost compensation-based query processing. In Gupta et al. [1906], pages 182– 193. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GrovlenHT98.html. Also known as

VLDB.

Weber:1998:QAP

[1298] Roger Weber, Hans-Jörg Schek, and Stephen Blott. A quantitative analysis and performance study for similaritysearch methods in high-dimensional spaces. In Gupta et al. [1906], pages 194–205. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
WeberSB98.html. Also known as

VLDB.

Ankerst:1998:IAS

[1299] Mihael Ankerst, Bernhard Braunmüller, Hans-Peter Kriegel, and Thomas Seidl. Improving adaptable similarity query processing by using approximations.

In Gupta et al. [1906], pages 206– 217. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AnkerstBKS98.html. Also known as

VLDB.

Ishikawa:1998:MQD

[1300] Yoshiharu Ishikawa, Ravishankar Subramanya, and Christos Faloutsos. MindReader: Querying databases through multiple examples. In Gupta et al. [1906], pages 218–227. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/IshikawaSF98.

html. Also known as VLDB.

Ganguly:1998:DAP

[1301] Sumit Ganguly. Design and analysis of parametric query optimization algorithms. In Gupta et al. [1906], pages 228–238. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ganguly98.html. Also known as

VLDB.

Cherniack:1998:IFS

[1302] Mitch Cherniack and Stanley B. Zdonik. Inferring function semantics to optimize queries. In Gupta et al. [1906], pages 239–250. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/CherniackZ98.

html. Also known as VLDB.

Nippl:1998:TCB

[1303] Clara Nippl and Bernhard Mitschang. TOPAZ: a cost-based, rule-driven, multi-phase parallelizer. In Gupta et al. [1906], pages 251–262. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/NipplM98.html.

Also known as VLDB.

Shivakumar:1998:FAP

[1304] Narayanan Shivakumar, Hector Garcia-Molina, and Chandra Chekuri. Filtering with approximate predicates. In Gupta et al. [1906], pages 263– 274. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShivakumarGC98.html. Also known as

VLDB.

Jagadish:1998:OHQ

[1305] H. V. Jagadish, Nick Koudas, S. Muthukrishnan, Viswanath Poosala, Kenneth C. Sevcik, and Torsten Suel. Optimal histograms with quality guarantees. In Gupta et al. [1906], pages 275–286. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JagadishKMPSS98.html. Also known

as VLDB.

Greco:1998:BPD

[1306] Sergio Greco. Binding propagation in disjunctive databases. In Gupta et al. [1906], pages 287–298. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

URL http://www.vldb.org/dblp/
db/conf/vldb/Greco98.html. Also

known as VLDB.

Fang:1998:CIQ

[1307] Min Fang, Narayanan Shivakumar, Hector Garcia-Molina, Rajeev Motwani, and Jeffrey D. Ullman. Computing Iceberg queries efficiently. In Gupta et al. [1906], pages 299– 310. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
FangSGMU98.html. Also known as

VLDB.

Gibson:1998:CCD

[1308] David Gibson, Jon M. Kleinberg, and Prabhakar Raghavan. Clustering categorical data: An approach based on dynamical systems. In Gupta et al. [1906], pages 311– 322. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GibsonKR98.html. Also known as

VLDB.

Ester:1998:ICM

[1309] Martin Ester, Hans-Peter Kriegel, Jörg Sander, Michael Wimmer, and Xiaowei Xu. Incremental clustering for mining in a data warehousing environment. In Gupta et al. [1906], pages 323–333. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
EsterKSWX98.html. Also known as

VLDB.

Garcia:1998:ONS

[1310] Yván J. García, Mario A. Lopez, and Scott T. Leutenegger. On optimal node splitting for R-trees. In Gupta et al. [1906], pages 334– 344. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GarciaLL98.html. Also known as

VLDB.

Bliujute:1998:RTB

[1311] Rasa Bliujute, Christian S. Jensen, Simonas Saltenis, and Giedrius Slivinskas. R-tree based indexing of nowrelative bitemporal data. In Gupta et al. [1906], pages 345–356. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/BliujuteJSS98.

html. Also known as VLDB.

Ooi:1998:FHD

[1312] Beng Chin Ooi, Cheng Hian Goh, and Kian-Lee Tan. Fast high-dimensional data search in incomplete databases. In Gupta et al. [1906], pages 357– 367. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

OoiGT98.html. Also known as VLDB.

Ramaswamy:1998:DIP

[1313] Sridhar Ramaswamy, Sameer Mahajan, and Abraham Silberschatz. On the discovery of interesting patterns in association rules. In Gupta et al. [1906], pages 368–379. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/RamaswamyMS98.

html. Also known as VLDB.

Morimoto:1998:AMA

[1314] Yasuhiko Morimoto, Takeshi Fukuda, Hirofumi Matsuzawa, Takeshi Tokuyama, and Kunikazu Yoda. Algorithms for mining association rules for binary segmentations of huge categorical databases. In Gupta et al. [1906], pages 380–391. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MorimotoFMTY98.html. Also known as

VLDB.

Knorr:1998:AMD

[1315] Edwin M. Knorr and Raymond T. Ng.

Algorithms for mining distance-based outliers in large datasets. In Gupta et al. [1906], pages 392–403. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

                                        1998. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KnorrN98.html.

Also known as VLDB.

Rastogi:1998:PDT

[1316] Rajeev Rastogi and Kyuseok Shim. PUBLIC: a decision tree classifier that integrates building and pruning. In Gupta et al. [1906], pages 404– 415. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        RastogiS98.html. Also known as

VLDB.

Gehrke:1998:RFF

[1317] Johannes Gehrke, Raghu Ramakrishnan, and Venkatesh Ganti. Rain-Forest — A framework for fast decision tree construction of large datasets. In Gupta et al. [1906], pages 416– 427. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GehrkeRG98.html. Also known as

VLDB.

Sheikholeslami:1998:WMR

[1318] Gholamhosein Sheikholeslami, Surojit Chatterjee, and Aidong Zhang. WaveCluster: a multi-resolution clustering approach for very large spatial databases. In Gupta et al. [1906], pages 428–439. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SheikholeslamiCZ98.html.  Also

known as VLDB.

Ozsu:1998:AAB

[1319] M. Tamer Özsu, Kaladhar Voruganti, and Ronald C. Unrau. An asynchronous avoidance-based cache consistency algorithm for client caching DBMSs. In Gupta et al. [1906], pages 440–451. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

OzsuVU98.html. Also known as VLDB.

Muth:1998:DIP

[1320] Peter Muth, Patrick E. O’Neil, Achim Pick, and Gerhard Weikum. Design, implementation, and performance of the LHAM log-structured history data access method. In Gupta et al. [1906], pages 452–463. ISBN 1-55860-566-5. LCCN QA76.9.D3

I559 1998. URL http://www.vldb.
org/dblp/db/conf/vldb/MuthOPW98.

html. Also known as VLDB.

George:1998:SBF

[1321] Binto George and Jayant R. Haritsa. Secure buffering in firm realtime database systems. In Gupta et al. [1906], pages 464–475. ISBN 1-55860-566-5. LCCN QA76.9.D3

I559 1998. URL http://www.vldb.
org/dblp/db/conf/vldb/GeorgeH98.

html. Also known as VLDB.

Moerkotte:1998:SMA

[1322] Guido Moerkotte. Small materialized aggregates: a light weight index structure for data warehousing. In Gupta et al. [1906], pages 476– 487. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Moerkotte98.html. Also known as

VLDB.

Shukla:1998:MVS

[1323] Amit Shukla, Prasad Deshpande, and Jeffrey F. Naughton. Materialized view selection for multidimensional datasets. In Gupta et al. [1906], pages 488– 499. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ShuklaDN98.html. Also known as

VLDB.

Garcia-Molina:1998:EDW

[1324] Hector Garcia-Molina, Wilburt J. Labio, and Jun Yang. Expiring data in a warehouse. In Gupta et al. [1906], pages 500–511. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Garcia-MolinaLY98.

html. Also known as VLDB.

Zou:1998:SEU

[1325] Chendong Zou and Betty Salzberg. Safely and efficiently updating references during on-line reorganization. In Gupta et al. [1906], pages 512– 522. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

ZouS98.html. Also known as VLDB. Zhang:1998:BRA

[1326] Weiye Zhang and Per-Åke Larson. Buffering and read-ahead strategies for external mergesort. In Gupta et al. [1906], pages 523–533. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

                                        1998. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ZhangL98.html.

Also known as VLDB.

Amer-Yahia:1998:BLT

[1327] Sihem Amer-Yahia, Sophie Cluet, and Claude Delobel. Bulk-loading techniques for object databases and an application to relational data. In Gupta et al. [1906], pages 534– 545. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Amer-YahiaCD98.html. Also known as

VLDB.

Papadias:1998:AQS

[1328] Dimitris Papadias, Nikos Mamoulis, and Vasilis Delis. Algorithms for querying by spatial structure. In Gupta et al. [1906], pages 546–557. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/PapadiasMD98.

html. Also known as VLDB.

Zimbrao:1998:RAP

[1329] Geraldo Zimbrao and Jano Moreira de Souza. A raster approximation for processing of spatial joins. In Gupta et al. [1906], pages 558– 569. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ZimbraoS98.html. Also known as

VLDB.

Arge:1998:SSB

[1330] Lars Arge, Octavian Procopiuc, Sridhar Ramaswamy, Torsten Suel, and Jeffrey Scott Vitter. Scalable sweeping-based spatial join. In Gupta et al. [1906], pages 570– 581. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ArgePRSV98.html. Also known as

VLDB.

Korn:1998:RRN

[1331] Flip Korn, Alexandros Labrinidis, Yannis Kotidis, and Christos Faloutsos. Ratio rules: a new paradigm for fast, quantifiable data mining. In Gupta et al. [1906], pages 582–593. ISBN 1-55860-566-5. LCCN QA76.9.D3

                                        I559 1998. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/KornLKF98.

html. Also known as VLDB.

Silverstein:1998:STM

[1332] Craig Silverstein, Sergey Brin, Rajeev Motwani, and Jeffrey D. Ullman. Scalable techniques for mining causal structures. In Gupta et al. [1906], pages 594–605. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SilversteinBMU98.html. Also known

as VLDB.

Chakrabarti:1998:MSP

[1333] Soumen Chakrabarti, Sunita Sarawagi, and Byron Dom. Mining surprising patterns using temporal description length. In Gupta et al. [1906], pages 606–617. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChakrabartiSD98.html. Also known

as VLDB.

Slutz:1998:MST

[1334] Donald R. Slutz. Massive stochastic testing of SQL. In Gupta et al. [1906], pages 618–622. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Slutz98.html. Also

known as VLDB.

Lakshmi:1998:SEE

[1335] M. Seetha Lakshmi and Shaoyu Zhou.

Selectivity estimation in extensible databases — A neural network approach. In Gupta et al. [1906], pages 623–627. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LakshmiZ98.html. Also known as

VLDB.

Boncz:1998:DB

[1336] Peter A. Boncz, Tim Rühl, and Fred Kwakkel. The drill down benchmark. In Gupta et al. [1906], pages 628– 632. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BonczRK98.html. Also known as

VLDB.

Do:1998:IDV

[1337] Lyman Do, Pamela Drew, Wei Jin, Vish Jumani, and David Van Rossum. Issues in developing very large data warehouses. In Gupta et al. [1906], pages 633–636. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

URL http://www.vldb.org/dblp/
db/conf/vldb/DoDJJR98.html. Also

known as VLDB.

Sterling:1998:NMK

[1338] Warren Sterling. The National Medical Knowledge Bank. In Gupta et al. [1906], pages 637– 640. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sterling98.html. Also known as

VLDB.

Carino:1998:BAC

[1339] Felipe Cariño and Mark Jahnke. Bank of America case study: The information currency advantage. In Gupta et al. [1906], pages 641–644. ISBN 1-55860-566-5. LCCN QA76.9.D3

                                        I559 1998. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/CarinoJ98.

html. Also known as VLDB.

Dar:1998:DDD

[1340] Shaul Dar, Gadi Entin, Shai Geva, and Eran Palmon. DTL’s DataSpot: Database exploration using plain language. In Gupta et al. [1906], pages 645–649. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

DarEGP98.html. Also known as VLDB.

Vieille:1998:DIK

[1341] Laurent Vieille. From data independence to knowledge independence: An on-going story. In Gupta et al. [1906], pages 650–654. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Vieille98.html. Also

known as VLDB.

Fankhauser:1998:EFD

[1342] Peter Fankhauser, Georges Gardarin, M. Lopez, J. Muñoz, and Anthony Tomasic. Experiences in federated databases: From IRO-DB to MIRO-Web. In Gupta et al. [1906], pages 655–658. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        FankhauserGLMT98.html. Also known

as VLDB.

Bello:1998:MVO

[1343] Randall G. Bello, Karl Dias, Alan Downing, James Feenan, Jr., William D. Norcott, Harry Sun, Andrew Witkowski, and Mohamed Ziauddin. Materialized views in Oracle. In Gupta et al. [1906], pages 659– 664. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BelloDDFNSWZ98.html. Also known as

VLDB.

Joshi:1998:CO

[1344] Ashok Joshi, William Bridge, Juan Loaiza, and Tirthankar Lahiri. Checkpointing in Oracle. In Gupta et al. [1906], pages 665–668. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

URL http://www.vldb.org/dblp/
db/conf/vldb/JoshiBLL98.html. Also

known as VLDB.

Bamford:1998:AOP

[1345] Roger Bamford, D. Butler, Boris Klots, and N. MacNaughton. Architecture of Oracle Parallel Server. In Gupta et al. [1906], pages 669–670. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/BamfordBKM98.

html. Also known as VLDB.

Arun:1998:KAI

[1346] Gopalan Arun and Ashok Joshi.

KODA — the architecture and interface of a data model independent kernel. In Gupta et al. [1906], pages 671–674. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

ArunJ98.html. Also known as VLDB.

Schoning:1998:ABP

[1347] Harald Schöning. The ADABAS buffer pool manager. In Gupta et al. [1906], pages 675–679. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Schoning98.html. Also

known as VLDB.

Baulier:1998:DSR

[1348] Jerry Baulier, Stephen Blott, Henry F. Korth, and Abraham Silberschatz. A database system for real-time event aggregation in telecommunication. In Gupta et al. [1906], pages 680– 684. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BaulierBKS98.html. Also known as

VLDB.

Venkataraman:1998:HDQ

[1349] Shivakumar Venkataraman and Tian Zhang. Heterogeneous database query optimization in DB2 universal Data-Joiner. In Gupta et al. [1906], pages 685–689. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        VenkataramanZ98.html. Also known

as VLDB.

Carino:1998:PPT

[1350] Felipe Cariño and William O’Connell. Plan-per-tuple optimization solution — parallel execution of expensive user-defined functions. In Gupta et al. [1906], pages 690–695. ISBN 1-55860-566-5. LCCN QA76.9.D3

                                        I559 1998. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/CarinoO98.

html. Also known as VLDB.

Florescu:1998:WSM

[1351] Daniela Florescu, Alon Y. Levy, and Dan Suciu. Is Web-site management a database problem? In Gupta et al. [1906], page 696. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559

1998. URL http://www.vldb.org/
dblp/db/conf/vldb/FlorescuLS98.

html. Also known as VLDB.

Blott:1998:ICM

[1352] Stephen Blott, Henry F. Korth, and Abraham Silberschatz. Information, communication, and money: For what can we charge and how can we meter it? In Gupta et al. [1906], page 697. ISBN 1-55860-566-5. LCCN QA76.9.D3

I559 1998. URL http://www.vldb.
org/dblp/db/conf/vldb/BlottKS98.

html. Also known as VLDB.

Orenstein:1998:SSE

[1353] Jack A. Orenstein. Starting (and sometimes ending) a database company. In Gupta et al. [1906], page 698. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Orenstein98.html. Also known as

VLDB.

deFerreiraRezende:1998:BHR

[1354] Fernando de Ferreira Rezende and Günter Sauter. Bridging heterogeneity: Research and practice of database middleware technology. In Gupta et al. [1906], page 699. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RezendeS98.html. Also known as

VLDB.

Roussopoulos:1998:CSO

[1355] Nick Roussopoulos and Yannis Kotidis. The cubetree storage organization. In Gupta et al. [1906], page 700. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RoussopoulosK98.html. Also known

as VLDB.

Baulier:1998:DHP

[1356] Jerry Baulier, Philip Bohannon, S. Gogate, S. Joshi, C. Gupta, A. Khivesera, Henry F. Korth, P. McIlroy, J. Miller, P. P. S. Narayan, M. Nemeth, Rajeev Rastogi, Abraham Silberschatz, and S. Sudarshan. DataBlitz: a high performance main-memory storage manager. In Gupta et al. [1906], page 701. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BaulierBGJGKKMMNNRSS98.html. Also

known as VLDB.

George:1998:DPD

[1357] Felicity George. DMS: a parallel data mining server. In Gupta et al. [1906], page 702. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

George98.html. Also known as VLDB.

Schiefer:1998:IDU

[1358] K. Bernhard Schiefer, Jim Kleewein, Karen Brannon, Guy M. Lohman, and Gene Fuh. IBM’s DB2 universal database demonstrations at VLDB’98. In Gupta et al. [1906], page 703. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

SchieferKBLF98.html. Also known as

VLDB.

Mina:1998:MSS

[1359] Chebel Mina. MapInfo Spatial-Ware: a spatial information server for RDBMS. In Gupta et al. [1906], page 704. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Mina98.html. Also known as VLDB.

Objectivity:1998:OIE

[1360] Objectivity, Inc. Objectivity industrial exhibit. In Gupta et al. [1906], page 705. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Objectivity98.html. Also known as

VLDB.

Pogue:1998:OIE

[1361] Amy Pogue. Oracle industrial exhibit. In Gupta et al. [1906], page 706. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Pogue98.html. Also known as VLDB.

Mihalisin:1998:SPC

[1362] Ted Mihalisin. A single pass computing engine for interactive analysis of VLDBs. In Gupta et al. [1906], page 707. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Mihalisin98.html. Also known as

VLDB.

Mohan:1999:RHB

[1363] C. Mohan. Repeating history beyond ARIES. In Atkinson et al. [1907], pages 1–17. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/Mohan99.html. Also

known as VLDB’99.

Tan:1999:OFN

[1364] Kian-Lee Tan, Cheng Hian Goh, and Beng Chin Ooi. Online feedback for nested aggregate queries with multithreading. In Atkinson et al. [1907], pages 18–29. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/TanGO99.html. Also

known as VLDB’99.

Kemper:1999:GHT

[1365] Alfons Kemper, Donald Kossmann, and Christian Wiesner. Generalised hash teams for join and group-by. In Atkinson et al. [1907], pages 30– 41. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KemperKW99.html. Also known as

VLDB’99.

Sarawagi:1999:EDM

[1366] Sunita Sarawagi. Explaining differences in multidimensional aggregates. In Atkinson et al. [1907], pages 42– 53. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Sarawagi99.html. Also known as

VLDB’99.

Boncz:1999:DAO

[1367] Peter A. Boncz, Stefan Manegold, and Martin L. Kersten. Database architecture optimized for the new bottleneck: Memory access. In Atkinson et al. [1907], pages 54–65. ISBN 1-55860-615-7. LCCN QA76.9.D3

I559 1999. URL http://www.vldb.
org/dblp/db/conf/vldb/BonczMK99.

html. Also known as VLDB’99.

Norvaag:1999:PCI

[1368] Kjetil Nørvåg. The persistent cache:

Improving OID indexing in temporal object-oriented database systems. In Atkinson et al. [1907], pages 66– 77. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Norvag99.html.  Also known as

VLDB’99.

Rao:1999:CCI

[1369] Jun Rao and Kenneth A. Ross. Cache conscious indexing for decision-support in main memory. In Atkinson et al. [1907], pages 78–89. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/RaoR99.html.  Also

known as VLDB’99.

Chawathe:1999:CHD

[1370] Sudarshan S. Chawathe. Comparing hierarchical data in external memory. In Atkinson et al. [1907], pages 90– 101. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Chawathe99.html. Also known as

VLDB’99.

Jagadish:1999:MDT

[1371] H. V. Jagadish, Nick Koudas, and S. Muthukrishnan. Mining deviants in a time series database. In Atkinson et al. [1907], pages 102– 113. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

                                        KoudasMJ99.html. Also known as

VLDB’99.

Pitoura:1999:EVH

[1372] Evaggelia Pitoura and Panos K. Chrysanthis. Exploiting versions for handling updates in broadcast disks. In Atkinson et al. [1907], pages 114– 125. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        PitouraC99.html. Also known as

VLDB’99.

Pacitti:1999:FAM

[1373] Esther Pacitti, Pascale Minet, and Eric Simon. Fast algorithms for maintaining replica consistency in lazy master replicated databases. In Atkinson et al. [1907], pages 126– 137. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        PacittMS99.html. Also known as

VLDB’99.

Abiteboul:1999:AVE

[1374] Serge Abiteboul, Bernd Amann, Sophie Cluet, Adi Eyal, Laurent Mignet, and Tova Milo. Active views for electronic commerce. In Atkinson et al. [1907], pages 138– 149. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AbiteboulACEMM99.html. Also known

as VLDB’99.

Voruganti:1999:AHS

[1375] Kaladhar Voruganti, M. Tamer Özsu, and Ronald C. Unrau. An adaptive hybrid server architecture for client caching ODBMSs. In Atkinson et al. [1907], pages 150–161. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/VorugantiOU99.

html. Also known as VLDB’99.

Tamura:1999:DLB

[1376] Masahisa Tamura and Masaru Kitsuregawa. Dynamic load balancing for parallel association rule mining on heterogeneous PC cluster systems. In Atkinson et al. [1907], pages 162–173. ISBN 1-55860-615-7. LCCN QA76.9.D3

I559 1999. URL http://www.vldb.
org/dblp/db/conf/vldb/TamuraK99.

html. Also known as VLDB’99.

Ioannidis:1999:HBA

[1377] Yannis E. Ioannidis and Viswanath Poosala. Histogram-based approximation of set-valued query-answers. In Atkinson et al. [1907], pages 174– 185. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
IoannidisP99.html. Also known as

VLDB’99.

Jagadish:1999:SCP

[1378] H. V. Jagadish, J. Madar, and Raymond T. Ng. Semantic compression and pattern extraction with fascicles. In Atkinson et al. [1907], pages 186– 198. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JagadishMN99.html. Also known as

VLDB’99.

Brodie:1999:INM

[1379] Michael L. Brodie and Surajit Chaudhuri. Issues in network management in the next millennium. In Atkinson et al. [1907], page 198. ISBN 1-55860-615-7. LCCN QA76.9.D3

                                        I559 1999. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/BrodieC99.

html. Also known as VLDB’99.

Bamford:1999:SHA

[1380] Roger Bamford, Rafiul Ahad, and Angelo Pruscino. A scalable and highly available networked database architecture. In Atkinson et al. [1907], pages 199–201. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        BamfordAP99.html. Also known as

VLDB’99.

Hamilton:1999:NDM

[1381] James Hamilton. Networked data management design points. In Atkinson et al. [1907], pages 202– 206. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        Hamilton99.html. Also known as

VLDB’99.

Pound:1999:CSN

[1382] Chris Pound. In cyber space no one can hear you scream. In Atkinson et al. [1907], pages 207–210. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Pound99.html. Also

known as VLDB’99.

Knorr:1999:FIK

[1383] Edwin M. Knorr and Raymond T. Ng. Finding intensional knowledge of distance-based outliers. In Atkinson et al. [1907], pages 211–222. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/KnorrN99.html.

Also known as VLDB’99.

Garofalakis:1999:SSP

[1384] Minos N. Garofalakis, Rajeev Rastogi, and Kyuseok Shim. SPIRIT: Sequential pattern mining with regular expression constraints. In Atkinson et al. [1907], pages 223– 234. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GarofalakisRS99.html. Also known

as VLDB’99.

Jermaine:1999:NIS

[1385] Chris Jermaine, Anindya Datta, and Edward Omiecinski. A novel index supporting high volume data warehouse insertion. In Atkinson et al. [1907], pages 235–246. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JermaineDO99.html. Also known as

VLDB’99.

Blum:1999:MEQ

[1386] Adam Blum. Microsoft English Query 7.5: Automatic extraction of semantics from relational databases and OLAP cubes. In Atkinson et al. [1907], pages 247–248. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/Blum99.html.  Also

known as VLDB’99.

Campbell:1999:NLL

[1387] David Campbell. The new locking, logging, and recovery architecture of Microsoft SQL Server 7.0. In Atkinson et al. [1907], page 249. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

Campbell99.html. Also known as

VLDB’99.

Graefe:1999:VMJ

[1388] Goetz Graefe. The value of merge-join and hash-join in SQL server. In Atkinson et al. [1907], pages 250–253. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/Graefe99.html.

Also known as VLDB’99.

Darmont:1999:VGD

[1389] Jérôme Darmont and Michel Schneider. VOODB: a generic discrete-event random simulation model to evaluate the performances of OODBs. In Atkinson et al. [1907], pages 254– 265. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DarmontS99.html. Also known as

VLDB’99.

Ailamaki:1999:DMP

[1390] Anastassia Ailamaki, David J. De-Witt, Mark D. Hill, and David A. Wood. DBMSs on a modern processor: Where does time go? In Atkinson et al. [1907], pages 266–277. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/AilamakiDHW99.

html. Also known as VLDB’99.

Johnson:1999:PMC

[1391] Theodore Johnson. Performance measurements of compressed bitmap indices. In Atkinson et al. [1907], pages 278–289. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Johnson99.html. Also known as

VLDB’99.

Dyreson:1999:CQM

[1392] Curtis E. Dyreson, Michael H. Böhlen, and Christian S. Jensen. Capturing and querying multiple aspects of semistructured data. In Atkinson et al. [1907], pages 290– 301. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DyresonBJ99.html. Also known as

VLDB’99.

Shanmugasundaram:1999:RDQ

[1393] Jayavel Shanmugasundaram, Kristin Tufte, Chun Zhang, Gang He, David J. DeWitt, and Jeffrey F. Naughton. Relational databases for querying XML documents: Limitations and opportunities. In Atkinson et al. [1907], pages 302–314. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShanmugasundaramGTZDN99.html. Also

known as VLDB’99.

McHugh:1999:QOX

[1394] Jason McHugh and Jennifer Widom.

Query optimization for XML. In Atkinson et al. [1907], pages 315–326. ISBN 1-55860-615-7. LCCN QA76.9.D3

I559 1999. URL http://www.vldb.
org/dblp/db/conf/vldb/McHughW99.

html. Also known as VLDB’99.

Bernstein:1999:CBP

[1395] Philip A. Bernstein, Shankar Pal, and David Shutt. Context-based prefetch for implementing objects on relations. In Atkinson et al. [1907], pages 327– 338. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

                                        BernsteinPS99.html. Also known as

VLDB’99.

Goh:1999:GFG

[1396] Cheng Hian Goh, Beng Chin Ooi, D. Sim, and Kian-Lee Tan. GHOST: Fine granularity buffering of indexes. In Atkinson et al. [1907], pages 339– 350. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GohOST99.html.  Also known as

VLDB’99.

Haas:1999:LCQ

[1397] Laura M. Haas, Donald Kossmann, and Ioana Ursu. Loading a cache with query results. In Atkinson et al. [1907], pages 351–362. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/HaasKU99.html.

Also known as VLDB’99.

Wang:1999:BHC

[1398] Ke Wang, Senqiang Zhou, and Shiang Chen Liew. Building hierarchical classifiers using class proximity. In Atkinson et al. [1907], pages 363– 374. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        WangZL99.html.  Also known as

VLDB’99.

Chakrabarti:1999:DHR

[1399] Soumen Chakrabarti, Martin van den Berg, and Byron Dom. Distributed hypertext resource discovery through examples. In Atkinson et al. [1907], pages 375–386. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/

ChakrabartiBD99.html. Also known

as VLDB’99.

Jagadish:1999:MDS

[1400] H. V. Jagadish, Olga Kapitskaia, Raymond T. Ng, and Divesh Srivastava. Multi-dimensional substring selectivity estimation. In Atkinson et al. [1907], pages 387–398. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/JagadishKNS99.

html. Also known as VLDB’99.

Chaudhuri:1999:ETK

[1401] Surajit Chaudhuri and Luis Gravano.

Evaluating top- k selection queries. In Atkinson et al. [1907], pages 397– 410. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChaudhuriG99.html. Also known as

VLDB’99.

Chaudhuri:1999:ETS

[1402] S. Chaudhuri and L. Gravano. Evaluating top-K selection queries. In Atkinson et al. [1907], pages 399–410. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. Also known as VLDB’99.

Donjerkovic:1999:POT

[1403] Donko Donjerkovic and Raghu Ramakrishnan. Probabilistic optimization of top N queries. In Atkinson et al. [1907], pages 411– 422. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DonjerkovicR99.html. Also known as

VLDB’99.

Konig:1999:CHP

[1404] Arnd Christian König and Gerhard Weikum. Combining histograms and parametric curve fitting for feedbackdriven query result-size estimation. In Atkinson et al. [1907], pages 423– 434. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KonigW99.html.  Also known as

VLDB’99.

Josifovski:1999:IHO

[1405] Vanja Josifovski and Tore Risch. Integrating heterogeneous overlapping databases through object-oriented transformations. In Atkinson et al. [1907], pages 435–446. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/JosifovskiR99.

html. Also known as VLDB’99.

Naumann:1999:QDI

[1406] Felix Naumann, Ulf Leser, and Johann Christoph Freytag. Qualitydriven integration of heterogeneous information systems. In Atkinson et al. [1907], pages 447–458. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/NaumannLF99.html. Also

known as VLDB’99.

Deutsch:1999:PDI

[1407] Alin Deutsch, Lucian Popa, and Val Tannen. Physical data independence, constraints, and optimization with universal plans. In Atkinson et al. [1907], pages 459– 470. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
DeutschPT99.html. Also known as

VLDB’99.

Lakshmanan:1999:EIS

[1408] Laks V. S. Lakshmanan, Fereidoon Sadri, and Subbu N. Subramanian. On efficiently implementing SchemaSQL on a SQL database system. In Atkinson et al. [1907], pages 471– 482. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LakshmananSS99.html. Also known as

VLDB’99.

Lee:1999:UCD

[1409] Sin Yeung Lee and Tok Wang Ling.

Unrolling cycles to decide trigger termination. In Atkinson et al. [1907], pages 483–493. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/LeeL99.html.  Also

known as VLDB’99.

Jaedicke:1999:UDT

[1410] Michael Jaedicke and Bernhard Mitschang.

User-defined table operators: Enhancing extensibility for ORDBMS. In Atkinson et al. [1907], pages 494– 505. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JaedickeM99.html. Also known as

VLDB’99.

Hinneburg:1999:OGC

[1411] Alexander Hinneburg and Daniel A.

Keim. Optimal grid-clustering: Towards breaking the curse of dimensionality in high-dimensional clustering. In Atkinson et al. [1907], pages 506–517. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/

KeimH99.html.  Also known as

VLDB’99.

Gionis:1999:SSH

[1412] Aristides Gionis, Piotr Indyk, and Rajeev Motwani. Similarity search in high dimensions via hashing. In Atkinson et al. [1907], pages 518– 529. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
GionisIM99.html. Also known as

VLDB’99.

Jagadish:1999:WCH

[1413] H. V. Jagadish, Laks V. S. Lakshmanan, and Divesh Srivastava. What can hierarchies do for data warehouses? In Atkinson et al. [1907], pages 530– 541. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JagadishLS99.html. Also known as

VLDB’99.

Carey:1999:WTD

[1414] Michael J. Carey, Donald D. Chamberlin, Srinivasa Narayanan, Bennet Vance, Doug Doole, Serge Rielau, Richard Swagerman, and Nelson Mendonça Mattos. O-O, what have they done to DB2? In Atkinson et al. [1907], pages 542– 553. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CareyCNVDRSM99.html. Also known as

VLDB’99.

Chen:1999:HLI

[1415] Weidong Chen, Jyh-Herng Chow, You-Chin Fuh, Jean Grandbois, Michelle Jou, Nelson Mendonça Mattos, Brian T. Tran, and Yun Wang.

High level indexing of user-defined types. In Atkinson et al. [1907], pages 554–564. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChenCFGJMTW99.html. Also known as

VLDB’99.

Fuh:1999:ISS

[1416] You-Chin Fuh, Stefan Deßloch, Weidong Chen, Nelson Mendonça Mattos, Brian T. Tran, Bruce G. Lindsay, Linda DeMichel, Serge Rielau, and Danko Mannhaupt. Implementation of SQL3 structured types with inheritance and value substitutability. In Atkinson et al. [1907], pages 565–574. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/FuhDCMTLDRM99.

html. Also known as VLDB’99.

Nink:1999:GCL

[1417] Udo Nink, Theo Härder, and Norbert Ritter. Generating call-level interfaces for advanced database application programming. In Atkinson et al. [1907], pages 575–586. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/NinkHR99.html.

Also known as VLDB’99.

Hosking:1999:POP

[1418] Antony L. Hosking and Jiawan Chen.

PM3: An orthogonal persistent systems programming language — design, implementation, performance. In Atkinson et al. [1907], pages 587– 598. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HoskingC99.html. Also known as

VLDB’99.

Roth:1999:CMD

[1419] Mary Tork Roth, Fatma Ozcan, and Laura M. Haas. Cost models DO matter: Providing cost information for diverse data sources in a federated system. In Atkinson et al. [1907], pages 599–610. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/RothOH99.html.

Also known as VLDB’99.

Carino:1999:ASH

[1420] Felipe Cariño, William O’Connell, John Burgess, and Joel H. Saltz. Active storage hierarchy, database systems and applications — socratic exegesis. In Atkinson et al. [1907], pages 611–614. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        CarinoOBS99.html. Also known as

VLDB’99.

Ceri:1999:DDO

[1421] Stefano Ceri, Piero Fraternali, and Stefano Paraboschi. Data-driven, oneto-one Web site generation for dataintensive applications. In Atkinson et al. [1907], pages 615–626. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/CeriFP99.html.

Also known as VLDB’99.

Florescu:1999:ORT

[1422] Daniela Florescu, Alon Y. Levy, Dan Suciu, and Khaled Yagoub. Optimization of run-time management of data intensive Web-sites. In Atkinson et al. [1907], pages 627–638. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/

dblp/db/conf/vldb/FlorescuLSY99.

html. Also known as VLDB’99.

Kumar:1999:ELS

[1423] S. Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. Extracting large-scale knowledge bases from the Web. In Atkinson et al. [1907], pages 639– 650. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KumarRRT99.html. Also known as

VLDB’99.

Li:1999:AAV

[1424] Jianzhong Li, Doron Rotem, and Jaideep Srivastava. Aggregation algorithms for very large compressed data warehouses. In Atkinson et al. [1907], pages 651–662. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

URL http://www.vldb.org/dblp/
db/conf/vldb/LiRS99.html.  Also

known as VLDB’99.

Pedersen:1999:EPP

[1425] Torben Bach Pedersen, Christian S.

Jensen, and Curtis E. Dyreson. Extending practical pre-aggregation in on-line analytical processing. In Atkinson et al. [1907], pages 663–674. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/PedersenJD99.

html. Also known as VLDB’99.

Chan:1999:HPC

[1426] Chee Yong Chan and Yannis E. Ioannidis. Hierarchical prefix cubes for range-sum queries. In Atkinson et al. [1907], pages 675–686. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/ChanI99.html. Also

known as VLDB’99.

Cheng:1999:ITS

[1427] Qi Cheng, Jarek Gryz, Fred Koo, T. Y. Cliff Leung, Linqi Liu, Xiaoyan Qian, and Berni Schiefer. Implementation of two semantic query optimization techniques in DB2 universal database. In Atkinson et al. [1907], pages 687–698. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ChengGKLLQS99.

html. Also known as VLDB’99.

Kornacker:1999:HPE

[1428] Marcel Kornacker. High-performance extensible indexing. In Atkinson et al. [1907], pages 699–708. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999.

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Kornacker99.html. Also

known as VLDB’99.

Raman:1999:ODR

[1429] Vijayshankar Raman, Bhaskaran Raman, and Joseph M. Hellerstein. Online dynamic reordering for interactive data processing. In Atkinson et al. [1907], pages 709–720. ISBN 1-55860-615-7. LCCN QA76.9.D3

                                        I559 1999. URL http://www.vldb.
                                        org/dblp/db/conf/vldb/RamanRH99.

html. Also known as VLDB’99.

DeRose:1999:WDW

[1430] Steven J. DeRose. What do those weird XML types want, anyway? In Atkinson et al. [1907], pages 721–724. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

                                        1999. URL http://www.vldb.org/

dblp/db/conf/vldb/DeRose99.html.

Also known as VLDB’99.

Dayal:1999:IPD

[1431] Umeshwar Dayal. Industrial panel on data warehousing technologies: Experiences, challenges, and directions. In Atkinson et al. [1907], page 725. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/Dayal99.html.

Also known as VLDB’99.

Zurek:1999:DMC

[1432] Thomas Zurek and Markus Sinnwell.

Datawarehousing has more colours than just black & white. In Atkinson et al. [1907], pages 726–729. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/ZurekS99.html.

Also known as VLDB’99.

Datta:1999:CNS

[1433] Anindya Datta, Krithi Ramamritham, and Helen M. Thomas. Curio: a novel solution for efficient storage and indexing in data warehouses. In Atkinson et al. [1907], pages 730–733. ISBN 1-55860-615-7. LCCN QA76.9.D3

I559 1999. URL http://www.vldb.
org/dblp/db/conf/vldb/DattaRT99.

html. Also known as VLDB’99.

Zirintsis:1999:HPJ

[1434] Evangelos Zirintsis, Graham N. C.

Kirby, and Ronald Morrison. Hyperprogramming in Java. In Atkinson et al. [1907], pages 734–737. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/ZirintsisKM99.

html. Also known as VLDB’99.

Sahuguet:1999:BLW

[1435] Arnaud Sahuguet and Fabien Azavant. Building light-weight wrappers for legacy Web data-sources using W4F. In Atkinson et al. [1907], pages 738–741. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SahuguetA99.html. Also known as

VLDB’99.

Abiteboul:1999:XRA

[1436] Serge Abiteboul, Vincent Aguilera, Sébastien Ailleret, Bernd Amann, Sophie Cluet, Brendan Hills, Frédéric Hubert, Jean-Claude Mamou, Amélie Marian, Laurent Mignet, Tova Milo, Cassio Souza dos Santos, Bruno Tessier, and Anne-Marie Vercoustre. XML repository and active views demonstration. In Atkinson et al. [1907], pages 742– 745. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AbiteboulAAACHHMMMMSTV99.html. Also

known as VLDB’99.

Baumann:1999:STR

[1437] Peter Baumann, Andreas Dehmel, Paula Furtado, Roland Ritsch, and Norbert Widmann. Spatio-temporal retrieval with RasDaMan. In Atkinson et al. [1907], pages 746–749. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559

1999. URL http://www.vldb.org/
dblp/db/conf/vldb/BaumannDFRW99.

html. Also known as VLDB’99.

Bouganim:1999:MWI

[1438] Luc Bouganim, Tatiana Chan-Sine-Ying, Tuyet-Tram Dang-Ngoc, Jean-Luc Darroux, Georges Gardarin, and Fei Sha. Miro Web: Integrating multiple data sources through semistructured data types. In Atkinson et al. [1907], pages 750– 753. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BouganimCDDGS99.html. Also known

as VLDB’99.

Acharya:1999:AFD

[1439] Swarup Acharya, Phillip B. Gibbons, and Viswanath Poosala. Aqua: a fast decision support systems using approximate query answers. In Atkinson et al. [1907], pages 754– 757. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AcharyaGP99.html. Also known as

VLDB’99.

deVries:1999:MMA

[1440] Arjen P. de Vries, Mark G. L. M.

van Doorn, Henk M. Blanken, and Peter M. G. Apers. The mirror MMDBMS architecture. In Atkinson et al. [1907], pages 758– 761. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:/

/www.vldb.org/dblp/db/conf/vldb/
VriesDBA99.html. Also known as

VLDB’99.

Chaudhuri:2000:RDS

[1441] Surajit Chaudhuri and Gerhard Weikum. Rethinking database system architecture: Towards a self-tuning RISC-style database system. In El Abbadi et al. [1908], pages 1–10. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChaudhuriW00.html.

Bobineau:2000:PSD

[1442] Christophe Bobineau, Luc Bouganim, Philippe Pucheral, and Patrick Valduriez. PicoDMBS: Scaling down database techniques for the Smartcard. In El Abbadi et al. [1908], pages 11–20. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChristopheBPV00.html.

Johnson:2000:MAU

[1443] Theodore Johnson, Laks V. S. Lakshmanan, and Raymond T. Ng. The 3W model and algebra for unified data mining. In El Abbadi et al. [1908], pages 21–32. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        LakshmananJN00.html.

Choenni:2000:DIG

[1444] Sunil Choenni. Design and implementation of a genetic-based algorithm for data mining. In El Abbadi et al. [1908], pages 33–42. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChoenniV00.html.

Wang:2000:MFI

[1445] Ke Wang, Yu He, and Jiawei Han. Mining frequent itemsets using support constraints. In El Abbadi et al. [1908], pages 43–52. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        WangHH00.html.

Altinel:2000:EFX

[1446] Mehmet Altinel and Michael J. Franklin. Efficient filtering of XML documents for selective dissemination of information. In El Abbadi et al. [1908], pages 53–64. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
AltinelF00.html.

Shanmugasundaram:2000:EPR

[1447] Jayavel Shanmugasundaram, Eugene J.

Shekita, Rimon Barr, Michael J. Carey, Bruce G. Lindsay, Hamid Pirahesh, and Berthold Reinwald. Efficiently publishing relational data as XML documents. In El Abbadi et al. [1908], pages 65–76. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ShanmugasundaramSBCLPR00.html.

Miller:2000:SMQ

[1448] Renée J. Miller, Laura M. Haas, and Mauricio A. Hernández. Schema mapping as query discovery. In El Abbadi et al. [1908], pages 77–88. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MillerHH00.html.

Chakrabarti:2000:LDR

[1449] Kaushik Chakrabarti and Sharad Mehrotra. Local dimensionality reduction: a new approach to indexing high dimensional spaces. In El Abbadi et al. [1908], pages 89–100. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChakrabartiM00.html.

Matias:2000:DMW

[1450] Yossi Matias, Jeffrey Scott Vitter, and Min Wang. Dynamic maintenance of wavelet-based histograms. In El Abbadi et al. [1908], pages 101– 110. ISBN 1-55860-715-3. LCCN ????

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/MatiasVW00.html.

Chakrabarti:2000:AQP

[1451] Kaushik Chakrabarti, Minos N. Garofalakis, Rajeev Rastogi, and Kyuseok Shim. Approximate query processing using wavelets. In El Abbadi et al. [1908], pages 111–122. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ChakrabartiGRS00.html.

Yu:2000:ENE

[1452] Haifeng Yu and Amin Vahdat. Efficient numerical error bounding for replicated network services. In El Abbadi et al. [1908], pages 123– 133. ISBN 1-55860-715-3. LCCN

                                        ???? URL http://www.vldb.org/
                                        dblp/db/conf/vldb/YuV00.html.

Kemme:2000:DLC

[1453] Bettina Kemme and Gustavo Alonso. Don’t be lazy, be consistent: Postgres-R, A new way to implement database replication. In El Abbadi et al. [1908], pages 134–143. ISBN 1-55860-715-3.

                                        LCCN ???? URL http://www.vldb.
                                        org/dblp/db/conf/vldb/KemmeA00.
                                        html.

Olston:2000:OPP

[1454] Chris Olston and Jennifer Widom. Offering a precision-performance tradeoff for aggregation queries over replicated data. In El Abbadi et al. [1908], pages 144–155. ISBN 1-55860-715-3.

                                        LCCN ???? URL http://www.vldb.
                                        org/dblp/db/conf/vldb/OlstonW00.
                                        html.

Grumbach:2000:MID

[1455] Stéphane Grumbach, Philippe Rigaux, and Luc Segoufin. Manipulating interpolated data is easier than you thought. In El Abbadi et al. [1908], pages 156– 165. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/GrumbachRS00.
html.

Wang:2000:USB

[1456] Haixun Wang and Carlo Zaniolo. Using SQL to build new aggregates and extenders for object- relational systems. In El Abbadi et al. [1908], pages 166– 175. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/WangZ00.html.

Ganti:2000:IST

[1457] Venkatesh Ganti, Mong-Li Lee, and Raghu Ramakrishnan. ICICLES: Selftuning samples for approximate query answering. In El Abbadi et al. [1908], pages 176–187. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/GantiLR00.
html.

Yagoub:2000:CSD

[1458] Khaled Yagoub, Daniela Florescu, Valérie Issarny, and Patrick Valduriez. Caching strategies for data-intensive Web sites. In El Abbadi et al. [1908], pages 188–199. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
YagoubFIV00.html.

Cho:2000:EWI

[1459] Junghoo Cho and Hector Garcia-Molina. The evolution of the Web and implications for an incremental crawler.

In El Abbadi et al. [1908], pages 200– 209. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/ChoG00.html.

Pardon:2000:CLT

[1460] Guy Pardon and Gustavo Alonso. CheeTah: a lightweight transaction server for plug-and-play Internet data management. In El Abbadi et al. [1908], pages 210–219. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PardonA00.html.

Balmin:2000:HQO

[1461] Andrey Balmin, Thanos Papadimitriou, and Yannis Papakonstantinou. Hypothetical queries in an OLAP environment. In El Abbadi et al. [1908], pages 220–231. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BalminPP00.html.

Lee:2000:HCC

[1462] Sin Yeung Lee, Tok Wang Ling, and Hua-Gang Li. Hierarchical compact cube for range-max queries. In El Abbadi et al. [1908], pages 232– 241. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/LingLL00.html.

Mendelzon:2000:TQO

[1463] Alberto O. Mendelzon and Alejandro A. Vaisman. Temporal queries in OLAP. In El Abbadi et al. [1908], pages 242–253. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/MendelzonV00.
html.

Ceri:2000:PAT

[1464] Stefano Ceri, Roberta Cochrane, and Jennifer Widom. Practical applications of triggers and constraints: Success and lingering issues (10-year award). In El Abbadi et al. [1908], pages 254– 262. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/CeriCW00.html.

Ramsak:2000:IUT

[1465] Frank Ramsak, Volker Markl, Robert Fenk, Martin Zirkel, Klaus Elhardt, and Rudolf Bayer. Integrating the UB-tree into a database system kernel. In El Abbadi et al. [1908], pages 263–272. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/RamsakMFZEB00.
html.

Stohr:2000:MDD

[1466] Thomas Stöhr, Holger Märtens, and Erhard Rahm. Multi-dimensional database allocation for parallel data warehouses. In El Abbadi et al. [1908], pages 273–284. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/StohrMR00.
html.

Srinivasan:2000:OIO

[1467] Jagannathan Srinivasan, Souripriya Das, Chuck Freiwald, Eugene Chong, Mahesh Jagannath, Aravind Yalamanchi, Ramkumar Krishnan, Anh-Tuan Tran, Samuel DeFazio, and Jayanta Banerjee. Oracle8i indexorganized table and its application to new domains. In El Abbadi et al. [1908], pages 285–296. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SrinivasanDFCJYKTDB00.html.

Wu:2000:FFA

[1468] Leejay Wu, Christos Faloutsos, Katia P. Sycara, and Terry R. Payne. FALCON: Feedback adaptive loop for content-based retrieval. In El Abbadi et al. [1908], pages 297– 306. ISBN 1-55860-715-3. LCCN

                                        ???? URL http://www.vldb.org/
                                        dblp/db/conf/vldb/WuFSP00.html.

Sarawagi:2000:UAE

[1469] Sunita Sarawagi. User-adaptive exploration of multidimensional data. In El Abbadi et al. [1908], pages 307– 316. ISBN 1-55860-715-3. LCCN ????

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/Sarawagi00.html.

Diao:2000:TLB

[1470] Yanlei Diao, Hongjun Lu, Songting Chen, and Zengping Tian. Toward learning based Web query processing. In El Abbadi et al. [1908], pages 317–328. ISBN 1-55860-715-3.

                                        LCCN ???? URL http://www.vldb.
                                        org/dblp/db/conf/vldb/DiaoLCT00.
                                        html.

Amer-Yahia:2000:OQC

[1471] Sihem Amer-Yahia and Theodore Johnson. Optimizing queries on compressed bitmaps. In El Abbadi et al. [1908], pages 329–338. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        JohnsonA00.html.

Manegold:2000:WHD

[1472] Stefan Manegold, Peter A. Boncz, and Martin L. Kersten. What happens during a join? Dissecting CPU and memory optimization effects. In El Abbadi et al. [1908], pages 339–350. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ManegoldBK00.html.

Ramasamy:2000:SCJ

[1473] Karthikeyan Ramasamy, Jignesh M.

Patel, Jeffrey F. Naughton, and Raghav Kaushik. Set containment joins: The good, the bad and the ugly. In El Abbadi et al. [1908], pages 351–362. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/RamasamyPNK00.
html.

Indyk:2000:IRT

[1474] Piotr Indyk, Nick Koudas, and S. Muthukrishnan. Identifying representative trends in massive time series data sets using sketches. In El Abbadi et al. [1908], pages 363–372. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
KoudasIM00.html.

Lu:2000:DTS

[1475] Hongjun Lu and Hongyan Liu. Decision tables: Scalable classification exploring RDBMS capabilities. In El Abbadi et al. [1908], pages 373– 384. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/LuL00.html.

Yi:2000:FTS

[1476] Byoung-Kee Yi and Christos Faloutsos. Fast time sequence indexing for arbitrary L_{p} norms. In El Abbadi et al. [1908], pages 385– 394. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/YiF00.html.

Pfoser:2000:NAQ

[1477] Dieter Pfoser, Christian S. Jensen, and Yannis Theodoridis. Novel approaches in query processing for moving object trajectories. In El Abbadi et al. [1908], pages 395–406. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        PfoserJT00.html.

Kriegel:2000:MIE

[1478] Hans-Peter Kriegel, Marco Pötke, and Thomas Seidl. Managing intervals efficiently in object-relational databases. In El Abbadi et al. [1908], pages 407–418. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        KriegelPS00.html.

Guntzer:2000:OMF

[1479] Ulrich Güntzer, Wolf-Tilo Balke, and Werner Kießling. Optimizing multifeature queries for image databases. In El Abbadi et al. [1908], pages 419– 428. ISBN 1-55860-715-3. LCCN ????

                                        URL http://www.vldb.org/dblp/
                                        db/conf/vldb/GuntzerBK00.html.

Goldstein:2000:CPP

[1480] Jonathan Goldstein and Raghu Ramakrishnan. Contrast plots and Psphere trees: Space vs. time in nearest neighbour searches. In El Abbadi et al. [1908], pages 429–440. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        GoldsteinR00.html.

Cowley:2000:TIC

[1481] Wes Cowley and Dimitris Plexousakis. Temporal integrity constraints with indeterminacy. In El Abbadi et al. [1908], pages 441–450. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/CowleyP00.
html.

Jiang:2000:BTB

[1482] Linan Jiang, Betty Salzberg, David B.

Lomet, and Manuel Barrena. The BTtree: a branched and temporal access method. In El Abbadi et al. [1908], pages 451–460. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
JiangSLB00.html.

Labio:2000:PII

[1483] Wilburt Labio, Jun Yang, Yingwei Cui, Hector Garcia-Molina, and Jennifer Widom. Performance issues in incremental warehouse maintenance. In El Abbadi et al. [1908], pages 461–472. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LabioYCGW00.html.

Jarke:2000:CPD

[1484] Matthias Jarke, Thomas List, and Jörg Köller. The challenge of process data warehousing. In El Abbadi et al. [1908], pages 473–483. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/JarkeKL00.
html.

Pottinger:2000:SAA

[1485] Rachel Pottinger and Alon Y. Levy.

A scalable algorithm for answering queries using views. In El Abbadi et al. [1908], pages 484–495. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PottingerL00.html.

Agrawal:2000:ASM

[1486] Sanjay Agrawal, Surajit Chaudhuri, and Vivek R. Narasayya. Automated selection of materialized views and indexes in SQL databases. In El Abbadi et al. [1908], pages 496–505. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        AgrawalCN00.html.

Hinneburg:2000:WNN

[1487] Alexander Hinneburg, Charu C. Aggarwal, and Daniel A. Keim. What is the nearest neighbor in high dimensional spaces? In El Abbadi et al. [1908], pages 506–515. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        HinneburgAK00.html.

Sakurai:2000:TIS

[1488] Yasushi Sakurai, Masatoshi Yoshikawa, Shunsuke Uemura, and Haruhiko Kojima. The A-tree: An index structure for high-dimensional spaces using relative approximation. In El Abbadi et al. [1908], pages 516–526. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        SakuraiYUK00.html.

Diligenti:2000:FCU

[1489] Michelangelo Diligenti, Frans Coetzee, Steve Lawrence, C. Lee Giles, and Marco Gori. Focused crawling using context graphs. In El Abbadi et al. [1908], pages 527–534. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DiligentiCLGG00.html.

Bar-Yossef:2000:AAQ

[1490] Ziv Bar-Yossef, Alexander Berg, Steve Chien, Jittat Fakcharoenphol, and Dror Weitz. Approximating aggregate queries about Web pages via random walks. In El Abbadi et al. [1908], pages 535–544. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bar-YossefBCFW00.html.

Ding:2000:CGS

[1491] Junyan Ding, Luis Gravano, and Narayanan Shivakumar. Computing geographical scopes of Web resources. In El Abbadi et al. [1908], pages 545– 556. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/DingGS00.html.

Panti:2000:CBA

[1492] Maurizio Panti, Luca Spalazzi, and Alberto Giretti. A case-based approach to information integration. In El Abbadi et al. [1908], pages 557–565. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PantiSG00.html.

Chang:2000:AQT

[1493] Chen-Chuan K. Chang and Hector Garcia-Molina. Approximate query translation across heterogeneous information sources. In El Abbadi et al. [1908], pages 566–577. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChangG00.html.

Pappas:2000:DDS

[1494] Nikos Pappas and Stavros Christodoulakis.

Design and development of a stream service in a heterogeneous client environment. In El Abbadi et al. [1908], pages 578–589. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/PappasC00.
html.

Rishe:2000:SAS

[1495] Naphtali Rishe, Jun Yuan, Rukshan Athauda, Shu-Ching Chen, Xiaoling Lu, Xiaobin Ma, Alexander Vaschillo, Artyom Shaposhnikov, and Dmitry Vasilevsky. Semantic access: Semantic interface for querying databases. In El Abbadi et al. [1908], pages 591–594. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RisheYACLMVSV00.html.

Pedersen:2000:TSR

[1496] Torben Bach Pedersen, Christian S. Jensen, and Curtis E. Dyreson. The TreeScape system: Reuse of precomputed aggregates over irregular OLAP hierarchies. In El Abbadi et al. [1908], pages 595–598. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
PedersenJD00.html.

Gu:2000:OPE

[1497] Junmin Gu, Torben Bach Pedersen, and Arie Shoshani. OLAP++: Powerful and easy-to-use federations of OLAP and object databases. In El Abbadi et al. [1908], pages 599– 602. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/GuPS00.html.

Chakrabarti:2000:MBA

[1498] Soumen Chakrabarti, Sandeep Srivastava, Mallela Subramanyam, and Mitul Tiwari. Memex: a browsing assistant for collaborative archiving and mining of surf trails. In El Abbadi et al. [1908], pages 603–606. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
ChakrabartiSST00.html.

Beneventano:2000:IIM

[1499] Domenico Beneventano, Sonia Bergamaschi, Silvana Castano, Alberto Corni, R. Guidetti, G. Malvezzi, Michele Melchiori, and Maurizio Vincini. Information integration: The MOMIS project demonstration. In El Abbadi et al. [1908], pages 611–614. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BeneventanoBCCGMMV00.html.

Caruso:2000:TDR

[1500] Francesco Caruso, Munir Cochinwala, Uma Ganapathy, Gail Lalk, and Paolo Missier. Telcordia’s database reconciliation and data quality analysis tool. In El Abbadi et al. [1908], pages 615– 618. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/CarusoCGLM00.
html.

Helal:2000:RDD

[1501] Abdelsalam Helal and Hua Li. Rainbow: Distributed database system for classroom education and experimental research. In El Abbadi et al. [1908], pages 619–622. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HelalL00.html.

Manolescu:2000:ALX

[1502] Ioana Manolescu, Daniela Florescu, Donald Kossmann, Florian Xhumari, and Don Olteanu. Agora: Living with XML and relational. In El Abbadi et al. [1908], pages 623–626. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        ManolescuFKXO00.html.

Pereira:2000:PSW

[1503] João Pereira, Françoise Fabret, François Llirbat, Radu Preotiuc-Pietro, Kenneth A. Ross, and Dennis Shasha. Publish/subscribe on the Web at extreme speed. In El Abbadi et al. [1908], pages 627–630. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        PereiraFLS00.html.

Dutta:2000:DES

[1504] Kaushik Dutta, Anindya Datta, Debra E. VanderMeer, Krithi Ramamritham, and Helen M. Thomas. Demonstration: Enabling scalable online personalization on the Web. In El Abbadi et al. [1908], pages 631–634. ISBN 1-55860-715-3. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/conf/vldb/
                                        DuttaDMRT00.html.

Shahabi:2000:ITI

[1505] Cyrus Shahabi, Adil Faisal, Farnoush Banaei Kashani, and Jabed Faruque. IN-SITE: a tool for interpreting users? Interaction with a Web space. In El Abbadi et al. [1908], pages 635– 638. ISBN 1-55860-715-3. LCCN

                                        ???? URL http://www.vldb.org/
                                        dblp/db/conf/vldb/ShahabiFKF00.
                                        html.

Ludascher:2000:MBI

[1506] Bertram Ludäscher, Amarnath Gupta, and Maryann E. Martone. Model-based information integration in a neuroscience mediator system. In El Abbadi et al. [1908], pages 639–642. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
LudasaherGM00.html.

Adi:2000:PTP

[1507] Asaf Adi, David Botzer, Opher Etzion, and Tali Yatzkar-Haham. Push technology personalization through event correlation. In El Abbadi et al. [1908], pages 643–645. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/AdiBEY00.
html.

Carey:2000:XMP

[1508] Michael J. Carey, Jerry Kiernan, Jayavel Shanmugasundaram, Eugene J. Shekita, and Subbu N. Subramanian. XPERANTO: Middleware for publishing object-relational data as XML documents. In El Abbadi et al. [1908], pages 646–648. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CareyKSSS00.html.

Basu:2000:BSI

[1509] Julie Basu, José Alberto Fernández, and Olga Peschansky. Building scalable Internet applications with Oracle8i server. In El Abbadi et al. [1908], page 649. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/BasuFP00.html.

Ramamritham:2000:SEG

[1510] Krithi Ramamritham, Yeha El Atfi, Carlo Batini, Michael Eitan, Valerie Gregg, and D. B. Phatak. Social, educational, and governmental change enabled through information technology. In El Abbadi et al. [1908], pages 651–654. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
RamamrithamABEGP00.html.

Deen:2000:VV

[1511] S. Misbah Deen, Anant Jhingran, Shamkant B. Navathe, Erich J. Neuhold, and Gio Wiederhold. A 20/20 vision of the VLDB-2020? In El Abbadi et al. [1908], pages 655– 659. ISBN 1-55860-715-3. LCCN ????

URL http://www.vldb.org/dblp/
db/conf/vldb/DeenJNNW00.html.

Bernstein:2000:PGM

[1512] Philip A. Bernstein, Laura M. Haas, Matthias Jarke, Erhard Rahm, and Gio Wiederhold. Panel: Is generic metadata management feasible? In El Abbadi et al. [1908], pages 660–662. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
BernsteinHJRW00.html.

Schek:2000:PFD

[1513] Hans-Jörg Schek. Panel: Future directions of database research — the VLDB broadening strategy, Part 1. In El Abbadi et al. [1908], pages 663– 664. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/Schek00.html.

Brodie:2000:PFD

[1514] Michael L. Brodie. Panel: Future directions of database research — the VLDB broadening strategy, Part 2. In El Abbadi et al. [1908], pages 665– 672. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/Brodie00.html.

Hvasshovd:2000:UHA

[1515] Svein-Olaf Hvasshovd, Svein Erik Bratsberg, and Øystein Torbjørnsen. An ultra highly available DBMS. In El Abbadi et al. [1908], page 673. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
HvasshovdBT00.html.

Liles:2000:ZLE

[1516] Dave Liles. The zero latency enterprise. In El Abbadi et al. [1908], pages 674–676. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/Ladin00.html.

Anonymous:2000:HPS

[1517] Anonymous. High-performance and scalability through application tier, in-memory data management. In El Abbadi et al. [1908], pages 677– 680. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/Team00.html.

Nori:2000:AEA

[1518] Anil Nori. Asera: Extranet architecture for B2B solutions. In El Abbadi et al. [1908], page 681. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Nori00.html.

Deshpande:2000:LBD

[1519] Anand Deshpande. Linking business to deliver value: a data management challenge. In El Abbadi et al. [1908], pages 682–683. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Deshpande00.html.

Mohan:2000:EGB

[1520] C. Mohan, Ron Barber, S. Watts, A. Somani, and Markos Zaharioudakis. Evolution of groupware for business applications: a database perspective on Lotus Domino/Notes. In El Abbadi et al. [1908], pages 684–687. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
MohanBWSZ00.html.

Casati:2000:PAF

[1521] Fabio Casati and Ming-Chien Shan. Process automation as the foundation for E-business. In El Abbadi et al. [1908], pages 688–691. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
CasatiS00.html.

Blecken:2000:MWI

[1522] Carsten Blecken. Media360 workflowimplementing a workflow engine inside a database. In El Abbadi et al. [1908], page 692. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/Blecken00.
html.

Bowker:2000:WIP

[1523] Geoffrey C. Bowker. Work and information practices in the sciences of biodiversity. In El Abbadi et al.

[1908], pages 693–696. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Bowker00.html.

Schnase:2000:RDB

[1524] John L. Schnase. Research directions in biodiversity informatics. In El Abbadi et al. [1908], pages 697–700. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Schnase00.html.

Cotter:2000:BII

[1525] Gladys A. Cotter and Barbara T.

Bauldock. Biodiversity informatics infrastructure: An information commons for the biodiversity community. In El Abbadi et al. [1908], pages 701–704. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/CotterB00.
html.

Banerjee:2000:DPB

[1526] Sandeepan Banerjee. A database platform for bioinformatics. In El Abbadi et al. [1908], pages 705–710. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Banerjee00.html.

Tsur:2000:DMB

[1527] Shalom Tsur. Data mining in the bioinformatics domain. In El Abbadi et al. [1908], pages 711–714. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Tsur00.html.

Schaefer:2000:CDW

[1528] Elvira Schaefer, Jan-Dirk Becker, Andreas Boehmer, and Matthias Jarke. Controlling data warehouses with knowledge networks. In El Abbadi et al. [1908], pages 715–718. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
SchaeferBBJ00.html.

Netz:2000:IDM

[1529] Amir Netz, Surajit Chaudhuri, Jeff Bernhardt, and Usama M. Fayyad. Integration of data mining with database technology. In El Abbadi et al. [1908], pages 719–722. ISBN 1-55860-715-3.

LCCN ???? URL http://www.vldb.
org/dblp/db/conf/vldb/NetzCBF00.
html.

Ramaswamy:2000:PET

[1530] Sridhar Ramaswamy. E.piphany epicenter technology overview. In El Abbadi et al. [1908], page 723. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Ramaswamy00.html.

Taylor:2000:CDW

[1531] Richard Taylor. Concurrency in the data warehouse. In El Abbadi et al. [1908], pages 724–727. ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
Taylor00.html.

Lane:2000:BIC

[1532] Meridith A. Lane, James L. Edwards, and Ebbe Nielsen. Biodiversity informatics: The challenge of rapid development, large databases, and complex data (keynote). In El Abbadi et al. [1908], pages 729– 732. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/LaneEN00.html.

Carey:2000:TWK

[1533] Michael J. Carey. Toto, we’re not in Kansas anymore: On transitioning from research to the real (invited industrial talk). In El Abbadi et al. [1908], page 733. ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/Carey00.html.

Anonymous:2000:AI

[1534] Anonymous. Author index. In El Abbadi et al. [1908], page ?? ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
X00a.html.

Anonymous:2000:OIC

[1535] Anonymous. Ordering information, conference organizers, program committees, additional reviewers, additional demonstrations reviewers, sponsors, VLDB endowment, preface, foreword. In El Abbadi et al. [1908], page ?? ISBN 1-55860-715-3. LCCN

???? URL http://www.vldb.org/
dblp/db/conf/vldb/X00.html.

Yagoub:2000:BCD

[1536] Khaled Yagoub, Daniela Florescu, Valérie Issarny, and Patrick Valduriez. Building and customizing dataintensive Web sites using weave. In El Abbadi et al. [1908], page ?? ISBN 1-55860-715-3. LCCN ???? URL http:/

/www.vldb.org/dblp/db/conf/vldb/
YagoubFIV00a.html.

Dayal:2001:BPC

[1537] Umeshwar Dayal, Meichun Hsu, and Rivka Ladin. Business process coordination: State of the art, trends, and open issues. In Apers et al.

[1909], pages 3–13. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001.

                                        URL http://www.vldb.org/conf/
                                        2001/P003.pdf.

Sol:2001:AIU

[1538] Egbert-Jan Sol. Ambient intelligence with the ubiquitous network, the embedded computer devices and the hidden databases (abstract). In Apers et al. [1909], page 14. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P014.pdf.

Wadler:2001:TXD

[1539] Philip Wadler. Et tu, XML? The downfall of the relational empire (abstract). In Apers et al. [1909], page 15. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P015.pdf.

Sondag:2001:SWP

[1540] Pierre-Paul Sondag. The semantic Web: Paving the way to the knowledge society. In Apers et al. [1909], page 16. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P016.pdf.

Stillger:2001:LDL

[1541] Michael Stillger, Guy M. Lohman, Volker Markl, and Mokhtar Kandil. LEO — DB2’s LEarning Optimizer. In Apers et al. [1909], pages 19–28. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P019.pdf.

Andrei:2001:UOC

[1542] Mihnea Andrei and Patrick Valduriez. User-optimizer communication using abstract plans in Sybase ASE. In Apers et al. [1909], pages 29–38. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P029.pdf.

VandenBercken:2001:XLA

[1543] Jochen Van den Bercken, Björn Blohsfeld, Jens-Peter Dittrich, Jürgen Krämer, Tobias Schäfer, Martin Schneider, and Bernhard Seeger. XXL — A library approach to supporting efficient implementations of advanced database queries. In Apers et al. [1909], pages 39–48. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P039.pdf.

Madhavan:2001:GSM

[1544] Jayant Madhavan, Philip A. Bernstein, and Erhard Rahm. Generic schema matching with cupid. In Apers et al. [1909], pages 49–58. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P049.pdf.

Chirkova:2001:FPV

[1545] Rada Chirkova, Alon Y. Halevy, and Dan Suciu. A formal perspective on the view selection problem. In Apers et al. [1909], pages 59–68. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P059.pdf.

Dill:2001:SSW

[1546] Stephen Dill, S. Ravi Kumar, Kevin S.

McCurley, Sridhar Rajagopalan, D. Sivakumar, and Andrew Tomkins. Selfsimilarity in the Web. In Apers et al. [1909], pages 69–78. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P069.pdf.

Gilbert:2001:SWS

[1547] Anna C. Gilbert, Yannis Kotidis, S. Muthukrishnan, and Martin Strauss. Surfing wavelets on streams: One-pass summaries for approximate aggregate queries. In Apers et al. [1909], pages 79–88. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P079.pdf.

Amir:2001:AQD

[1548] Amihood Amir, Reuven Kashi, and Nathan S. Netanyahu. Analyzing quantitative databases: Image is everything. In Apers et al. [1909], pages 89–98. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P089.pdf.

Stanoi:2001:DIS

[1549] Ioana Stanoi, Mirek Riedewald, Divyakant Agrawal, and Amr El Abbadi. Discovery of influence sets in frequently updated databases. In Apers et al. [1909], pages 99–108. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P099.pdf.

Crescenzi:2001:RTA

[1550] Valter Crescenzi, Giansalvatore Mecca, and Paolo Merialdo. RoadRunner: Towards automatic data extraction from large Web sites. In Apers et al. [1909], pages 109–118. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P109.pdf.

Baumgartner:2001:VWI

[1551] Robert Baumgartner, Sergio Flesca, and Georg Gottlob. Visual Web information extraction with Lixto. In Apers et al. [1909], pages 119–128. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P119.pdf.

Raghavan:2001:CHW

[1552] Sriram Raghavan and Hector Garcia-Molina. Crawling the hidden Web. In Apers et al. [1909], pages 129– 138. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P129.pdf.

Hunt:2001:DIL

[1553] Ela Hunt, Malcolm P. Atkinson, and Robert W. Irving. A database index to large biological sequences. In Apers et al. [1909], pages 139–148. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P139.pdf.

Agrawal:2001:SQC

[1554] Rakesh Agrawal, Amit Somani, and Yirong Xu. Storage and querying of E-commerce data. In Apers et al. [1909], pages 149–158. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P149.pdf.

Grigori:2001:IBP

[1555] Daniela Grigori, Fabio Casati, Umeshwar Dayal, and Ming-Chien Shan. Improving business process quality through exception understanding, prediction, and prevention. In Apers et al. [1909], pages 159–168. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P159.pdf.

Ailamaki:2001:WRC

[1556] Anastassia Ailamaki, David J. DeWitt, Mark D. Hill, and Marios Skounakis. Weaving relations for cache performance. In Apers et al. [1909], pages 169–180. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P169.pdf.

Cha:2001:CCC

[1557] Sang K. Cha, Sangyong Hwang, Kihong Kim, and Keunjoo Kwon. Cacheconscious concurrency control of mainmemory indexes on shared-memory multiprocessor systems. In Apers et al. [1909], pages 181–190. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P181.pdf.

Luo:2001:FBP

[1558] Qiong Luo and Jeffrey F. Naughton. Form-based proxy caching for databasebacked Web sites. In Apers et al. [1909], pages 191–200. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P191.pdf.

Bartolini:2001:FNA

[1559] Ilaria Bartolini, Paolo Ciaccia, and Florian Waas. FeedbackBypass: a new approach to interactive similarity query processing. In Apers et al. [1909], pages 201–210. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P201.pdf.

Bohm:2001:FET

[1560] Klemens Böhm, Michael Mlivoncic, Hans-Jörg Schek, and Roger Weber. Fast evaluation techniques for complex similarity queries. In Apers et al. [1909], pages 211–220. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P211.pdf.

Lin:2001:EHM

[1561] Shu Lin, M. Tamer Özsu, Vincent Oria, and Raymond T. Ng. An extendible hash for multi-precision similarity querying of image databases. In Apers et al. [1909], pages 221– 230. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P221.pdf.

Sakurai:2001:SSA

[1562] Yasushi Sakurai, Masatoshi Yoshikawa, Ryoji Kataoka, and Shunsuke Uemura. Similarity search for adaptive ellipsoid queries using spatial transformation. In Apers et al. [1909], pages 231–240. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P231.pdf.

Manolescu:2001:AXQ

[1563] Ioana Manolescu, Daniela Florescu, and Donald Kossmann. Answering XML queries on heterogeneous data sources. In Apers et al. [1909], pages 241–250. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P241.pdf.

Fegaras:2001:QEW

[1564] Leonidas Fegaras and Ramez Elmasri.

Query engines for Web-accessible XML data. In Apers et al. [1909], pages 251–260. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P251.pdf.

Shanmugasundaram:2001:QXV

[1565] Jayavel Shanmugasundaram, Jerry Kiernan, Eugene J. Shekita, Catalina Fan, and John Funderburk. Querying XML views of relational data. In Apers et al. [1909], pages 261–270. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P261.pdf.

Cluet:2001:VLS

[1566] Sophie Cluet, Pierangelo Veltri, and Dan Vodislav. Views in a large scale XML repository. In Apers et al. [1909], pages 271–280. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P271.pdf.

Natsev:2001:SIJ

[1567] Apostol Natsev, Yuan-Chi Chang, John R. Smith, Chung-Sheng Li, and Jeffrey Scott Vitter. Supporting incremental join queries on ranked inputs. In Apers et al. [1909], pages 281–290. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P281.pdf.

Chien:2001:EMM

[1568] Shu-Yao Chien, Vassilis J. Tsotras, and Carlo Zaniolo. Efficient management of multiversion documents by object referencing. In Apers et al. [1909], pages 291–300. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P291.pdf.

Tan:2001:EPS

[1569] Kian-Lee Tan, Pin-Kwang Eng, and Beng Chin Ooi. Efficient progressive skyline computation. In Apers et al. [1909], pages 301–310. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P301.pdf.

Margaritis:2001:NST

[1570] Dimitris Margaritis, Christos Faloutsos, and Sebastian Thrun. NetCube: a scalable tool for fast data mining and compression. In Apers et al. [1909], pages 311–320. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P311.pdf.

Dong:2001:MMD

[1571] Guozhu Dong, Jiawei Han, Joyce M. W. Lam, Jian Pei, and Ke Wang. Mining multi-dimensional constrained gradients in data cubes. In Apers et al. [1909], pages 321–330. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P321.pdf.

Nanopoulos:2001:CCB

[1572] Alexandros Nanopoulos, Yannis Theodoridis, and Yannis Manolopoulos. C^{2}P: Clustering based on closest pairs. In Apers et al. [1909], pages 331– 340. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P331.pdf.

Cooper:2001:FIS

[1573] Brian Cooper, Neal Sample, Michael J.

Franklin, Gisli R. Hjaltason, and Moshe Shadmon. A fast index for semistructured data. In Apers et al. [1909], pages 341–350. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P341.pdf.

Kahveci:2001:EIS

[1574] Tamer Kahveci and Ambuj K. Singh. Efficient index structures for string databases. In Apers et al. [1909], pages 351–360. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P351.pdf.

Li:2001:IQX

[1575] Quanzhong Li and Bongki Moon. Indexing and querying XML data for regular path expressions. In Apers et al. [1909], pages 361–370. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P361.pdf.

Galhardas:2001:DDC

[1576] Helena Galhardas, Daniela Florescu, Dennis Shasha, Eric Simon, and Cristian-Augustin Saita. Declarative data cleaning: Language, model, and algorithms. In Apers et al. [1909], pages 371–380. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P371.pdf.

Raman:2001:PWI

[1577] Vijayshankar Raman and Joseph M. Hellerstein. Potter’s Wheel: An interactive data cleaning system. In Apers et al. [1909], pages 381–390. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P381.pdf.

Labrinidis:2001:UPS

[1578] Alexandros Labrinidis and Nick Roussopoulos. Update propagation strategies for improving the quality of data on the Web. In Apers et al. [1909], pages 391–400. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P391.pdf.

Shou:2001:WTV

[1579] Lidan Shou, Jason Chionh, Zhiyong Huang, Yixin Ruan, and Kian-Lee Tan. Walking through a very large virtual environment in real-time. In Apers et al. [1909], pages 401–410. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P401.pdf.

An:2001:AEB

[1580] Ning An, Anand Sivasubramaniam, Narayanan Vijaykrishnan, Mahmut T. Kandemir, Mary Jane Irwin, and Sudhanva Gurumurthi. Analyzing energy behavior of spatial access methods for memory-resident data. In Apers et al. [1909], pages 411–420. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P411.pdf.

Yu:2001:IDE

[1581] Cui Yu, Beng Chin Ooi, Kian-Lee Tan, and H. V. Jagadish. Indexing the distance: An efficient method to KNN processing. In Apers et al. [1909], pages 421–430. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P421.pdf.

Tao:2001:MTS

[1582] Yufei Tao and Dimitris Papadias. MV3R-tree: a spatio-temporal access method for timestamp and interval queries. In Apers et al. [1909], pages 431–440. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P431.pdf.

Jensen:2001:TTT

[1583] Christian S. Jensen and David B. Lomet. Transaction timestamping in (temporal) databases. In Apers et al. [1909], pages 441–450. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P441.pdf.

Bratsberg:2001:OSH

[1584] Svein Erik Bratsberg and Rune Humborstad. Online scaling in a highly available database. In Apers et al. [1909], pages 451–460. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P451.pdf.

VandenBercken:2001:EGB

[1585] Jochen Van den Bercken and Bernhard Seeger. An evaluation of generic bulk loading techniques. In Apers et al. [1909], pages 461–470. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P461.pdf.

Cui:2001:LTG

[1586] Yingwei Cui and Jennifer Widom. Lineage tracing for general data warehouse transformations. In Apers et al. [1909], pages 471–480. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P471.pdf.

Kotidis:2001:DWA

[1587] Yannis Kotidis. A data warehousing architecture for enabling service provisioning process. In Apers et al. [1909], pages 481–490. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P481.pdf.

Gravano:2001:ASJ

[1588] Luis Gravano, Panagiotis G. Ipeirotis, H. V. Jagadish, Nick Koudas, S. Muthukrishnan, and Divesh Srivastava. Approximate string joins in a database (almost) for free. In Apers et al. [1909], pages 491–500. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P491.pdf.

Urhan:2001:DPS

[1589] Tolga Urhan and Michael J. Franklin.

Dynamic pipeline scheduling for improving interactive query performance. In Apers et al. [1909], pages 501– 510. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P501.pdf.

Sawai:2001:FPI

[1590] Rie Sawai, Masahiko Tsukamoto, Yin-Huei Loh, Tsutomu Terada, and Shojiro Nishio. Functional properties of information filtering. In Apers et al. [1909], pages 511–520. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P511.pdf.

Chun:2001:DUC

[1591] Seok-Ju Chun, Chin-Wan Chung, Ju-Hong Lee, and Seok-Lyong Lee. Dynamic update cube for range-sum queries. In Apers et al. [1909], pages 521–530. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P521.pdf.

Sathe:2001:IRM

[1592] Gayatri Sathe and Sunita Sarawagi. Intelligent rollups in multidimensional OLAP data. In Apers et al. [1909], pages 531–540. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001.

                                        URL http://www.vldb.org/conf/
                                        2001/P531.pdf.

Gibbons:2001:DSH

[1593] Phillip B. Gibbons. Distinct sampling for highly-accurate answers to distinct values queries and event reports. In Apers et al. [1909], pages 541–550. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P541.pdf.

Kemper:2001:HDD

[1594] Alfons Kemper and Christian Wiesner. Hyperqueries: Dynamic distributed query processing on the Internet. In Apers et al. [1909], pages 551–560. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P551.pdf.

Yang:2001:CHP

[1595] Beverly Yang and Hector Garcia-Molina. Comparing hybrid peerto-peer systems. In Apers et al. [1909], pages 561–570. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P561.pdf.

Aksoy:2001:DSD

[1596] Demet Aksoy, Michael J. Franklin, and Stanley B. Zdonik. Data staging for on-demand broadcast. In Apers et al. [1909], pages 571–580. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P571.pdf.

Marian:2001:CCM

[1597] Amélie Marian, Serge Abiteboul, Gregory Cobena, and Laurent Mignet. Change-centric management of versions in an XML warehouse. In Apers et al. [1909], pages 581–590. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P581.pdf.

Aboulnaga:2001:ESX

[1598] Ashraf Aboulnaga, Alaa R. Alameldeen, and Jeffrey F. Naughton. Estimating the selectivity of XML path expressions for Internet scale applications. In Apers et al. [1909], pages 591–600. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P591.pdf.

Marron:2001:PXL

[1599] Pedro José Marrón and Georg Lausen.

On processing XML in LDAP. In Apers et al. [1909], pages 601–610. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P601.pdf.

OConnell:2001:CAP

[1600] William O’Connell, Andrew Witkowski, and Goetz Graefe. Collaborative analytical processing — dream or reality? (panel abstract). In Apers et al. [1909], page 613. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P613.pdf.

Tsur:2001:WSN

[1601] Shalom Tsur, Serge Abiteboul, Rakesh Agrawal, Umeshwar Dayal, Johannes Klein, and Gerhard Weikum. Are Web services the next revolution in e-commerce? (panel). In Apers et al. [1909], pages 614–617. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P614.pdf.

Ozden:2001:SSP

[1602] Banu Özden, Eran Gabber, Bruce Hillyer, Wee Teck Ng, Elizabeth A. M. Shriver, David J. DeWitt, Bruce Gordon, Jim Gray, and John Wilkes. Storage service providers: a solution for storage management? (panel). In Apers et al. [1909], pages 618– 619. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P618.pdf.

Novaretti:2001:FGA

[1603] Serge Novaretti. French government activity in the conservation of data and electronic documents. In Apers et al. [1909], pages 621–624. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P621.pdf.

Duranti:2001:LTP

[1604] Luciana Duranti. The long-term preservation of authentic electronic records. In Apers et al. [1909], pages 625–628. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P625.pdf.

Gentili:2001:IEI

[1605] Mario Gentili. Italian electronic identity card — principle and architecture. In Apers et al. [1909], pages 629–632. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P629.pdf.

Bouguettaya:2001:OBS

[1606] Athman Bouguettaya, Ahmed K. Elmagarmid, Brahim Medjahed, and M. Ouzzani. Ontology-based support for digital government. In Apers et al. [1909], pages 633–636. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P633.pdf.

Draluk:2001:DWS

[1607] Vadim Draluk. Discovering Web services: An overview. In Apers et al. [1909], pages 637–640. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P637.pdf.

Shmueli:2001:AIW

[1608] Oded Shmueli. Architectures for internal Web services deployment. In Apers et al. [1909], pages 641–644. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P641.pdf.

Huyn:2001:SOB

[1609] Nam Huyn. Scientific OLAP for the biotech domain. In Apers et al. [1909], pages 645–648. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P645.pdf.

Bonifati:2001:WWD

[1610] Angela Bonifati, Fabio Casati, Umeshwar Dayal, and Ming-Chien Shan. Warehousing workflow data: Challenges and opportunities. In Apers et al. [1909], pages 649–652. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P649.pdf.

Sadri:2001:SPQ

[1611] Reza Sadri, Carlo Zaniolo, Amir M. Zarkesh, and Jafar Adibi. A sequential pattern query language for supporting instant data mining for e-services. In Apers et al. [1909], pages 653– 656. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P653.pdf.

Bunker:2001:AMD

[1612] Craig J. Bunker, Latha S. Colby, Richard L. Cole, William J. McKenna, Gopal Mulagund, and David Wilhite. Aggregate maintenance for data warehousing in Informix Red Brick Vista. In Apers et al. [1909], pages 659– 662. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P659.pdf.

Abrol:2001:NLS

[1613] Man Abrol, Neil Latarche, Uma Mahadevan, Jianchang Mao, Rajat Mukherjee, Prabhakar Raghavan, Michel Tourn, John Wang, and Grace Zhang. Navigating large-scale semistructured data in business portals. In Apers et al. [1909], pages 663– 666. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P663.pdf.

Datta:2001:CSA

[1614] Anindya Datta, Kaushik Dutta, Helen M. Thomas, Debra E. VanderMeer, Krithi Ramamritham, and Dan Fishman. A comparative study of alternative middle tier caching solutions to support dynamic Web content acceleration. In Apers et al. [1909], pages 667–670. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P667.pdf.

Carey:2001:PDS

[1615] Michael J. Carey, Steve Kirsch, Mary Roth, Bert Van der Linden, Nicolas Adiba, Michael Blow, Daniela Florescu, David Li, Ivan Oprencak, Rajendra Panwar, Runping Qi, David Rieber, John C. Shafer, Brian Sterling, Tolga Urhan, Brian Vickery, Dan Wineman, and Kuan Yee. The Propel Distributed Services Platform. In Apers et al. [1909], pages 671– 674. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P671.pdf.

Narayanan:2001:TSA

[1616] Srinivasa Narayanan and Subbu N.

Subramanian. Tavant system architecture for sell-side channel management. In Apers et al. [1909], pages 675–678. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P675.pdf.

Catozzi:2001:OSE

[1617] John Catozzi and Sorana Rabinovici.

Operating system extensions for the Teradata Parallel VLDB. In Apers et al. [1909], pages 679–682. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P679.pdf.

Lahiri:2001:CFE

[1618] Tirthankar Lahiri, Vinay Srihari, Wilson Chan, N. MacNaughton, and Sashikanth Chandrasekaran. Cache fusion: Extending shared-disk clusters with shared caches. In Apers et al. [1909], pages 683–686. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P683.pdf.

Adler:2001:DSE

[1619] David W. Adler. DB2 spatial extender — spatial data within the RDBMS. In Apers et al. [1909], pages 687– 690. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P687.pdf.

Brown:2001:PCB

[1620] Leonard Brown and Le Gruenwald. A prototype content-based retrieval system that uses virtual images to save space. In Apers et al. [1909], pages 693–694. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P693.pdf.

Maeda:2001:AAD

[1621] Yoko Maeda and Motomichi Toyama. ACTIVIEW: Adaptive data presentation using SuperSQL. In Apers et al. [1909], pages 695–696. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P695.pdf.

Zeng:2001:AAB

[1622] Liangzhao Zeng, Boualem Benatallah, Puong Nguyen, and Anne H. H. Ngu. AgFlow: Agent-based crossenterprise workflow management system. In Apers et al. [1909], pages 697–698. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

www.vldb.org/conf/2001/P697.pdf.

Li:2001:CPT

[1623] Wen-Syan Li, K. Selçuk Candan, Wang-Pin Hsiung, Oliver Po, Divyakant Agrawal, Qiong Luo, Wei-Kuang Waine Huang, Yusuf Akca, and Cemal Yilmaz. Cache portal: Technology for accelerating database-driven e-commerce Web sites. In Apers et al. [1909], pages 699–700. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P699.pdf.

Sundara:2001:DIS

[1624] Seema Sundara, Ying Hu, Timothy Chorma, Nipun Agarwal, and Jagannathan Srinivasan. Developing an indexing scheme for XML document collection using the Oracle8i Extensibility Framework. In Apers et al. [1909], pages 701–702. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P701.pdf.

Morishima:2001:EEU

[1625] Atsuyuki Morishima, Seiichi Koizumi, Hiroyuki Kitagawa, and Satoshi Takano. Enabling end-users to construct data-intensive Web-sites from XML repositories: An example-based approach. In Apers et al. [1909], pages 703–704. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P703.pdf.

Blok:2001:FSD

[1626] Henk Ernst Blok, Menzo Windhouwer, Roelof van Zwol, Milan Petkovic, Peter M. G. Apers, Martin L. Kersten, and Willem Jonker. Flexible and scalable digital library search. In Apers et al. [1909], pages 705–706. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P705.pdf.

May:2001:LSX

[1627] Wolfgang May. LoPiX: a system for XML data integration and manipulation. In Apers et al. [1909], pages 707–708. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001. URL http://

                                        www.vldb.org/conf/2001/P707.pdf.

Anciaux:2001:PVE

[1628] Nicolas Anciaux, Christophe Bobineau, Luc Bouganim, Philippe Pucheral, and Patrick Valduriez. PicoDBMS: Validation and experience. In Apers et al. [1909], pages 709–710. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P709.pdf.

Sindoni:2001:SRL

[1629] Giuseppe Sindoni, Leonardo Tininini, Amedea Ambrosetti, Cristina Bedeschi, Stefano De Francisci, Orietta Gargano, Rossella Molinaro, Mario Paolucci, Paola Patteri, and Pina Ticca. SIT-IN: a real-life spatiotemporal information system. In Apers et al. [1909], pages 711–712. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P711.pdf.

Kosch:2001:SDM

[1630] Harald Kosch, László Böszörményi, Alexander Bachlechner, Christian Hanin, Christian Hofbauer, Margit Lang, Carmen Riedler, and Roland Tusch. SMOOTH — A distributed multimedia database system. In Apers et al. [1909], pages 713–714. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P713.pdf.

Baumgartner:2001:SWG

[1631] Robert Baumgartner, Sergio Flesca, and Georg Gottlob. Supervised wrapper generation with Lixto. In Apers et al. [1909], pages 715–716. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P715.pdf.

Ambite:2001:WAS

[1632] José Luis Ambite, Craig A. Knoblock, Mohammad R. Kolahdouzan, Maria Muslea, Cyrus Shahabi, and Snehal Thakkar. The WorlInfo Assistant: Spatio-temporal information integration on the Web. In Apers et al. [1909], pages 717–718. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/P717.pdf.

Zhou:2001:VVX

[1633] Aoying Zhou, Hongjun Lu, Shihui Zheng, Yuqi Liang, Long Zhang, Wenyun Ji, and Zengping Tian. VXMLR: a visual XML-relational database system. In Apers et al.

[1909], pages 719–720. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P719.pdf.

Stohr:2001:WDA

[1634] Thomas Stöhr and Erhard Rahm. WARLOCK: a data allocation tool for parallel warehouses. In Apers et al. [1909], pages 721–722. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P721.pdf.

Pereira:2001:WHT

[1635] João Pereira, Françoise Fabret, Hans-Arno Jacobsen, François Llirbat, and Dennis Shasha. WebFilter: a highthroughput XML-based publish and subscribe system. In Apers et al. [1909], pages 723–724. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/P723.pdf.

Chaudhuri:2001:SRX

[1636] Surajit Chaudhuri and Kyuseok Shim. Storage and retrieval of XML data using relational databases. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/tut1.pdf.

Cherniack:2001:DMP

[1637] Mitch Cherniack, Michael J. Franklin, and Stanley B. Zdonik. Data management for pervasive computing. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

                                        2001. URL http://www.vldb.org/
                                        conf/2001/tut2.pdf.

Garofalakis:2001:AQP

[1638] Minos N. Garofalakis and Phillip B.

Gibbon. Approximate query processing: Taming the TeraBytes. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/tut4.pdf.

Leymann:2001:MBPa

[1639] Frank Leymann. Managing business processes via workflow technology. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/tut6.pdf.

Leymann:2001:MBPb

[1640] Frank Leymann. Managing business processes via workflow technology. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/tut6.pdf.

Mohan:2001:CTW

[1641] C. Mohan. Caching technologies for Web applications. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001.

URL http://www.almaden.ibm.com/
u/mohan/Caching_VLDB2001.pdf.

Paton:2001:IMG

[1642] Norman W. Paton and Carole A.

Goble. Information management for genome level bioinformatics. In Apers et al. [1909], page ?? ISBN 1-55860-804-4. LCCN QA76.9.D3 I559

2001. URL http://www.vldb.org/
conf/2001/tut5.pdf.

Imielinski:2002:WGD

[1643] Tomasz Imielinski and B. R. Badrinath. Wireless graffiti — data, data everywhere matters. In Bernstein et al. [1910], pages 9–19. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S02P01.pdf.

Weikum:2002:STD

[1644] Gerhard Weikum, Axel Mönkeberg, Christof Hasse, and Peter Zabback. Self-tuning database technology and information services: from wishful thinking to viable engineering. In Bernstein et al. [1910], pages 20– 31. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S02P02.pdf.

Cosley:2002:ROF

[1645] Dan Cosley, Steve Lawrence, and David M. Pennock. REFEREE: An open framework for practical testing of recommender systems using ResearchIndex. In Bernstein et al. [1910], pages 35–46. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S03P01.pdf.

Ashwin:2002:ASS

[1646] T. V. Ashwin, Rahul Gupta, and Sugata Ghosal. Adaptable similarity search using non-relevant information. In Bernstein et al. [1910], pages 47–58. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S03P02.pdf.

Ozsoyoglu:2002:SVA

[1647] Gultekin Özsoyoglu, Abdullah Al-Hamdani, Ismail Sengör Altingövde, Selma Ayse Özel, Özgür Ulusoy, and Z. Meral Özsoyoglu. Sideway value algebra for object-relational databases. In Bernstein et al. [1910], pages 59– 70. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S03P03.pdf.

Conrad:2002:DSU

[1648] Jack G. Conrad, Xi S. Guo, Peter Jackson, and Monem Meziou. Database selection using actual physical and acquired logical collection resources in a massive domain-specific operational environment. In Bernstein et al. [1910], pages 71–82. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S03P04.pdf.

Park:2002:SFI

[1649] Chang-Won Park, Jun-Ki Min, and Chin-Wan Chung. Structural function inlining technique for structurally recursive XML queries. In Bernstein et al. [1910], pages 83–94. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S04P01.
pdf.

Gottlob:2002:EAP

[1650] Georg Gottlob, Christoph Koch, and Reinhard Pichler. Efficient algorithms for processing XPath queries. In Bernstein et al. [1910], pages 95– 106. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S04P02.pdf.

Moerkotte:2002:IXP

[1651] Guido Moerkotte. Incorporating XSL processing into database engines. In Bernstein et al. [1910], pages 107– 118. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S04P03.pdf.

Bohannon:2002:OVQ

[1652] Philip Bohannon, S. Ganguly, Henry F. Korth, P. P. S. Narayan, and Pradeep Shenoy. Optimizing view queries in ROLEX to support navigable result trees. In Bernstein et al. [1910], pages 119–130. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S04P04.pdf.

Bouganim:2002:CSD

[1653] Luc Bouganim and Philippe Pucheral. Chip-secured data access: Confidential data on untrusted servers. In Bernstein et al. [1910], pages 131– 142. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S05P01.pdf.

Agrawal:2002:HD

[1654] Rakesh Agrawal, Jerry Kiernan, Ramakrishnan Srikant, and Yirong Xu. Hippocratic databases. In Bernstein et al. [1910], pages 143–154. ISBN 1-55860-869-9. LCCN ???? URL http:

                                        //www.vldb.org/conf/2002/S05P02.
                                        pdf.

Agrawal:2002:WRD

[1655] Rakesh Agrawal and Jerry Kiernan. Watermarking relational databases. In Bernstein et al. [1910], pages 155– 166. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S05P03.pdf.

Hulgeri:2002:PQO

[1656] Arvind Hulgeri and S. Sudarshan. Parametric query optimization for linear and piecewise linear cost functions.

In Bernstein et al. [1910], pages 167– 178. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S06P01.pdf.

Ghosh:2002:PSB

[1657] Antara Ghosh, Jignashu Parikh, Vibhuti S. Sengar, and Jayant R. Haritsa. Plan selection based on query clustering. In Bernstein et al. [1910], pages 179–190. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S06P02.pdf.

Manegold:2002:GDC

[1658] Stefan Manegold, Peter A. Boncz, and Martin L. Kersten. Generic database cost models for hierarchical memory systems. In Bernstein et al. [1910], pages 191–202. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S06P03.pdf.

Chandrasekaran:2002:SQS

[1659] Sirish Chandrasekaran and Michael J.

Franklin. Streaming queries over streaming data. In Bernstein et al. [1910], pages 203–214. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S07P01.pdf.

Carney:2002:MSN

[1660] Donald Carney, Ugur Çetintemel, Mitch Cherniack, Christian Convey, Sangdon Lee, Greg Seidman, Michael Stonebraker, Nesime Tatbul, and Stanley B. Zdonik. Monitoring streams — A new class of data management applications. In Bernstein et al. [1910], pages 215–226. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S07P02.pdf.

Ludascher:2002:TBX

[1661] Bertram Ludäscher, Pratik Mukhopadhyay, and Yannis Papakonstantinou. A transducer-based XML query processor. In Bernstein et al. [1910], pages 227–238. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S07P03.pdf.

Kaushik:2002:USI

[1662] Raghav Kaushik, Philip Bohannon, Jeffrey F. Naughton, and Pradeep Shenoy. Updates for structure indexes. In Bernstein et al. [1910], pages 239– 250. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S08P01.pdf.

Chan:2002:RTE

[1663] Chee Yong Chan, Minos N. Garofalakis, and Rajeev Rastogi. RE-Tree: An efficient index structure for regular expressions. In Bernstein et al. [1910], pages 263–274. ISBN 1-55860-869-9.

                                        LCCN ???? URL http://www.vldb.
                                        org/conf/2002/S08P02.pdf.

Chien:2002:ESJ

[1664] Shu-Yao Chien, Zografoula Vagena, Donghui Zhang, Vassilis J. Tsotras, and Carlo Zaniolo. Efficient structural joins on indexed XML documents. In Bernstein et al. [1910], pages 263– 274. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S08P03.pdf.

Kossmann:2002:SSS

[1665] Donald Kossmann, Frank Ramsak, and Steffen Rost. Shooting stars in the sky: An online algorithm for skyline queries. In Bernstein et al. [1910], pages 275–286. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S09P01.pdf.

Tao:2002:CNN

[1666] Yufei Tao, Dimitris Papadias, and Qiongmao Shen. Continuous nearest neighbor search. In Bernstein et al. [1910], pages 287–298. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S09P02.pdf.

Dittrich:2002:PMJ

[1667] Jens-Peter Dittrich, Bernhard Seeger, David Scot Taylor, and Peter Widmayer. Progressive merge join: a generic and non-blocking sort-based join algorithm. In Bernstein et al. [1910], pages 299–310. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S09P03.pdf.

Kiessling:2002:FPD

[1668] Werner Kießling. Foundations of preferences in database systems. In Bernstein et al. [1910], pages 311– 322. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S09P04.pdf.

Chen:2002:MDR

[1669] Yixin Chen, Guozhu Dong, Jiawei Han, Benjamin W. Wah, and Jianyong Wang. Multi-dimensional regression analysis of time-series data streams. In Bernstein et al. [1910], pages 323– 334. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S10P01.pdf.

Cormode:2002:CDS

[1670] Graham Cormode, Mayur Datar, Piotr Indyk, and S. Muthukrishnan.

Comparing data streams using Hamming norms (how to zero in). In Bernstein et al. [1910], pages 335– 345. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S10P02.pdf.

Manku:2002:AFC

[1671] Gurmeet Singh Manku and Rajeev Motwani. Approximate frequency counts over data streams. In Bernstein et al. [1910], pages 346–357. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S10P03.
pdf.

Zhu:2002:SSM

[1672] Yunyue Zhu and Dennis Shasha. Stat-Stream: Statistical monitoring of thousands of data streams in real time. In Bernstein et al. [1910], pages 358– 369. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S10P04.pdf.

Lempel:2002:ORP

[1673] Ronny Lempel and Shlomo Moran. Optimizing result prefetching in Web search engines with segmented indices. In Bernstein et al. [1910], pages 370– 381. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S11P01.pdf.

Lifantsev:2002:CDP

[1674] Maxim Lifantsev and Tzi cker Chiueh. I/O-conscious data preparation for large-scale Web search engines. In Bernstein et al. [1910], pages 382– 393. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S11P02.pdf.

Ipeirotis:2002:DSH

[1675] Panagiotis G. Ipeirotis and Luis Gravano. Distributed search over the hidden Web: Hierarchical database sampling and selection. In Bernstein et al. [1910], pages 394–405. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S11P03.pdf.

Keogh:2002:EID

[1676] Eamonn J. Keogh. Exact indexing of dynamic time warping. In Bernstein et al. [1910], pages 406– 417. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S12P01.pdf.

Tao:2002:AIS

[1677] Yufei Tao and Dimitris Papadias.

Adaptive index structures. In Bernstein et al. [1910], pages 418–429. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S12P02.
pdf.

Ramamurthy:2002:CFM

[1678] Ravishankar Ramamurthy, David J.

DeWitt, and Qi Su. A case for fractured mirrors. In Bernstein et al. [1910], pages 430–441. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S12P03.pdf.

Lim:2002:XLS

[1679] Lipyeow Lim, Min Wang, Sriram Padmanabhan, Jeffrey Scott Vitter, and Ronald Parr. XPathLearner: An online self-tuning Markov histogram for XML path selectivity estimation. In Bernstein et al. [1910], pages 442– 453. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S13P01.pdf.

Gilbert:2002:HSU

[1680] Anna C. Gilbert, Yannis Kotidis, S. Muthukrishnan, and Martin Strauss. How to summarize the universe: Dynamic maintenance of quantiles. In Bernstein et al. [1910], pages 454– 465. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S13P02.pdf.

Polyzotis:2002:SVS

[1681] Neoklis Polyzotis and Minos N. Garofalakis. Structure and value synopses for XML data graphs. In Bernstein et al. [1910], pages 466– 477. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S13P03.pdf.

Yu:2002:CAM

[1682] Ting Yu, Divesh Srivastava, Laks V. S. Lakshmanan, and H. V. Jagadish. Compressed accessibility map: Efficient access control for XML. In Bernstein et al. [1910], pages 478– 489. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S14P01.pdf.

Cho:2002:OSE

[1683] SungRan Cho, Sihem Amer-Yahia, Laks V. S. Lakshmanan, and Divesh Srivastava. Optimizing the secure evaluation of twig queries. In Bernstein et al. [1910], pages 490– 501. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S14P02.pdf.

Bettini:2002:POP

[1684] Claudio Bettini, Sushil Jajodia, Xiaoyang Sean Wang, and Duminda Wijesekera. Provisions and obligations in policy management and security applications. In Bernstein et al. [1910], pages 502–513. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S14P03.pdf.

Cho:2002:ECD

[1685] Junghoo Cho and Alexandros Ntoulas.

Effective change detection using sampling. In Bernstein et al. [1910], pages 514–525. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S15P01.pdf.

Shah:2002:MCD

[1686] Shetal Shah, Krithi Ramamritham, and Prashant J. Shenoy. Maintaining coherency of dynamic data in cooperating repositories. In Bernstein et al. [1910], pages 526–537. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S15P02.pdf.

Uhl:2002:BMI

[1687] Axel Uhl. A bandwidth model for Internet search. In Bernstein et al. [1910], pages 538–549. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S15P03.pdf.

Bright:2002:ULR

[1688] Laura Bright and Louiqa Raschid. Using latency-recency profiles for data delivery on the Web. In Bernstein et al. [1910], pages 550–561. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S16P01.pdf.

Candan:2002:VID

[1689] K. Selçuk Candan, Divyakant Agrawal, Wen-Syan Li, Oliver Po, and Wang-Pin Hsiung. View invalidation for dynamic content caching in multitiered architectures. In Bernstein et al. [1910], pages 562–573. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S16P02.pdf.

Kounev:2002:IDA

[1690] Samuel Kounev and Alejandro P. Buchmann. Improving data access of J2EE applications by exploiting asynchronous messaging and caching services. In Bernstein et al. [1910], pages 574–585. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S16P03.pdf.

Ananthakrishna:2002:EFD

[1691] Rohit Ananthakrishna, Surajit Chaudhuri, and Venkatesh Ganti. Eliminating fuzzy duplicates in data warehouses. In Bernstein et al. [1910], pages 586–597. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S17P01.pdf.

Popa:2002:TWD

[1692] Lucian Popa, Yannis Velegrakis, Renée J. Miller, Mauricio A. Hernández, and Ronald Fagin. Translating Web data. In Bernstein et al. [1910], pages 598–609. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S17P02.pdf.

Do:2002:CSF

[1693] Hong Hai Do and Erhard Rahm. COMA — A system for flexible combination of schema matching approaches. In Bernstein et al. [1910], pages 610– 621. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S17P03.pdf.

Stolte:2002:EEL

[1694] Etzard Stolte and Gustavo Alonso. Efficient exploration of large scientific databases. In Bernstein et al. [1910], pages 622–633. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S18P01.pdf.

Hammel:2002:SSS

[1695] Laurie Hammel and Jignesh M. Patel.

Searching on the secondary structure of protein sequences. In Bernstein et al. [1910], pages 634–645. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S18P02.pdf.

Nierman:2002:PPD

[1696] Andrew Nierman and H. V. Jagadish.

ProTDB: Probabilistic data in XML. In Bernstein et al. [1910], pages 646– 657. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S18P03.pdf.

Chakrabarti:2002:FAT

[1697] Soumen Chakrabarti, Shourya Roy, and Mahesh V. Soundalgekar. Fast and accurate text classification via multiple linear discriminant projections. In Bernstein et al. [1910], pages 658– 669. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S19P01.pdf.

Hristidis:2002:DKS

[1698] Vagelis Hristidis and Yannis Papakonstantinou. DISCOVER: Keyword search in relational databases. In Bernstein et al. [1910], pages 670– 681. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S19P02.pdf.

Rizvi:2002:MDP

[1699] Shariq Rizvi and Jayant R. Haritsa. Maintaining data privacy in association rule mining. In Bernstein et al. [1910], pages 682–693. ISBN 1-55860-869-9.

                                        LCCN ???? URL http://www.vldb.
                                        org/conf/2002/S19P03.pdf.

Senkul:2002:LFS

[1700] Pinar Senkul, Michael Kifer, and Ismail H. Toroslu. A logical framework for scheduling workflows under resource allocation constraints. In Bernstein et al. [1910], pages 694– 705. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S20P01.pdf.

Blott:2002:ASP

[1701] Stephen Blott and Henry F. Korth. An almost-serial protocol for transaction execution in main-memory database systems. In Bernstein et al. [1910], pages 706–717. ISBN 1-55860-869-9.

                                        LCCN ???? URL http://www.vldb.
                                        org/conf/2002/S20P02.pdf.

Daynes:2002:LFI

[1702] Laurent Daynès and Grzegorz Czajkowski. Lightweight flexible isolation for language-based extensible systems. In Bernstein et al. [1910], pages 718– 729. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S20P03.pdf.

Karayannidis:2002:PSQ

[1703] Nikos Karayannidis, Aris Tsois, Timos K. Sellis, Roland Pieringer, Volker Markl, Frank Ramsak, Robert Fenk, Klaus Elhardt, and Rudolf Bayer. Processing star queries on hierarchically-clustered fact tables. In Bernstein et al. [1910], pages 730– 741. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S21P01.pdf.

Kang:2002:EVL

[1704] Heum-Geun Kang and Chin-Wan Chung. Exploiting versions for on-line data warehouse maintenance in MO-LAP servers. In Bernstein et al. [1910], pages 742–753. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S21P02.pdf.

Rohm:2002:FFS

[1705] Uwe Röhm, Klemens Böhm, Hans-Jörg Schek, and Heiko Schuldt. FAS — A freshness-sensitive coordination middleware for a cluster of OLAP components. In Bernstein et al. [1910], pages 754–765. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S21P03.pdf.

Lakshmanan:2002:GMA

[1706] Laks V. S. Lakshmanan, Raymond T.

Ng, Christine Xing Wang, Xiaodong Zhou, and Theodore Johnson. The generalized MDL approach for summarization. In Bernstein et al. [1910], pages 766–777. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S22P01.pdf.

Lakshmanan:2002:QCH

[1707] Laks V. S. Lakshmanan, Jian Pei, and Jiawei Han. Quotient cube: How to summarize the semantics of a data cube. In Bernstein et al. [1910], pages 778–789. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S22P02.pdf.

Lee:2002:OPA

[1708] Young-Koo Lee, Kyu-Young Whang, Yang-Sae Moon, and Il-Yeol Song. A one-pass aggregation algorithm with the optimal buffer size in multidimensional OLAP. In Bernstein et al. [1910], pages 790–801. ISBN 1-55860-869-9.

                                        LCCN ???? URL http://www.vldb.
                                        org/conf/2002/S22P03.pdf.

Palpanas:2002:IMN

[1709] Themistoklis Palpanas, Richard Sidle, Roberta Cochrane, and Hamid Pirahesh. Incremental maintenance for nondistributive aggregate functions. In Bernstein et al. [1910], pages 802– 813. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S22P04.pdf.

Korn:2002:RNN

[1710] Flip Korn, S. Muthukrishnan, and Divesh Srivastava. Reverse nearest neighbor aggregates over data streams. In Bernstein et al. [1910], pages 814– 825. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S23P01.pdf.

Chan:2002:TPA

[1711] Chee Yong Chan, Wenfei Fan, Pascal Felber, Minos N. Garofalakis, and Rajeev Rastogi. Tree pattern aggregation for scalable XML data dissemination. In Bernstein et al. [1910], pages 826–837. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S23P02.pdf.

Benedikt:2002:DDP

[1712] Michael Benedikt, Chee Yong Chan, Wenfei Fan, Rajeev Rastogi, Shihui Zheng, and Aoying Zhou. DTDdirected publishing with attribute translation grammars. In Bernstein et al. [1910], pages 838–849. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S23P03.
pdf.

Seifert:2002:MVC

[1713] Andrè Seifert and Marc H. Scholl. A multi-version cache replacement and prefetching policy for hybrid data delivery environments. In Bernstein et al. [1910], pages 850–861. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S23P04.pdf.

Fox:2002:TRO

[1714] Armando Fox. Toward recoveryoriented computing. In Bernstein et al. [1910], pages 873–876. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S25P01.pdf.

Lohman:2002:SMD

[1715] Guy M. Lohman and Sam Lightstone.

SMART: Making DB2 (more) autonomic. In Bernstein et al. [1910], pages 877–879. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S25P02.pdf.

Sayal:2002:BPC

[1716] Mehmet Sayal, Fabio Casati, Umeshwar Dayal, and Ming-Chien Shan. Business process cockpit. In Bernstein et al. [1910], pages 880–883. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S26P01.
pdf.

Vingralek:2002:GSF

[1717] Radek Vingralek. GnatDb: a smallfootprint, secure database system. In Bernstein et al. [1910], pages 884– 893. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S26P02.pdf.

Zeller:2002:ERE

[1718] Bernhard Zeller and Alfons Kemper. Experience report: Exploiting advanced database optimization features for large-scale SAP R/3 installations. In Bernstein et al. [1910], pages 894– 905. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S26P03.pdf.

Selinger:2002:IIX

[1719] Patricia G. Selinger. Information integration and XML in IBM’s DB2. In Bernstein et al. [1910], pages 906– 907. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S27P01.pdf.

Goto:2002:NPS

[1720] Koichi Goto and Yahiko Kambayashi. A new passenger support system for public transport using mobile database access. In Bernstein et al. [1910], pages 908–919. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S28P01.pdf.

Alexandrov:2002:OOH

[1721] I. Alexandrov, A. Amorim, E. Badescu, M. Barczyk, D. Burckhart-Chromek, M. Caprini, M. Dobson, J. Flammer, R. Hart, R. Jones, A. Kazarov, S. Kolos, V. Kotov, D. Liko, L. Lucio, L. Mapelli, M. Mineev, L. Moneta, I. Papadopoulos, M. Nassiakou, N. Parrington, L. Pedro, A. Ribeiro, Yu. Ryabov, D. Schweiger, I. Soloviev, and H. Wolters. OBK — an online high energy physics’ meta-data repository. In Bernstein et al. [1910], pages 920–927. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S28P02.pdf.

Laud:2002:GHP

[1722] Amey V. Laud, Sourav S. Bhowmick, Pedro Cruz, Dadabhai T. Singh, and George Rajesh. The gRNA: a highly programmable infrastructure for prototyping, developing and deploying genomics-centric applications. In Bernstein et al. [1910], pages 928– 939. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S28P03.pdf.

Mohan:2002:EMP

[1723] C. Mohan. An efficient method for performing record deletions and updates using index scans. In Bernstein et al. [1910], pages 940–949. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S29P01.pdf.

Ilyas:2002:JRI

[1724] Ihab F. Ilyas, Walid G. Aref, and Ahmed K. Elmagarmid. Joining ranked inputs in practice. In Bernstein et al. [1910], pages 950–961. ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/conf/2002/S29P02.pdf.

Dageville:2002:SMM

[1725] Benoît Dageville and Mohamed Zaït.

SQL memory management in oracle9i. In Bernstein et al. [1910], pages 962– 973. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S29P03.pdf.

Schmidt:2002:XBX

[1726] Albrecht Schmidt, Florian Waas, Martin L. Kersten, Michael J. Carey, Ioana Manolescu, and Ralph Busse. XMark: a benchmark for XML data management. In Bernstein et al. [1910], pages 974–985. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S30P01.pdf.

Pan:2002:DDI

[1727] Alberto Pan, Juan Raposo, Manuel Álvarez, Paula Montoto, Vicente Orjales, Justo Hidalgo, Lucía Ardao, Anastasio Molano, and Ángel Viña. The denodo data integration platform. In Bernstein et al. [1910], pages 986– 989. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S30P02.pdf.

Kiessling:2002:PSD

[1728] Werner Kießling and Gerhard Köstler. Preference SQL — design, implementation, experiences. In Bernstein et al. [1910], pages 990–1001. ISBN 1-55860-869-9. LCCN ???? URL http://www.

                                        vldb.org/conf/2002/S30P03.pdf.

King:2002:RSD

[1729] Roger King. The rubicon of smart data. In Bernstein et al. [1910], pages 1002– 1005. ISBN 1-55860-869-9. LCCN

                                        ???? URL http://www.vldb.org/
                                        conf/2002/S31P01.pdf.

Sripada:2002:IMC

[1730] Suryanarayana M. Sripada. Information management challenges from the aerospace industry. In Bernstein et al.

[1910], pages 1006–1007. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S31P02.
pdf.

Ohura:2002:EQE

[1731] Yusuke Ohura, Katsumi Takahasii, Iko Pramudiono, and Masaru Kitsuregawa. Experiments on query expansion for Internet Yellow Page services using Web log mining. In Bernstein et al. [1910], pages 1008–1018. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S32P01.pdf.

Li:2002:IEC

[1732] Wen-Syan Li, Wang-Pin Hsiung, Dmitri V. Kalashnikov, Radu Sion, Oliver Po, Divyakant Agrawal, and K. Selçuk Candan. Issues and evaluations of caching solutions for Web application acceleration. In Bernstein et al. [1910], pages 1019–1030. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S32P02.
pdf.

Aggarwal:2002:ASW

[1733] Charu C. Aggarwal and Philip S. Yu.

An automated system for Web portal personalization. In Bernstein et al. [1910], pages 1031–1040. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S32P03.
pdf.

Shegalov:2002:EEO

[1734] German Shegalov, Gerhard Weikum, Roger S. Barga, and David B. Lomet. EOS: Exactly-once E-service middleware. In Bernstein et al. [1910], pages 1043–1046. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S33P01.pdf.

Keidl:2002:SDS

[1735] Markus Keidl, Stefan Seltzsam, Konrad Stocker, and Alfons Kemper. ServiceGlobe: Distributing E-services across the Internet. In Bernstein et al. [1910], pages 1047–1050. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S33P02.
pdf.

Sheng:2002:SSP

[1736] Quan Z. Sheng, Boualem Benatallah, Marlon Dumas, and Eileen Oi-Yan Mak. SELF-SERV: a platform for rapid composition of Web services in a peer-to-peer environment. In Bernstein et al. [1910], pages 1051– 1054. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P03.pdf.

Agrawal:2002:DTE

[1737] Rakesh Agrawal, Ramakrishnan Srikant, and Yirong Xu. Database technologies for electronic commerce. In Bernstein et al. [1910], pages 1055–1058. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S33P04.
pdf.

Hsu:2002:ADT

[1738] Wynne Hsu, Mong-Li Lee, Beng Chin Ooi, Pranab Kumar Mohanty, Keng Lik Teo, and Chenyi Xia. Advanced database technologies in a diabetic healthcare system. In Bernstein et al. [1910], pages 1059–1062. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S33P05.
pdf.

Lam:2002:RRT

[1739] Kam yiu Lam, Alan Kwan, and Krithi Ramamritham. RTMonitor: Real-time data monitoring using mobile agent technologies. In Bernstein et al. [1910], pages 1063–1066. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S33P06.pdf.

Heymann:2002:VTF

[1740] Stephan Heymann, Katja Tham, Axel Kilian, Gunnar Wegner, Peter Rieger, Dieter Merkel, and Johann Christoph Freytag. Viator — A tool family for graphical networking and data view creation. In Bernstein et al. [1910], pages 1067–1070. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S33P07.pdf.

Cilia:2002:PIC

[1741] Mariano Cilia, P. Hasselmayer, and Alejandro P. Buchmann. Profiling and Internet connectivity in automotive environments. In Bernstein et al. [1910], pages 1071–1074. ISBN 1-55860-869-9.

LCCN ???? URL http://www.vldb.
org/conf/2002/S33P08.pdf.

Santini:2002:GGM

[1742] Simone Santini and Amarnath Gupta.

GeMBASE: a geometric mediator for brain analysis with surface ensembles. In Bernstein et al. [1910], pages 1075– 1078. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P09.pdf.

Mahnke:2002:EOS

[1743] Wolfgang Mahnke, Christian Mathis, and Hans-Peter Steiert. Extending an ORDBMS: The StateMachine module.

In Bernstein et al. [1910], pages 1079– 1082. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P10.pdf.

Aditya:2002:BBK

[1744] B. Aditya, Gaurav Bhalotia, Soumen Chakrabarti, Arvind Hulgeri, Charuta Nakhe, Parag, and S. Sudarshan. BANKS: Browsing and keyword searching in relational databases. In Bernstein et al. [1910], pages 1083– 1086. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P11.pdf.

Abiteboul:2002:AXP

[1745] Serge Abiteboul, Omar Benjelloun, Ioana Manolescu, Tova Milo, and Roger Weber. Active XML: Peer-to-peer data and Web services integration. In Bernstein et al. [1910], pages 1087– 1090. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P12.pdf.

Bohannon:2002:LCR

[1746] Philip Bohannon, Juliana Freire, Jayant R. Haritsa, Maya Ramanath, Prasan Roy, and Jérôme Siméon. LegoDB: Customizing relational storage for XML documents. In Bernstein et al. [1910], pages 1091–1094. ISBN 1-55860-869-9. LCCN ???? URL http:

//www.vldb.org/conf/2002/S33P13.
pdf.

Herlekar:2002:ESF

[1747] Anandi Herlekar, Atul Deopujari, Krithi Ramamritham, Shaymsunder Gopale, and Shridhar Shukla. enTrans: a system for flexible consistency maintenance in directory applications. In Bernstein et al. [1910], pages 1095– 1098. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P14.pdf.

Rantzau:2002:CDC

[1748] Ralf Rantzau, Carmen Constantinescu, Uwe Heinkel, and Holger Meinecke. Champagne: Data change propagation for heterogeneous information systems. In Bernstein et al. [1910], pages 1099– 1102. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P15.pdf.

Sarawagi:2002:AAL

[1749] Sunita Sarawagi, Anuradha Bhamidipaty, Alok Kirpal, and Chandra Mouli. ALIAS: An active learning led interactive deduplication system. In Bernstein et al. [1910], pages 1103– 1106. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P16.pdf.

Liu:2002:TPS

[1750] Haifeng Liu and Hans-Arno Jacobsen.

A-TOPSS — A publish/subscribe system supporting approximate matching. In Bernstein et al. [1910], pages 1107– 1110. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
conf/2002/S33P17.pdf.

Chakrabarti:2002:TSM

[1751] Soumen Chakrabarti. Tutorial: Searching and mining fine-grained semistructured data. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
c/Chakrabarti:Soumen.html.

Faloutsos:2002:TSD

[1752] Christos Faloutsos. Tutorial: Sensor data mining: Similarity search and pattern analysis. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        f/Faloutsos:Christos.html.

Garofalakis:2002:TQM

[1753] Minos N. Garofalakis and Johannes Gehrke. Tutorial: Querying and mining data streams: You only get one look. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9.

                                        LCCN ????  URL http://www.
                                        vldb.org/dblp/db/indices/a-tree/
                                        g/Garofalakis:Minos_N=.html.

Mohan:2002:TAS

[1754] C. Mohan. Tutorial: Application servers and associated technologies. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/m/Mohan:C=.html.

Nori:2002:TES

[1755] Anil Nori. Tutorial: eBusiness standards and architectures. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        n/Nori:Anil.html.

Sarawagi:2002:TAI

[1756] Sunita Sarawagi. Tutorial: Automation in information extraction and data integration. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        s/Sarawagi:Sunita.html.

Shasha:2002:TDT

[1757] Dennis Shasha and Philippe Bonnet.

Tutorial: Database tuning: Principles, experiments, and troubleshooting techniques. In Bernstein et al. [1910], page ?? ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/s/Shasha:
Dennis.html.

Eleftheriou:2003:NBA

[1758] E. Eleftheriou, P. Bächtold, G. Cherubini, A. Dholakia, C. Hagleitner, T. Loeliger, A. Pantazi, H. Pozidis, T. R. Albrecht, Gerd K. Binnig, Michel Despont, Ute Drechsler, Urs Dürig, B. Gotsmann, D. Jubin, Walter Häberle, M. A. Lantz, Hugo E. Rothuizen, Richard Stutz, Peter Vettiger, and D. Wiesmann. A nanotechnology-based approach to data storage. In Freytag et al. [1911], pages 3–7. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
e/Eleftheriou:E=.html.

Mattos:2003:IID

[1759] Nelson Mendonça Mattos. Integrating information for on demand computing. In Freytag et al. [1911], pages 8–14. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/m/Mattos:
Nelson_Mendon=ccedil=a.html.

Shenker:2003:DCR

[1760] Scott Shenker. The data-centric revolution in networking. In Freytag et al. [1911], pages 15–15 (??). ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/s/Shenker:Scott.html.

Ioannidis:2003:HHA

[1761] Yannis E. Ioannidis. The history of histograms (abridged). In Freytag et al. [1911], pages 19–30. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        i/Ioannidis:Yannis_E=.html.

Raghavan:2003:CQW

[1762] Sriram Raghavan and Hector Garcia-Molina. Complex queries over Web repositories. In Freytag et al. [1911], pages 33–44. ISBN 0-12-722442-4.

                                        LCCN ????  URL http://www.
                                        vldb.org/dblp/db/indices/a-tree/
                                        r/Raghavan:Sriram.html.

Cohen:2003:XSS

[1763] Sara Cohen, Jonathan Mamou, Yaron Kanza, and Yehoshua Sagiv. XSEarch: a semantic search engine for XML. In Freytag et al. [1911], pages 45– 56. ISBN 0-12-722442-4. LCCN

                                        ???? URL http://www.vldb.org/
                                        dblp/db/indices/a-tree/c/Cohen:
                                        Sara.html.

Shah:2003:ERA

[1764] Shetal Shah, Shyamshankar Dharmarajan, and Krithi Ramamritham. An efficient and resilient approach to filtering and disseminating streaming data. In Freytag et al. [1911], pages 57–68. ISBN 0-12-722442-4.

                                        LCCN ????  URL http://www.
                                        vldb.org/dblp/db/indices/a-tree/
                                        s/Shah:Shetal.html.

Yang:2003:EMX

[1765] Liang Huai Yang, Mong-Li Lee, and Wynne Hsu. Efficient mining of XML query patterns for caching. In Freytag et al. [1911], pages 69–80. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/y/Yang:Liang_Huai.html.

Aggarwal:2003:FCE

[1766] Charu C. Aggarwal, Jiawei Han, Jianyong Wang, and Philip S. Yu. A framework for clustering evolving data streams. In Freytag et al. [1911], pages 81–92. ISBN 0-12-722442-4.

LCCN ????  URL http://www.
vldb.org/dblp/db/indices/a-tree/
a/Aggarwal:Charu_C=.html.

Teng:2003:RBT

[1767] Wei-Guang Teng, Ming-Syan Chen, and Philip S. Yu. A regression-based temporal pattern mining scheme for data streams. In Freytag et al. [1911], pages 93–104. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
t/Teng:Wei=Guang.html.

Kumaran:2003:CMD

[1768] A. Kumaran and Jayant R. Haritsa.

On the costs of multilingualism in database systems. In Freytag et al. [1911], pages 105–116. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/k/Kumaran:A=.html.

Yu:2003:DTQ

[1769] Clement T. Yu, George Philip, and Weiyi Meng. Distributed top-N query processing with possibly uncooperative local systems. In Freytag et al. [1911], pages 117–128. ISBN 0-12-722442-4.

LCCN ???? URL http://www.vldb.
org/dblp/db/indices/a-tree/y/Yu:
Clement_T=.html.

Long:2003:OQE

[1770] Xiaohui Long and Torsten Suel. Optimized query execution in large search engines with global page ordering. In Freytag et al. [1911], pages 129–140. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/Long:Xiaohui.html.

Buneman:2003:PQC

[1771] Peter Buneman, Martin Grohe, and Christoph Koch. Path queries on compressed XML. In Freytag et al. [1911], pages 141–152. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        b/Buneman:Peter.html.

Flesca:2003:MXQ

[1772] Sergio Flesca, Filippo Furfaro, and Elio Masciari. On the minimization of Xpath queries. In Freytag et al. [1911], pages 153–164. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/f/Flesca:Sergio.html.

Ramanan:2003:CIX

[1773] Prakash Ramanan. Covering indexes for XML queries: Bisimulation — simulation = negation. In Freytag et al. [1911], pages 165–176. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/r/Ramanan:Prakash.html.

Amer-Yahia:2003:PMX

[1774] Sihem Amer-Yahia, Mary F. Fernandez, Divesh Srivastava, and Yu Xu. Phrase matching in XML. In Freytag et al. [1911], pages 177–188. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/a/Amer=Yahia:Sihem.html.

Chen:2003:RRR

[1775] Yi Chen, Susan B. Davidson, Carmem S.

Hara, and Yifeng Zheng. RRXF: Redundancy reducing XML storage in relations. In Freytag et al. [1911], pages 189–200. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
c/Chen:Yi.html.

Deutsch:2003:MSP

[1776] Alin Deutsch and Val Tannen. MARS:

a system for publishing XML from mixed and redundant storage. In Freytag et al. [1911], pages 201–212. ISBN 0-12-722442-4. LCCN ???? URL http:

//www.vldb.org/dblp/db/indices/
a-tree/d/Deutsch:Alin.html.

Marian:2003:PXD

[1777] Amélie Marian and Jérôme Siméon.

Projecting XML documents. In Freytag et al. [1911], pages 213– 224. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/m/Marian:
Am=eacute=lie.html.

Halverson:2003:MMX

[1778] Alan Halverson, Josef Burger, Leonidas Galanis, Ameet Kini, Rajasekar Krishnamurthy, Ajith Nagaraja Rao, Feng Tian, Stratis Viglas, Yuan Wang, Jeffrey F. Naughton, and David J. De-Witt. Mixed mode XML query processing. In Freytag et al. [1911], pages 225–236. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
h/Halverson:Alan.html.

Chen:2003:TPG

[1779] Zhimin Chen, H. V. Jagadish, Laks V. S. Lakshmanan, and Stelios Paparizos. From tree patterns to generalized tree patterns: On efficient evaluation of XQuery. In Freytag et al. [1911], pages 237–248. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/c/Chen:Zhimin.html.

Koch:2003:EPE

[1780] Christoph Koch. Efficient processing of expressive node-selecting queries on XML data in secondary storage: a tree automata-based approach. In Freytag et al. [1911], pages 249–260. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/k/Koch:Christoph.html.

Diao:2003:QPH

[1781] Yanlei Diao and Michael J. Franklin. Query processing for high-volume XML message brokering. In Freytag et al. [1911], pages 261–272. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/d/Diao:Yanlei.html.

Jiang:2003:HTJ

[1782] Haifeng Jiang, Wei Wang, Hongjun Lu, and Jeffrey Xu Yu. Holistic twig joins on indexed XML documents. In Freytag et al. [1911], pages 273–284. ISBN 0-12-722442-4. LCCN ???? URL http:

                                        //www.vldb.org/dblp/db/indices/
                                        a-tree/j/Jiang:Haifeng.html.

Viglas:2003:MOR

[1783] Stratis Viglas, Jeffrey F. Naughton, and Josef Burger. Maximizing the output rate of multi-way join queries over streaming information sources. In Freytag et al. [1911], pages 285– 296. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/v/Viglas:
Stratis.html.

Hammad:2003:SSW

[1784] Moustafa A. Hammad, Michael J.

Franklin, Walid G. Aref, and Ahmed K. Elmagarmid. Scheduling for shared window joins over data streams. In Freytag et al. [1911], pages 297– 308. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/h/Hammad:
Moustafa_A=.html.

Tatbul:2003:LSD

[1785] Nesime Tatbul, Ugur Çetintemel, Stanley B. Zdonik, Mitch Cherniack, and Michael Stonebraker. Load shedding in a data stream manager. In Freytag et al. [1911], pages 309–320. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/t/Tatbul:Nesime.html.

Huebsch:2003:QIP

[1786] Ryan Huebsch, Joseph M. Hellerstein, Nick Lanham, Boon Thau Loo, Scott Shenker, and Ion Stoica. Querying the Internet with PIER. In Freytag et al. [1911], pages 321–332. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/h/Huebsch:Ryan.html.

Tian:2003:TRS

[1787] Feng Tian and David J. DeWitt. Tuple routing strategies for distributed eddies. In Freytag et al. [1911], pages 333–344. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        t/Tian:Feng.html.

Lerner:2003:AQL

[1788] Alberto Lerner and Dennis Shasha. AQuery: Query language for ordered data, optimization techniques, and experiments. In Freytag et al. [1911], pages 345–356. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/Lerner:Alberto.html.

He:2003:WIA

[1789] Hai He, Weiyi Meng, Clement T. Yu, and Zonghuan Wu. WISE-integrator: An automatic integrator of Web search interfaces for E-commerce. In Freytag et al. [1911], pages 357–368. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/h/He:Hai.html.

Lim:2003:SSA

[1790] Lipyeow Lim, Min Wang, and Jeffrey Scott Vitter. SASH: a self-adaptive histogram set for dynamically changing workloads. In Freytag et al. [1911], pages 369–380. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/Lim:Lipyeow.html.

Lin:2003:VVL

[1791] Chi-Chun Lin and Ming-Syan Chen. VIPAS: Virtual Link Powered Authority Search in the Web. In Freytag et al. [1911], pages 381–392. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/l/Lin:Chi=Chun.html.

Labrinidis:2003:BPD

[1792] Alexandros Labrinidis and Nick Roussopoulos. Balancing performance and data freshness in Web database servers. In Freytag et al. [1911], pages 393–404. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
l/Labrinidis:Alexandros.html.

Zhou:2003:BAM

[1793] Jingren Zhou and Kenneth A. Ross.

Buffering accesses to memory-resident index structures. In Freytag et al. [1911], pages 405–416. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/z/Zhou:Jingren.html.

Hankins:2003:DMA

[1794] Richard A. Hankins and Jignesh M.

Patel. Data morphing: An adaptive, cache-conscious storage technique. In Freytag et al. [1911], pages 417–428. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
h/Hankins:Richard_A=.html.

Hinneburg:2003:COD

[1795] Alexander Hinneburg, Wolfgang Lehner, and Dirk Habich. COMBI-operator: Database support for data mining applications. In Freytag et al. [1911], pages 429–439. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
h/Hinneburg:Alexander.html.

Shi:2003:SBA

[1796] Yong Shi, Yuqing Song, and Aidong Zhang. A shrinking-based approach for multi-dimensional data analysis. In Freytag et al. [1911], pages 440–451. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
s/Shi:Yong.html.

Zhou:2003:DBN

[1797] Jianjun Zhou and Jörg Sander. Data bubbles for non-vector data: Speedingup hierarchical clustering in arbitrary metric spaces. In Freytag et al. [1911], pages 452–463. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
z/Zhou:Jianjun.html.

Cormode:2003:FHH

[1798] Graham Cormode, Flip Korn, S. Muthukrishnan, and Divesh Srivastava. Finding hierarchical heavy hitters in data streams. In Freytag et al. [1911], pages 464–475. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
c/Cormode:Graham.html.

Xin:2003:SCC

[1799] Dong Xin, Jiawei Han, Xiaolei Li, and Benjamin W. Wah. Star-cubing: Computing Iceberg cubes by top-down and bottom-up integration. In Freytag et al. [1911], pages 476–487. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/x/Xin:Dong.html.

Kraft:2003:CGO

[1800] Tobias Kraft, Holger Schwarz, Ralf Rantzau, and Bernhard Mitschang. Coarse-grained optimization: Techniques for rewriting SQL statement sequences. In Freytag et al. [1911], pages 488–499. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
k/Kraft:Tobias.html.

Golab:2003:PSW

[1801] Lukasz Golab and M. Tamer Özsu.

Processing sliding window multijoins in continuous queries over data streams. In Freytag et al. [1911], pages 500–511. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
g/Golab:Lukasz.html.

Iwerks:2003:CKN

[1802] Glenn S. Iwerks, Hanan Samet, and Ken Smith. Continuous K-nearest neighbor queries for continuously moving points with updates. In Freytag et al. [1911], pages 512–523. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/i/Iwerks:Glenn_S=.html.

Grust:2003:SJT

[1803] Torsten Grust, Maurice van Keulen, and Jens Teubner. Staircase join: Teach a relational DBMS to watch its (axis) steps. In Freytag et al. [1911], pages 524–525. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
g/Grust:Torsten.html.

Korn:2003:CBM

[1804] Flip Korn, S. Muthukrishnan, and Yunyue Zhu. Checks and balances: Monitoring data quality problems in network traffic databases. In Freytag et al. [1911], pages 536–547. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/k/Korn:Flip.html.

Lubbers:2003:SDD

[1805] Dominik Lübbers, Udo Grimmer, and Matthias Jarke. Systematic development of data mining-based data quality tools. 548-559. In Freytag et al. [1911], pages 548–559. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/L=uuml=bbers:Dominik.html.

Papadimitriou:2003:AHS

[1806] Spiros Papadimitriou, Anthony Brockwell, and Christos Faloutsos. Adaptive, hands-off stream mining. In Freytag et al. [1911], pages 560–571. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/p/Papadimitriou:Spiros.html.

Madhavan:2003:CMA

[1807] Jayant Madhavan and Alon Y. Halevy. Composing mappings among data sources. In Freytag et al. [1911], pages 572–583. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        m/Madhavan:Jayant.html.

Velegrakis:2003:MAU

[1808] Yannis Velegrakis, Renée J. Miller, and Lucian Popa. Mapping adaptation under evolving schemas. In Freytag et al. [1911], pages 584–595. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/v/Velegrakis:Yannis.html.

Luo:2003:LPM

[1809] Gang Luo, Jeffrey F. Naughton, Curt Ellmann, and Michael Watzke. Locking protocols for materialized aggregate join views. In Freytag et al. [1911], pages 596–607. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
l/Luo:Gang.html.

Lee:2003:SFU

[1810] Mong-Li Lee, Wynne Hsu, Christian S.

Jensen, Bin Cui, and Keng Lik Teo. Supporting frequent updates in Rtrees: a bottom-up approach. In Freytag et al. [1911], pages 608–619. ISBN 0-12-722442-4. LCCN ???? URL http:

//www.vldb.org/dblp/db/indices/
a-tree/l/Lee:Mong=Li.html.

Qian:2003:NTD

[1811] Gang Qian, Qiang Zhu, Qiang Xue, and Sakti Pramanik. The ND-tree: a dynamic indexing technique for multidimensional non-ordered discrete data spaces. In Freytag et al. [1911], pages 620–631. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
q/Qian:Gang.html.

Gao:2003:TSE

[1812] Dengfeng Gao and Richard T. Snodgrass. Temporal slicing in the evaluation of XML queries. In Freytag et al. [1911], pages 632–643. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/g/Gao:Dengfeng.html.

Tsois:2003:GPG

[1813] Aris Tsois and Timos K. Sellis. The generalized pre-grouping transformation: Aggregate-query optimization in the presence of dependencies. In Freytag et al. [1911], pages 644–655. ISBN 0-12-722442-4. LCCN ???? URL http:

//www.vldb.org/dblp/db/indices/
a-tree/t/Tsois:Aris.html.

Helmer:2003:EOC

[1814] Sven Helmer, Thomas Neumann, and Guido Moerkotte. Estimating the output cardinality of partial preaggregation with a measure of clusteredness. In Freytag et al. [1911], pages 656– 667. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/h/Helmer:
Sven.html.

Brown:2003:BAD

[1815] Paul Brown and Peter J. Haas. BHUNT: Automatic discovery of fuzzy algebraic constraints in relational data. In Freytag et al. [1911], pages 668– 679. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/b/Brown:
Paul.html.

Yu:2003:TPR

[1816] Hailing Yu, Divyakant Agrawal, and Amr El Abbadi. Tabular placement of relational data on MEMS-based storage devices. In Freytag et al. [1911], pages 680–693. ISBN 0-12-722442-4.

LCCN ???? URL http://www.vldb.
org/dblp/db/indices/a-tree/y/Yu:
Hailing.html.

Anciaux:2003:MRQ

[1817] Nicolas Anciaux, Luc Bouganim, and Philippe Pucheral. Memory requirements for query execution in highly constrained devices. In Freytag et al. [1911], pages 694–705. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/a/Anciaux:Nicolas.html.

Schindler:2003:LRD

[1818] Jiri Schindler, Anastassia Ailamaki, and Gregory R. Ganger. Lachesis: Robust database storage management based on device-specific performance characteristics. In Freytag et al. [1911], pages 706–717. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
s/Schindler:Jiri.html.

Altinel:2003:CTP

[1819] Mehmet Altinel, Christof Bornhövd, Sailesh Krishnamurthy, C. Mohan, Hamid Pirahesh, and Berthold Reinwald. Cache tables: Paving the way for an adaptive database cache. In Freytag et al. [1911], pages 718–729. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/a/Altinel:Mehmet.html.

Chaudhuri:2003:PWS

[1820] Surajit Chaudhuri, Prasanna Ganesan, and Vivek R. Narasayya. Primitives for workload summarization and implications for SQL. In Freytag et al. [1911], pages 730–741. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/c/Chaudhuri:Surajit.html.

Vieira:2003:DBO

[1821] Marco Vieira and Henrique Madeira.

A dependability benchmark for OLTP application environments. In Freytag et al. [1911], pages 742–753. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/v/Vieira:Marco.html.

Ilyas:2003:STK

[1822] Ihab F. Ilyas, Walid G. Aref, and Ahmed K. Elmagarmid. Supporting top-k join queries in relational databases. In Freytag et al. [1911], pages 754–765. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
i/Ilyas:Ihab_F=.html.

Hulgeri:2003:AAN

[1823] Arvind Hulgeri and S. Sudarshan. AniPQO: Almost non-intrusive parametric query optimization for nonlinear cost functions. In Freytag et al. [1911], pages 766–777. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/h/Hulgeri:Arvind.html.

Guha:2003:EAO

[1824] Sudipto Guha, Dimitrios Gunopulos, Nick Koudas, Divesh Srivastava, and Michail Vlachos. Efficient approximation of optimization queries under parametric aggregation constraints. In Freytag et al. [1911], pages 778–789. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
g/Guha:Sudipto.html.

Tao:2003:TTO

[1825] Yufei Tao, Dimitris Papadias, and Jimeng Sun. The TPR*-tree: An optimized spatio-temporal access method for predictive queries. In Freytag et al. [1911], pages 790–801. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/t/Tao:Yufei.html.

Papadias:2003:QPS

[1826] Dimitris Papadias, Jun Zhang, Nikos Mamoulis, and Yufei Tao. Query processing in spatial network databases. In Freytag et al. [1911], pages 802–813. ISBN 0-12-722442-4.

LCCN ????  URL http://www.
vldb.org/dblp/db/indices/a-tree/
p/Papadias:Dimitris.html.

Lin:2003:MHS

[1827] Xuemin Lin, Qing Liu, Yidong Yuan, and Xiaofang Zhou. Multiscale histograms: Summarizing topological relations in large spatial datasets. In Freytag et al. [1911], pages 814–825. ISBN 0-12-722442-4. LCCN ????

URL http://www.vldb.org/dblp/
db/indices/a-tree/l/Lin:Xuemin.
html.

Wang:2003:AOG

[1828] Xiaoyu Wang and Mitch Cherniack.

Avoiding ordering and grouping in query processing. In Freytag et al. [1911], pages 826–837. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/w/Wang:Xiaoyu.html.

Pottinger:2003:MMB

[1829] Rachel Pottinger and Philip A. Bernstein. Merging models based on given correspondences. In Freytag et al. [1911], pages 826–873. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/p/Pottinger:Rachel.html.

Carney:2003:OSD

[1830] Donald Carney, Ugur Çetintemel, Alex Rasin, Stanley B. Zdonik, Mitch Cherniack, and Michael Stonebraker. Operator scheduling in a data stream manager. In Freytag et al. [1911], pages 838–849. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/c/Carney:
Donald.html.

Hristidis:2003:EIS

[1831] Vagelis Hristidis, Luis Gravano, and Yannis Papakonstantinou. Efficient IRstyle keyword search over relational databases. In Freytag et al. [1911], pages 850–861. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
h/Hristidis:Vagelis.html.

Galanis:2003:LDS

[1832] Leonidas Galanis, Yuan Wang, Shawn R. Jeffery, and David J. DeWitt. Locating data sources in large distributed systems. In Freytag et al. [1911], pages 874–885. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
g/Galanis:Leonidas.html.

Jermaine:2003:RES

[1833] Chris Jermaine. Robust estimation with sampling and approximate preaggregation. In Freytag et al. [1911], pages 886–897. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
j/Jermaine:Chris.html.

Miklau:2003:CAP

[1834] Gerome Miklau and Dan Suciu. Controlling access to published data using cryptography. In Freytag et al. [1911], pages 898–909. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/m/Miklau:Gerome.html.

Meek:2003:OOA

[1835] Colin Meek, Jignesh M. Patel, and Shruti Kasetty. OASIS: An online and accurate technique for localalignment searches on biological sequences. In Freytag et al. [1911], pages 910–921. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
m/Meek:Colin.html.

Bawa:2003:PPI

[1836] Mayank Bawa, Roberto J. Bayardo, Jr., and Rakesh Agrawal. Privacypreserving indexing of documents on the network. In Freytag et al. [1911], pages 922–933. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
b/Bawa:Mayank.html.

Poss:2003:DCO

[1837] Meikel Pöss and Dmitry Potapov. Data compression in oracle. In Freytag et al. [1911], pages 937–947. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/p/P=ouml=ss:Meikel.html.

An:2003:IPB

[1838] Ning An, Kothuri Venkata Ravi Kanth, and Siva Ravada. Improving performance with bulk-inserts in Oracle R-trees. In Freytag et al. [1911], pages 948–951. ISBN 0-12-722442-4.

LCCN ???? URL http://www.vldb.
org/dblp/db/indices/a-tree/a/An:
Ning.html.

Galindo-Legaria:2003:SV

[1839] César A. Galindo-Legaria, Milind Joshi, Florian Waas, and Ming-Chuan Wu. Statistics on views. In Freytag et al. [1911], pages 952–962. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/g/Galindo=Legaria:C=eacute=
sar_A=.html.

Bhattacharjee:2003:EQP

[1840] Bishwaranjan Bhattacharjee, Sriram Padmanabhan, Timothy Malkemus, Tony Lai, Leslie Cranston, and Matthew Huras. Efficient query processing for multi-dimensionally clustered tables in DB2. In Freytag et al. [1911], pages 963–974. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/b/Bhattacharjee:Bishwaranjan.
html.

Srinivasa:2003:PBM

[1841] Srinath Srinivasa and Sujit Kumar. A platform based on the multidimensional data model for analysis of bio-molecular structures. In Freytag et al. [1911], pages 975–986. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/s/Srinivasa:Srinath.html.

Liu:2003:CGT

[1842] Chengfei Liu, Bruce G. Lindsay, Serge Bourbonnais, Elizabeth Hamel, Tuong C. Truong, and Jens Stankiewitz. Capturing global transactions from multiple recovery log files in a partitioned database system. In Freytag et al. [1911], pages 987–996. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/l/Liu:Chengfei.html.

Florescu:2003:BXS

[1843] Daniela Florescu, Chris Hillery, Donald Kossmann, Paul Lucas, Fabio Riccardi, Till Westmann, Michael J. Carey, Arvind Sundararajan, and Geetika Agrawal. The BEA/XQRL streaming XQuery processor. In Freytag et al. [1911], pages 997–1008. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/f/Florescu:Daniela.html.

Murthy:2003:XSO

[1844] Ravi Murthy and Sandeepan Banerjee. XML schemas in oracle XML DB. In Freytag et al. [1911], pages 1009– 1018. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/m/Murthy:
Ravi.html.

Hage:2003:IDM

[1845] Christian Hage, Christian S. Jensen, Torben Bach Pedersen, Laurynas Speicys, and Igor Timko. Integrated data management for mobile services in the real world. In Freytag et al. [1911], pages 1019–1030. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
h/Hage:Christian.html.

Cabrera:2003:WSI

[1846] Felipe Cabrera. Web services (industrial session). In Freytag et al. [1911], pages 1031–1032. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
c/Cabrera:Felipe.html.

Leymann:2003:GAI

[1847] Frank Leymann. Grid and applications (industrial session). In Freytag et al. [1911], page 1033. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/Leymann:Frank.html.

Schoning:2003:CUD

[1848] Harald Schöning. Commercial use of database technology. In Freytag et al. [1911], page 1034. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        s/Sch=ouml=ning:Harald.html.

Kuhn:2003:ZDD

[1849] eva Kühn. The zero-delay data warehouse: Mobilizing heterogeneous databases. In Freytag et al. [1911], pages 1035–1040. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        k/K=uuml=hn:eva.html.

Kersten:2003:DSH

[1850] Martin L. Kersten, Gerhard Weikum, Michael J. Franklin, Daniel A. Keim, Alejandro P. Buchmann, and Surajit Chaudhuri. A database striptease or how to manage your personal databases. In Freytag et al. [1911], pages 1043–1044. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        k/Kersten:Martin_L=.html.

Cluet:2003:WNX

[1851] Sophie Cluet. Who needs XML databases? In Freytag et al. [1911], page 1045. ISBN 0-12-722442-4. LCCN

                                        ???? URL http://www.vldb.org/

dblp/db/indices/a-tree/c/Cluet:
Sophie.html.

Brodie:2003:IDS

[1852] Michael L. Brodie. Illuminating the dark side of Web services. In Freytag et al. [1911], pages 1046–1049. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/b/Brodie:Michael_L=.html.

Berger:2003:XVP

[1853] Sacha Berger, François Bry, Sebastian Schaffert, and Christoph Wieser. Xcerpt and visXcerpt: From patternbased to visual querying of XML and semistructured data. In Freytag et al. [1911], pages 1053–1056. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/b/Berger:Sacha.html.

Meng:2003:OSB

[1854] Xiaofeng Meng, Daofeng Luo, Mong-Li Lee, and Jing An. OrientStore: a schema based native XML storage system. In Freytag et al. [1911], pages 1057–1060. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
m/Meng:Xiaofeng.html.

Abiteboul:2003:MDW

[1855] Serge Abiteboul, Jérôme Baumgarten, Angela Bonifati, Gregory Cobena, Cosmin Cremarenco, Florin Dragan, Ioana Manolescu, Tova Milo, and Nicoleta Preda. Managing distributed workspaces with Active XML. In Freytag et al. [1911], pages 1061–1064. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        a/Abiteboul:Serge.html.

Arion:2003:XPQ

[1856] Andrei Arion, Angela Bonifati, Gianni Costa, Sandra D’Aguanno, Ioana Manolescu, and Andrea Pugliese. Xquec: Pushing queries to compressed XML data. In Freytag et al. [1911], pages 1065–1068. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        a/Arion:Andrei.html.

Balmin:2003:SKP

[1857] Andrey Balmin, Vagelis Hristidis, Nick Koudas, Yannis Papakonstantinou, Divesh Srivastava, and Tianqiu Wang. A system for keyword proximity search on XML databases. In Freytag et al. [1911], pages 1069–1072. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/b/Balmin:Andrey.html.

Harding:2003:XRX

[1858] Philip J. Harding, Quanzhong Li, and Bongki Moon. XISS/R: XML indexing and storage system using RDBMS. In Freytag et al. [1911], pages 1073–1076. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        h/Harding:Philip_J=.html.

Fernandez:2003:IXG

[1859] Mary F. Fernandez, Jérôme Siméon, Byron Choi, Amélie Marian, and Gargi Sur. Implementing Xquery 1.0: The Galax experience. In Freytag et al. [1911], pages 1077–1080. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/f/Fernandez:Mary_F=.html.

Weber:2003:WCO

[1860] Roger Weber, Christoph Schuler, Patrick Neukomm, Heiko Schuldt, and Hans-Jörg Schek. WebService composition with O’GRAPE and OSIRIS. In Freytag et al. [1911], pages 1081– 1084. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/w/Weber:
Roger.html.

Torlone:2003:CEC

[1861] Riccardo Torlone and Paolo Atzeni.

Chameleon: an extensible and customizable tool for Web data translation. In Freytag et al. [1911], pages 1085–1088. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
t/Torlone:Riccardo.html.

Nicklas:2003:NAL

[1862] Daniela Nicklas, Matthias Großmann, and Thomas Schwarz. NexusScout: An advanced location-based application on a distributed, open mediation platform. In Freytag et al. [1911], pages 1089–1092. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
n/Nicklas:Daniela.html.

Abiteboul:2003:SDC

[1863] Serge Abiteboul, Bernd Amann, Jérôme Baumgarten, Omar Benjelloun, Frederic Dang Ngoc, and Tova Milo. Schema-driven customization of Web services. In Freytag et al. [1911], pages 1093–1096. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        a/Abiteboul:Serge.html.

Nie:2003:BSE

[1864] Zaiqing Nie, Subbarao Kambhampati, and Thomas Hernandez. BibFinder/ StatMiner: Effectively mining and using coverage and overlap statistics in data integration. In Freytag et al. [1911], pages 1097–1100. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/n/Nie:Zaiqing.html.

Petrovic:2003:TST

[1865] Milenko Petrovic, Ioana Burcea, and Hans-Arno Jacobsen. S-ToPSS: Semantic Toronto Publish/Subscribe System. In Freytag et al. [1911], pages 1101–1104. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        p/Petrovic:Milenko.html.

Sizov:2003:FCE

[1866] Sergej Sizov, Jens Graupmann, and Martin Theobald. From focused crawling to expert information: an application framework for Web exploration and portal generation. In Freytag et al. [1911], pages 1105–1108. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/s/Sizov:Sergej.html.

Li:2003:CIA

[1867] Wen-Syan Li, Oliver Po, Wang-Pin Hsiung, K. Selçuk Candan, Divyakant Agrawal, Yusuf Akca, and Kunihiro Taniguchi. CachePortal II: Acceleration of very large scale data centerhosted database-driven Web applications. In Freytag et al. [1911], pages 1109–1112. ISBN 0-12-722442-4.

LCCN ???? URL http://www.vldb.
org/dblp/db/indices/a-tree/l/Li:
Wen=Syan.html.

Wang:2003:ASC

[1868] Haixun Wang, Carlo Zaniolo, and Chang Richard Luo. ATLAS: a small but complete SQL extension for data mining and data streams. In Freytag et al. [1911], pages 1113–1116. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/w/Wang:Haixun.html.

Witkowski:2003:BMU

[1869] Andrew Witkowski, Srikanth Bellamkonda, Tolga Bozkaya, Nathan Folkert, Abhinav Gupta, Lei Sheng, and Sankar Subramanian. Business modelling using SQL spreadsheets. In Freytag et al. [1911], pages 1117–1120. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
w/Witkowski:Andrew.html.

Wang:2003:IVD

[1870] Tianqiu Wang, Simone Santini, and Amarnath Gupta. An interpolated volume data model. In Freytag et al. [1911], pages 1121–1124. ISBN 0-12-722442-4. LCCN ???? URL http:/

/www.vldb.org/dblp/db/indices/a-
tree/w/Wang:Tianqiu.html.

Lakshmanan:2003:EDC

[1871] Laks V. S. Lakshmanan, Jian Pei, and Yan Zhao. Efficacious data cube exploration by semantic summarization and compression. In Freytag et al. [1911], pages 1125–1128. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        l/Lakshmanan:Laks_V=_S=.html.

Sattler:2003:QCQ

[1872] Kai-Uwe Sattler, Ingolf Geist, and Eike Schallehn. QUIET: Continuous querydriven index tuning. In Freytag et al. [1911], pages 1129–1132. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/s/Sattler:Kai=Uwe.html.

Bouganim:2003:CSD

[1873] Luc Bouganim, François Dang Ngoc, Philippe Pucheral, and Lilan Wu. Chip-secured data access: Reconciling access rights with data encryption. In Freytag et al. [1911], pages 1133–1136. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        b/Bouganim:Luc.html.

Nath:2003:IAI

[1874] Suman Nath, Amol Deshpande, Yan Ke, Phillip B. Gibbons, Brad Karp, and Srinivasan Seshan. IrisNet: An architecture for Internet-scale sensing services. In Freytag et al. [1911], pages 1137–1140. ISBN 0-12-722442-4. LCCN ???? URL http://www.

                                        vldb.org/dblp/db/indices/a-tree/
                                        n/Nath:Suman.html.

Baumann:2003:LSS

[1875] Peter Baumann. Large-scale, standardsbased Earth observation imagery and Web mapping services. In Freytag et al. [1911], pages 1141–1144. ISBN 0-12-722442-4. LCCN ???? URL http:/

                                        /www.vldb.org/dblp/db/indices/a-
                                        tree/b/Baumann:Peter.html.

Clifton:2003:PED

[1876] Chris Clifton, Irini Fundulaki, Richard Hull, Bharat Kumar, Daniel F. Lieuwen, and Arnaud Sahuguet. Privacy-enhanced data management for next-generation e-commerce. In Freytag et al. [1911], pages 1147– 1148. ISBN 0-12-722442-4. LCCN ????

URL http://www.vldb.org/dblp/
db/indices/a-tree/c/Clifton:Chris.
html.

Decker:2003:SWS

[1877] Stefan Decker and Vipul Kashyap. The Semantic Web: Semantics for data on the Web. In Freytag et al. [1911], page 1148. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/d/Decker:
Stefan.html.

Koudas:2003:DSQ

[1878] Nick Koudas and Divesh Srivastava.

Data stream query processing: a tutorial. In Freytag et al. [1911], page 1149. ISBN 0-12-722442-4. LCCN

???? URL http://www.vldb.org/
dblp/db/indices/a-tree/k/Koudas:
Nick.html.

Jagatheesan:2003:GDM

[1879] Arun Jagatheesan, Reagan Moore, Norman W. Paton, and Paul Watson. Grid data management systems & services. In Freytag et al. [1911], page 1150. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
j/Jagatheesan:Arun.html.

Ceri:2003:CID

[1880] Stefano Ceri and Ioana Manolescu.

Constructing and integrating datacentric Web applications: Methods, tools, and techniques. In Freytag et al. [1911], page 1151. ISBN 0-12-722442-4. LCCN ???? URL http://www.

vldb.org/dblp/db/indices/a-tree/
c/Ceri:Stefano.html.

Kerr:1975:PIC

[1881] Douglas S. Kerr, editor. Proceedings of the International Conference on Very Large Data Bases, Framingham, MA, USA, September 22–24, 1975. ACM Press, New York, NY 10036, USA, 1975. ISSN 0278-2596. LCCN QA76.9.D3 I55 1975. US$15.00. ACM SIGMOD v. 1, no. 1, September 1975.

Lockemann:1976:SLD

[1882] Peter C. Lockemann and Erich J. Neuhold, editors. Systems for large data bases: proceedings of the 2nd International Conference on Very Large Date [sic] Bases (Brussels, Belgium, 1976). North-Holland Publishing Co., Amsterdam, The Netherlands, 1976. ISBN 0-7204-0463-0, 0-7204-0546-7. LCCN QA76.9.D3 I56 1976. FL60.00.

ACM:1977:VLD

[1883] ACM, editor. Very large data bases: proceedings, Third International Conference on Very Large Data Bases, Tokyo, Japan, October 6–8, 1977, SIGMOD record; v. 9, no. 4. ACM Press, New York, NY 10036, USA, 1977. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I55 1977.

Yao:1978:VLD

[1884] S. Bing Yao, editor. Very large data bases: Fourth International Conference on Very Large Data Bases, West Berlin, Germany, September 13–15, 1978. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. ISBN ???? LCCN QA76.9.D3 I55 1978. US$25.00.

Furtado:1979:VLD

[1885] Antonio L. Furtado and Howard L.

Morgan, editors. Very large data bases: Fifth International Conference on Very Large Data Bases, Rio de Janeiro, Brazil, October 3–5, 1979. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979. LCCN QA76.9.D3 I55 1979.

Weber:1979:IDB

[1886] Herbert Weber and Anthony I. Wasserman, editors. Issues in data base management: proceedings of the Fourth International Conference on Very Large Data Bases, September 13–15, 1978, West-Berlin, Germany. Elsevier North-Holland, Amsterdam, 1979. ISBN 0-444-85316-2. LCCN QA76.9 .D3 I61 1978.

CIPS:1980:VLD

[1887] Canadian Information Processing Society, editor. Very large data bases: Sixth International Conference on Very Large Data Bases (Montréal, PQ, Canada). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. LCCN QA76.9.D3 .I559 1980. Sponsored by IEEE Computer Society, Association for Computing Machinery, and Canadian Information Processing Society. IEEE Catalog Number 80CH1534-7C. ACM Order Number 471800. Long Beach Order Number 322.

Zaniolo:1981:VLD

[1888] Carlo Zaniolo and C. Delobel, editors. Very large data bases: proceedings: seventh International Conference on Very Large Data Bases, Cannes, France, September 9–11, 1981. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA76.9.D3 I559 1981. IEEE catalog no. 81CH1701-2.

VLDB:1982:VLD

[1889] VLDB Endowment, editor. Very large data bases: eighth International Conference on Very Large Data Bases, Mexico City, Mexico, September 8–10, 1982. VLDB Endowment, P.O. Box 2245, Saratoga, CA, USA, 1982. ISBN 0-934613-14-1. LCCN QA76.9.D3 I559 1982.

Schkolnick:1983:ICV

[1890] Mario Schkolnick and C. (Costantino) Thanos, editors. 9th International Conference on Very Large Data Bases (Florence, Italy, October 31–November 2, 1983). VLDB Endowment, P.O. Box 2245, Saratoga, CA, USA, 1983. ISBN 0-934613-15-X. LCCN QA 76.9 D3 I61 1983. This conference is sponsored by VLDB Endowment and co-sponsored by IFIP et al.

Dayal:1984:VLD

[1891] Umeshwar Dayal, Gunter Schlageter, and Huat Seng Lim, editors. Very large data bases: proceedings tenth International Conference on Very Large Data Bases, Singapore, August 27–31, 1984. VLDB Endowment, P.O. Box 2245, Saratoga, CA, USA, 1984. ISBN 0-934613-16-8. LCCN QA76.9.D3 I559 1984.

Pirotte:1985:VLD

[1892] Alain Pirotte and Yannis Vassiliou, editors. Very large data bases, Stockholm 1985: 11th International Conference on Very Large Data Bases, Stockholm, August 21–23, 1985. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1985. ISBN 0-934613-17-6. LCCN QA 76.9 D3 I61 1985.

Kambayashi:1986:TIC

[1893] Yahiko Kambayashi, Wesley Chu, Georges Gardarin, and Setsuo Ohsuga, editors. Twelfth international conference on very large data bases, proceedings (VLDB ’86). Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1986. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

Stocker:1987:VLD

[1894] Peter M. Stocker, William Kent, and Peter Hammersley, editors. Very large data bases: 1987, 13th VLDB, Brighton, England, September 1–4: proceedings of the Thirteenth International Conference on Very Large Data Bases, Brighton, England, 1987. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1987. ISBN 0-934613-46-X. LCCN QA 76.9 D3 I61 1987.

Bancilhon:1988:VLD

[1895] François Bancilhon and David J. De-Witt, editors. Very large data bases: 1988, 14th VLDB, Los Angeles, USA, August 29–September 1: proceedings of the Fourteenth International Conference on Very Large Data Bases. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1988. ISBN 0-934613-75-3 (paperback). LCCN QA 76.9 D3 I61 1988. Sponsored by VLDB Endowment, ACM-SIGMOD, and the Computer Society of the IEEE.

Apers:1989:VLD

[1896] P. M. G. (Petrus Maria Gerardus) Apers and Gio Wiederhold, editors. Very large data bases: proceedings: proceedings of the Fifteenth International Conference on Very Large Data Bases, August 22–25, 1989, Amsterdam, The Netherlands. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1989. ISBN 1-55860-101-5. LCCN QA 76.9 D3 I61 1989.

McLeod:1990:VLD

[1897] Dennis McLeod, Ron Sacks-Davis, and H.-J Schek, editors. Very large data bases: 16th International Conference on Very Large Data Bases; August 13– 16, 1990, Brisbane, Australia. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1990. ISBN 1-55860-149-X. LCCN QA76.9.D3I559 1990.

Camps:1991:PSI

[1898] Rafael Camps, Guy M. Lohman, and Amílcar Sernadas, editors. Proceedings of the seventeenth International Conference on Very Large Data Bases, September 3–6, 1991, Barcelona (Catalonia, Spain). Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1991. ISBN 1-55860-150-3. LCCN QA 76.9 D3 I61 1991.

Yuan:1992:VLD

[1899] Li-Yan Yuan, editor. Very large data bases: VLDB ’92, proceedings of the 18th International Conference on Very Large Data Bases, August 23– 27, 1992, Vancouver, Canada. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1992. ISBN 1-55860-151-1. LCCN QA76.9.D3 I61 1992.

Agrawal:1993:VLD

[1900] Rakesh Agrawal, Sean Baker, and David Bell, editors. Very large data bases, VLDB ’93: proceedings of the 19th International Conference on Very Large Data Bases, August 24–27, 1993, Dublin, Ireland. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1993. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Bocca:1994:ICV

[1901] Jorge B. Bocca, Matthias Jarke, and Carlo Zaniolo, editors. 20th International Conference on Very Large Data Bases, September 12–15, 1994, Santiago, Chile proceedings. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

Dayal:1995:VPI

[1902] Umeshwar Dayal, Peter M. D. Gray, and Shojiro Nishio, editors. VLDB ’95: proceedings of the 21st International Conference on Very Large Data Bases, Zurich, Switzerland, Sept. 11–15, 1995. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1995. ISBN 1-55860-379-4. LCCN QA76.9.D3 I61

Svensson:1996:PEI

[1903] P. (Per) Svensson and J. C. (James Cornelius) French, editors. Proceedings: Eighth International Conference on Scientific and Statistical Database Systems, June 18–20, 1996, Stockholm, Sweden. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7264-1, 0-8186-7266-8. LCCN QA 76.9 D3 I576 1996. IEEE catalog number 96TB100051.

Vijayaraman:1996:PTS

[1904] T. M. Vijayaraman, Alejandro P. Buchmann, C. Mohan, and Nandlal L. Sarda, editors. Proceedings of the twenty-second international Conference on Very Large Data Bases, September 3–6, 1996, Mumbai (Bombay), India. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

Jarke:1997:PTT

[1905] Matthias Jarke, Michael J. Carey, Klaus R. Dittrich, Frederick H. Lochovsky, Pericles Loucopoulos, and Manfred A. Jeusfeld, editors. Proceedings of the Twenty-third International Conference on Very Large Data Bases, Athens, Greece, 26–29 August 1997. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1997. ISBN 1-55860-470-7. LCCN QA76.9 D3 I61 1997. Also known as VLDB’97.

Gupta:1998:PTF

[1906] Ashish Gupta, Oded Shmueli, and Jennifer Widom, editors. Proceedings of the Twenty-fourth International Conference on Very Large Databases, New York, NY, USA, 24–27 August, 1998. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-566-5. LCCN QA76.9.D3 I559 1998. Also known as VLDB.

Atkinson:1999:PTF

[1907] Malcolm P. Atkinson, Maria E. Orlowska, Patrick Valduriez, Stanley B. Zdonik, and Michael L. Brodie, editors. Proceedings of the Twentyfifth International Conference on Very Large Databases, Edinburgh, Scotland, UK, 7–10 September, 1999. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. Also known as VLDB’99.

ElAbbadi:2000:VPI

[1908] Amr El Abbadi, Michael L. Brodie, Sharma Chakravarthy, Umeshwar Dayal, Nabil Kamel, Gunter Schlageter, and Kyu-Young Whang, editors. VLDB 2000, Proceedings of 26th International Conference on Very Large Data Bases, September 10–14, 2000, Cairo, Egypt. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-715-3. LCCN ????

Apers:2001:PTS

[1909] Peter M. G. Apers, Paolo Atzeni, Stefano Ceri, Stefano Paraboschi, Kotagiri Ramamohanarao, and Richard T. Snodgrass, editors. Proceedings of the Twenty-seventh International Conference on Very Large Data Bases: Roma, Italy, 11–14th September, 2001. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-804-4. LCCN QA76.9.D3 I559 2001.

Bernstein:2002:VPT

[1910] Philip A. Bernstein et al., editors. VLDB 2002: proceedings of the Twenty-Eighth International Conference on Very Large Data Bases, Hong Kong SAR, China, 20–23 August 2002. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-869-9. LCCN ????

Freytag:2003:VPI

[1911] Johann Christoph Freytag, Peter C. Lockemann, Serge Abiteboul, Michael J. Carey, Patricia G. Selinger, and Andreas Heuer, editors. VLDB 2003: Proceedings of 29th International Conference on Very Large Data Bases, September 9–12, 2003, Berlin, Germany. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-12-722442-4. LCCN ????